A Complete Bibliography of Publications in Fisheries Oceanography

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

27 December 2023 Version 1.15

Title word cross-reference

1 [CW98, ODMRM98]. 3 [EHW08, PJD14]. ^{13} [WP93]. ^{137} [MFS^{+}17]. ^{15} [WP93]. ^{90} [MFS^{+}17]. ^{∘} [Jes22]. _{2} [HLH^{+}17, KTO^{+}11]. β [LPCG23]. : [FKUY16, YOY00]. δ [WP93]. $US [Gre99].

-D [ODMRM98]. -diversity [LPCG23].

0-12-487570-X [Gre99]. 0-group [KSAF13]. 06 [Aut08].

120^{o} [KEJK00]. 1980s [DHMT96]. 1990s [DHMT96, ZHL^{+}03]. 1996 [BBS99]. 1997 [CP03]. 1999 [REM02].

20 [Jes22]. 2000 [CP03]. 2009 [JMP^{+}14]. 2011 [KKK^{+}17, MTT^{+}17, OKU17]. 2012/2013 [óóSV18]. 20th [SLM13, SB04]. 21st [BEiI^{+}23]. 22^{o} [CG18]. 25^{o} [CG18].

30th [Kim23]. 30 [BEiI^{+}23]. 32-year [CDG^{+}19].

60^{o} [KEJK00].

abalone [KTO^{+}11, TWK13, TKW^{+}17]. Abiotic [FYK^{+}13, CDG^{+}19, HVHC10, KSAF13, REG^{+}13]. Abukuma [SAO^{+}17]. Abundance [LSW^{+}03, AOVAG22, BJV^{+}17, Bea03, BHM02, BT99, BWS^{+}01, CSFC05, CP92, CP03, Coy05, DHC^{+}07, DP01, DHMT96, GTB10, GDM^{+}17, GVRC04, GEGHPCC17, HJ99, HEG08, HCWF21, HCC^{+}09, JCH05, JHK^{+}15, JCCB15, LCCdS^{+}19, LYT^{+}20, LC95, LP10, LéEPW^{+}12, LBSS^{+}92, LS15, LA05, MESMM18, MDKS93, MFH05, MLRS07, MSC^{+}17, MTLL^{+}16, MRHL09, MWR^{+}98, NHM94, Oda94, OFS^{+}16, PP01, PLSO98, PDD03, Pol96, RSF13, RAT^{+}02, SRR07, SHG^{+}22, SSW^{+}17, SGN^{+}05, SCKJ^{+}18, SFL16, SNV^{+}12, SNL19, TID^{+}96, TAN^{+}17b, TBB^{+}03, TCC^{+}98, TTH15, UTMS06, VCB^{+}98, VHLM15, WK03, WSC05, YWM^{+}00, YOIW21, YLA13, éSMB20]. abundance-biomass [GEGHPCC17]. abundances [JYH^{+}18, RS92]. Academic [Gre99]. acanthias [SPM02]. accident [MFS^{+}17]. Accuracy [PSC05, WSP^{+}07, BFF15]. accurately [WM06]. Acknowledgements [Ano95a]. Acoustic [AI92, Hor00, MAS^{+}98, MIY^{+}09, BH97, GCF^{+}21, HHK^{+}10, RMM02]. Acoustical [Gre99]. across [AM18, GS99, HGG^{+}17, KBB^{+}20, KEWDA18, LSW^{+}03, MTZG23, RKD^{+}20, SGW^{+}21, SFL16, TNK^{+}16]. actions [JPHA^{+}16]. active [KSY^{+}23]. activities [WLWZ98]. activity [FRS^{+}05, HSLP19, HTP14, MFS^{+}17, PVBV19, SAT^{+}18]. acutorostrata [MTK^{+}07]. adaptation [JPHA^{+}16, SMS^{+}23]. Adapting [OTIK20]. ADCP [TKH08]. Additive [HHF09, MTP07, FODCN00, YOK^{+}17]. address [JPHA^{+}16]. adjacent [DWH11, LLCJ16, MBH^{+}99, NSGL^{+}22, TCS^{+}09, XWL^{+}23]. adjoint [MLM^{+}98]. adjust [Jes22]. Adriatic [CLM^{+}21, CMB^{+}15, DG00, VZP98, ZVKŠ13]. adult [BYM16, FKUY16, FKSA21, RWLP12, SKHN11, SSR13, Tan17a, WTK^{+}16, WSF^{+}14, ZSY^{+}21]. Advection [SSP^{+}07, ÅGN^{+}04, ASK99, BHH98, DPL^{+}20, Dd95, ESTJ03, ETB^{+}17, MAHG94, MGHS14, WPL^{+}93]. advective [BSF01a, GP94, HBO^{+}01]. aeglefinus [BCL04, HG98, LOS^{+}14, LSK^{+}18]. aestivation [TY04]. affected [YCS^{+}19]. affecting [FYK^{+}13, HQH^{+}06, INM^{+}18, LAG^{+}11, NKS00, OWK04, Spe08]. affects [VCKH05]. affiliation [SWAAB20]. Africa [BJV^{+}17, DBRSC16, JHC^{+}15, MRL^{+}14, MHM^{+}20, SGFR^{+}21, TAN^{+}17b, VCB^{+}98]. African [LéEPW^{+}12, LRBJ21]. after [KKK^{+}17, KYSM11, MFS^{+}17, NSH^{+}17, OK17]. Age [HHK^{+}10, BMH^{+}21, FYA^{+}21, FFF^{+}18, HFF^{+}19, HAS^{+}19, IFF^{+}18, MSS12, OTIK20, OH23, SYT^{+}09, SSW^{+}17, SADA^{+}23, TMMM20, TY04, WSC05, XDP^{+}20, YCH^{+}15]. age-0 [FFF^{+}18, HFF^{+}19, IFF^{+}18, MSS12, SYT^{+}09, SADA^{+}23, TMMM20, WSC05]. age-1 [YCH^{+}15]. Age-dependent [HHK^{+}10]. age-specific [FYA^{+}21]. age-structured [SSW^{+}17]. ages [Jes22]. aggregating [DBFW13, GCF^{+}21, GAH^{+}19]. aggregation [GSBB07]. aggregations [CLKP19, OE17, VPRG13]. Agulhas [VCB^{+}98]. al [Sim96]. alalunga [AAKMG06, BML11, CLT05, CSK11, DSPH07, Dom09, KNS97, NPS^{+}23, SA10, ZSS08]. Alaska [LA05, WGS^{+}08, APL^{+}96, APL^{+}08, ADAHL10, BBMY93, BPZR19, BG01, BWKM15, BT99, CAB^{+}01, CCSS01, CP03, ECM^{+}01, FYA^{+}21, GV01, HAS^{+}19, KNE^{+}04, KPHG14, LK21, LDAWM10, MSS12, MWGK92, MM03, NBF^{+}01, RBBG12, RFM^{+}21, RTK01, RKZHC19, SGW^{+}21, SMF^{+}05, TGRS^{+}19, TMM^{+}07, VIS92, WJP^{+}01, WS08, WCP^{+}01, YCS^{+}19, ZP21a]. Alaskan [CL05, CP92, NBH99, RZM^{+}03]. albacares [BCR20, DWH11, GCF^{+}21, MSST16, Nis92, NdLOO23, SFA14, SF22, SZX^{+}08]. Albacore [NPS^{+}23, ZSS08, AAKMG06, BML11, CLT05, CSK11, CH16, CGI^{+}19, DSPH07, Dom09, Gla11, KNS97, SA10, ZHT14, ZHX^{+}20]. albatross [MJH14]. albatrosses [HKA^{+}06, XTC^{+}04]. albidus [HKLG07]. Alboran [BGM^{+}18, VYGT^{+}20]. Aleutian [BRO18, BRR05, CCL^{+}05, Coy05, aTCK05, FRS^{+}05, HWS^{+}05, HS05, JCH05, LJH^{+}05, LHM^{+}05, LAB^{+}05, MSL^{+}05, ROB05, SMF^{+}05, SCDA10, SPV96, SHM05, SKKS05, ZP21b]. alfredi [AAG11]. along [BPLC11, BUE02, FKH^{+}17, FRHMAM^{+}06, GNP^{+}19, HA07, HT99, HONH04, IWK^{+}21, JHC^{+}15, KFS22, KN08, KSC^{+}10, KBS^{+}16, KMM^{+}06, LPCA15, LJBR20, LRBJ21, MBY^{+}18, MSL^{+}05, Mor11, MSVY^{+}13, NYI11, PDER10, PKP^{+}00, SSP^{+}07, SME^{+}14, SS19, TSK^{+}92, Tan99, Tan02, TKM^{+}22, TDE09, UIU^{+}99, WTK^{+}16, WZK^{+}98, WKN^{+}95]. Alopias [HRB^{+}18]. Alosa [LAFF15]. alpinus [RDE^{+}07]. Alternating [NFN00]. alternations [NTIO18]. Alternative [APL^{+}96, SP93]. alternatives [CLKP19]. alters [LéEPW^{+}12]. alutus [KPHG14, RBBG12, Sco95]. Amazon [JMP^{+}14]. amberjack [TNC^{+}22]. ambient [III^{+}06, WJT97, ZHX^{+}20]. Amblyraja [GHM21, SB06]. America [HFC01, PS06]. American [DDS^{+}17, BMOT17, CCC^{+}23, CHM^{+}94, DSPH07, Dom09, DHMT96, DTC06, MFMG20, PTS^{+}24, QCR22, SCTB19]. americana [NH06]. americanus [AOVAG22, BMOT17, DTC06, HDH^{+}05, IN00, IXW^{+}10, PWML12, SCTB19]. americanust [DHMT96]. Ammodytes [KKNY92, NNOU20, TY04]. among [BDVS^{+}19, CHF^{+}04, DAW^{+}23, ERR^{+}21, LPH^{+}19, NH01, PEKL14, QLB^{+}05, RAT^{+}02, Rog94, RS92, WQI00]. amphipod [VPRG13]. Amundsen [KEJK00]. anadromous [AHAM03]. analyses [DDS^{+}17, HCC^{+}09, KM93, áRÁSG^{+}16, YAM^{+}18]. Analysis [GPS22, BHV^{+}06, BM99a, BSF^{+}20, BEF^{+}12, CPM^{+}15, DWHdP21, FPBDC11, HHK^{+}17, HP02, HPG^{+}20, HHH^{+}18, IMO^{+}12, KKNY92, LRBJ21, MMBC07, MMMS14, OK17, PHH13, PCR^{+}18, SB94, SMB03b, TCS^{+}09, VIS92, YOK^{+}17]. anchoita [DBS^{+}19, HMM01, LC95, MSM^{+}13]. Anchoveta [GNP^{+}19, RPG^{+}22, CRVL^{+}17]. Anchovy [CDG^{+}19, GSBB07, RR18, AB02, ACT^{+}10, ACG^{+}16, APL01, APGL03, APLG07, APL07, BH97, BGP^{+}06, BBP^{+}13, BPP07, BBB^{+}16, BUE^{+}98, BFSV08, BRC^{+}03, BPC^{+}16, CMB^{+}15, CH95, Cur04, CCP07, DBGW04, ESA^{+}16, FYK^{+}21, GIT^{+}13, Gla11, GöEIOS16, GFO14, HMM01, HJR^{+}03, HSLP19, HCC^{+}09, HBG^{+}16, ICB^{+}08, IK97, IYN^{+}09, ISN^{+}11, KL01, LGM^{+}02, LVC^{+}05, LC95, LPSS04, LBSS^{+}92, MSM^{+}13, MYHvdL15, Mul97, MFP^{+}03, NFN00, NTIO18, PHH^{+}98, PVMP03, PBL07, RCB08, RGQPN09, SSP^{+}07, SGFR^{+}21, SLL19, TWKW01, TW05, TCL^{+}12, TA06, TMN^{+}15, TCC^{+}98, TTC^{+}12, WMD^{+}06, ZKT07, ZYY^{+}21, ZYT^{+}22, ZHL^{+}03, ZVKŠ13]. Anguilla [AM18, BCR08, BBT^{+}09, CSS^{+}21, HZTS12, HXC^{+}17, KSY^{+}23, SOTM^{+}18]. animals [LPG^{+}06]. anniversary [Kim23]. Annual [BAB^{+}06, CP03, HL98, KTH^{+}15, Kas97, Kas98, Kas99, Liv00, RCS98, Woo95, Woo97, AYK03, ETB^{+}17, GFG98, LP10, LAPL21, MBY^{+}18, OE17, SCTB19, TAN^{+}17b, VYGT^{+}20]. Anomalies [OBA01, BMHW13, KJZ97, LJM^{+}10]. Anomalous [BBS99, NH01, SWZ^{+}01, TCL^{+}12]. Anomaly [MM94a]. Anoplopoma [GJR18, KMB00, SC06, SE19]. Antarctic [BCA^{+}18, LPCA15, MMI^{+}22, MKH^{+}13, TBB^{+}03]. Antarctica [MKH^{+}13, SRCV09, BCA^{+}18]. antennatus [CLPC18]. anthropogenic [CH16]. antipredator [VN97]. appears [Jan16]. Application [BHM02, BGM^{+}18, AB02]. Applications [CH99]. applied [LPS19, LBW^{+}05]. appraisal [GPA^{+}21]. appreciation [BD93]. approach [BHV^{+}06, BBY08, CC03, CH95, CMS16, HVHC10, LVPK11, LMBL03, MLM^{+}98, MMBC07, MCB^{+}16, NH06, OIA^{+}12, PVMP03, PLP^{+}11, PQH16, RBPCR^{+}22, SP15, WKR^{+}18]. approaches [CIS20, GNP^{+}19, Hor00]. April [JMP^{+}14]. aquaculture [HSEH16]. aquatic [SAO^{+}17]. Aransas [BHJ^{+}04]. Arc [SPV96]. Archipelago [SFA14, FKH^{+}17, SPS^{+}20, aTCK05, HS05, MSL^{+}05, SHM05]. archival [AMD^{+}16, APR^{+}08, CÅP^{+}13, DPM^{+}11, GJR18, HLG^{+}11, HKLG07, MKK13, MLR10, MBB^{+}03, PECG08, RHG^{+}13, SF22, SMB03b, WSP^{+}07]. Arctic [ÅGN^{+}04, HPL13, LOS^{+}14, LSK^{+}18, LS21, MFRR96, RDE^{+}07, SB07]. Arcto [OS95, VSÅO07, LOS^{+}14]. arcto-boreal [LOS^{+}14]. Arcto-Norwegian [OS95, VSÅO07]. area [AOVAG22, AM18, BPZR19, BCJ^{+}13, CLM^{+}21, CLKP19, CAR^{+}10, Dom04, FHD98, GSNFL99, HQW^{+}99, ISN^{+}11, KKNY04, KVR^{+}18, KHB02, Mar01, NSH^{+}17, NHS^{+}07, STI^{+}09, SHK^{+}19, TTI^{+}20, WZK97, Yam04, YKH^{+}21, SAH^{+}18]. areas [BJCS12, BSG^{+}13, BBR^{+}05, BHJ^{+}04, DWHdP21, FIDC00, FKUY16, GGF17, IWK^{+}21, KY17, MBH^{+}99, NBH99, OM10, RHRL12, RRF^{+}21, RF07, SF22, SLL19, UYF92, WJM15]. Argentina [ASCM12, TMMM20]. Argentine [HMM01, JMLG06, MSM^{+}13, MMSL19, PVHT01, TMMM20]. argentinus [ABI^{+}21, CAB12, WRTP01]. argo [ZWL21]. argo-based [ZWL21]. Argopecten [LCCS15]. Arguin [FIDC00]. argus [EF10]. Ariake [SKNT14]. Aristeus [CLPC18]. armorhead [LRS^{+}23]. arrowtooth [RKZHC19]. ascent [Hea99b]. ash [PW12]. Asia [HZTS12]. Asian [RZM^{+}03]. Aspects [MBJ^{+}07, SPM^{+}19]. aspera [BMHW13, Por22]. assemblage [DTO^{+}23, MHG^{+}11, SKM04, SSM^{+}10, TTH15]. Assemblages [SKKW02, ADAHL10, BDAMD14, CCK^{+}22, DABM^{+}06, ESA09, FGGDSMF08, FBRB12, FRHMAM^{+}06, FRZVHM^{+}11, GHV95, GDM^{+}17, HFC01, HLWL12, JMLG06, JMP^{+}14, KN08, KYA^{+}15, KGW13, LLCJ16, MBY^{+}18, MTZG23, MBKP08, MSVY^{+}13, MMB^{+}11, OKU17, OK17, OEV^{+}10, SKHI04, XMW^{+}23]. assess [MLVO05, MDR^{+}16]. assessed [GCF^{+}21]. Assessing [DDS^{+}17, ESTJ03, LVC^{+}05, LPH^{+}19, MFG99, RR18, VCB^{+}98, BFF15, PDD03, TMM^{+}07]. assessment [BJCS12, KSAF13, KYA^{+}15, OTIK20, OTH09, SC05, SSP^{+}11]. assessments [Bri94]. assimilation [MLM^{+}98]. associated [EBO04, GCF^{+}21, GAH^{+}19, LPCG23, MSST16, MMRH^{+}16, MBB^{+}03, PM95, TID^{+}96, TCC^{+}98, WFRS93, YIT^{+}22]. Association [LLCV18, BGM^{+}18, HMS16]. Associations [GBAD^{+}17, CJ04, GPS22, JJBCW09, KR14, Mar01, MTSH15, PFAM96, PWML12, PMFC10, RMH^{+}19, SPM02, SB06]. asynchronous [SPM^{+}24]. at-sea [PLSO98]. at-sea-sampling [FCJ^{+}15]. Atka [MFH05]. Atlantic [APLG07, APL07, ADPC21, CBdSF^{+}23, FC04, FMG^{+}22, HKLG07, MSM^{+}13, OCH99, SPM^{+}19, SPS^{+}20, SCS05, AUOGMM19, And03, AAKMG06, BC97, BC04, Bea03, BBR^{+}05, BBT^{+}09, BUE02, BSF01a, BB07, BvDSDC18, BCL04, BDTR23, BPS^{+}14, áCGNGC19, CTWS08, CJ04, CMMK^{+}15, COW^{+}99, CRC11, CGI^{+}19, CIS20, CWCM14, DHC^{+}07, DH11, DPM^{+}11, DB93, DDS^{+}17, DBS^{+}19, DGB^{+}16, DDZ09, DB03, Erz05, FDT^{+}99, FHD98, FRBB14, GI13, GHV95, GRT^{+}07, GCW17, GVRC04, HB99, HT18, HA07, HBPC15, HKWL17, HLG^{+}11, HBR^{+}15, HDJ15, IIS^{+}07, IHS97, ISS02, Jan16, KSP^{+}22, KVR^{+}18, KR10, LLCV18, LPS19, LJR^{+}22, LC95, Mar01, MMSL19, MDVB^{+}20, MHRC18, MM94a, MTSH15, MSL^{+}20, MMMS14, MHB^{+}14, MLR10, MMB93, NdLOO23, PLT09, PL03, PGL^{+}15, PLG^{+}10, QBMW99, QC99, QCM^{+}16, RF04, RFD^{+}04, RDF^{+}11, RQN^{+}99, RCPS09]. Atlantic [RSZ^{+}03, RBB^{+}21, RF07, SA10, SHS^{+}23, SGL22, SR02, SLZ^{+}23, SGHW05, SQW^{+}99, SNL19, SRM^{+}18, Swa99, VHCN14, VGPL^{+}11, WRTP01, WKN^{+}95, WJ93, XMH^{+}18, ZJH^{+}22]. Atlantis [OCCF^{+}18]. atmosphere [SCS05]. Atmospheric [OBA01, Sha13, BBS99, MCG^{+}14, PWML12]. Atoll [HK06]. audax [APMRH17, APMVOGMR19, GSNFL99, SDHB07]. auratus [Fra93]. aurita [MBE^{+}15]. australasicus [NK08]. Australia [BYM16, CB93, Cap08, DWH11, DBGW04, FML^{+}14, FvPH^{+}16, FHK^{+}10, FHK^{+}12, HHK^{+}10, LJM^{+}10, MDR^{+}16, MCS^{+}06, NK08, RHG^{+}13, RHP^{+}15, RRF^{+}21, SWS^{+}19, SBD^{+}19]. Australian [MMB^{+}11, MGHS14, NK08, RHP^{+}15, WMD^{+}06, DBGW04, JPHA^{+}16, KN08, MBKP08, PECG08, SHG^{+}22]. australis [DBGW04, WMD^{+}06]. Author [Ano01a, Ano03b, Ano04a, Ano05a]. autumn [FM93, FMG^{+}22, IMS^{+}04, SDHB07]. autumn-spawned [FM93]. Availability [ISS02, CMMK^{+}15, Jan16, NZI95, OEV^{+}10, PBF00, RJHC99, RBB^{+}21, SHS^{+}23, SBY^{+}15, TW05, Tan99]. average [RMM02, WGFR06]. avid [BZ21]. axis [TNK^{+}16]. Azores [SPS^{+}20, APR^{+}08, SPM^{+}19]. aztecus [MCB^{+}16].

B. [SMK^{+}13]. back [MTH^{+}04]. backscatter [TKH08]. backscatterings [MIY^{+}09]. Backward [GGQF22]. Baird [MIK07]. bairdii [MIK07]. Baja [AGSSL^{+}22, FRHMAM^{+}06, GPCGdlT^{+}22, HT99]. Balaenoptera [MTK^{+}07, MKH^{+}13, SMK^{+}13]. balanced [Gre13]. Balancing [PVBV19]. Balearic [CAGPC21]. Baltic [MKF^{+}03, AMK08, BML^{+}14, BSG^{+}13, BHV^{+}06, HBO^{+}01, HLMS03, HVHC10, Neu02, NHNA07, SHG12, SHB^{+}11, TLS98, VHJ99, VDHF08, WJT97]. balticus [SHB^{+}11]. Bank [FIDC00, MATL98, RAT^{+}02, VCB^{+}98, BSF01b, BCL04, LBW^{+}05, Lou10, MLM^{+}98, MLC^{+}98, NGGJ09, PSN^{+}99, PJD14, TCS^{+}09, WPL^{+}93]. banks [HDH^{+}05]. barbatus [GGF17]. barcoding [ARM16, BBB^{+}19, KBB^{+}20]. Barents [NFO^{+}23, ESTJ03, FGS95, HEG08, HCFP20, OÅL00, OH23, SPLY23, WPN12]. Barotropic [LHF^{+}99]. Barrier [LHF^{+}99, MSVY^{+}13]. bartramii [ASM^{+}15, FCC^{+}19, IMS^{+}04, ISI^{+}18, NII^{+}14, NTM^{+}15, YWM^{+}00]. based [ACT^{+}10, AMK08, BC04, BRC04, BJCS12, BHV^{+}06, BLH98, BHM02, CAB12, DPK^{+}08, DMH16, FGS95, GNP^{+}19, HHK^{+}17, HP02, HBC07, HHB^{+}15, ITH23, KMM^{+}06, MLVO05, MCHSNEO13, MPM19, MKK13, MLC^{+}98, MMMS14, NK08, NBMS06, Nis92, OTIK20, PG06, PLG^{+}10, QBMW99, RHRL12, RBPCR^{+}22, RWLP12, RWP11, SYT^{+}09, VN97, VFS^{+}24, ZWL21, ZSY^{+}21]. baseline [Yam04]. Basin [BHH98, HBLC22, SGHW05, Neu02, SHG12, TLS98, CAGPC21]. Basin-scale [BHH98, SGHW05]. basins [NSGL^{+}22]. basis [TR11, Tan17a]. Basking [Wil04, CSFC05, SR02]. bass [EHW08, NASTF10, NH06, SFK^{+}20]. bathymetric [JYH^{+}18]. bathymetry [OR12]. Bay [APL^{+}96, APL^{+}08, COW^{+}99, KKK^{+}17, KTH^{+}15, KU95, KUO^{+}17, LA05, MBY^{+}17, MWN^{+}23, QLB^{+}05, RTK01, SFK^{+}20, TNM^{+}02, TKH08, TKMS11, TY04, YIT^{+}22, LCCS15, ACT^{+}10, ACG^{+}16, APL01, APGL03, APLG07, APL07, BPP07, BUE^{+}98, BFSV08, BBA^{+}21, BPC^{+}16, DPL^{+}20, GHG^{+}19, HBG^{+}16, ICB^{+}08, JR07, LOGLD^{+}15, OKU17, PLT09, PBL07, SPM02, VGPL^{+}11]. bay-mouth [KKK^{+}17]. Bayesian [RGQPN09]. bays [GV01, SBT20]. BC [MFG99]. be [Jan16]. beach [TSG^{+}20]. beach-seine [TSG^{+}20]. beaches [XMW^{+}23]. beaked [MIK07]. bearded [SBY^{+}15]. Beaufort [BAL^{+}99]. bed [FKH^{+}17, MTT^{+}17, VPRG13]. before [OK17]. beginnings [Sha95]. Behavior [GCF^{+}21, BMOT17, CSK11, EHW08, KKNY04, MKK13, MSST16, SRCV09, SFA14, TNC^{+}22, TTC^{+}12, WPL^{+}93]. behavioral [CCM^{+}08, HKM^{+}21]. behavioral-physical [CCM^{+}08]. behaviors [DPM^{+}11, SAH^{+}18]. Behaviour [FDT^{+}99, KSMY00, BGH09, FMYN06, FCL93, FHK^{+}12, HT18, HQW^{+}99, KFH00, MIK07, OA06, PECG08, SSR13, VN97, Wil01]. Behavioural [RDF^{+}11]. Bellingshausen [KEJK00]. belone [ABG19]. Belt [SMF96]. benefit [MTL^{+}22]. Bengal [GHG^{+}19]. Benguela [IMO^{+}12, AJ15, Col99, JHK^{+}15, KYA^{+}15, KYSM11, KYS15, LRL^{+}06, MYHvdL15, MFP^{+}03, OCCF^{+}18, PHH^{+}98, PVMP03, Sko05, SSSB03, WJM15, YMK^{+}15]. Benthic [DMF^{+}17, BPZR19, JYH^{+}18, QM01, SFL16, TKM^{+}22, VPRG13]. benthic/demersal [QM01]. bentincki [GMH^{+}12]. Berardius [MIK07]. Bergen [LJR^{+}22]. Bering [WSC05, AYMK01, BCBDA10, BHC^{+}01, BH18, BRO18, BO05, BMO^{+}99, BDAMD14, CRW20, CEM^{+}11, DABM^{+}06, KEWDA18, MSS12, MW92, Mor11, NKS00, NH01, Por22, Ree95, SGW^{+}21, SS94, SCDA10, SADA^{+}23, Spe08, SMF96, SBK^{+}01, SWZ^{+}01, ST97, SP13, TID^{+}96, UMK20, Wat17, WQI00, WQ00, WEW98, YCH^{+}15]. best [TSG^{+}20]. between [And03, Ano99, BEF^{+}12, BBR^{+}05, BUE^{+}98, BBB^{+}19, RPG^{+}22, CSB94, EBFF17, GGF17, GI13, GPS22, GBAD^{+}17, GEGHPCC17, HMM01, HA07, HBO^{+}01, HMS16, HCC^{+}09, IMS^{+}04, KSP^{+}22, KSAF13, LLCV18, LS21, MEK^{+}09, Mal20, MHM^{+}20, MDR^{+}16, MKF^{+}03, NZI95, NTIO18, Nis19, NdLOO23, OM10, OHS06, QM01, RRF^{+}21, RZM^{+}03, SMK^{+}13, SPM02, SPT^{+}17, TKM^{+}22, WTK^{+}16, WMKR09, Wat17, WGFR06, WGS^{+}08, YW94, ZLTM11, ZKT07, ZHT14]. bicoloratus [YTY96, YOY00]. bifurcation [BF07, KFS22]. Bigeye [APR^{+}08, BHM02, Dom23, GCF^{+}21, HKM^{+}19, HKM^{+}21, HK06, LLCV18, MKK13, MSST16, MBB^{+}03, RWI^{+}16, SMB03b, SLZ^{+}23, ZSY^{+}21, ZWC^{+}21]. Bight [FMG^{+}22, HSS19, RHP^{+}15, SGL22, BK94a, BK94b, CTWS08, CG18, MDKS93, OCH99, SKNLD10, SGN^{+}05, SCS05, WMD^{+}06]. bilinearis [RPC^{+}19]. billfish [HBLC22]. billfishes [PLG^{+}10]. Bio [HG98, HZTS12, LAB^{+}98]. bio-geochemical [LAB^{+}98]. Bio-physical [HG98]. bio-tracers [HZTS12]. biochemical [ITH23, ODMRM98]. biochronologies [BMHW13]. biodiversity [JHK^{+}15, LS15]. bioenergetics [GiIW^{+}20, IKK^{+}04]. biogeochemical [AGK^{+}08, LCH03, MEK^{+}09, SMDM98]. biogeochemical-populations [LCH03]. Biogeography [KOWM16, PAS^{+}18]. Biological [Har92, LOGLD^{+}15, SPM^{+}19, BLD^{+}03, CH95, JGS93, KO95, LLCJ16, LSD^{+}21, MTL^{+}22, MLM^{+}98, MIY^{+}09, MWR^{+}98, MMB93, NKS00, PHH^{+}98, PMG^{+}94, TR11, Tan17a]. biologists [Tyl92]. Biology [NH01, DLCQ22, LJR^{+}22, Tan02]. Biomass [HKT^{+}03, BKvdP^{+}22, BW92, CP03, Coy05, GEGHPCC17, HH99, KSC^{+}10, KL01, LP10, MM03, NKM01, NY03, OS95, RFM^{+}21, RCD^{+}99, ST97, ST98, TCO^{+}05, UMK20]. Biophysical [APLG07, CLKP19, Ols01, APL07, BTGM07, HRS^{+}21, IXW^{+}10, LBW^{+}05, MDR^{+}16, PML06, PJD14, RRF^{+}21]. Biosphere [SFA14]. Biotic [REG^{+}13, FYK^{+}13, HVHC10]. biovolume [CC03]. bird [SBT20]. birds [BG01, BWKM15, CCL^{+}05, LH96, SPV96]. Biscay [APLG07, APL07, ACT^{+}10, ACG^{+}16, APL01, APGL03, BPP07, BUE^{+}98, BFSV08, BPC^{+}16, HBG^{+}16, ICB^{+}08, LOGLD^{+}15, PLT09, PBL07, VGPL^{+}11]. bivalve [MPM19]. black [EHW08, MJH14, GöEIOS16, GFO14, ODMRM98, Shi98, Zai92]. black-footed [MJH14]. blackbelly [SPS^{+}20]. Blackspot [SFGE21, GEGHPCC17]. Blob [RWDA^{+}21, YCS^{+}19]. block [RMM02]. bloom [CP92, FYKSP07, KSYT97, KWO^{+}18, MRHL09, SFL16]. Blue [OHF12, BC97, BBH99, CKA^{+}17, CIS20, CWCM14, ERE^{+}10, GPCGdlT^{+}22, HEG08, MMRS16, MP18, NK08, OFS^{+}16, REL07, RCPS09, SSPY08, SSP^{+}11, TDE09]. Bluefin [RF07, SGL04, AUOGMM19, AMD^{+}16, BGH09, DGB^{+}16, FRBB14, FHK^{+}10, FHK^{+}12, FFF^{+}18, GCQ^{+}13, HKWL17, HFF^{+}19, HHTF10, HHK^{+}10, IFF^{+}18, KKNY04, KBF^{+}07, Mat06, MLR10, PECG08, Pol96, RF04, RSZ^{+}03, RBB^{+}21, RMH^{+}19, SL09, SAT^{+}18, TTI^{+}20, VHCN14, WMD^{+}06]. bluefish [CTWS08, VHLM15]. bluemouth [MBJ^{+}07]. bocaccio [ZLTM11]. Body [Mor11, AGSSL^{+}22, AOVAG22, AI05, BMHW13, CHPT20, HKM^{+}19, HKM^{+}21, IMS^{+}04, KHN^{+}22, OFS^{+}16, PGL^{+}15, REG^{+}13, TB92]. bogaraveo [GEGHPCC17, NSGL^{+}22, SFGE21]. Bohai [GFG98, TJW^{+}03, WLWZ98]. bonaerensis [MKH^{+}13]. bonasus [CGMM10]. Bonga [BDE^{+}19]. bongo [MM03, PSC05]. Bonnaterre [NdLOO23]. Book [Ano94, Gra98, Gre99, Par99]. boosted [MCB^{+}16]. boreal [LOS^{+}14]. borealis [FYKSP07, KFYP07, OA06, SMK^{+}13]. Boreogadus [MFRR96]. Bornholm [Neu02, SHG12, TLS98]. both [DBB^{+}18, TAN^{+}17b]. bottlenose [KFS22]. Bottom [SBD^{+}19, TMM^{+}07, AAI16, AJ15, ESA09, FMM^{+}20, HAS^{+}19, JHC^{+}15, KCW^{+}15, LA05, Lou10, OUKH04, SYT^{+}09, SCTB19, TMMM20]. Bottom-up [TMM^{+}07, HAS^{+}19, TMMM20]. Boundary [Esc98, EvST^{+}17, LOS^{+}14, SES^{+}20, SBD^{+}19, WMD^{+}06]. brachyuran [éSMB20]. Brama [QCM^{+}16]. Brandt [ESA^{+}16]. Bransfield [LLCJ16]. Brazil [CG18, ABI^{+}21, MHS^{+}21]. Brazilian [AG99, CMM06, LC95, MDKS93, SS98]. break [CMM06, SHS^{+}23]. breakwater [KKK^{+}17]. bream [YOYK20]. breeding [BRR05, HKA^{+}06, XTC^{+}04]. Bregmacerotidae [MDKS93]. Brevoortia [FDT^{+}99, HT18, QBMW99]. brief [Sch23]. Bright [HMT07]. Bristol [APL^{+}96, DPL^{+}20, LA05, RTK01]. Britain [CSFC05]. British [APL^{+}96, GDM^{+}17, JTYB18, PHWM96, PMT^{+}94, HTL^{+}00, SME^{+}14, Tan17a, WWSE00]. Broad [RHG^{+}13, VPRG13, MTSH15]. Broad-scale [RHG^{+}13, VPRG13, MTSH15]. Brown [MCB^{+}16, DST11, HTP14, HSS19, SGN^{+}05]. Browns [BSF01b]. Bryde [MTK^{+}07, SMK^{+}13]. brydei [SMK^{+}13]. building [MLR10]. bungii [TSK04]. Buoyancy [PSS^{+}21, HBG^{+}16, PVMP03, VJ99]. buoys [MBB^{+}03]. bust [SFL16]. Buzzards [LCCS15]. by-catches [LAFF15]. bycatch [AUOGMM19, BMH^{+}21, CIS20].

C [Sim96, Jes22, WP93]. Ca [FKUY16, YOY00]. Cádiz [RR18]. Calanoida [TSK04]. Calanus [Ano99, BM99a, BHH98, CW98, Esc98, GMH^{+}99, HTE^{+}03, Hea99b, HBR^{+}99, HJ99, HDF^{+}99, IHHH99, IH03, Jan16, Jón99, LSW^{+}03, MBH^{+}99, MLC^{+}98, MTLL^{+}16, Mul94, Mul97, NGGJ09, PHH13, RCS98, RJHC99, RD96, SGHW05, TDT03, VJ99]. calibration [HDF^{+}99].

California [AGSSL^{+}22, ERR^{+}21, FRHMAM^{+}06, GPCGdlT^{+}22, HT99, JCCB15, JJBCW09, KGW13, SKNLD10, ARL93, Aut08, BRFRJRLC18, BDSM07, CC03, CCP07, EBFF17, ESA^{+}16, FRZVHM^{+}11, Gla11, HTLJ20, HCWF21, HKA^{+}06, KCW^{+}15, KBS^{+}16, LBLCLC05, Lyn03, MRRN05, MJH14, MLRS07, MWB^{+}00, Mul94, NPS^{+}23, PM95, PDER10, PMFC10, PMG^{+}23, PCR^{+}18, RCB08, RMH^{+}19, SRR99, SRR07, SC06, SWAAB20, SCKJ^{+}18, Sim92b, TCL^{+}12, THL^{+}18, VFS^{+}24, VMT^{+}23, WGW07, WGS^{+}08]. Californian [Mul97]. called [GSNFL99]. Callinectes [CWCM14, ERE^{+}10, OHF12, REL07, TDE09]. Callorhinus [HMS16, YKB08]. camtschaticus [LA05]. can [BBT^{+}09]. Canada [GDM^{+}17, RDF^{+}11, War92, éSMB20, DTC06, JR07, PBF00, XDP^{+}20]. canadum [CBdSF^{+}23]. Canaria [BAB^{+}06, MRHL09]. Canary [BAB^{+}06, MRHL09, HL98, MRBBHL14, SGFR^{+}21]. Cancer [MAHG94, Sha13]. candidate [HTP14]. cannibalism [NGGJ09]. Cantabrian [GQPGA04, RBPCR^{+}22]. canyons [CCK^{+}22]. capacities [VAFG95]. capacity [Mat06]. cape [BKvdP^{+}22, GS99, KvdPBW17, KYS15]. capelin [APL^{+}08, CP92, FGS95, HWSS07, IHS97, LDAWM10, OR12, OR13, WPN12]. capensis [BKvdP^{+}22, IMO^{+}12, KvdPBW17, KYS15, MFP^{+}03, PVMP03, WJM15]. capensis/encrasicolus [MFP^{+}03]. capture [HHTF10]. captures [BCR20]. carangid [MSC^{+}17, RS15]. carbon [JCH04, Ste98, VZP98]. Carcharhinus [RHG^{+}13]. carcharias [MCHSNEO13]. Carcharodon [MCHSNEO13]. Caretta [PKP^{+}00, PBH^{+}04]. Caribbean [JMP^{+}14]. Carolina [GP94, COW^{+}99, OHF12, QLB^{+}05, WBQL99]. Carrying [Mat06, VAFG95]. cascade [BRO18]. case [BML^{+}14, BSG^{+}13, BFSV08, CIS20, DWHdP21, FH94, GEGHPCC17, HLMS03, HBN^{+}21, KU95, LOS^{+}14, LVM^{+}18, PVBV19, RF07, SNV^{+}12, TSG^{+}20, TAN^{+}17b, TFB^{+}17, VGPL^{+}11]. Castellanos [CAB12]. Catalan [OEV^{+}10, SSP^{+}07]. catch [ARL93, BBH99, BML11, CIS20, DWH11, Dom23, DTC06, FML^{+}14, GHG^{+}19, HBLC22, HK06, HBR^{+}15, KvdPBW17, LLCV18, MDR^{+}16, MMBC07, MMRH^{+}16, MHB^{+}14, NLN^{+}21, NNOU20, RMH^{+}19, VHCN14, WMKR09, Wat17, YOYK20, ZHX^{+}20]. catch-per-unit-effort [NLN^{+}21]. catchability [SBD^{+}19]. catches [BRN^{+}95, FCJ^{+}15, GPCGdlT^{+}22, HSLP19, HDJ15, IHS97, LAFF15, SA10, SR02]. catching [TSG^{+}20]. caught [NFN00, YAM^{+}18]. causality [NTIO18]. cause [McK13]. caused [OKU17]. Causes [FCL93, Fun11, KHN^{+}22, SGN^{+}05]. cavalla [WMKR09]. cavity [AI05]. Celtic [PLT09, PSJF93]. Central [NdLOO23, TR11, AYMK01, ASM^{+}15, Aut08, BHV^{+}06, BS94, CCK^{+}22, Coy05, ESA^{+}16, FGGDSMF08, FYK^{+}21, GMH^{+}12, GQPGA04, HJ10, INM^{+}18, JCH05, KNE^{+}04, KTPM17, LAFF15, LHM^{+}05, LTL^{+}22, Lyn03, MRRN05, MSST16, MJH14, NPY^{+}15, PDER10, PMG^{+}94, PKP^{+}00, PBH^{+}04, QM01, SRR99, SRR07, SHG12, SF22, SLL19, SMF^{+}05, SHB^{+}11, TID^{+}96, TY04, WMKR09, AMK08, CKA^{+}17, GGQF22, MKF^{+}03, VFS^{+}24]. central-northern [SLL19]. central-south [QM01]. central-southern [NPY^{+}15]. Centropristis [EHW08]. century [BEiI^{+}23, REB^{+}03, SLM13, SB04, War95]. cephalopod [PQH16, áRÁSG^{+}16]. Cephalopoda [OKT^{+}23]. cesium [Kae17]. cetacean [SMF^{+}05]. Cetorhinus [CSFC05, SR02, Wil04]. chaetognath [TSK^{+}95]. chaetognaths [BT99]. chakograrnma [BBMY93]. chalcogramma [AYMK01, BCBDA10, Fun07, Fun11, FYK^{+}13, HYW04, HWSS07, HONH04, IST^{+}04, LDAWM10, MTH^{+}04, NKS00, NHS^{+}07, SB94, WSC05, Yam04]. chalcogrammus [KNS^{+}22, KTH^{+}15, KEWDA18, LK21, OTIK20, SADA^{+}23, YCH^{+}15]. Challenges [McK13, BEiI^{+}23]. chamaeleonticeps [NLN^{+}21]. chance [KWB^{+}16]. Change [KNE^{+}04, SB06, TID^{+}96, BML^{+}14, BBA^{+}21, BMO^{+}99, BB07, CCL^{+}05, CEM^{+}11, DG00, DMH16, FMM^{+}20, FvPH^{+}16, GHM21, GVRC04, Han11, HGG^{+}17, HB92, JPHA^{+}16, LPHM21, LVM^{+}18, MSS12, NTIO18, NPLS22, OCCF^{+}18, PRDC^{+}13, Pol96, RPE98, SMS^{+}23, Swa99, TMN^{+}15, War95, WK03, YW07]. changed [MYHvdL15]. Changes [AS08, CH16, HKM^{+}19, LBLCLC05, MSR20, OHS06, SADA^{+}23, SBBB03, Swa99, TB92, AGS^{+}04, ABS^{+}11, AOVAG22, Bea03, BCR08, BDSM07, CGI^{+}19, FRBB14, GHV95, GöEIOS16, HYW04, HK06, IHHH99, IFF^{+}18, JGS93, KYSM11, KNO^{+}04, LA05, LMBL03, MFMG20, MHG^{+}11, MTH^{+}04, NIIS04, OTH09, PP01, RF04, REB^{+}03, SFGE21, SHG12, SW05, SC06, SPG^{+}16, Shi98, Spe08, TAS04, TBB^{+}03, YOYK20, YKB08, Zai92, ZP21a]. changing [DB93, FCC^{+}19, FPFL13, LOS^{+}14, LBC23, SCTB19, SMS^{+}19]. Changjiang [IK97, XWL^{+}23]. Channel [Hea99b, HJ99, IH03, Jón99, LGM^{+}02, LVC^{+}05, NPLS22, OUKH04, RJHC99]. Characterising [DWHdP21]. Characteristics [RPG^{+}22, CCK^{+}22, LOGLD^{+}15, MSNK10, MHvD^{+}24, PSJF93, RQN^{+}99, SHK^{+}19]. Characterization [RD96, GR98, MIY^{+}09]. Characterizing [DTO^{+}23, GIT^{+}13, MMRH^{+}16, SRM^{+}18, BPLC11]. Charlotte [JTYB18]. charr [RDE^{+}07]. chemistry [ACT^{+}10, RSZ^{+}03]. Chikugo [SKNT14]. Chile [REM02, CRVL^{+}17, FYC22, GMH^{+}12, HSLP19, NPY^{+}15, QM01, SLL19]. Chilean [Esc98, GNP^{+}19, LPCG23]. chilensis [Esc98]. China [IK97, KKH^{+}20, KKNY04, KMK^{+}18, LTL^{+}22, MTLL^{+}16, OTH09, SKM06, SYT^{+}09, TTC^{+}12, LJBR20, LSW^{+}03]. Chinook [BRN^{+}95, BRPC08, DDB17, HHH^{+}16, HFHW19, HTT^{+}16, HMT07, LMB^{+}19, MRRN05, PMFC10, RAK^{+}17, SMB^{+}03a, SW05, SVEW^{+}13, VFS^{+}24, WS08, WGFR06, WGW07, WGS^{+}08, XDP^{+}20]. Chionoecetes [KBS^{+}16, SP13]. Chlorophyll [YW07, ST97]. Chlorophyll-a [YW07]. Choice [ZYY^{+}21]. chokka [DBRSC16, MRL^{+}14]. chronology [SMB^{+}03a]. Chub [GiIW^{+}20, KOS^{+}19, KM93, PVHT01, TYO21, YWI^{+}05]. Chum [YCH^{+}15, AI04, AI05, FYA^{+}21, MWN^{+}23, Mor11, PHWM96, SKHN11, Sai22, TID^{+}96, WTK^{+}16, Wat17, ZZ93]. ciliates [KT93, ST95]. Circulation [CFL^{+}99, HB99, AYK03, BK94b, EHW08, HQW^{+}99, MLC^{+}98, RPT^{+}00, SNV^{+}12, TDE09, VSÅO07, WJP^{+}01]. Citharichthys [SRR99]. Clarence [Gre99]. Clarifying [YOIW21]. class [ASCM12, Fra93, GPS22, KMB00, LK21, MSS12, NDC05, RTK01, TY04, YCH^{+}15]. classification [CL05, MLR10]. Clay [Gre99]. clear [BBB^{+}19]. clearly [Bow11]. Climate [ASCM12, BB02, BH18, BFSV08, CEM^{+}11, CCHL23, DG00, GCQ^{+}13, GVRC04, HBR^{+}99, HAS^{+}19, HDJ15, KHB02, MTLL^{+}16, NTIO18, PCR^{+}18, SLM13, VOB^{+}19, XDP^{+}20, AH97, BML^{+}14, BCGB14, BYM16, BBY08, BBA^{+}21, BMO^{+}99, BB07, CSFC05, CHF^{+}04, DLCQ22, DHC^{+}07, DMH16, ERR^{+}21, FHHW98, FvPH^{+}16, GPCGdlT^{+}22, GHM21, GFO14, Han11, HA07, HGG^{+}17, HCWF21, HB92, JHK^{+}15, JPHA^{+}16, KNE^{+}04, Kae23, KWB^{+}16, KGW13, LBC23, LPHM21, LCH03, LYT^{+}20, LVM^{+}18, LS15, MTL^{+}22, MKF^{+}03, NH01, NPLS22, OTIK20, OCCF^{+}18, OHS06, OH23, PRDC^{+}13, PL03, PMG^{+}94, Pol96, RPE98, Rob94, ROB05, RR18, RCD^{+}99, SBY^{+}15, SGFR^{+}21, SW05, SVEW^{+}13, SDRL96, SNL19, SPT^{+}17, TMN^{+}15, TGRS^{+}19, TMM^{+}07, TTH15, War95, WWSE00, YSW^{+}99, ZLTM11, ZHT14]. Climate-driven [MTLL^{+}16, OTIK20]. Climate-Fisheries [BB02]. Climate-induced [ASCM12, GCQ^{+}13, SLM13, VOB^{+}19, MTL^{+}22, Pol96, SW05]. Climate-influenced [CCHL23]. climate-oceanological [SDRL96]. climate-related [LCH03]. climate-to-fish [SGFR^{+}21]. Climatic [BCR20, MMBC07, APL^{+}96, HQH^{+}06, PSM00, RR18, TAS04]. climatically [LOS^{+}14]. close [HTP14]. Cloudy [KWB^{+}16]. Clupea [BML^{+}14, BG01, BWKM15, BDTR23, CAB^{+}01, FPBDC11, FM93, FBRB12, GPA^{+}21, LYT^{+}20, MLVO05, NDC05, Neu02, REG^{+}13, SNV^{+}12, Tan17a, WQI00, WQ00, óóSV18]. clupeiform [BAB^{+}06]. Clupeoid [Sko05, Col99, TTY^{+}23]. co [AOVAG22, BH97, EPG^{+}16, HSH^{+}22, PMG^{+}23, HLH^{+}17]. co-occurrence [EPG^{+}16, PMG^{+}23]. co-occurring [AOVAG22, BH97, HSH^{+}22]. Coast [KNK^{+}18, AGSSL^{+}22, AG99, ASK99, ABS^{+}11, BJV^{+}17, BPLC11, DDS^{+}17, FYK^{+}13, FRHMAM^{+}06, GNP^{+}19, GPCGdlT^{+}22, HYW04, HA07, HT99, HFF^{+}19, HONH04, ISI^{+}18, IST^{+}23, JHC^{+}15, KBS^{+}16, KK00, KB08, KY17, LRBJ21, MRRN05, MAS^{+}98, MTH^{+}04, MBKP08, MTT^{+}17, OK17, OEV^{+}10, PDER10, SK03, TMN^{+}15, TTI^{+}20, Tan99, Tan02, WTK^{+}16, YKB08, DAW^{+}23, DWH11, KSC^{+}10, SME^{+}14, SMS^{+}21, XB09]. Coastal [Col00, FM93, SHG^{+}22, BSG^{+}13, BBB^{+}16, CHPT20, CRVL^{+}17, DCLC15, DLD^{+}23, FvPH^{+}16, GPS22, Han11, HCC^{+}09, IXW^{+}10, IHS97, IWK^{+}21, JPMH20, JMLG06, LJBR20, LML^{+}03, MBY^{+}18, MAHG94, MWP02, NFN00, NASTF10, NHS^{+}07, OM10, QM01, RFD^{+}04, Rob94, RHRL12, REM02, RMM02, áRÁSG^{+}16, RAK^{+}17, SSW^{+}17, STI^{+}09, SLL19, Sim92b, SNL19, TKM^{+}22, TCS^{+}09, TDE09, TCC^{+}98, WTK^{+}16, WZK97, WL21, ZYY^{+}21, ZYT^{+}22]. coastal-offshore [SSW^{+}17]. coastal-pelagic [CHPT20]. coasts [BUE02, CSS^{+}21, PS06]. Cobb [DP01]. cobia [CBdSF^{+}23]. coccolithophore [HGH93]. Cod [HBPC15, HMP92, MMB93, PSN^{+}99, AHKP16, AMK08, BCGB14, BSF01a, BTGM07, BCL04, CSB94, CRC11, D’A93, Dd95, DB93, DB03, ETB^{+}17, FUA^{+}98, FODCN00, GRT^{+}07, GCW17, HL07, HBO^{+}01, HCS^{+}09, IHS97, JCA^{+}16, KSAF13, KR10, LS21, SL95, LBW^{+}05, Lou10, MATL98, MFRR96, MRD^{+}19, NSH^{+}17, Neu02, NHNA07, OS95, OHS06, OH23, PA14, RKD^{+}20, SHG12, SP93, SC05, SB07, SB04, Swa99, TNM^{+}02, TLS98, VSÅO07, VHJ99, WPL^{+}93, WJT97, WKN^{+}95]. codlet [MDKS93]. Coexistence [AHAM03]. Coherence [DAW^{+}23, PWML12, RAK^{+}17]. coherent [Pol96]. Coho [BNM^{+}00, BRPC08, BDSM07, Col00, DAW^{+}23, KHB02, LML^{+}03, PMFC10, RWLP12, RWP11, SMB^{+}03a, WS08, WGFR06]. Cohort [CTWS08, NII^{+}14, NTM^{+}15, VFS^{+}24]. cohorts [IMS^{+}04]. Coilia [SKNT14]. coincidence [SS94]. Cold [FMG^{+}22, YKH^{+}21, APL^{+}08, KEWDA18, MPW^{+}99, OUKH04, Por22, REM02, SADA^{+}23]. collapse [KKK^{+}17, Kaw93, MRD^{+}19]. collected [KBB^{+}20]. collection [KSM^{+}20, Sch23]. Cololabis [FKSA21, INM^{+}18, IST^{+}04, III^{+}06, KHN^{+}22, KNO^{+}04, MVK^{+}20, OWK^{+}03, OWK04, OTO^{+}09, SK04, TKO^{+}14, TNK^{+}16, YW07, YOIW21]. colony [PLSO98, SAG^{+}09]. colour [RR18]. Columbia [GDM^{+}17, APL^{+}96, EBO04, HTL^{+}00, HMT07, JTYB18, PHWM96, PMT^{+}94, SME^{+}14, SMB^{+}03a, Tan17a, WWSE00, WSF^{+}14]. columns [OA06]. combination [DST11]. Combined [SPLY23, CC03, RGQPN09]. combining [DLD^{+}23, HVHC10, VHJ99]. come [GGQF22, GJR18]. comment [Bau95]. Comments [Sim96]. commercial [BSF^{+}20, DWH11, FML^{+}14, HKLG07, HHH^{+}18, KMD^{+}09, NLN^{+}21, PBF00, SRR05, WKN^{+}95]. commercially [KTO^{+}11, SLM13]. common [GMH^{+}12, KFS22, KYU^{+}06, MTK^{+}07, ST95]. communities [CCSS01, DDZ09, FvPH^{+}16, MTT^{+}17, PFSL09, SDRL96]. community [APM^{+}12, Aut08, CAGPC21, FKH^{+}17, GR98, HT99, KKH^{+}20, KMK^{+}18, LéEPW^{+}12, LAG^{+}11, áRÁSG^{+}16, Shi98, UIU^{+}99]. commuting [HKA^{+}06]. Comparative [SB94, APGL03, BB03, KYA^{+}15, ZSY^{+}21]. compared [LVF12]. Comparing [CIS20, DB03, RKZHC19]. Comparison [IMO^{+}12, MSST16, MWGK92, NBH99, RMM02, SLZ^{+}23, BRC04, CHF^{+}04, PSC05, TF08]. compass [CSS^{+}21, CLH^{+}22, DLTI95, Sim96]. Competition [RZM^{+}03, LDAWM10]. competitive [WP93]. complete [DST11]. complex [PRDC^{+}13]. complexity [SPLS15]. components [BDVS^{+}19]. Composition [CAGPC21, ARM16, CMM06, GDM^{+}17, HKT^{+}03, KPHG14, KMK^{+}18, NKM01, OTIK20, PJO99]. compression [PG06, PLG^{+}10]. Computer [DLTI95, HTL^{+}00, Sim96]. Concentration [PTS^{+}24, RSC96, BBR^{+}05, HSLP19, KKK^{+}17, MWGK92, ST97]. Concentrations [MFS^{+}17, SS94, TDE09, WZK^{+}98, ZKT07]. concept [BNM^{+}00]. Concurrent [FYC22]. Condition [CLPC18, ADPC21, CHPT20, DDB17, DDB^{+}20, DBS^{+}19, LDDC06, MMMS14, NNOU20, PM95, PHWM96, PTS^{+}24, PGL^{+}15, TMMM20, TGRS^{+}19, VHLM15]. conditions [AGSSL^{+}22, AMK08, BGP^{+}06, BBP^{+}13, BFF15, Col99, CRVL^{+}17, DDB17, DAW^{+}23, DH11, DGB^{+}16, DHM^{+}15, ESA^{+}16, ECM^{+}01, GCQ^{+}13, HBLC22, HTT^{+}16, HWSS07, IFF^{+}18, KBF^{+}07, KYSM11, KB08, LLSF01, Mal20, MMSL19, MFMG20, MSC^{+}17, Mul97, NH01, NPS^{+}23, OEV^{+}10, PVHT01, PWML12, SC06, SWZ^{+}01, SK04, TSG^{+}20, TAN^{+}17b, TH11, TCC^{+}98, VYGT^{+}20, WMKR09, WGS^{+}08, WSF^{+}14, YWM^{+}00, ZSS08, ZHX^{+}20, ZVKŠ13]. conducive [ZVKŠ13]. conducted [WSP^{+}07]. configuration [TCS^{+}09]. Confirmation [GSNFL99]. Confluence [ABI^{+}21]. conger [LJBR20, LJBR20]. Congruent [SR02]. connection [SDRL96]. connections [MMI^{+}22]. Connectivity [CLM^{+}21, IXW^{+}10, KSP^{+}22, LCCS15, BJCS12, BCA^{+}18, CLKP19, CBdSF^{+}23, GGF17, HSH^{+}22, LPCA15, LPH^{+}19, LPHM21, MLP22, MHM^{+}20, NSGL^{+}22, POA^{+}17, PEKL14, QCR22, RRF^{+}21, RWI^{+}16, SGW^{+}21]. consequences [MM03, PMG^{+}94, WEW98, ZHL^{+}03]. Conservation [SAH^{+}18, CL05, HRS^{+}21, PFB^{+}16]. conservation/management [CL05]. Considerations [Nis92]. consistent [GPS22]. consumption [BWKM15, SFL16, WSC05]. contamination [SAO^{+}17]. content [DDS^{+}17, Jón99, NKM01]. Contents [Ano06, TID^{+}96]. context [Tyl92, VHLM15]. Continental [FMM^{+}20, MHRC18, CGMM10, EHW08, GMH^{+}99, GI13, GP94, HB99, HZTS12, HHK^{+}10, HCWF21, LPHM21, LP10, MPM^{+}13, RHP^{+}15, SSP^{+}07, SME^{+}14, SFL16, WBQL99, WKN^{+}95]. Continuous [BM99a, BM99b, RPE98, YCS^{+}15, COSC97, LVF12, PSC05, VCB^{+}98]. contraction [HGS^{+}21]. Contrasted [DBB^{+}18]. contrasting [SPT^{+}17, TNK^{+}16, WSC05]. contrasts [CHHS05]. contribute [PW12]. contribution [DBRSC16, LK21]. Contributions [IST^{+}04, YOY00]. Control [KEWDA18, CEM^{+}11, MEK^{+}09, TJW^{+}03, VGPL^{+}11]. controlling [CRC11, SHM05]. controls [BDVS^{+}19, CAB^{+}01, HGH93, HAS^{+}19, LVPK11, MLM^{+}98, PLP^{+}11, REG^{+}13, XWL^{+}23]. convection [MMRS16]. Convergence [ARM16, HJR^{+}03]. convergences [PTS^{+}24]. convergent [NTIO18, TDE09]. conversion [HBC07]. cooling [SBD^{+}19]. Copepod [KEWDA18, PL03, Bea03, GTB10, Jan16, MKF^{+}03, Mul94, TDT03, WZK^{+}98, ZKT07]. Copepoda [HT99, TSK04]. copepodite [IHHH99]. copepodites [BWJ03]. Copepods [BPLC11, HL98, NKM01, NIIS04, RAT^{+}02, STI^{+}09, UYF92, YCS^{+}15]. Coral [VOB^{+}19]. Corals [HWS^{+}05]. core [AI92, GSNFL99]. cormorant [ESA^{+}16]. correlated [CHPT20]. correlates [SRR05]. Correlation [YOYK20, MWP02]. correlations [Tyl92]. correspondence [BBR^{+}05]. Corridor [LJR^{+}22]. Corrigendum [Ano11a, Ano11b, Ano12, Ano15, Ano17, Ano19a, JJBCW17]. Coryphaena [MESMM18]. Counter [HDF^{+}99, GTB10, GR98]. coupled [CW98, CCM^{+}08, EHW08, HQW^{+}99, IKK^{+}04, ITH23, LAB^{+}98, LCH03, MEK^{+}09, SMDM98, TTC^{+}12, IXW^{+}10]. Coupling [CMB^{+}15, DPK^{+}08, MLC^{+}98, RHRL12, TKM^{+}22]. Covariability [RFM^{+}21]. covariates [HBN^{+}21]. Covariation [RCD^{+}99, WGFR06]. cover [Gre99, WEW98]. cownose [CGMM10]. CPUE [FCC^{+}19]. crab [CWCM14, DPL^{+}20, ERE^{+}10, KBS^{+}16, MAHG94, OHF12, SPM^{+}19, Sha13, SBD^{+}19, SP13, TDE09, YTIS95]. crabs [HSH^{+}22, LA05, REL07, RTK01, éSMB20]. Crangon [DST11, HTP14, HSS19, SGN^{+}05, TD02]. Crassostrea [KSM^{+}20, YIT^{+}22]. cristatus [TSK04]. critical [HSS19, PFSL09, REG^{+}13, ROH16]. croaker [ASCM12, HT18, HGS^{+}21, HA07, KJZ97, XWL^{+}23]. cross [BBT^{+}09, HWSS07, NTIO18, QLB^{+}05, RCG^{+}15, WJM15]. cross-shelf [HWSS07, QLB^{+}05, RCG^{+}15, WJM15]. Crustacea [HTP14]. crustacean [BBMY93]. cryopreserved [OK17]. Cs [MFS^{+}17]. Cs/ [MFS^{+}17]. Ctenolabrus [CLH^{+}22]. ctenophore [Shi98]. Ctenophores [CH92]. Cuba [CLKP19, KBB^{+}20]. cucumber [HMTG^{+}05]. cues [HALO00]. CUFES [PSC05]. cultural [DL94]. curl [WGW07]. Current [AJ15, BRFRJRLC18, CCP07, HKA^{+}06, JCCB15, KYA^{+}15, LLB^{+}20, MLRS07, NPS^{+}23, PMG^{+}23, SC06, SCKJ^{+}18, VMT^{+}23, AW92, EvST^{+}17, Gla11, HZTS12, HP02, HLWL12, JYH^{+}18, SES^{+}20, SLL19, Sim92b, TKH08, TDE09, WMD^{+}06, Aut08, AS08, BF07, BDSM07, Cap08, CC03, EBFF17, Esc98, FM93, FHK^{+}12, FRZVHM^{+}11, GSBB07, HTLJ20, HZW^{+}98, HCWF21, HXC^{+}17, IST^{+}23, IMO^{+}12, IWK^{+}21, JJBCW09, KFS22, KKS92, KCW^{+}15, KIS01, KMK^{+}18, KGW13, LBLCLC05, MCM^{+}17, MRBBHL14, MMB^{+}11, MGHS14, NKM01, NK08, PMFC10, PCR^{+}18, RCB08, RMH^{+}19, SGFR^{+}21, SMK02, SKM06, TCL^{+}12, TKO^{+}14, TYO21, THL^{+}18, TTH15, WZK^{+}98, YMK^{+}15]. currents [ABI^{+}21, AI04, FKH^{+}17, GV01, GP94, TIH^{+}92, Zam01]. Cushing [BD93]. cycle [BAB^{+}06, CP03, DST11, HL98, KU95, LVC^{+}05, OE17, TD02, TAS04]. cycles [GFG98, MMB93, PRDC^{+}13]. Cyclic [MMRS16]. cygnus [CB93, Cap08].

D [CW98, EHW08, ODMRM98, PJD14]. dab [LDDC06]. dactylopterus [MBJ^{+}07]. Dai [MFS^{+}17]. Dai-ichi [MFS^{+}17]. Daily [SK04, FML^{+}14, HPG^{+}20, KNO^{+}04, SPG^{+}16, SGS^{+}06, ZKT07]. Dall [OM10]. dalli [OM10]. damage [MMF95]. data [BH97, BRC04, BFF15, BM99a, BM99b, BHS^{+}15, DWHdP21, DWH11, FCJ^{+}15, GYS14, HBLC22, HLG^{+}11, KSMY00, LJBR20, LPG^{+}06, MPM19, MKK13, MFH05, MLM^{+}98, MMMS14, MIK07, MLR10, MBB^{+}03, NHNA07, Nis92, OFS^{+}16, PH11, ROH16, RDE^{+}07, SL09, Sch23, SDRL96, SMB03b, SSPY08, SRR05, WMD^{+}00, WSP^{+}07, ZSS08, ZWL21, ZSY^{+}21]. data-recording [KSMY00]. date [ACG^{+}16, FYK^{+}21, KNO^{+}04]. David [BD93]. day [HKM^{+}19]. Decadal [FH94, HYW04, KMB00, NH03, Pol96, TJW^{+}03, WK03, YKB08, ABS^{+}11, CHHS05, Gar97, LSK^{+}18, MM03, SNL19]. Decadal-scale [FH94, KMB00, NH03, TJW^{+}03, MM03]. decade [NNOU20]. decades [KK00]. decapod [CAGPC21]. Decapoda [HTP14, MHS^{+}21]. decision [HSEH16]. decline [CHM^{+}94, Fun11, JCA^{+}16, NNOU20, SR02, TMM^{+}07, ZHL^{+}03]. Declines [BRN^{+}95]. decrease [KY17, NNOU20]. decreased [SSW^{+}17]. Decreasing [KFYP07]. deep [CAGPC21, DBRSC16, GTB10, GGQF22, GJR18, HJ10, KvdPBW17, LTL^{+}22, MHG^{+}11, SPM^{+}19]. deep-spawned [DBRSC16]. deep-water [GTB10, GJR18, KvdPBW17, MHG^{+}11, SPM^{+}19]. define [Sco95]. defining [NBH99, SQW^{+}99]. Delaroche [MBJ^{+}07]. delayed [KHN^{+}22]. delays [KWO^{+}18]. delineation [BBB^{+}19]. delta [LPSS04]. Demersal [KSC^{+}10, KMD^{+}09, KCW^{+}15, KYA^{+}15, LVF12, LAB^{+}05, OKU17, PLT09, QM01, TTH15, YMK^{+}15]. demographic [GNP^{+}19]. demography [Mul97, SGHW05, WB93]. demonstrates [KBB^{+}20]. dendrochronological [BBY08]. dense [VPRG13]. Density [FYA^{+}21, Spe08, TYO21, KKCL06, KM93, MCB^{+}16, POA^{+}17, SB06, TKW^{+}17, TKO^{+}14, TB92, WZK97, XB09, ZLTM11]. Density-dependent [FYA^{+}21, TYO21, KKCL06, Spe08, SB06]. Density-independent [Spe08]. dentex [MTP07, MTP07]. deoxygenation [FKF^{+}22]. Departure [FHK^{+}12]. Dependence [EF10, XB09, YKI98]. Dependency [HLMS03]. dependent [AW92, CH92, CLH^{+}22, FYA^{+}21, Fun07, HHK^{+}10, IUY10, KKCL06, MW92, QCR22, SB07, Spe08, SB06, TYO21, Wil01]. depleted [JHC^{+}15, LBC23]. deposition [BDTR23]. depressed [JTYB18]. Depth [YMB99, AW92, CJ04, KN08, NY08, RDE^{+}07, SAG^{+}09, WM06, ZP21a]. depth-dependent [AW92]. depths [CLPC18, Hea99b]. derived [HLG^{+}11, Kae17, WKN^{+}95]. description [Yam04, ZP21b]. design [BPZR19, BH97, IKK^{+}04, SNV^{+}12]. Designing [PH11]. destination [KPW19]. detailed [ZP21b]. Detection [NTIO18]. deterioration [LRBJ21]. determinants [TA06]. determination [NDC05]. determine [Fra93, HEG08, TFB^{+}17]. determined [APR^{+}08, FODCN00, HHK^{+}10, MTP07, OFS^{+}16, PECG08, RHG^{+}13, SMB03b]. determining [DH11]. Development [HKM^{+}21, BWJ03, DL94, KD98, KTO^{+}11, LDH14, MPM19, QCR22, WJT97]. developmental [BMOT17]. devices [DBFW13, GCF^{+}21, GAH^{+}19]. dFADs [GCF^{+}21]. diagnosis [MLM^{+}98]. diamond [OKT^{+}23, OHM^{+}10]. diapause [TDT03]. Diatom [WB93]. Did [PW12]. Diego [Gre99]. Diel [GJR18, MTH^{+}04, SRR99, SE19, WMK^{+}99, BM99a, CCM^{+}08, HRB^{+}18, HHF09, SMB^{+}03a]. diet [DDS^{+}17, ESA^{+}16, HFF^{+}19, LK21, SKT21, SMF^{+}05, YKB08]. Diets [MLRS07, BDSM07, JCH05]. difference [LCC15, MTH^{+}04, Spr92]. Differences [MAH12, NZI95, OM10, ACG^{+}16, BWJ03, CP92, MCHSNEO13, OKT^{+}23, PJD14, SGL22, SCF^{+}20]. different [BDVS^{+}19, DDZ09, GFG98, GIT^{+}13, KT93, MATL98, QCR22, SLZ^{+}23, TA06, WQI00]. differential [FCL93]. differentials [AW92]. differentiate [GEGHPCC17]. differentiation [SMK^{+}13]. Differing [HGG^{+}17, IMS^{+}04]. diffusion [ÅGN^{+}04]. dimensional [APL01, HQW^{+}99, HNHP09, KU95, PML06]. Diomedea [XTC^{+}04]. direct [AMK08, BDBP93, HBC07]. direction [DLTI95, Sim96]. direction-finding [DLTI95, Sim96]. discontinuity [FKH^{+}17]. discovery [TTI^{+}20]. discrete [SF22]. discriminate [KN08]. discus [KTO^{+}11, TWK13, TKW^{+}17]. disentangle [RBPCR^{+}22]. Dispersal [EHW08, EvST^{+}17, REL07, SCAG^{+}21, GGF17, HZW^{+}98, KR10, MLP22, NSGL^{+}22, POA^{+}17, PHH13, PDER10, PEKL14, SES^{+}20, SMA14, THH12]. dispersion [BK94a, BK94b, BC97, ETB05, HLMS03, Kae17]. dissociate [FCJ^{+}15]. dissolved [JCCB15, KKK^{+}17]. Dissostichus [MMI^{+}22, PSS^{+}21]. distance [SAG^{+}09]. distinct [JJBCW09]. Distribution [AAI16, APL^{+}08, Aut08, BRR05, CLT05, CG18, DDZ09, HJ99, IK97, KEJK00, LC95, MDKS93, Mul94, OFS^{+}16, OKT^{+}23, QLB^{+}05, SME^{+}14, SKHI04, SKM06, SYT^{+}09, SCDA10, TMS^{+}08, Tak04, APL07, AAG11, AOVAG22, AS08, BJV^{+}17, BH18, BRFRJRLC18, BRPC08, BPLC11, BBB^{+}16, BRC^{+}03, BT99, BvDSDC18, Cap08, CAGPC21, CKA^{+}17, CDG^{+}19, RPG^{+}22, CGI^{+}19, Coy05, CMM06, D’A93, EBO04, FKF^{+}22, FKSA21, GP94, HT18, HGS^{+}21, Han11, HMM01, HDH^{+}05, HHH^{+}16, HJ10, HSLP19, HHF09, HGH93, HWSS07, HHK^{+}10, HMS16, HCWF21, ISI^{+}18, JCH05, JHK^{+}15, JCCB15, KvdPBW17, KMD^{+}09, KYU^{+}06, KIS01, KMM^{+}06, KM94, LLCJ16, LOS^{+}14, LS21, LJBR20, LTL^{+}22, LS15, LH96, LA05, LVPK11, LSD^{+}21, MBH^{+}99, MBJ^{+}07, MTP07, MFMG20, MDVB^{+}20, MP18, MTH^{+}04, MSC^{+}17, MCB^{+}16, MRHL09, MRBBHL14, MKH^{+}13, OTIK20, OHM^{+}10]. distribution [OA06, PLSO98, PMFC10, PLP^{+}11, Por22, RS15, RCG^{+}15, REM02, SA10, SRR99, SMK02, SHG12, SGL04, SL09, SAG^{+}09, SMS^{+}21, SADA^{+}23, SMH^{+}92, SSSB03, SHB^{+}11, SBBB03, SSPY08, Swa99, SB06, TSK^{+}92, TNM^{+}02, TKH08, TSK^{+}95, TDT03, TLS98, TTH15, VCB^{+}98, WRTP01, WM06, WMD^{+}06, WPL^{+}93, WJT97, WL21, WKN^{+}95, XH95, XWL^{+}23, YOYK20, YOK^{+}17, YLA13]. distributional [Neu02]. distributions [ACG^{+}16, AW92, BCBDA10, FCL93, HP02, IIS^{+}07, KTPM17, KWB^{+}16, LAB^{+}98, LBLCLC05, MCS^{+}06, MKH^{+}13, PP01, PML06, SF22, SLL19, Spe08, SRR05, TF08, WKR^{+}18, WEW98, YCS^{+}19]. Diurnal [WMD^{+}00, XMW^{+}23]. dive [FRS^{+}05, MIK07]. divergent [HSH^{+}22]. diverse [MWN^{+}23]. diversion [MFG99]. Diversity [RS15, ARM16, Bea03, FGGDSMF08, LPCG23, LéEPW^{+}12, PL03, SSM^{+}10, YMK^{+}15]. Diving [KKNY04, MIK07]. DNA [ARM16, BEF^{+}12, BBB^{+}19, KBB^{+}20, MWGK92, OK17]. Do [Gla11, MBE^{+}15, SMF^{+}05, WM06, DBFW13, GGQF22, HBLC22, Spr92]. Does [FPBDC11, Fra93, TFB^{+}17, HLH^{+}17]. dogfish [SPM02, YOK^{+}17]. dolphinfish [KR14, MESMM18]. dolphins [KFS22]. domains [MAH12, SMF^{+}05]. dominance [NFN00]. dominant [DTO^{+}23, éSMB20]. dominated [CFL^{+}99]. Doryteuthis [PS16]. Dosidicus [DLCQ22]. Doto [Yam04]. down [Gla11, GJR18]. downscaling [NFO^{+}23]. downwelling [MAH12]. dramatic [LK21]. drift [APLG07, EBFF17, HDH^{+}05, SCDA10, VHJ99]. drifting [GCF^{+}21, MSST16, UTMS06]. driftnet [YWM^{+}00]. drive [FRBB14, HSH^{+}22, NTIO18]. driven [ASK99, CRW20, HLWL12, Jan16, JR07, LHF^{+}99, MTLL^{+}16, OTIK20, REL07, SBD^{+}19]. drivers [AMDM12, BSF^{+}20, BDVS^{+}19, CMMK^{+}15, FPFL13, GPA^{+}21, HTLJ20, HPG^{+}20, HGG^{+}17, LSD^{+}21, MSL^{+}20, NLN^{+}21, RS15, SFL16, TSK^{+}22, THL^{+}18, VMT^{+}23, éSMB20]. drives [RBB^{+}21, Sha13]. driving [BBB^{+}16]. drum [GPS22]. Dual [KOS^{+}19]. due [MMF95]. dumerili [TNC^{+}22]. Dungeness [MAHG94, Sha13]. duration [BWJ03, HKLG07, MM94b]. during [AI05, BHC^{+}01, BPP07, BWKM15, CRVL^{+}17, CP92, DGB^{+}16, DHMT96, DTC06, ETB05, FDT^{+}99, FM93, FKSA21, FRZVHM^{+}11, HMM01, HQW^{+}99, HMS16, IUY10, JMP^{+}14, Jón99, KSM^{+}20, KSY^{+}23, KYU^{+}06, KK00, KB08, KNO^{+}04, LMB^{+}19, MRRN05, Mor11, MRHL09, MRBBHL14, Mul94, Mul97, MRD^{+}19, NASTF10, NFKY21, PSJF93, REB^{+}03, REG^{+}13, RCG^{+}15, REM02, SBT20, SDHB07, SCDA10, SLM13, SADA^{+}23, SSM^{+}10, SB04, TW05, TKO^{+}14, TLS98, VMG01, VDHF08, WBQL99, YWM^{+}00, YOIW21]. dusky [RHG^{+}13]. DVM [SSR13]. dwell [GS96]. Dynamic [BCJ^{+}13, HHTF10, MFMG20, HHB^{+}15, KFH00, MJH14, MLC^{+}98, RG97, XWL^{+}23]. dynamical [LAB^{+}98, SMDM98]. Dynamics [ABI^{+}21, Har92, SGFR^{+}21, SS19, APL^{+}08, APM^{+}12, BB03, BML^{+}14, BLH98, BPC^{+}16, CWCM14, DH11, DSHL18, Esc98, ECM^{+}01, FPBDC11, FBRB12, FRBB14, GSBB07, HMTG^{+}05, IXW^{+}10, KNE^{+}04, Kae23, KEWDA18, KKNY92, LCH03, LMB^{+}19, MWN^{+}23, NDC05, NK08, Ols01, PHH13, RCS98, REL07, RQN^{+}99, RKD^{+}20, RR18, SBY^{+}15, SOTM^{+}18, SK03, SKNT14, SP13, TAS04, UMK20, YKI98, ZZ93, ZYY^{+}21].

Earlier [CGI^{+}19]. Early [BCA^{+}18, HHH^{+}16, WSC05, ACT^{+}10, ACG^{+}16, ADPC21, BC04, BSF01b, CAR^{+}10, DHMT96, FYK^{+}21, GPS22, HMM01, HG98, HBO^{+}01, IUY10, KTO^{+}11, KR10, LPCA15, LGM^{+}02, LLB^{+}20, LMB^{+}19, LCC15, MBH^{+}99, MLVO05, MW92, MFP^{+}03, NFKY21, NHS^{+}07, NH06, Oda94, PSS^{+}21, RS15, ROH16, RWDA^{+}21, RAK^{+}17, RD96, SKHN11, SS19, SB94, SCDA10, SK03, SCF^{+}20, TWKW01, TTY^{+}23, THH12, WPL^{+}93, XWL^{+}23, YK96]. Early- [WSC05]. early-life [NH06]. earthquake [ONK17, TWK13, TKW^{+}17, KKK^{+}17, MTT^{+}17, NSH^{+}17, OKU17, ONK17]. East [DWH11, Jan16, MTT^{+}17, NSH^{+}17, OKU17, ONK17, PLT09, ÅGN^{+}04, Bea03, BUE02, BB07, DLTI95, GHV95, HA07, IIS^{+}07, MWP02, Sim96, SR02, SGHW05, WQI00, WQ00, HZTS12, IK97, KKH^{+}20, KKNY04, KMK^{+}18, MTLL^{+}16, MMI^{+}22, MMB^{+}11, MGHS14, NK08, OTH09, SKM06, SYT^{+}09, SBD^{+}19, TTC^{+}12, FH94, SB07]. Eastern [Esc98, APMRH17, APMVOGMR19, AOVAG22, BCBDA10, BHC^{+}01, BH18, BC97, BC04, BDAMD14, CSB94, CRW20, Coy05, DL94, DTC06, DABM^{+}06, FRS^{+}05, FMYN06, FYA^{+}21, GSNFL99, HB99, HFC01, HBO^{+}01, HLG^{+}11, ISI^{+}18, ISS02, JCH05, JPHA^{+}16, KMD^{+}09, KN08, KEWDA18, KKNY92, KBF^{+}07, LHM^{+}05, MPW^{+}99, MSS12, MDKS93, MAS^{+}98, MTH^{+}04, MSL^{+}05, NKS00, NH01, NK08, PJO99, Por22, SF22, SS94, SWS^{+}19, SDRL96, SCDA10, SGS^{+}06, Spe08, SBK^{+}01, SWZ^{+}01, SP13, TNC^{+}22, UMK20, WFRS93, YOYK20, YCH^{+}15, EvST^{+}17, HBLC22, HHH^{+}18, Kaw93, KO95, MMRH^{+}16, QCM^{+}16]. Ebre [LPSS04]. Ebro [LPSS04]. Ecological [KK00, WCP^{+}01, BBA^{+}21, CL05, MM94a, SPM^{+}19, ZHL^{+}03, ZWC^{+}21]. ecology [CC03, Hea93, HTT^{+}16, HS05, KNE^{+}04, LCCQ^{+}22, NPS^{+}23, NBF^{+}01, RDF^{+}11, WMD^{+}06, XTC^{+}04]. Economic [Dom09]. economically [FYC22]. economy [RKZHC19]. Ecosystem [AS08, CAB^{+}01, EBFF17, HTLJ20, PFB^{+}16, BO05, BBA^{+}21, CW98, CGMM10, aTCK05, CMS16, DPK^{+}08, FPFL13, GSBB07, HHK^{+}17, HHH^{+}16, HHB^{+}15, IMO^{+}12, IKK^{+}04, KTS15, KCW^{+}15, LRL^{+}06, MTL^{+}22, MWR^{+}98, NH03, ODMRM98, OUKH04, PCR^{+}18, RD96, RKZHC19, SGFR^{+}21, SPLY23, SMF96, SHM05, SMS^{+}19, SP15, TJW^{+}03, TB92, YMK^{+}15, AAI16, AJ15, FMM^{+}20, GAH^{+}19, KYA^{+}15, NPS^{+}23, PMG^{+}23, VMT^{+}23]. ecosystem-based [HHK^{+}17, HHB^{+}15]. Ecosystems [FC04, Har92, CHHS05, DDZ09, FH94, FHHW98, Gre13, PO03, TFB^{+}17]. Ecuador [HMTG^{+}05]. eddies [ADAHL10, HBR^{+}15, KBB^{+}20, LS01, SS94]. eddy [HTL^{+}00]. edeni [MTK^{+}07]. edge [PKP^{+}00, RHP^{+}15, SMF96, TDE09, WKN^{+}95]. editor [CW94]. eDNA [MWN^{+}23]. edulis [ITH23, YAM^{+}18]. Edwards [SCTB19]. edwardsii [FML^{+}14, HGG^{+}17, LJM^{+}10]. eel [AM18, BCR08, BBT^{+}09, CHM^{+}94, HZTS12, HXC^{+}17, KSY^{+}23, KIS01, SOTM^{+}18]. eels [AM18, CSS^{+}21, KMM^{+}06]. Effect [BGP^{+}06, HSLP19, HWSS07, IST^{+}23, KNS^{+}22, PGL^{+}15, SSW^{+}17, TAN^{+}17b, TY04, ASK99, BCR20, BMOT17, CB93, ETB05, HBPC15, JMP^{+}14, KJZ97, KIS01, OKU17, PVMP03, PW14, SPG^{+}16, SB07, SSSB03, THH12, WL21]. effective [BHM02]. effectiveness [LVF12]. Effects [AYK03, DB93, FHHW98, HCS^{+}09, KvdPBW17, KTO^{+}11, LRS^{+}23, MCM^{+}17, MTT^{+}17, NSH^{+}17, OS95, PLSO98, RTK01, SKHN11, SS98, TW05, TKW^{+}17, TGRS^{+}19, TTH15, APL^{+}96, AHKP16, AMK08, BJV^{+}17, BB03, BH18, BBH99, BYM16, CSFC05, DHC^{+}07, Dom09, FYA^{+}21, GEGHPCC17, HKWL17, HTL^{+}00, HP02, HHF09, HFF^{+}19, HAS^{+}19, HCWF21, HK06, JCH04, JHK^{+}15, KOS^{+}19, KK00, KKCL06, KM93, LAFF15, LDH14, LS15, MAH12, OR13, ONK17, OCCF^{+}18, Par95, PJB05, SFGE21, SPLY23, SNV^{+}12, SSPY08, SP15, Swa99, TDE09, TB92, UMK20, VFS^{+}24, WHT92, WMD^{+}06, WGW07, XB09, XDP^{+}20, YWI^{+}05, YOIW21, ZHX^{+}20]. efficiencies [Bau95]. efficiency [FCJ^{+}15]. effort [BHM02, Dom23, MTSH15, NLN^{+}21, SSW^{+}17, VHCN14, Wat17]. Egg [IIS^{+}07, AMK08, BCBDA10, BDTR23, COSC97, ICB^{+}08, KNS^{+}22, KBB^{+}20, KL01, LVF12, MMI^{+}22, PSC05, PML06, RJHC99, SGS^{+}06, TYO21, TMN^{+}15, VCB^{+}98, WZK97]. eggs [ÅGN^{+}04, BBMY93, BRC^{+}03, BSF01a, CAR^{+}10, Cur04, CCP07, Dd95, HJR^{+}03, HBG^{+}16, IK97, IYN^{+}09, III^{+}06, KBB^{+}20, LS21, LVF12, MOE06, MHM^{+}20, NLS^{+}24, NYI^{+}13, PVMP03, PSS^{+}21, SBBB03, SFK^{+}20, TF08, TKMS11, VCB^{+}98, WJT97]. Eighth [Liv00]. electronic [AMD^{+}16, KSMY00, NHNA07]. elegans [BT99, TSK^{+}95]. Eleginops [QM01]. elemental [LCC15]. Elephant [SRCV09]. Elevated [HLH^{+}17, KTO^{+}11]. elongatus [ARL93, MKF^{+}03]. elver [Jes22]. embayment [CP92]. embedded [AYK03]. emergence [TDT03]. Emiliania [HGH93]. Emperor [LRS^{+}23]. emphasis [MBY^{+}17, YKI98]. Empirical [JPMH20, NY03]. encouraged [KSY^{+}23]. encrasicolus [AB02, ACT^{+}10, APL01, APGL03, APLG07, APL07, BGP^{+}06, BBP^{+}13, BPP07, BUE^{+}98, BFSV08, BRC^{+}03, CPM^{+}15, GIT^{+}13, GöEIOS16, HBG^{+}16, ICB^{+}08, LVC^{+}05, LPSS04, MYHvdL15, MFP^{+}03, PBL07, RGQPN09, SSP^{+}07, ZVKŠ13]. endogenous [DDB^{+}20]. energetic [SPLS15]. Energy [LMB^{+}19, CHF^{+}04, PSM00]. England [XMH^{+}18, PWML12]. English [Bow11, IH03]. Engraulidae [SKNT14]. Engraulis [AB02, ACT^{+}10, APL01, APGL03, APLG07, APL07, BGP^{+}06, BBP^{+}13, BPP07, BUE^{+}98, BFSV08, BRC^{+}03, RPG^{+}22, CRVL^{+}17, Cur04, CCP07, CPM^{+}15, DBGW04, DBS^{+}19, FYK^{+}21, GNP^{+}19, GIT^{+}13, GöEIOS16, GSBB07, HMM01, HJR^{+}03, HSLP19, HBG^{+}16, ICB^{+}08, IK97, IYN^{+}09, ISN^{+}11, KL01, LVC^{+}05, LC95, LPSS04, MSM^{+}13, MYHvdL15, MFP^{+}03, PVMP03, PBL07, RCB08, RGQPN09, SSP^{+}07, SLL19, TWKW01, TW05, TCL^{+}12, TA06, TMN^{+}15, TTC^{+}12, WMD^{+}06, ZKT07, ZYY^{+}21, ZYT^{+}22, ZHL^{+}03, ZVKŠ13]. enhance [SBD^{+}19]. Enhancing [HHB^{+}15]. enrichment [LRL^{+}06]. Ensemble [WB93, CW98]. Enshu [NFN00]. Enshu-nada [NFN00]. ENSO [FYC22, HSLP19, LBLCLC05, OBA01]. Entrainment [MMB^{+}11, MGHS14]. entropy [WKR^{+}18]. environment [APL07, AAKMG06, BDE^{+}19, BRN^{+}95, Bea03, BBB^{+}16, BUE^{+}98, BSF01a, BvDSDC18, Buc92, CB93, CHM^{+}94, DBFW13, GPS22, HBG^{+}16, KFYP07, LHM^{+}05, LLCV18, LOS^{+}14, LCCdS^{+}19, NKS00, NDC05, NII^{+}14, Nis19, QCM^{+}16, Ree95, RBB^{+}21, RGQPN09, RWLP12, SA10, SBK^{+}01, SPLS15, ST97, ST98, WMD^{+}00, WGFR06]. environment-based [RWLP12]. environment-recruitment [GPS22]. Environmental [BJV^{+}17, BB03, BBH99, BBB^{+}16, BUE02, BDVS^{+}19, Col99, Dom09, DHM^{+}15, EPG^{+}16, FML^{+}14, HMP92, ISN^{+}11, MEK^{+}09, MESMM18, MTSH15, NLN^{+}21, OWK04, PHH^{+}98, PBF00, RF07, RMH^{+}19, SFGE21, SHK^{+}19, SZX^{+}08, SSPY08, TA06, VHCN14, VGPL^{+}11, VDHF08, YWI^{+}05, ZVKŠ13, AUOGMM19, ADPC21, AGS^{+}04, AMDM12, ABS^{+}11, AS08, BKvdP^{+}22, BFF15, BHV^{+}06, BSF^{+}20, BLH98, BCR08, BDSM07, CLPC18, CLW^{+}19, CLT05, CH95, CAB12, DPK^{+}08, DH11, DBB^{+}18, DGB^{+}16, DPL02, ERE^{+}10, Erz05, FCJ^{+}15, FYK^{+}21, GCQ^{+}13, GEGHPCC17, HKWL17, HBLC22, HHF09, HPG^{+}20, HGG^{+}17, HVHC10, HBN^{+}21, HCC^{+}09, HALO00, IFF^{+}18, IYN^{+}09, JCCB15, KvdPBW17, KEJK00, KYSM11, LAFF15, LPCG23, LLSF01, LML^{+}03, MTP07, MSM^{+}13, MMSL19, MPM19, MBY^{+}18, MMRH^{+}16, MHB^{+}14, MWP02, NPS^{+}23, NYI11, OHF12, PM95, PJB05, PGL^{+}15, RF04]. environmental [RS15, RPC^{+}19, SME^{+}14, SGFR^{+}21, SC05, SFL16, SEM^{+}14, SCF^{+}20, SRR05, Swa99, SB06, TSK^{+}22, TKO^{+}14, TSG^{+}20, TAN^{+}17b, TCC^{+}98, WMKR09, WQ00, WJW20, XB09, YOIW21, YIT^{+}22, ZWL21]. environmental/physiographic [KEJK00]. Environmentally [CRW20, HBPC15]. environmentally-explicit [HBPC15]. environments [FMM^{+}20, HLMS03, TNK^{+}16]. environs [AI92]. Epinephelus [OE17]. epipelagic [PFAM96, TSK^{+}95]. epiplanktonic [HL98]. episodic [BKvdP^{+}22, BO05, IHS97, ZLTM11]. Equatorial [HXC^{+}17, KIS01, Dom23, HJ10, LAB^{+}98, LCCdS^{+}19, MSST16, MHB^{+}14, SMDM98]. Errata [Ano00a, Ano02]. Erratum [Ano00b, Ano14, Woo97]. error [AW92]. Essential [DWHdP21, CLM^{+}21]. Establishing [BBY08]. estimate [BFF15, BHM02, CC03]. estimated [APL01, IYN^{+}09, MTH^{+}04, YOY00]. estimates [CCM^{+}08, PP01, PS06, QCR22, RMM02, ZHT14]. Estimating [FKSA21, MFH05, PH11, Gla11]. Estimation [DWH11, III^{+}06, KOKM15, SP93]. estimations [GiIW^{+}20]. estuaries [BWK^{+}99, RS92]. estuarine [BHJ^{+}04, DMF^{+}17, DHM^{+}15, FKUY16, HSH^{+}22, MLVO05, MW92, NH06, SS19, SHG^{+}22, SGL22, SKNT14, YOY00]. estuarine-dependent [MW92]. estuary [ASCM12, CFL^{+}99, MW92, QBMW99, REL07, SAO^{+}17, SQW^{+}99, SKNT14, YLA13, XWL^{+}23]. Ethmalosa [BDE^{+}19]. Etrumeus [VCB^{+}98]. Eucalanus [TSK04]. Eulerian [GP94]. Eumetopias [CL05, FRS^{+}05, SMF^{+}05, TMM^{+}07]. Euphausia [MAS^{+}98, SRCV09, Tak04, TBB^{+}03]. euphausiid [RMM02, Tan02]. euphausiids [PMG^{+}23, Tan99]. Europe [Ano99, BUE02]. European [AB02, ACT^{+}10, ACG^{+}16, AH97, AM18, BGP^{+}06, BBP^{+}13, BCR08, BBT^{+}09, CHM^{+}94, CSS^{+}21, DWHdP21, GIT^{+}13, GI13, HB99, Jes22, LCCQ^{+}22, MOE06, PVBV19, PWE98]. eutrophic [UIU^{+}99]. evaluate [OIA^{+}12]. evaluated [VFS^{+}24]. Evaluating [DDB^{+}20, GCW17, HHF09, OCCF^{+}18, PS06, XMH^{+}18, HBPC15, JPHA^{+}16]. Evaluation [SSP^{+}11, AJ15, AI04, CWCM14]. event [MPW^{+}99, PMG^{+}94, REM02]. events [BO05, KNE^{+}04, LBLCLC05, MHG^{+}11, SES^{+}20, Sim92b]. Evidence [BMO^{+}99, JTYB18, KKCL06, KMM^{+}06, MTP07, NNOU20, SCS05, MPM^{+}13, SWS^{+}19]. exacerbate [HLH^{+}17]. examined [DPM^{+}11]. example [AB02, FIDC00, SHB^{+}11]. exceptional [ARL93]. Exchange [GS99, HBO^{+}01, KKK^{+}17, QLB^{+}05, SHS^{+}23]. Exclusive [Dom09]. exert [Gla11]. exhibit [RAK^{+}17]. exogenous [DDB^{+}20]. expansion [HGS^{+}21, TKW^{+}17]. expansion/contraction [HGS^{+}21]. experienced [FHD98, RFD^{+}04, WMD^{+}00]. Experiment [OCH99, BAL^{+}99, MEK^{+}09, OA06, ZWL21]. experiments [IYN^{+}09, NYI^{+}13, YAM^{+}18]. explain [ABI^{+}21, BMPC16, FKH^{+}17]. explaining [HA07]. explicit [FGS95, GYS14, HBPC15, MLVO05, PDD03]. exploitation [DH11, FCJ^{+}15, FRBB14, RR18]. exploited [BEF^{+}12, HMTG^{+}05, HRS^{+}21, PFSL09]. Exploring [GGF17, BM99b]. export [CAR^{+}10, NLS^{+}24, TKM^{+}22]. exposed [YOY00]. extant [MPM^{+}13]. Extended [SPM^{+}24, RP93]. extending [MRL^{+}14]. Extension [NIIS04, NY08, NY03, SHK^{+}19, YW07]. extensive [AM18]. extent [BEF^{+}12]. extremes [MCG^{+}14]. exulans [XTC^{+}04]. Ezo [KTO^{+}11].

factor [DHMT96, NNOU20, FCC^{+}19]. factors [ABS^{+}11, AS08, BUE02, BDTR23, CLPC18, CLT05, EPG^{+}16, FYK^{+}21, FYK^{+}13, HQH^{+}06, INM^{+}18, LPCG23, LAPL21, LAG^{+}11, MTP07, MHB^{+}14, OWK04, PM95, Spe08, TKO^{+}14, VDHF08, WKB^{+}05, YOIW21, YIT^{+}22]. FAD [GAH^{+}19]. FADs [DBFW13, MSST16]. failure [VGPL^{+}11]. Falkland [AGS^{+}04]. fall [ESA^{+}16, HMT07, WBQL99]. fallacy [Bau98]. fallax [LAFF15]. False [ZP21a]. Family [WMK^{+}99]. far [HKA^{+}06, SDRL96, Kaw93, KO95]. far-eastern [SDRL96]. far-ranging [HKA^{+}06]. Farfantepenaeus [MCB^{+}16]. farm [KNK^{+}18]. Faroe [Hea99b, HJ99, Jón99, RJHC99]. fast [BBT^{+}09]. fatness [HFF^{+}19]. faunal [LBLCLC05]. favorable [YKH^{+}21]. features [CG18, DDB^{+}20, FRS^{+}05, HSH^{+}22, LJH^{+}05, MJH14, MFB^{+}09, Sco95, SHB^{+}11, WFRS93]. Feeding [FBRB12, MATL98, MFRR96, WLWZ98, BT99, CC03, DDB17, DPL02, HTT^{+}16, KNE^{+}04, KKNY04, KNO^{+}04, MVK^{+}20, NKS00, NII^{+}14, PHWM96, RAT^{+}02, SMB^{+}03a, SSR13, SK04, SKNT14, TNM^{+}02, VDHF08, YKH^{+}21]. female [BMOT17]. ferruginea [SCS05]. fertilised [PSS^{+}21]. fertilization [KTS15]. fictitious [BWK^{+}99]. fidelity [CLH^{+}22]. Field [HDF^{+}99, BRC04, FMYN06, IUY10, JR07, OA06, PP01, TKH08, VHJ99]. Fifth [Kas97]. filter [SMB03b]. fimbria [GJR18, KMB00, SC06, SE19]. fimbriata [BDE^{+}19]. finding [DLTI95, Sim96]. Fine [Cur04, SKNLD10]. Fine-scale [SKNLD10]. finmarchicus [Ano99, BM99a, BHH98, CW98, GMH^{+}99, HTE^{+}03, Hea99b, HBR^{+}99, HJ99, HDF^{+}99, IHHH99, Jón99, MLC^{+}98, NGGJ09, PHH13, RCS98, RJHC99, SGHW05, TDT03, VJ99]. finned [DHC^{+}07, KOKM15]. First [Jan16, ZP21b, AHKP16, ABG19, BMPC16, BEiI^{+}23, MIK07]. First-year [Jan16]. Fish [DWHdP21, JMLG06, KGW13, Nak98, REM02, AI92, ASK99, ABS^{+}11, ARM16, BB03, BH18, BML^{+}14, BJCS12, BCJ^{+}13, BRFRJRLC18, BEF^{+}12, BS94, BB07, Buc92, CLM^{+}21, CHPT20, COSC97, CÅP^{+}13, CFL^{+}99, CH92, CAR^{+}10, DBFW13, DLD^{+}23, DPL02, ESA09, ERR^{+}21, FRP^{+}99, FCL93, FvPH^{+}16, FKSA21, FKH^{+}17, FRHMAM^{+}06, FRZVHM^{+}11, GQPGA04, GCF^{+}21, GAH^{+}19, GDM^{+}17, HHF09, HPG^{+}20, HNHP09, HLMS03, HPL13, HLWL12, HCFP20, IIS^{+}07, IKK^{+}04, JMP^{+}14, KN08, KSC^{+}10, KBB^{+}20, LLCJ16, LVF12, LVM^{+}18, LéEPW^{+}12, LH96, LSD^{+}21, MBY^{+}17, MBY^{+}18, MSR20, MTZG23, MHG^{+}11, MCS^{+}06, MRHL09, MRBBHL14, MBKP08, MSVY^{+}13, MMB^{+}11, NLS^{+}24, OKU17, OEV^{+}10, PP01, PJO99, PST03, PDD03, PSC05, PLT09, PML06, PRDC^{+}13, PFSL09, PJB05, PLP^{+}11, QM01, Ree95, RPT^{+}00, RAT^{+}02, REG^{+}13, Rob94, RCG^{+}15, RSC96, Rog94]. fish [RG97, SBY^{+}15, SGFR^{+}21, SS19, SKKW02, SKHI04, SKM04, SBT20, SES^{+}20, SHG^{+}22, SCKJ^{+}18, SFL16, SC97, SRR05, SPT^{+}17, Tan02, TAN^{+}17b, TGRS^{+}19, TFB^{+}17, THH12, TTC^{+}12, TTH15, VN97, VCB^{+}98, VAFG95, WHT92, WKR^{+}18, WEW98, XMW^{+}23, YMK^{+}15, Zam01, óóSV18]. fished [OHS06]. fisherie [SMS^{+}23]. Fisheries [BB02, CAR^{+}10, FC04, ONK17, Par95, RBPCR^{+}22, War92, AAI16, Bau98, Bri94, BHS^{+}15, CIS20, CMS16, DSHL18, DTC06, ERR^{+}21, Erz05, EPG^{+}16, FMV03, HA07, HHK^{+}17, HSEH16, JCH04, JPHA^{+}16, KD98, KPW19, LAG^{+}11, MKF^{+}03, Par96, Ric96, RS92, SHG^{+}22, Sch23, Sha95, Sim92a, SSPY08, SR93, SP15, Tyl92, VOB^{+}19, XTC^{+}04, dBdOJdO^{+}22, KYY00, BEiI^{+}23, Kim23]. Fisheries-based [RBPCR^{+}22]. Fishery [CMB^{+}15, DL94, AG99, And03, BBH99, BLG^{+}16, Cap08, CMMK^{+}15, CSB94, CCHL23, DWHdP21, DLCQ22, Dom09, DMH16, FCJ^{+}15, GYS14, GEGHPCC17, HGG^{+}17, HHTF10, HBR^{+}15, HDJ15, KB08, MPM19, MDR^{+}16, MMRH^{+}16, NFN00, Nis92, PVHT01, SR02, SS98, VIS92, ZWL21, ZSY^{+}21, ZHX^{+}20]. fishes [BBB^{+}19, EBO04, GP94, GS99, HALO00, KCW^{+}15, MTL^{+}22, MFS^{+}17, MSC^{+}17, MFB^{+}09, PM95, PG06, QLB^{+}05, RS15, SMK02, SNL19, WM06, WMK^{+}99, WK03]. fishing [ASM^{+}15, BSF^{+}20, BHM02, DSPH07, EBFF17, GAH^{+}19, HKLG07, ITH23, KFYP07, KY17, LPS19, LVM^{+}18, LAPL21, MHS^{+}21, Par95, PVBV19, PVHT01, PBF00, PKP^{+}00, RKZHC19, SSW^{+}17, SPLY23, SLZ^{+}23, SNL19, YW94, YK96, YWM^{+}00, YOIW21, ZSS08]. fitness [FGS95]. fitness-based [FGS95]. fixed [NH06, SRR07]. fixed-location [NH06]. fjord [ASK99, KR10]. fjords [APM^{+}12, GV01, VAFG95]. flatfish [DMF^{+}17, HLH^{+}17, NBH99, SLM13, Spe08]. flights [HKA^{+}06]. float [YW07]. floating [DBFW13]. Florida [CMMK^{+}15, Dom04, EF10, KBB^{+}20, RCPS09, WMKR09]. flounder [DCLC15, KUO^{+}17, RKZHC19, SSW^{+}17, SCS05, XMH^{+}18, YTY96, YOY00]. Flow [JR07, BEF^{+}12, KM94, RSF13, SAG^{+}09]. Flow-field [JR07]. flowing [SAO^{+}17]. fluctuating [DDB17]. Fluctuation [KIS01, TCC^{+}98, KJZ97, OE17, TID^{+}96]. Fluctuations [BCR08, LLSF01, ASCM12, ABI^{+}21, Bea03, BPP07, BAL^{+}99, FYC22, Gar97, HBR^{+}99, HEG08, KO95, LLB^{+}20, LBSS^{+}92, MMRS16, RF04]. flux [GS99, JCH04, Ste98]. fluxes [VZP98]. flying [ASM^{+}15, IMS^{+}04, ISI^{+}18, LCC15, NII^{+}14, NTM^{+}15, YWM^{+}00]. focus [BB03]. folk [FvPH^{+}16]. following [MTZG23]. Food [WS08, BCL04, DMF^{+}17, HLMS03, NHM94, NZI95, NNOU20, PDD03, PAS^{+}18, RJHC99, SPV96, SP15, TW05]. food-limited [BCL04, NNOU20]. foods [YKH^{+}21]. footed [MJH14]. Forage [PBH^{+}04, Dom09, LPCA15, LAB^{+}98, LVM^{+}18, PJO99, SBT20, SPT^{+}17, TGRS^{+}19, Zam01]. Foraging [Wil01, XTC^{+}04, JPMH20, LJH^{+}05, MCHSNEO13, NPS^{+}23, SRCV09, SAG^{+}09]. forced [DST11, TF08]. Forcing [BBA^{+}21, AH97, ADPC21, ABS^{+}11, ERE^{+}10, GQPGA04, OHF12, PA14, RGQPN09, SGFR^{+}21, Sha13, SCKJ^{+}18, SCS05, TMM^{+}07]. Forecast [SMS^{+}23, CH95, MPM19]. Forecast-ready [SMS^{+}23]. Forecasting [BML^{+}14, NPY^{+}15, SW05, HSEH16, KWB^{+}16, PHH^{+}98, RWLP12, WQ00, YW94]. forecasts [CRW20, GYS14, HBN^{+}21, PST03]. forest [KM94]. Foreword [Ano03a, CHPA98]. form [KOKM15]. Four [Bow11, LBSS^{+}92, MHS^{+}21, PLSO98]. Fourth [Woo97]. framework [LPG^{+}06, OCCF^{+}18]. franciscanus [MWB^{+}00]. Fraser [Sim96, APL^{+}96, DLTI95, McK13, MCG^{+}14, MFG99, PW12, PW14, RFM^{+}21, SMH^{+}92, TIH^{+}92, TH11, XDP^{+}20]. frequency [PP01, PHH13, SRR07]. fresh [HQH^{+}06]. freshwater [HQH^{+}06, LPSS04, QM01, WSF^{+}14]. front [GS99, HJR^{+}03, LSW^{+}03, MSM^{+}13, VCKH05, YW94, KT93, MMB^{+}11, MGHS14]. frontal [BBR^{+}05, ISN^{+}11, MIY^{+}09, NZI95]. fronts [AAI16, BGM^{+}18, CMB^{+}15, KFH00, OR12, PKP^{+}00, RSC96, SGL04, UYF92]. FRS [BD93]. fry [ZZ93]. fuegensis [ADPC21]. Fuegian [ADPC21]. Fukushima [Kae17, MFS^{+}17, SSW^{+}17, SAO^{+}17]. Fukushima-derived [Kae17]. Fulmar [BMH^{+}21]. function [TD02, Zam01]. functions [QCR22]. Fundamentals [Gre99]. Fundy [JR07, SPM02]. Funka [KTH^{+}15]. furnieri [ASCM12]. further [SWS^{+}19]. fuscus [HMTG^{+}05]. future [JYH^{+}18, MDVB^{+}20, NFO^{+}23, SLL19, Sim92a].

G [Sim96]. gadid [WL21]. gadoid [LOS^{+}14]. Gadus [AHKP16, AMK08, BCGB14, BSF01a, BTGM07, BCL04, CRC11, D’A93, DB03, FODCN00, GRT^{+}07, GCW17, HBPC15, HL07, HBO^{+}01, HCS^{+}09, KNS^{+}22, KTH^{+}15, KEWDA18, KR10, LK21, LBW^{+}05, Lou10, MRD^{+}19, NSH^{+}17, Neu02, NHNA07, OTIK20, OHS06, RKD^{+}20, SHG12, SC05, SB07, SADA^{+}23, SB04, Swa99, TNM^{+}02, TLS98, VSÅO07, VHJ99, WJT97, WKN^{+}95, YCH^{+}15]. Gadusmorhua [IHS97]. gahi [AGS^{+}04]. gain [NFO^{+}23]. Galápagos [HMTG^{+}05]. Galician [LCCQ^{+}22]. gastropod [KTO^{+}11, SPM^{+}24]. gauntlet [JPMH20]. gear [PBF00, SB94]. gears [HKLG07]. GEE [CIS20]. gelatinous [BMO^{+}99, GBAD^{+}17]. gene [BEF^{+}12]. general [AYK03, LBW^{+}05]. Generalized [HHF09, MTP07, FODCN00, YOK^{+}17]. generated [BWK^{+}99, MHG^{+}11]. Generation [RP93, BZ21]. genetic [CPM^{+}15, KPHG14]. genetics [HRS^{+}21]. geochemical [LAB^{+}98]. Geographic [KMK^{+}18, LAB^{+}05, MVK^{+}20, Mar01, OKT^{+}23, SB06]. Geographical [UIU^{+}99, FKSA21, Sim92a]. geography [BvDSDC18]. Geolocation [GRT^{+}07, NBMS06]. geomagnetic [BA12]. George [RRF^{+}21]. Georges [MLM^{+}98, BCL04, LBW^{+}05, Lou10, MLC^{+}98, NGGJ09, PSN^{+}99, PJD14, TCS^{+}09, WPL^{+}93]. Georgia [GDM^{+}17, PMT^{+}94, WKB^{+}05, BRN^{+}95, MWR^{+}98, SMA14, TBB^{+}03, XTC^{+}04]. geostatistical [RMM02]. geostrophic [RPT^{+}00]. German [BK94a, BK94b, HSS19, SGN^{+}05]. giant [POA^{+}17]. Gibraltar [GEGHPCC17, NSGL^{+}22, SFGE21]. gigas [DLCQ22, KSM^{+}20, YIT^{+}22]. gillnet [EBFF17]. glacial [APM^{+}12]. Glacier [APL^{+}08]. gladius [SKNLD10, SAH^{+}18]. Glass [SOTM^{+}18, CSS^{+}21, KSY^{+}23]. glauca [GPCGdlT^{+}22, HRB^{+}18]. Glaucosoma [BEF^{+}12]. GLM [CIS20]. GLMM [CIS20]. Global [HB92, LMBL03, SMS^{+}23, Sim92a, XH95]. GLOBEC [Ano03a, CHPA98]. Globicephala [KOKM15]. go [HBLC22]. goby [SBY^{+}15]. goes [GJR18]. going [RSF13]. golden [NLN^{+}21]. Goldsinny [CLH^{+}22]. good [Sha95, UYF92]. gorbuscha [BWS^{+}01, CAB^{+}01, FYA^{+}21, MAH12, RZM^{+}03, TID^{+}96, WCP^{+}01, Wil01]. gradient [SS19]. gradients [APM^{+}12, MBY^{+}18, Mor11]. Gran [BAB^{+}06, MRHL09]. gray [BASS11]. grazing [RP93]. grazing-extended [RP93]. Great [MTT^{+}17, OKU17, CLPC18, KUO^{+}17, MM94a, NSH^{+}17, ONK17, RHP^{+}15, TR11, WMD^{+}06]. greater [TNC^{+}22]. Green [SMF96]. Greenland [MFRR96, ÅGN^{+}04, DDS^{+}17, SL95, SCDA10, SB04, YLA13]. grey [KSAF13]. gridded [ZSY^{+}21]. grooved [BYM16, KBS^{+}16]. gross [RS92]. ground [ASM^{+}15, ABI^{+}21, FKSA21, HONH04, IK97, III^{+}06, LSK^{+}18, MHB^{+}14, PVHT01, RCPS09, SAT^{+}18, TTI^{+}20, WZK97, YW94, YKH^{+}21, ZSS08]. Groundfish [JJBCW09, SSM^{+}10, DTO^{+}23, GHV95, MSS12, YCS^{+}19]. groundfishes [HCWF21]. grounds [DSPH07, GöEIOS16, ITH23, KUO^{+}17, PVMP03, PKP^{+}00, QBMW99, RRF^{+}21, SHK^{+}19, SLZ^{+}23, WZK^{+}98, YTY96, YOY00, YW94, YK96]. group [KSAF13]. grouper [OE17]. Growth [ACT^{+}10, AM18, CRVL^{+}17, IUY10, MHS^{+}21, OWK^{+}03, RBBG12, TWKW01, TNK^{+}16, AHKP16, AYMK01, ACG^{+}16, APGL03, APLG07, BC04, BMPC16, BHV^{+}06, BBY08, BASS11, BCL04, BWS^{+}01, DPK^{+}08, DBS^{+}19, DPL02, DB03, ERR^{+}21, FYA^{+}21, FYKSP07, GHBM99, GCQ^{+}13, HFHW19, HPG^{+}20, HBC07, HVHC10, HFF^{+}19, HAS^{+}19, ISN^{+}11, JTYB18, KNS^{+}22, LDH14, LDDC06, LMB^{+}19, LBW^{+}05, MRRN05, MBJ^{+}07, MSS12, MSL^{+}20, MMMS14, MRD^{+}19, NFKY21, NHS^{+}07, OTH09, OIA^{+}12, OWK04, PDD03, PA14, SKHN11, SKT21, SPG^{+}16, Sko05, SCF^{+}20, TW05, TCL^{+}12, TSK^{+}22, TA06, TTY^{+}23, Tan17a, TKM^{+}22, TY04, TB92, VSÅO07, WGW07, WGS^{+}08, WSC05, XDP^{+}20, YCH^{+}15, ZJH^{+}22, ZNI96]. Growth-dependent [IUY10]. Guam [KPW19]. guide [PST03]. Gulf [MCB^{+}16, TMMM20, DBGW04, AUOGMM19, ADAHL10, AOVAG22, BBMY93, BPZR19, BASS11, BT99, BDVS^{+}19, BBB^{+}19, BPS^{+}14, CM10, CRC11, CP03, D’A93, DCLC15, DGB^{+}16, ERR^{+}21, GS99, GBAD^{+}17, GRT^{+}07, GCW17, HDH^{+}05, HBPC15, IN00, IXW^{+}10, KNE^{+}04, KPHG14, KSP^{+}22, KR14, LK21, LDAWM10, MSS12, MTZG23, MLM^{+}98, MM03, MSL^{+}20, MSC^{+}17, MLR10, NLS^{+}24, PGL^{+}15, ROH16, RBBG12, RFM^{+}21, RBB^{+}21, RR18, RD96, RCD^{+}99, RKZHC19, SGW^{+}21, SCAG^{+}21, SGL04, SCTB19, SMS^{+}19, SJB^{+}22, Swa99, SB06, TGRS^{+}19, VHCN14, WFRS93, Wil04, XMH^{+}18, YCS^{+}19, YLA13, éSMB20]. gulfs [RRF^{+}21, LSD^{+}21]. Gunnerus [IHHH99]. gurnard [KSAF13]. gut [DDS^{+}17, NKM01].

habit [SK04]. Habitat [APMRH17, CGMM10, DWHdP21, FFF^{+}18, GPL^{+}11, HTE^{+}03, HKLG07, HHH^{+}18, KR14, Mar01, MSNK10, PLT09, SMK^{+}13, AB02, BPZR19, BGP^{+}06, BLH98, BRPC08, BHM02, CHPT20, CLW^{+}19, COW^{+}99, CH16, DWH11, DMF^{+}17, DSPH07, EBFF17, FRS^{+}05, FYC22, FKF^{+}22, FHK^{+}10, GIT^{+}13, GCW17, HHK^{+}17, HLG^{+}11, HCWF21, HHB^{+}15, ISI^{+}18, KOKM15, KMD^{+}09, KSAF13, LOS^{+}14, LPS19, LDAWM10, LMBL03, LPG^{+}06, Lyn03, MCHSNEO13, MSR20, MFMG20, MDVB^{+}20, MHRC18, MYHvdL15, MJH14, NASTF10, Nis19, PMFC10, PBL07, PBH^{+}04, PG06, PLG^{+}10, RFD^{+}04, RCB08, RHG^{+}13, RHP^{+}15, SFA14, Sco95, SLL19, SDHB07, SGS^{+}06, SSP^{+}11, SRM^{+}18, SB06, WM06, ZSY^{+}21, ZWC^{+}21]. habitat-based [BHM02]. habitats [BHS^{+}15, CLM^{+}21, GTB10, HCFP20, JJBCW09, KYS15, LPHM21, RBBG12, SPV96, SJB^{+}22, TFB^{+}17, VOB^{+}19, VPRG13]. habits [TNM^{+}02, WS08]. haddock [BSF01b, BCL04, GHBM99, HG98, LOS^{+}14, LSK^{+}18, PSN^{+}99, PJD14, PA14, WPL^{+}93]. Haimovici [LAPL21]. hairtail [SCF^{+}20]. hake [BKvdP^{+}22, CC03, DDB^{+}20, GI13, IMO^{+}12, LCCQ^{+}22, MMSL19, MOE06, PVBV19, PMG^{+}23, RPC^{+}19, SRR07, SMA14, Tan99, TMMM20, VMT^{+}23]. hakes [KvdPBW17]. half [SB04, War95]. halibut [ÅGN^{+}04, FKF^{+}22, HAS^{+}19, RKZHC19, SME^{+}14, SGW^{+}21, SCDA10, YLA13]. Haliotis [KTO^{+}11, TWK13, TKW^{+}17]. Halocyprididae [LTL^{+}22]. hannai [KTO^{+}11, TWK13, TKW^{+}17]. Hansen [MAS^{+}98]. hard [Gre99]. harengus [BML^{+}14, BDTR23, FPBDC11, FM93, GPA^{+}21, MLVO05, NDC05, Neu02, SNV^{+}12, óóSV18]. harvested [SPM^{+}24]. Harvey [MTZG23]. hatch [ACG^{+}16, FYK^{+}21, KNO^{+}04]. hatch-date [ACG^{+}16]. hatchery [MAH12, Sai22, ZZ93]. hatching [KVR^{+}18, NHS^{+}07]. Hatteras [GS99]. hauls [LVF12]. Hawaii [SMB03b]. Hawaiian [HKA^{+}06, MBB^{+}03]. Heather [Hea99a]. heatwave [RWDA^{+}21]. heavily [OHS06]. hebraicum [BEF^{+}12]. height [WGW07]. heights [LRBJ21]. helgolandicus [IH03]. Helicolenus [MBJ^{+}07]. help [Bow11]. hemisphere [WTR04]. Henry [BD93]. herbivorous [UYF92]. Herman [Gre99]. Herring [CMMK^{+}15, AH97, BML^{+}14, BMPC16, BSG^{+}13, BG01, BWKM15, BDVS^{+}19, BDTR23, CAB^{+}01, CP92, FPBDC11, FUA^{+}98, FM93, FBRB12, GPA^{+}21, JGS93, LYT^{+}20, MLVO05, Mar01, MWGK92, NDC05, Neu02, NBF^{+}01, REG^{+}13, SMA14, SMH^{+}92, SNV^{+}12, SPLS15, Tan17a, VCB^{+}98, WQI00, WQ00, óT10, óóSV18]. HF [HP02]. High [DP01, SRR07, TDE09, LSD^{+}21, LRBJ21, MHM^{+}20, PHH13, SZX^{+}08]. High-frequency [SRR07]. high-resolution [LRBJ21, MHM^{+}20]. higher [NFO^{+}23]. Highlights [Kas99, Liv00, DAW^{+}23]. highly [BBB^{+}16, CGMM10, CCHL23]. Hilsa [GHG^{+}19]. Hindcast [ZWC^{+}21]. hippoglossoides [ÅGN^{+}04, SCDA10, YLA13]. Hippoglossus [HAS^{+}19, SME^{+}14, SGW^{+}21]. historic [ZP21a]. Historical [BPP07, LA05, SFGE21, FH94, QM01]. histories [AHAM03, BHV^{+}06, BASS11, ISN^{+}11, TSK04]. history [BC04, BCA^{+}18, MLVO05, MW92, NDC05, NBF^{+}01, PSS^{+}21, QBMW99, QC99, REG^{+}13, RG97, SS19, Sch23, Tak04, THH12]. Hiuchi [YTIS95]. Hiuchi-Nada [YTIS95]. Hokkaido [KTH^{+}15, FYK^{+}13, HONH04, KSYT97, KY17, MTH^{+}04, NHS^{+}07, SKHN11, TKM^{+}22]. Homarus [DHMT96, BMOT17, DTC06, HDH^{+}05, IN00, IXW^{+}10, PWML12, SCTB19]. homeward [DLTI95, Sim96]. homeward-migrating [DLTI95, Sim96]. homing [AI05, DHM^{+}15]. Honshu [NSH^{+}17]. Horizontal [KBF^{+}07, SMK02, SF22, SMB03b, TKH08, TSK^{+}95, FDT^{+}99, KSY^{+}23, SWAAB20, SSSB03]. horse [ISS02, KVR^{+}18, KYS15]. hot [MESMM18]. hubbsi [MMSL19, TMMM20]. human [PO03]. humans [CCL^{+}05]. Humboldt [Esc98, AS08, GSBB07, LLB^{+}20]. Hurricane [MTZG23]. huxleyi [HGH93]. hydroclimatic [Bea03]. Hydrodynamic [NSGL^{+}22, PST03, APL01, BHV^{+}06, BEF^{+}12, HB99, QBMW99, RQN^{+}99, RHRL12, TCS^{+}09, TTC^{+}12]. Hydrographic [LJH^{+}05, LGM^{+}02, AMK08, CRVL^{+}17, LVC^{+}05, MFB^{+}09, RS92, SPM02, SHB^{+}11]. hydrographical [MHvD^{+}24]. Hydrography [TSK^{+}92, GV01, HFC01, HEG08, TSK^{+}95, UTMS06]. hydrological [LOGLD^{+}15]. Hypothesis [KEWDA18, BA12, CEM^{+}11, IMS^{+}04, IUY10, McK13, MCG^{+}14, NGGJ09, PJD14, TMM^{+}07, MRL^{+}14]. Hypoxia [PG06, PLG^{+}10, YLA13]. Hypoxia-based [PG06, PLG^{+}10]. hypoxic [CGMM10, KSC^{+}10].

Iberian [áRÁSG^{+}16, áCGNGC19, GVRC04, PVBV19, RCG^{+}15, SOTM^{+}18]. IBM [MHM^{+}20, PVMP03]. ice [WEW98]. Icelandic [BTGM07, JGS93, OR12, OR13, SP93, SSM^{+}10, óT10, óóSV18]. ichi [MFS^{+}17]. ichthyofauna [DG00, LAB^{+}05]. Ichthyoplankton [CCK^{+}22, JCCB15, NK08, ADAHL10, Aut08, BDAMD14, CMM06, DDZ09, DABM^{+}06, FGGDSMF08, HFC01, HP02, LPCG23, SB94]. Ichthyoplankton-based [NK08]. ICOS [Ano99]. idealized [BLD^{+}03]. ideas [Sha95]. Identification [RSZ^{+}03, Hor00, MAS^{+}98]. identify [ROH16, SRR05]. Identifying [ISI^{+}18, Erz05, LCC15]. II [IXW^{+}10]. ilisha [GHG^{+}19]. illecebrosus [DHC^{+}07, SHS^{+}23]. Illex [ABI^{+}21, CAB12, DHC^{+}07, SHS^{+}23, WRTP01]. imagery [BDBP93, Col99, LVC^{+}05]. images [KYY00]. Immature [FRS^{+}05, AI04, KSMY00, KKNY04]. immigrating [RQN^{+}99]. immigration [SP93]. Impact [Cap08, FKF^{+}22, GMH^{+}12, KUO^{+}17, LJM^{+}10, LPSS04, NYI11, NII^{+}14, QCR22, CCC^{+}23, GFO14, LPHM21, SL95, MM94a, SMS^{+}21, TMN^{+}15, YWM^{+}00]. Impacts [BBA^{+}21, FCC^{+}19, PRDC^{+}13, RWDA^{+}21, TWK13, VSÅO07, CLW^{+}19, CEM^{+}11, GHM21, GAH^{+}19, JPHA^{+}16, Kae23, NPLS22]. implication [YW07]. Implications [BMH^{+}21, DPL^{+}20, HT18, KEWDA18, Tan02, dBdOJdO^{+}22, ACG^{+}16, BH97, Cap08, CEM^{+}11, ETB^{+}17, Fun11, HFC01, HXC^{+}17, Mul97, PSS^{+}21, PHH13, PVHT01, PKHF98, QLB^{+}05, RCG^{+}15, VPRG13, WBQL99, WQ00]. Importance [FKUY16, RCPS09, BO05, CMB^{+}15, DAW^{+}23, DBGW04, ESA^{+}16, Erz05, FIDC00, FMYN06, FBRB12, LCCdS^{+}19, LJR^{+}22, LMB^{+}19, Lou10]. important [BHJ^{+}04, FYC22, KTO^{+}11, LPCA15, LJBR20, MHS^{+}21, SLM13]. imprinting [BA12]. improve [FCJ^{+}15]. Improvement [KKK^{+}17, KWO^{+}18]. improves [SL09]. Improving [HBN^{+}21, MPM19, NBMS06, Sai22]. in-pot [BLG^{+}16]. incidental [MMRH^{+}16]. incidentally [NSH^{+}17]. incidents [óóSV18]. including [NBMS06]. Inclusion [SL09]. Incorporating [MTL^{+}22, SC05, SSW^{+}17, SSP^{+}11]. increase [BMO^{+}99, NFKY21, PW12]. increased [DHMT96, SES^{+}20]. increases [CSS^{+}21, LéEPW^{+}12]. increasing [DAW^{+}23]. increment [KTH^{+}15, KNO^{+}04, SPG^{+}16]. independent [Spe08]. Index [Ano01a, Ano01b, Ano03b, Ano03c, Ano04a, Ano04b, Ano05a, Ano05b, WTR04, BLH98, Bez00, CRC11, RWP11, XMH^{+}18]. Indexes [Ano03d]. India [KB08]. Indian [BCR20, BGH09, CLT05, GCF^{+}21, HRB^{+}18, HBN^{+}21, MMBC07, MTSH15, Nis92, Rog94, SZX^{+}08, WSP^{+}07, ZSY^{+}21]. indicate [MLRS07, SHG^{+}22]. indicated [WMD^{+}00]. indicates [ESA^{+}16]. indicator [HMS^{+}23, KT93]. indicators [BMHW13, BGM^{+}18, MCHSNEO13, RS92, WP93, YCH^{+}15]. Indices [FMG^{+}22, BMPC16, LCCdS^{+}19, MSS12, OTIK20, YCH^{+}15, ZHT14]. indirect [AMK08]. individual [BC04, BRC04, DPK^{+}08, FMYN06, HBC07, MLVO05, MLC^{+}98, NHNA07, PJB05, QBMW99, RHRL12, VN97, VFS^{+}24]. individual-based [BC04, BRC04, DPK^{+}08, HBC07, MLVO05, MLC^{+}98, QBMW99, RHRL12, VFS^{+}24]. individuals [MFP^{+}03]. induce [BRO18]. induced [ASCM12, BSG^{+}13, GCQ^{+}13, MTL^{+}22, NTM^{+}15, Pol96, SW05, SLM13, VOB^{+}19, XH95]. induction [TY04]. inferences [HKLG07, QC99, RQN^{+}99]. inferred [BCBDA10, GP94, KO95, PDER10, SSPY08]. inflow [REB^{+}03]. Influence [AGSSL^{+}22, ADAHL10, BWS^{+}01, CKA^{+}17, Dom23, GQPGA04, IHS97, KM94, MMSL19, MSC^{+}17, RKD^{+}20, SAG^{+}09, WSF^{+}14, XMW^{+}23, AUOGMM19, APL01, APL07, AGS^{+}04, AI04, BSF^{+}20, BSF01a, BvDSDC18, CCM^{+}08, DPK^{+}08, DHM^{+}15, DTC06, FPBDC11, FUA^{+}98, KSAF13, LDAWM10, Mar01, MJH14, NDC05, Neu02, PDD03, PS16, PMG^{+}23, QCM^{+}16, RCS98, Rob94, SHS^{+}23, SHB^{+}11, TIH^{+}92, TAS04, VYGT^{+}20, WTR04, óT10]. influenced [CCHL23, HTP14, OUKH04]. Influences [FRHMAM^{+}06, HTT^{+}16, OR12, WPL^{+}93, FML^{+}14, HMP92, HDJ15, KB08, LLB^{+}20, MRD^{+}19, OH23, PBF00, SRCV09, SAT^{+}18, TLS98, VHCN14, dBdOJdO^{+}22]. influencing [BDTR23, GCQ^{+}13, LGM^{+}02, LVC^{+}05, VDHF08, WKB^{+}05, WCP^{+}01]. Information [Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, GRT^{+}07, Sim92a, ZWL21]. Informing [BPZR19]. ingestion [FUA^{+}98]. Ingress [SOTM^{+}18, BAL^{+}99]. inhabiting [SPM^{+}19]. Initial [IKK^{+}04, SCTB19]. initiation [KHN^{+}22, TH11]. Inland [FYK^{+}21, KKNY92, YOYK20, ZKT07, OUKH04]. Inlet [BAL^{+}99, LHF^{+}99, FRP^{+}99, BHJ^{+}04]. inlets [RMM02]. inner [HSS19, MMB^{+}11]. innermost [SFK^{+}20]. input [BBB^{+}16, LPSS04, QM01]. Inshore [KSY^{+}23, BSF01a, CSB94, DBRSC16, YOY00]. Insights [DLD^{+}23, GNP^{+}19, SWAAB20, EvST^{+}17, MMI^{+}22, áRÁSG^{+}16]. Institute [KYY00]. instrumental [Sch23]. insularis [LAPL21]. Integrated [PFB^{+}16, Sch23]. integrative [NH06]. intensity [AAI16, TFB^{+}17]. Inter [OE17, ETB^{+}17, LP10, LAPL21, MIY^{+}09, TAN^{+}17b, VYGT^{+}20]. Inter-annual [OE17, ETB^{+}17, LP10, LAPL21, TAN^{+}17b, VYGT^{+}20]. inter-frontal [MIY^{+}09]. interaction [RD96, ZYT^{+}22]. Interactions [Har92, NdLOO23, ZLTM11, GPCGdlT^{+}22, LLCJ16, LAG^{+}11, NTIO18, PDD03, PO03, REM02, Wat17, XTC^{+}04]. Interannual [AYMK01, ACG^{+}16, BDSM07, CP92, DDB17, FGGDSMF08, FHK^{+}10, GDM^{+}17, HFF^{+}19, IH03, KPHG14, MAHG94, MWR^{+}98, NKM01, NNOU20, NHS^{+}07, OUKH04, PJD14, RSF13, SRCV09, SKT21, SC06, ST97, TCO^{+}05, Tan99, WL21, YWM^{+}00, BMPC16, GCQ^{+}13, HQH^{+}06, HSS19, INM^{+}18, IST^{+}23, KB08, LLCV18, MLP22, SGN^{+}05, Tan02]. Intercalibrating [MM03]. intercomparison [GTB10]. Interdecadal [KY17, ST98, YSW^{+}99, BDSM07, FHHW98, ST97]. Internet [KYY00]. interpolation [RMM02]. interpret [QBMW99]. interpretation [CAB12, LCCdS^{+}19]. interpreting [MM03]. interspecific [KM93, LDAWM10, NTIO18]. Intra [MBY^{+}18, KM93, SCTB19]. intra- [KM93]. Intra-annual [MBY^{+}18, SCTB19]. Introduction [Ano01c, Hea99c, SHM05, OCH99]. intrusion [OUKH04, STI^{+}09]. Invasion [Shi98, HBR^{+}99, MBJ^{+}07]. invertebrate [KSC^{+}10]. invertebrates [BBMY93, JYH^{+}18, SNL19]. investigate [BGM^{+}18, RRF^{+}21, TTC^{+}12]. investigated [APGL03, APL07]. Investigating [FMV03, MHM^{+}20, BCGB14]. Investigation [Ano99, BA12, DSPH07, TR11, Tan17a]. IPRC [BB02]. Ireland [MLP22, SR02]. IRI [BB02]. Irish [BCGB14, BSF^{+}20, ETB05, FODCN00, FMYN06, LDDC06, PA14]. iron [KTS15]. irradians [LCCS15]. isada [MAS^{+}98]. ISBN [Gre99]. Ise [TY04]. Iskenderun [MBY^{+}17]. Island [BAB^{+}06, LHF^{+}99, MSL^{+}05, MRHL09, PHWM96, PBF00, SRCV09, JR07, Coy05, HL98, LH96, TMN^{+}15, Tan99, Tan02, TR11]. Islands [HMTG^{+}05, SFA14, WSC05, Zam01, MBB^{+}03, APR^{+}08, BRO18, BAB^{+}06, BRR05, CCL^{+}05, FRS^{+}05, HWS^{+}05, JCH05, LHM^{+}05, LAB^{+}05, MRHL09, MBB^{+}03, SMF^{+}05, SCDA10, TSK^{+}92, ZP21b]. isolated [DP01, SPM^{+}19]. Isostichopus [HMTG^{+}05]. isotope [DDS^{+}17, IMO^{+}12, MCHSNEO13, OM10, OKT^{+}23]. Issue [Ano03a, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, CHPA98]. issues [PO03]. istiophorid [PLG^{+}10]. Istiophorus [HLG^{+}11, MHB^{+}14, RCPS09]. Isurus [MCHSNEO13, RHP^{+}15]. Iwate [OK17].

jack [DSHL18, IST^{+}23, IWK^{+}21, NPY^{+}15, SKM06, SYT^{+}09, SKT21, TSK^{+}22]. Japan [MTT^{+}17, NSH^{+}17, OKU17, ONK17, War92, CHHS05, FYK^{+}21, Fun07, Fun11, FYK^{+}13, HYW04, HFF^{+}19, HH99, HONH04, HMS16, ISI^{+}18, IST^{+}23, IFF^{+}18, KKK^{+}17, KNK^{+}18, KTH^{+}15, KSYT97, KKNY92, KMM^{+}06, KU95, KM93, KWO^{+}18, KYY00, KNO^{+}04, KUO^{+}17, KY17, MWN^{+}23, MAS^{+}98, MTH^{+}04, NSH^{+}17, NNOU20, OTH09, OFS^{+}16, OHM^{+}10, OUKH04, SKT21, SK03, SK04, SKNT14, SFK^{+}20, TSK^{+}22, TWK13, TKW^{+}17, TNM^{+}02, TMN^{+}15, Tak04, TKH08, TKMS11, TTI^{+}20, TY04, TTH15, WTK^{+}16, YAM^{+}18, YOYK20, Yam04, YTIS95, YIT^{+}22, YKB08, ZKT07]. Japanese [FYA^{+}21, FYK^{+}21, FKH^{+}17, HZTS12, HZW^{+}98, HXC^{+}17, IST^{+}23, IK97, IWK^{+}21, IYN^{+}09, ISN^{+}11, KSY^{+}23, KKS92, KYU^{+}06, KKCL06, KIS01, KWO^{+}18, KUO^{+}17, MTSH15, NHM94, NZI95, NFKY21, NY08, NYI11, NYI^{+}13, Nis19, NY03, OTH09, OKT^{+}23, OIA^{+}12, SKT21, SSW^{+}17, SHK^{+}19, SK03, SFK^{+}20, TWKW01, TW05, TSK^{+}22, TA06, TMN^{+}15, TF08, TY04, TTC^{+}12, WZK97, WZK^{+}98, YSW^{+}99, YKH^{+}21, YWI^{+}05, ZKT07, ZYY^{+}21, ZYT^{+}22]. japonica [HZTS12, HXC^{+}17, KSY^{+}23]. japonicus [AGSSL^{+}22, FKUY16, FYK^{+}21, GiIW^{+}20, HJR^{+}03, IST^{+}23, IK97, IWK^{+}21, IUY10, IYN^{+}09, ISN^{+}11, KOS^{+}19, KL01, NNOU20, PVHT01, SKM06, SYT^{+}09, SKT21, SCF^{+}20, SFK^{+}20, TWKW01, TW05, TSK^{+}22, TA06, TMN^{+}15, TTC^{+}12, YWI^{+}05, ZKT07, ZYY^{+}21, ZYT^{+}22, ZHL^{+}03]. Jasus [FML^{+}14, HGG^{+}17, LJM^{+}10]. jellyfish [SFL16]. jet [NYI11]. Johnstone [JTYB18]. Joint [War92]. jordani [Han11, PBF00]. Jorge [TMMM20]. journal [BZ21]. Juan [Zam01]. jubatus [CL05, FRS^{+}05, SMF^{+}05, TMM^{+}07]. July [MVK^{+}20]. Jumbo [LCC15]. June [MVK^{+}20]. Just [GAH^{+}19]. Juvenile [Lou10, MRRN05, NPS^{+}23, ARL93, ACT^{+}10, ACG^{+}16, APGL03, AI04, BGH09, BPLC11, BS94, BPC^{+}16, BDSM07, CCC^{+}23, CSK11, CAB^{+}01, DST11, EBO04, FMYN06, FHK^{+}10, FHK^{+}12, FFF^{+}18, Gla11, HHH^{+}16, HL07, HTT^{+}16, HKM^{+}19, HKM^{+}21, HONH04, HHK^{+}10, IST^{+}23, ICB^{+}08, IWK^{+}21, JPMH20, JTYB18, KBF^{+}07, KUO^{+}17, LDAWM10, MSS12, MLRS07, MWN^{+}23, MSC^{+}17, MRD^{+}19, NHS^{+}07, NBF^{+}01, PHWM96, PMT^{+}94, RSF13, RHP^{+}15, RWP11, SKHN11, SMB^{+}03a, SMH^{+}92, SSR13, TWKW01, TSK^{+}22, TKW^{+}17, UMK20, UTMS06, VFS^{+}24, WS08, WCP^{+}01, Wil01, WL21, YKH^{+}21]. juveniles [BDTR23, GPL^{+}11, LCCQ^{+}22, MOE06, NII^{+}14, SKM06, SKT21, SKNT14, TKO^{+}14, VSÅO07].

Kajikia [APMRH17, APMVOGMR19]. Kalman [SMB03b]. Kamchatka [FYA^{+}21]. Kareius [YTY96]. Karnataka [KB08]. Kasatoshi [McK13, PW12, PW14]. Katsuwonus [And03, GCF^{+}21, LPS19, LMBL03, MSST16, MSNK10, NPLS22]. Kattegat [FCJ^{+}15, JCA^{+}16]. kelp [MTT^{+}17, YKI98]. kelts [RFD^{+}04, RDF^{+}11]. keta [AI04, AI05, FYA^{+}21, SKHN11, TID^{+}96, WTK^{+}16, YCH^{+}15]. Key [WKB^{+}05, HVHC10, SBY^{+}15]. Kii [OUKH04]. kill [MPW^{+}99, óóSV18]. King [EvST^{+}17, DPL^{+}20, LA05, WMKR09, RRF^{+}21]. kisutch [BRPC08, BDSM07, KHB02, LML^{+}03, PMFC10, RWLP12, RWP11, SMB^{+}03a, WGFR06]. Korea [KL01, KKCL06]. Korean [KK00]. krill [MAS^{+}98, MKH^{+}13, MWR^{+}98, SRCV09, TBB^{+}03]. Kurile [TSK^{+}92]. Kuroshio [AI92, AGK^{+}08, FFF^{+}18, HZW^{+}98, IST^{+}23, IWK^{+}21, ISN^{+}11, KFS22, KKH^{+}20, KKS92, KKNY04, KMK^{+}18, MTL^{+}22, MIY^{+}09, NHM94, NZI95, NFN00, NKM01, NH03, NIIS04, NY08, NYI11, NY03, OWK^{+}03, SMK02, SKM04, SKM06, SHK^{+}19, TWKW01, TW05, TMS^{+}08, TKO^{+}14, TNK^{+}16, TYO21, UTMS06, WZK^{+}98, WK03, YW07, ZNI96]. Kuwait [YMB99]. Kyushu [TMN^{+}15].

L. [ACT^{+}10, BK94a, BK94b, BUE02, DPK^{+}08, DDS^{+}17, FM93, FODCN00, FMYN06, FHD98, GGF17, GI13, HBO^{+}01, HVHC10, HRS^{+}21, KVR^{+}18, NDC05, NHNA07, PGL^{+}15, RFD^{+}04, RDF^{+}11, SGN^{+}05, VHJ99, WJT97]. laboratory [OA06]. Labrador [FYKSP07, CSB94, GHV95, HMP92, KFYP07, LPH^{+}19, LPHM21, TDT03, WKN^{+}95]. lacustrine [AHAM03]. Lagrangian [APGL03, CW98, GGQF22, TF08, WB93]. lakes [TR11]. Laminaria [YKI98]. Lamna [CJ04]. lance [KKNY92, MW92, MWGK92, NNOU20, SJB^{+}22]. landfall [TIH^{+}92]. landing [CSB94, MAHG94, SFGE21]. landings [BGM^{+}18, CMMK^{+}15, Erz05, HBN^{+}21, LLSF01, LPSS04, NPY^{+}15, NLN^{+}21, QM01, SMS^{+}21, SRR05, VYGT^{+}20]. landscapes [LOGLD^{+}15]. Large [AAI16, AJ15, KYA^{+}15, NPS^{+}23, PWML12, FH94, HL07, HALO00, KCW^{+}15, KNS97, LTL^{+}22, LH96, LPG^{+}06, McK13, PW14, PECG08, QCR22, STI^{+}09, YMK^{+}15, ZHT14]. Large-scale [PWML12, HL07, QCR22, ZHT14]. large-sized [LTL^{+}22]. largehead [SCF^{+}20]. largely [Jes22]. largest [MDR^{+}16]. Larimichthys [HGS^{+}21, XWL^{+}23]. larvae [ÅGN^{+}04, APL07, AGSSL^{+}22, ABS^{+}11, ARM16, BBMY93, BBS99, BK94a, BK94b, BC97, BRFRJRLC18, BAB^{+}06, BSS94, BS94, BWK^{+}99, BBT^{+}09, BSF01b, BTGM07, BHJ^{+}04, CH92, CAR^{+}10, DST11, Dd95, DCLC15, DMF^{+}17, DBS^{+}19, DGB^{+}16, EHW08, ETB^{+}17, EvST^{+}17, FDT^{+}99, FRP^{+}99, FM93, FRHMAM^{+}06, GQPGA04, HLH^{+}17, IN00, IYN^{+}09, III^{+}06, KNS^{+}22, KTH^{+}15, KKS92, KPW19, KR14, LCCQ^{+}22, LDH14, LDDC06, LS01, MDKS93, MOE06, MWGK92, MCS^{+}06, MFRR96, MLR10, NHM94, NZI95, NYI^{+}13, OWK^{+}03, OTO^{+}09, PP01, Por22, RQN^{+}99, RCG^{+}15, REM02, SSP^{+}07, SKM06, SMA14, SSSB03, SNV^{+}12, SBBB03, SKNT14, SFK^{+}20, TKO^{+}14, TNK^{+}16, TNM^{+}02, TKMS11, TTI^{+}20, TCS^{+}09, TDE09, VSÅO07, WHT92, WKB^{+}05, ZNI96]. Larval [CPM^{+}15, HZTS12, HDH^{+}05, HQW^{+}99, HLWL12, KN08, LHF^{+}99, MRHL09, MRBBHL14, MBKP08, MSVY^{+}13, PEKL14, SJB^{+}22, YIT^{+}22, APGL03, APLG07, AM18, BCBDA10, BJCS12, BCJ^{+}13, BSG^{+}13, BEF^{+}12, BAL^{+}99, BHJ^{+}04, BCL04, CAGPC21, CC03, CM10, CFL^{+}99, CRVL^{+}17, DPK^{+}08, DPL^{+}20, DDB^{+}20, Dom04, DP01, DPL02, EHW08, FPBDC11, FUA^{+}98, FCL93, FBRB12, FRZVHM^{+}11, GHBM99, GCQ^{+}13, GP94, GS99, GDM^{+}17, HT18, HFC01, HZW^{+}98, HL07, HHF09, HMS^{+}23, HNHP09, HLMS03, HVHC10, HCC^{+}09, HXC^{+}17, HCS^{+}09, IIS^{+}07, ISN^{+}11, JMP^{+}14, JCA^{+}16, KSM^{+}20, KIS01, LLCJ16, LBW^{+}05, MBY^{+}18, MTZG23, MLP22, MHRC18, MAHG94, MATL98, MDR^{+}16, MSC^{+}17, MMI^{+}22, MMB^{+}11, MGHS14, MHvD^{+}24, NKS00, NGGJ09, Nis19, OHF12, OEV^{+}10, OWK04, OA06, POA^{+}17, PST03, PDD03, PDER10, PJD14, PA14, PWE98, QLB^{+}05, QCR22, RPT^{+}00, RAT^{+}02, REL07]. larval [RHRL12, RKD^{+}20, RD96, SRR99, SRR07, SMK02, SKHI04, SKM04, SKM06, SES^{+}20, SHG^{+}22, SS94, Sko05, SPLS15, SRM^{+}18, TWKW01, TW05, TCL^{+}12, TA06, TMN^{+}15, TFB^{+}17, TCC^{+}98, VIS92, VHJ99, VDHF08, WBQL99, YTY96, ZKT07, éSMB20]. laser [GTB10]. last [KK00, NNOU20]. Late [SKM04, HMM01, LS01, MCS^{+}06, MRHL09, PSJF93, TW05, TH11, WSC05]. late-run [TH11]. late-stage [MCS^{+}06]. late-summer [WSC05]. Lateolabrax [FKUY16, IUY10, SFK^{+}20]. latitude [PSM00, Sim92b, TIH^{+}92]. Latitudinal [BWJ03, SCF^{+}20]. Lawrence [éSMB20, BDVS^{+}19, CM10, D’A93, PGL^{+}15, RD96, RCD^{+}99, Swa99, SB06, VHCN14, YLA13]. layer [CCSS01, NIIS04, NY08, SBD^{+}19, YW07]. layered [AW92, GP94]. layers [AI92, HJ10]. learning [SLZ^{+}23]. Leatherback [SAH^{+}18, EBFF17, HHB^{+}15]. Leeuwin [Cap08, FHK^{+}12]. legislation [SFGE21]. leidyi [Shi98]. Leite [LAPL21]. Length [SPM02, OFS^{+}16, PP01, TGRS^{+}19]. lengths [WGFR06]. Lepidochelys [MMRH^{+}16, PBH^{+}04]. Lepidopsetta [CRW20, LDH14]. leptocephali [KMM^{+}06, TMS^{+}08]. less [Jes22]. Lessepsian [MBY^{+}17]. Letter [CW94]. level [CHHS05, D’A93, MCHSNEO13, WGW07]. levels [JCCB15, KCW^{+}15, NFO^{+}23]. LiDAR [JYH^{+}18]. Life [áRÁSG^{+}16, TD02, TSK04, AHKP16, AHAM03, BC04, BSF01b, BCA^{+}18, CAR^{+}10, DST11, GIT^{+}13, HG98, HBO^{+}01, IUY10, KR10, LPCA15, LGM^{+}02, LVC^{+}05, LLB^{+}20, LCCdS^{+}19, MLVO05, MW92, NDC05, NBF^{+}01, NH06, PSS^{+}21, PRDC^{+}13, QBMW99, QC99, RS15, REG^{+}13, ROH16, RWDA^{+}21, RG97, SGW^{+}21, SS19, SB94, SCDA10, SK03, Tak04, TTY^{+}23, TAS04, THH12, WPL^{+}93, XWL^{+}23, ZZ93]. life-history [SS19]. life-stage [SGW^{+}21]. light [BKvdP^{+}22, FUA^{+}98, HCS^{+}09, LS21, NBMS06]. light-based [NBMS06]. likely [HTP14]. Limanda [BMHW13, LDDC06, Por22, SCS05]. Limited [OKU17, BCL04, HLMS03, LJBR20, NNOU20]. limits [DB03]. lingcod [ARL93]. link [DPL02, GPS22, HTP14, LS21, OHS06, OH23]. Linkages [WMKR09, KKH^{+}20, NH01]. linked [BBS99, HFHW19, MMRS16, MFMG20, MFB^{+}09, QBMW99, REB^{+}03, SSR13]. Linking [BHV^{+}06, BCGB14, ESA09, SEM^{+}14, TSK^{+}22, HLWL12, KN08]. Links [GI13, OBA01, BMO^{+}99, Han11, HA07, NK08]. lion [CL05, FRS^{+}05, SMF^{+}05]. lions [TMM^{+}07, RBB^{+}21]. Lipid [Jón99, YKH^{+}21]. lipid-rich [YKH^{+}21]. Lipids [VJ99]. Lis [SOTM^{+}18]. List [Ano07, Ano10]. literature [DLD^{+}23]. Litopenaeus [WKB^{+}05]. Living [RHP^{+}15]. Lloyd [Bez00]. Lobster [CM10, BMOT17, BLG^{+}16, CB93, Cap08, CCC^{+}23, DHMT96, DTC06, EF10, FCJ^{+}15, FML^{+}14, GBAD^{+}17, HDH^{+}05, HGG^{+}17, IN00, IXW^{+}10, LJM^{+}10, MFMG20, MLP22, PWML12, PTS^{+}24, QCR22, SCTB19]. local [BJCS12, HBLC22, KMM^{+}06, MBE^{+}15]. local-scale [BJCS12]. Location [HHF09, BPP07, KYSM11, NH06, PLSO98]. Locations [YK96, HDJ15, III^{+}06, YW94]. Lofoten [ETB^{+}17]. logger [MIK07]. loggerhead [PKP^{+}00, PBH^{+}04]. logistic [RP93]. logit [BM99b]. loliginid [CG18]. Loligo [AGS^{+}04, DHC^{+}07, DBRSC16, MRL^{+}14]. Long [AH97, Bea03, BW92, BB07, Buc92, DLCQ22, IFF^{+}18, LYT^{+}20, MLP22, OTH09, OH23, SGN^{+}05, éSMB20, AS08, DHC^{+}07, RF04, RPE98, RHRL12, RSC96, RS92, SR02, VYGT^{+}20, YW07]. long-finned [DHC^{+}07]. Long-term [AH97, Bea03, BW92, BB07, Buc92, DLCQ22, IFF^{+}18, LYT^{+}20, MLP22, OTH09, OH23, SGN^{+}05, éSMB20, AS08, RF04, RPE98, RHRL12, RS92, SR02, VYGT^{+}20, YW07]. longevity [MHS^{+}21]. longiceps [HBN^{+}21, XB09]. longitudinal [WJM15]. longline [BBH99, BHM02, BML11, DSPH07, Dom09, Dom23, GHM21, HHTF10, HBR^{+}15, MTSH15, OFS^{+}16, PKP^{+}00, SSPY08, ZSY^{+}21, ZHX^{+}20]. longliners [AUOGMM19]. longlining [SZX^{+}08]. look [Tyl92]. loophole [BB03]. Lopholatilus [NLN^{+}21]. loricae [ST95]. Loss [MMF95, BSF01a]. low [GYS14, KIS01, Nis19]. low-salinity [KIS01]. low-stock [Nis19]. lower [CHHS05, IKK^{+}04]. lucens [TKMS11]. lucetia [LLB^{+}20]. lunar [CSS^{+}21, GHG^{+}19, OE17, SAT^{+}18]. Lutjanus [BASS11].

M [Ano01d, CLPC18]. maccoyii [BGH09, FHK^{+}10, FHK^{+}12, HHTF10, HHK^{+}10, PECG08, WMD^{+}06]. machine [SLZ^{+}23]. Mackerel [GiIW^{+}20, PGL^{+}15, BC04, BRC04, BUE02, BvDSDC18, DSHL18, HDJ15, IST^{+}23, IWK^{+}21, ISS02, Jan16, KOS^{+}19, KM93, KVR^{+}18, KYS15, MDVB^{+}20, MHRC18, MFH05, NPY^{+}15, NK08, PVHT01, RBPCR^{+}22, RCD^{+}99, SKM06, SYT^{+}09, SKT21, TSK^{+}22, TYO21, VGPL^{+}11, WMKR09, YWI^{+}05, ZYT^{+}22]. mackerels [SHK^{+}19]. maclovinus [QM01]. macroalgal [TKW^{+}17]. macrocephalus [HCS^{+}09, NSH^{+}17, SC05, TNM^{+}02]. macrorhynchus [KOKM15]. Macroscale [MSL^{+}20]. macrotidal [SKNT14]. Madden [Hea99a]. magellanicus [TCS^{+}09, ZJH^{+}22]. magister [MAHG94, Sha13]. magnetic [CLH^{+}22]. magnitude [KSYT97]. main [ABI^{+}21, AGK^{+}08, FKSA21, MBB^{+}03]. Maine [MLM^{+}98, BPS^{+}14, CRC11, DCLC15, GRT^{+}07, GCW17, HDH^{+}05, HBPC15, IN00, IXW^{+}10, ROH16, SGL04, SCTB19, SMS^{+}19, SJB^{+}22, Wil04]. mainly [WJ93]. maintaining [CLH^{+}22]. major [YOYK20]. majority [TNK^{+}16]. Makaira [CKA^{+}17, RCPS09, SSPY08, SSP^{+}11]. make [Spr92]. making [DWH11]. mako [MCHSNEO13, OFS^{+}16]. makos [RHP^{+}15]. Malabar [KB08, XB09]. Maldives [AAG11]. Mallotus [APL^{+}08, HWSS07, LDAWM10, OR12, OR13, WPN12]. Mallotusvillosus [IHS97]. Malvinas [ABI^{+}21]. mammals [JR07]. man [RGQPN09]. manage [HHTF10]. Management [GNP^{+}19, BEF^{+}12, CL05, CLM^{+}21, CLKP19, CH99, CMS16, Fun11, HHK^{+}17, HRS^{+}21, HHB^{+}15, JPHA^{+}16, LPH^{+}19, MPM19, Par96, PVHT01, PKHF98, YWI^{+}05, dBdOJdO^{+}22]. Mangalore [KB08]. Manta [AAG11]. mantas [AAG11]. manuscript [BZ21]. mapping [NTIO18, NH06]. maps [BPZR19]. March [RJHC99]. margin [SOTM^{+}18]. mariculture [KU95]. Marine [AAI16, AGK^{+}08, AJ15, FHD98, GPCGdlT^{+}22, Har92, HQH^{+}06, KYA^{+}15, LHM^{+}05, NPS^{+}23, RWLP12, SBT20, Woo93, BJCS12, BCJ^{+}13, BRN^{+}95, BNM^{+}00, BEF^{+}12, BWKM15, BWS^{+}01, CCL^{+}05, CLKP19, CH92, CAR^{+}10, DAW^{+}23, Dom04, ERR^{+}21, FYC22, FH94, Gre13, HSEH16, HKA^{+}06, JHK^{+}15, JR07, KCW^{+}15, KMM^{+}06, KHB02, LJR^{+}22, LMB^{+}19, LBLCLC05, LS15, LML^{+}03, MCG^{+}14, MFS^{+}17, MAH12, MMMS14, MKF^{+}03, MWR^{+}98, NH03, NSH^{+}17, PFB^{+}16, PO03, PFSL09, PEKL14, RDE^{+}07, RWDA^{+}21, RAK^{+}17, SKHN11, Sim92a, SC97, SPV96, THH12, Tyl92, VCKH05, WKR^{+}18, WS08, YMK^{+}15]. Marine-climate [GPCGdlT^{+}22]. marine-protected [NSH^{+}17]. Marini [TMMM20]. market [PS16]. marlin [APMRH17, APMVOGMR19, CKA^{+}17, GSNFL99, HKLG07, RCPS09, SDHB07, SSPY08, SSP^{+}11]. maroccanus [MTP07]. Mass [BHC^{+}01, MBKP08, óóSV18]. Massachusetts [LCCS15, CCC^{+}23, NASTF10]. masses [Coy05, ESA09, GNP^{+}19, KT93, KN08, SL95, MATL98, QLB^{+}05]. masses-impact [SL95]. massive [OKU17]. Match [MM94b, MBE^{+}15]. Match/mismatch [MM94b]. Mathematical [YKI98]. matrix [QC99]. Matsushima [YIT^{+}22]. matter [TH11]. matters [BH18, MLP22]. maturation [FKSA21, WGW07]. mature [WGFR06]. maturity [KBS^{+}16, OR13]. Mauritania [FIDC00, TFB^{+}17]. Mauritanian [BJV^{+}17, MBE^{+}15]. Maurolicus [RG97, SSR13]. mawsoni [MMI^{+}22, PSS^{+}21]. Maxent [SLL19]. maximum [MPM^{+}13, NH06, RP93, SKNT14, WKR^{+}18]. maximus [CSFC05, HRS^{+}21, SR02, Wil04]. may [Jes22, Aut08, BBS99]. maya [AOVAG22]. mean [WPL^{+}93]. meander [NHM94, NFN00]. measurements [ESTJ03, GiIW^{+}20]. measures [RAT^{+}02]. Measuring [GTB10, Par95]. Mechanism [AI05, DLTI95, Gar97, SHG^{+}22, Sim96]. Mechanisms [ETB^{+}17, AB02, HKM^{+}21, IUY10, KO95, NH06, TJW^{+}03, YTY96]. Mechanistic [HA07, PCR^{+}18]. mediated [HFHW19, HNHP09, VZP98]. mediator [MKF^{+}03]. Mediterranean [CAGPC21, GGQF22, PQH16, AMD^{+}16, AB02, ABG19, BGM^{+}18, CLPC18, GCQ^{+}13, GGF17, GPL^{+}11, GIT^{+}13, KMD^{+}09, LAFF15, LLSF01, LPSS04, MTP07, MMRS16, MBY^{+}17, MBY^{+}18, MOE06, MSR20, OEV^{+}10, RS15, SSP^{+}07, SGS^{+}06, VHLM15]. Medwin [Gre99]. Meeting [Kas98, Woo97, PFB^{+}16, Kas97, Kas99, Liv00, Woo95]. mega [TWK13]. mega-earthquake [TWK13]. megafauna [EPG^{+}16]. Mejillones [REM02]. Melanogrammus [BCL04, HG98, LOS^{+}14, LSK^{+}18]. melanosticta [KKCL06]. melanostictus [HZW^{+}98, IYN^{+}09, ISN^{+}11, NY08, NYI11, NYI^{+}13, Nis19, NY03, OTH09, OIA^{+}12, SK03, TF08, WZK^{+}98, YWI^{+}05]. Memoriam [Per23, Hea99a]. Menhaden [MSL^{+}20, COW^{+}99, FDT^{+}99, HT18, QBMW99, QC99, RQN^{+}99, SQW^{+}99, WBQL99]. Meridional [HJ10]. Merlangius [LVPK11]. merlangus [LVPK11]. Merluccius [BKvdP^{+}22, CC03, GI13, IMO^{+}12, KvdPBW17, LCCQ^{+}22, MMSL19, MOE06, RPC^{+}19, SRR07, Tan99, TMMM20, VMT^{+}23, WJM15]. meso [IST^{+}04, RWP11, SHB^{+}11]. meso-scale [SHB^{+}11]. meso-zooplankton [IST^{+}04, RWP11]. Mesoamerican [MSVY^{+}13]. Mesopelagic [FRZVHM^{+}11, FRHMAM^{+}06, SKKW02, SKM04]. Mesoscale [CMM06, HSH^{+}22, KFS22, LS01, APL01, ADAHL10, DPL02, GQPGA04, HLWL12, HBR^{+}15, KN08, KBB^{+}20, MCS^{+}06, MSVY^{+}13, RSC96, WRTP01, ZHX^{+}20]. mesotrophic [UIU^{+}99]. mesozooplankton [KKH^{+}20, KMK^{+}18]. Metabolic [GiIW^{+}20]. meteorological [VYGT^{+}20]. meteorology [SS98]. method [LPG^{+}06, MTH^{+}04, SSP^{+}11, WB93, YW94]. methods [RMM02, SP93]. Mexican [AUOGMM19, FGGDSMF08]. Mexico [BASS11, AUOGMM19, AOVAG22, BBB^{+}19, DGB^{+}16, ERR^{+}21, GPCGdlT^{+}22, GSNFL99, GBAD^{+}17, HT99, KSP^{+}22, KR14, MESMM18, MTZG23, MSC^{+}17, MCB^{+}16, MLR10, NLS^{+}24, SCAG^{+}21, SFA14]. Michael [Ano01d]. microbially [VZP98]. microcomputer [WHT92]. Microdistribution [SKNT14]. microdon [AHAM03]. Micromesistius [BC97, HEG08, MMRS16, MP18]. micronekton [HKT^{+}03]. microplankton [RD96]. Micropogonias [ASCM12, HT18, HA07]. Microsatellite [BEF^{+}12]. microstructure [ACT^{+}10, BHV^{+}06]. Mid [PSM00, SPM^{+}19, XMH^{+}18, CTWS08, MSM^{+}13, Sim92b, SGL22]. Mid-Atlantic [SPM^{+}19, CTWS08, SGL22]. Mid-latitude [PSM00, Sim92b]. mid-shelf [MSM^{+}13]. middle [Bau98, FMG^{+}22, SCS05]. Mie [KYY00]. migrating [BK94b, BGH09, DLTI95, Sim96, YKB08]. Migration [BPS^{+}14, CSK11, KNS97, TNC^{+}22, AMD^{+}16, AYK03, AI04, AI05, BM99a, CCM^{+}08, CGI^{+}19, CHF^{+}04, DST11, ETB05, GMH^{+}99, GS96, GJR18, HTL^{+}00, Hea99b, HQH^{+}06, HXC^{+}17, HTP14, HALO00, KHN^{+}22, KSY^{+}23, KYU^{+}06, KNO^{+}04, MESMM18, OR12, OR13, OHM^{+}10, Pol96, PBH^{+}04, QCM^{+}16, RBPCR^{+}22, RCG^{+}15, SYT^{+}09, SWAAB20, SE19, SHB^{+}11, SK04, TIH^{+}92, TH11, TTC^{+}12, VJ99, WMK^{+}99]. Migrations [HEG08, Ano99, AGS^{+}04, CÅP^{+}13, FGS95, NHNA07, PMT^{+}94, SSW^{+}17, WJM15]. migratory [CCHL23, HT18, LJBR20, SKKW02, WK03, YAM^{+}18]. millennia [CCL^{+}05]. Milne [SCTB19]. minke [KEJK00, MTK^{+}07, MKH^{+}13]. minority [TNK^{+}16]. mismatch [MM94b]. missing [Bau98, HTP14]. mixed [CMB^{+}15, JPMH20, NY08, STI^{+}09, YW07]. mixed-layer [YW07]. mixing [MCS^{+}06, RCG^{+}15, SF22, TF08]. Miyagi [TWK13]. Mnemiopsis [Shi98]. MOCNESS [CC03]. mode [APL01]. Model [BPZR19, BJCS12, AYK03, AMK08, AI04, BK94b, BC97, BC04, BRC04, BM99b, BLH98, BHM02, BBA^{+}21, BSF01b, BTGM07, BPS^{+}14, CW98, CMB^{+}15, CCM^{+}08, CÅP^{+}13, EHW08, FGS95, Fun07, GMH^{+}99, GGF17, GYS14, HQW^{+}99, HBPC15, HNHP09, HBC07, HHB^{+}15, IKK^{+}04, ITH23, KFH00, KU95, LAB^{+}98, LCH03, LVPK11, LBW^{+}05, MLVO05, MDR^{+}16, MLC^{+}98, MLR10, NY03, OTIK20, PST03, PJD14, PA14, PLP^{+}11, PCR^{+}18, QC99, RQN^{+}99, RRF^{+}21, RG97, RP93, RGQPN09, RWP11, SGFR^{+}21, SSW^{+}17, SLL19, SMDM98, SSP^{+}11, SK03, TF08, TAS04, TCS^{+}09, TTC^{+}12, VN97, VFS^{+}24, VZP98, Yam04, ZWC^{+}21]. Model-based [BJCS12, OTIK20, RWP11]. Modeled [DCLC15]. Modeling [AUOGMM19, AHKP16, GHM21, GFO14, KKS92, LAFF15, MMI^{+}22, NGGJ09, TAS04, WKR^{+}18, ZJH^{+}22, CLM^{+}21, CIS20, HKWL17, HVHC10, MESMM18, OHF12, OIA^{+}12, SCAG^{+}21, SB04, Yam04]. modelled [ECM^{+}01, LRBJ21]. Modelling [ÅGN^{+}04, BSS94, BRC^{+}03, BSF01b, BHS^{+}15, CLW^{+}19, Dd95, DBRSC16, DSHL18, FUA^{+}98, GFG98, HZW^{+}98, IN00, JYH^{+}18, LCH03, MTP07, MRL^{+}14, MDVB^{+}20, NPLS22, POA^{+}17, PVMP03, PP01, PHH13, PBL07, PWE98, SMA14, SGHW05, TDT03, APLG07, BHV^{+}06, BEF^{+}12, CAB12, DST11, DLD^{+}23, ESTJ03, FODCN00, FMYN06, GPL^{+}11, GiIW^{+}20, HB99, HG98, HRS^{+}21, LMBL03, MEK^{+}09, MFP^{+}03, OCCF^{+}18, PML06, SSSB03, SP15, VHJ99]. Models [HHF09, APL^{+}96, AAKMG06, Bri94, CMB^{+}15, DPK^{+}08, KWB^{+}16, NBH99, ODMRM98, QBMW99, RHRL12, SMS^{+}23, SLZ^{+}23, UMK20, WM06, YOK^{+}17, YKI98, ZSY^{+}21]. Modern [Sch23]. modify [DBFW13]. modifying [Sai22]. modulation [VZP98]. module [HHK^{+}17]. Moller [MWGK92, MW92]. molt [SCTB19]. monitoring [HHK^{+}10, LPS19, PHH^{+}98, PHH13]. monopterygius [MFH05]. monsoon [HLWL12, MHG^{+}11, SFK^{+}20]. monsoon-driven [HLWL12]. monsoon-generated [MHG^{+}11]. monsoons [AAG11]. moorings [SKKS05]. mordax [Cur04, CCP07, RCB08, TCL^{+}12]. morhua [AHKP16, AMK08, BCGB14, BSF01a, BTGM07, BCL04, CRC11, D’A93, DB03, FODCN00, GRT^{+}07, GCW17, HBPC15, HL07, HBO^{+}01, KR10, LBW^{+}05, Lou10, MRD^{+}19, Neu02, NHNA07, OHS06, RKD^{+}20, SHG12, SB07, SB04, Swa99, TLS98, VSÅO07, VHJ99, WJT97, WKN^{+}95]. Morocco [MTP07]. Morone [NASTF10, NH06]. morphology [ARM16]. morphometric [CPM^{+}15]. mortality [AMK08, BHC^{+}01, BC04, BMH^{+}21, BLG^{+}16, CRVL^{+}17, DBS^{+}19, FPBDC11, FCL93, Gla11, IUY10, Lou10, MHS^{+}21, NGGJ09, NY08, OTO^{+}09, SPLY23, WCP^{+}01]. most [DWH11]. motivate [MCG^{+}14]. mouth [KKK^{+}17]. Movement [ABG19, HONH04, PECG08, ACT^{+}10, Bri94, FFF^{+}18, HRB^{+}18, HPL13, KFS22, MFH05, OIA^{+}12, PKP^{+}00, SPS^{+}20]. Movements [DPM^{+}11, SFA14, SAH^{+}18, APR^{+}08, BYM16, DHM^{+}15, GRT^{+}07, HKM^{+}19, HKLG07, HCS^{+}09, KBF^{+}07, MBB^{+}03, RHG^{+}13, SF22, SKNLD10, SWAAB20, SMB03b, SDHB07, WKN^{+}95]. Mozambique [NPLS22]. Mt. [PW12]. much [DBFW13, Spr92]. muelleri [RG97, SSR13]. Mullin [Ano01d]. Mullus [GGF17]. Multi [LSK^{+}18, SNL19, Wat17, HHTF10, LVPK11, PLP^{+}11]. Multi-decadal [LSK^{+}18, SNL19]. multi-model [LVPK11, PLP^{+}11]. multi-species [HHTF10]. Multi-timescale [Wat17]. Multidecadal [BASS11, BMHW13]. multinet [GTB10]. multinomial [BM99b]. Multiple [SGW^{+}21, SES^{+}20, GNP^{+}19, OTIK20, SWS^{+}19, WSP^{+}07]. multiple-tagging [WSP^{+}07]. Multispecies [UMK20, DMH16]. murphyi [NPY^{+}15]. must [GJR18]. Mutsu [TNM^{+}02]. myctophid [SMK02, SKHI04, WMK^{+}99, WK03]. Myctophidae [WMK^{+}99]. mykiss [AMDM12, WWSE00]. myriaster [LJBR20].

N [WP93]. Nada [YTIS95, NFN00]. Namibia [IMO^{+}12, KvdPBW17, SBY^{+}15]. NansClim [LS15]. NAO [SB07]. nasus [CJ04, SKNT14]. Natal [RWI^{+}16, LCC15]. native [Bow11]. natural [DL94, FBRB12]. natural-cultural [DL94]. nature [VN97]. Naupliar [WZK^{+}98]. nauplii [ZKT07]. NC [BAL^{+}99]. near [ISN^{+}11, KCW^{+}15, MBB^{+}03, SRCV09, SMB03b, UTMS06, WSC05, ZWL21]. near-bottom [KCW^{+}15]. near-real-time [ZWL21]. nearshore [GPS22, JCCB15, KVR^{+}18, NBH99]. nekton [PLSO98, PFAM96, SDRL96]. NEMURO [AYK03]. Neocalanus [BWJ03, LP10, TCO^{+}05, TSK04]. neon [ASM^{+}15, IMS^{+}04, ISI^{+}18, NII^{+}14, NTM^{+}15, YWM^{+}00]. Nephrops [CLM^{+}21, FCJ^{+}15, MLP22]. Neritic [MTZG23, BBB^{+}19]. nerka [APL^{+}96, BWS^{+}01, CHF^{+}04, HQH^{+}06, McK13, PW12, PW14, PMT^{+}94, RZM^{+}03, TR11, TH11, WSF^{+}14]. net [CCSS01]. nets [MM03, PSC05]. network [BJCS12, CLM^{+}21, PEKL14]. networks [NPY^{+}15]. neural [NPY^{+}15]. Newfoundland [CSB94, Dd95, FYKSP07, GHV95, HMP92, IHS97, KR10, KFYP07, LPH^{+}19, LPHM21, PHH13, RFD^{+}04, RDF^{+}11, WKN^{+}95]. Newfoundland/Labrador [GHV95, HMP92]. newly [YMB99]. niche [ABG19, BBA^{+}21, MTL^{+}22, ZWC^{+}21]. nigricans [CKA^{+}17, RCPS09, SSPY08, SSP^{+}11]. nigripes [MJH14]. Niña [MRRN05]. Niño [Dom23, PS16, TCC^{+}98, BB03, FRHMAM^{+}06, FRZVHM^{+}11, HT99, HK06, KK00, MRRN05, Mul97]. Niño-southern [FRZVHM^{+}11]. niphonius [ZYT^{+}22]. nitrogen [KU95]. NOAA [KYY00]. noise [RF07, RR18]. Non [HKWL17, Bow11, ICB^{+}08, KN08]. non-depth-discriminate [KN08]. non-native [Bow11]. Non-parametric [HKWL17]. non-upwelling [ICB^{+}08]. nonlinear [GYS14]. Nonlocal [ASK99]. Nordic [NFO^{+}23]. NORPAC [MM03]. North [BJV^{+}17, BBH99, COW^{+}99, CHHS05, HXC^{+}17, HSS19, PLT09, QLB^{+}05, Woo93, ÅGN^{+}04, Ano99, Bea03, BUE02, BB07, Col00, DLTI95, GHV95, HB99, HFC01, IIS^{+}07, ISS02, LLSF01, LPSS04, MPW^{+}99, MDVB^{+}20, MAS^{+}98, MWP02, PJO99, PWE98, Sim96, SR02, SGHW05, WFRS93, WQI00, WQ00, YW94, AHKP16, ASM^{+}15, AGK^{+}08, AAKMG06, AMDM12, AI05, BC97, BC04, BF07, BSS94, CSK11, CSS^{+}21, DPK^{+}08, DL94, DB93, DDS^{+}17, DB03, ESA09, FPBDC11, FH94, FC04, FHD98, GMH^{+}99, GHBM99, Gar97, Gla11, GP94, HB99, Hea93, HG98, HBR^{+}99, HGH93, HKLG07, HLWL12, IMS^{+}04, INM^{+}18, IYN^{+}09, ISS02, III^{+}06, Jan16, JCA^{+}16, Kae23, KTPM17, KOKM15, KOWM16, KT93, KYU^{+}06, KSAF13, KNS97, KIS01, LRS^{+}23, LVF12, LYT^{+}20, LVM^{+}18, LVPK11, MBH^{+}99, MLVO05, MCM^{+}17, MBJ^{+}07, MM94a, MIK07, MVK^{+}20]. North [MSNK10, MFB^{+}09, MHvD^{+}24, MTK^{+}07, MIY^{+}09, MMB93, NTIO18, NDC05, NFKY21, NTM^{+}15, Oda94, OHF12, OM10, OBA01, PSM00, PFAM96, PAS^{+}18, PS06, PMG^{+}94, PKP^{+}00, PBH^{+}04, QCM^{+}16, REB^{+}03, RJHC99, RKD^{+}20, RZM^{+}03, SMK^{+}13, SKKW02, SKHI04, SKM04, SAT^{+}18, Sim92b, SB07, Spr92, SRM^{+}18, ST98, SPT^{+}17, TCO^{+}05, TMS^{+}08, TD02, TAN^{+}17b, WMD^{+}00, WMK^{+}99, WBQL99, WJ93, YOK^{+}17, YWM^{+}00, YOIW21, YCS^{+}15, ZSS08, ZHT14]. North-East [PLT09, ÅGN^{+}04, Bea03, BUE02, BB07, DLTI95, GHV95, IIS^{+}07, MWP02, Sim96, SR02, SGHW05, WQI00, WQ00, FH94, SB07]. north-eastern [HFC01, ISS02, MPW^{+}99, MAS^{+}98, PJO99, WFRS93, QCM^{+}16]. North-West [BJV^{+}17, Ano99, Col00, GHV95, HB99, MDVB^{+}20, PWE98, TAN^{+}17b, MM94a, MMB93]. north-western [LLSF01, LPSS04, YW94]. Northeast [FMM^{+}20, BBY08, BvDSDC18, CH16, HDJ15, MAH12, WL21, CGI^{+}19, DP01, DDZ09, FKF^{+}22, GHM21, LOS^{+}14, LSK^{+}18, LS21, MFMG20, MHRC18, MFRR96, SEM^{+}14, SP15, VGPL^{+}11, ZJH^{+}22]. Northeastern [MBY^{+}17, JMP^{+}14, KMM^{+}06, LTL^{+}22, MBY^{+}18, NSH^{+}17, NLS^{+}24, OFS^{+}16, RS92, SA10, TKW^{+}17, Tak04, TTI^{+}20, YKB08]. Northerly [YCS^{+}15]. Northern [BMH^{+}21, MRD^{+}19, RCB08, Aut08, BYM16, BS94, BASS11, BT99, BDSM07, Col99, CRVL^{+}17, CRW20, CP03, Cur04, ESA^{+}16, Fun07, Fun11, GHBM99, GHG^{+}19, Gla11, GI13, GBAD^{+}17, HYW04, HSLP19, HMS16, HCWF21, JMLG06, KYSM11, KYS15, LDH14, LPHM21, MBH^{+}99, MBJ^{+}07, MFG99, MM94a, MSC^{+}17, MWB^{+}00, MLR10, MMB93, NASTF10, Oda94, PVBV19, PMFC10, Pol96, ROH16, RJHC99, REM02, RD96, SHS^{+}23, SLL19, TCL^{+}12, TB92, VYGT^{+}20, VZP98, WMD^{+}06, WJM15, WKN^{+}95, Yam04, YKB08, HTE^{+}03, IMO^{+}12, JCCB15, JJBCW09, MCB^{+}16, SSSB03]. northward [KYU^{+}06, KNO^{+}04]. northwest [CJ04, DHC^{+}07, FCC^{+}19, HBR^{+}15, MMMS14, DH11, SHS^{+}23, SVEW^{+}13]. northwestern [HKM^{+}19, IST^{+}04, KBB^{+}20, MMRS16, MKK13, MTZG23, OWK04, OTO^{+}09, RS15, TAS04, YK96, CLW^{+}19, MSR20]. norvegicus [CLM^{+}21, FCJ^{+}15, MLP22]. Norway [FCJ^{+}15, HTE^{+}03, MLP22]. Norwegian [GTB10, HTE^{+}03, BS94, FM93, GPA^{+}21, OS95, SNV^{+}12, VAFG95, VSÅO07]. Note [Ano16]. notothenioid [LLCJ16]. NPZ [HNHP09]. Nuclear [MFS^{+}17]. nudus [TWK13]. Numerical [OHM^{+}10, WJP^{+}01, BC97, IYN^{+}09, KKNY92, KU95, LPG^{+}06, OHF12, PDER10, TKMS11]. nurseries [MLVO05, RSZ^{+}03]. nursery [BHJ^{+}04, CAB12, DMF^{+}17, FKUY16, GGF17, HONH04, KUO^{+}17, NBH99, PVMP03, RHRL12, RRF^{+}21, RBBG12, SHK^{+}19, WJM15, YTY96, YOY00]. nutrient [KNK^{+}18, OUKH04]. nutrient-rich [OUKH04]. Nutrients [MSL^{+}05, SWZ^{+}01]. Nutritional [ADPC21, DDB^{+}20, DBS^{+}19, HLH^{+}17]. NW [LCCQ^{+}22, MOE06, MCS^{+}06, OEV^{+}10, RCG^{+}15, áRÁSG^{+}16, SSP^{+}07, SGFR^{+}21].

O. [BWS^{+}01, FYA^{+}21, PMFC10, RZM^{+}03, SMB^{+}03a, TID^{+}96, WGFR06]. obesus [APR^{+}08, BHM02, HKM^{+}19, HKM^{+}21, HK06, LLCV18, MKK13, MSST16, MBB^{+}03, SMB03b]. Obituary [Ano95b]. object [DBFW13]. objectives [JPHA^{+}16]. obscurus [RHG^{+}13]. Observation [VIS92, óóSV18, RKD^{+}20]. Observations [RPT^{+}00, SKKS05, AI92, BT99, DLD^{+}23, HP02, JR07, MPM19, OA06, SMH^{+}92, SSSB03, TF08, VHJ99, WKN^{+}95]. observed [ECM^{+}01, KSMY00, OFS^{+}16, RPE98, VN97]. Observing [CMB^{+}15, Sch23]. obsesus [GCF^{+}21]. obtained [RMM02]. occasion [Kim23]. occurred [NSH^{+}17]. Occurrence [ARL93, KCW^{+}15, TKO^{+}14, TTI^{+}20, EPG^{+}16, HBLC22, IWK^{+}21, MESMM18, MLR10, PMG^{+}23, SMF^{+}05, YIT^{+}22]. occurring [AOVAG22, BH97, HSH^{+}22]. Ocean [DHC^{+}07, FC04, Hea93, LéEPW^{+}12, PMFC10, PLG^{+}10, SPS^{+}20, APL01, BBS99, BRPC08, CCC^{+}23, CHM^{+}94, CHF^{+}04, DBFW13, DDB17, ESA^{+}16, Han11, HHH^{+}16, HFHW19, HTT^{+}16, HKM^{+}19, HWSS07, HB92, HMT07, KFH00, LCH03, Mal20, NH01, Rob94, RWP11, Sch23, SVEW^{+}13, Sim96, SCS05, SPT^{+}17, TGRS^{+}19, TIH^{+}92, TH11, TMM^{+}07, UMK20, WWSE00, WGFR06, WSF^{+}14, YSW^{+}99, APMRH17, APMVOGMR19, AB02, ADPC21, AMDM12, AI05, BCR20, BGH09, BBT^{+}09, BML11, BW92, CLW^{+}19, CLT05, CH16, CBdSF^{+}23, DLTI95, DHC^{+}07, FYC22, FC04, GCF^{+}21, HRB^{+}18, HKM^{+}19, HPL13, HKLG07, HHH^{+}18, Kae23, KPHG14, KOWM16, KSP^{+}22, KYU^{+}06, KTS15, KNS97, KBF^{+}07, LLCV18, LCCdS^{+}19, LJR^{+}22, MESMM18, MSM^{+}13, MMSL19, MKK13, MSST16, MMBC07, MAH12, MVK^{+}20, MHB^{+}14, MWP02, Nis92, NTM^{+}15, Oda94, OWK04, OUKH04, PSM00, PFAM96, PL03]. Ocean [PBH^{+}04, Rog94, RWI^{+}16, RBBG12, RZM^{+}03, SF22, Sco95, SDHB07, SZX^{+}08, SLZ^{+}23, SSPY08, SSP^{+}11, TSK^{+}92, TSK^{+}95, TSK04, WQI00, WQ00, WSP^{+}07, WL21, YWM^{+}00, YOIW21, ZSS08, ZSY^{+}21, ZWC^{+}21]. ocean-mediated [HFHW19]. ocean/climate [YSW^{+}99]. Oceanic [Jes22, Kae17, MCG^{+}14, Ano99, AI04, BHS^{+}15, BBB^{+}19, FHHW98, GR98, KNS97, LLCV18, OM10, PKP^{+}00, QBMW99, REB^{+}03, TAS04, WZK97, WGS^{+}08, ZWL21]. oceanic-climatic [TAS04]. Oceanographic [APM^{+}12, CHPT20, DSPH07, FRBB14, HTLJ20, INM^{+}18, MP18, Sco95, THL^{+}18, TLS98, TBB^{+}03, AGSSL^{+}22, BPZR19, BBP^{+}13, CMB^{+}15, CMMK^{+}15, CG18, GBAD^{+}17, HSH^{+}22, HK06, IMS^{+}04, JYH^{+}18, JJBCW09, KOWM16, KBF^{+}07, KB08, LC95, LAPL21, MFMG20, MSC^{+}17, MSVY^{+}13, MP94, NdLOO23, OEV^{+}10, SC06, SMF^{+}05, SOTM^{+}18, SK04, VYGT^{+}20, YWM^{+}00, ZSS08, ZHX^{+}20]. Oceanography [HS05, War92, BGH09, BFSV08, Bri94, HCWF21, KD98, LPS19, LRS^{+}23, LJR^{+}22, Sch23, SR93, SS98, WRTP01, WBQL99, Gre99, BEiI^{+}23, Kim23]. oceanological [SDRL96]. Oceans [Har92, LBSS^{+}92, DPM^{+}11, HKWL17, MTSH15]. October [CP03, RJHC99]. octopoda [SCAG^{+}21]. octopus [AOVAG22, AOVAG22, FIDC00, LAPL21]. odontocetes [KOWM16]. off [ARL93, ADPC21, AG99, Ano99, ABS^{+}11, AS08, BJV^{+}17, BRPC08, CCC^{+}23, CDG^{+}19, Col00, CG18, CSFC05, DDB17, DLCQ22, DDS^{+}17, DBRSC16, DTC06, FYC22, FYK^{+}13, GMH^{+}99, GMH^{+}12, GSNFL99, GP94, HTE^{+}03, HYW04, HFC01, HFF^{+}19, HHK^{+}10, ISI^{+}18, IK97, KvdPBW17, KSYT97, KBB^{+}20, KK00, KKCL06, KFYP07, KB08, KNO^{+}04, KY17, LP10, SL95, LH96, Lyn03, MESMM18, MPW^{+}99, MRL^{+}14, MHM^{+}20, MDR^{+}16, MAS^{+}98, MTH^{+}04, MRHL09, MBKP08, Mul94, NSH^{+}17, OEV^{+}10, PHWM96, PBF00, PS06, RHG^{+}13, REM02, SRR99, SRR07, SBY^{+}15, SGFR^{+}21, SWS^{+}19, SSW^{+}17, SLL19, SR02, SBBB03, SK03, SK04, TMN^{+}15, Tak04, TTI^{+}20, TAN^{+}17b, TCC^{+}98, VFS^{+}24, WMKR09, WFRS93, XB09, YKB08]. Offshore [FKH^{+}17, BYM16, HDH^{+}05, KSC^{+}10, NZI95, OFS^{+}16, SSW^{+}17, TDE09, YKH^{+}21, ZNI96]. Ofunato [KKK^{+}17]. oglinum [CMMK^{+}15]. Oil [XB09, HBN^{+}21]. Okhotsk [MMF95, TKM^{+}22]. old [Sha95]. oligotrophic [MBKP08, UIU^{+}99]. oligotrophication [OUKH04]. olivacea [MMRH^{+}16, PBH^{+}04]. olivaceus [KUO^{+}17, SSW^{+}17]. olive [MMRH^{+}16, PBH^{+}04]. Ommastrephes [ASM^{+}15, FCC^{+}19, IMS^{+}04, ISI^{+}18, NII^{+}14, NTM^{+}15, YWM^{+}00]. oncaeid [NIIS04]. Onchorynchus [CAB^{+}01]. Oncorhynchus [APL^{+}96, AMDM12, AI04, AI05, BRPC08, BDSM07, BWS^{+}01, CHF^{+}04, EBO04, FYA^{+}21, HTT^{+}16, HQH^{+}06, HMT07, JTYB18, KNE^{+}04, KHB02, LML^{+}03, MRRN05, McK13, MAH12, PW12, PW14, PMT^{+}94, PMFC10, RZM^{+}03, RWLP12, RWP11, SKHN11, SMB^{+}03a, SW05, SVEW^{+}13, TID^{+}96, TR11, TH11, WTK^{+}16, WP93, WWSE00, WGFR06, WGW07, WGS^{+}08, WCP^{+}01, Wil01, WSF^{+}14, XDP^{+}20, YCH^{+}15]. One [PML06]. One-dimensional [PML06]. ongus [OE17]. onset [CHF^{+}04]. onshore [BYM16]. Onslow [COW^{+}99, QLB^{+}05]. ontogenetic [AYK03, AGS^{+}04, HHF09, IMO^{+}12, LCC15]. Ontogeny [ADPC21, BH18, FUA^{+}98, HCS^{+}09]. opalescens [PS16]. OPC [CC03]. OPC/MOCNESS [CC03]. open [MMRS16]. open-sea [MMRS16]. Operational [LPS19]. operations [BDBP93]. Ophiodon [ARL93]. opilio [SP13]. Opisthonema [CMMK^{+}15]. opportunities [BSF^{+}20]. opposing [LH96]. Optical [HDF^{+}99, GTB10, GR98]. optimal [DBB^{+}18, Gar97]. optimisation [KFH00]. Optimized [BTGM07]. Optimizing [BFF15, PH11]. Optimum [BCL04, Sai22, RPG^{+}22]. Oregon [BRPC08, ABS^{+}11, BPLC11, DDB17, DAW^{+}23, KHB02, LP10, LML^{+}03]. organisms [JHK^{+}15, LS15, RSC96, SAO^{+}17]. Organization [Woo93]. orientalis [FFF^{+}18, HFF^{+}19, IFF^{+}18, KKNY04, KBF^{+}07, Mat06, RMH^{+}19, SAT^{+}18, TTI^{+}20]. orientation [DLTI95, Sim96]. origin [BMOT17, RWI^{+}16]. originating [Dom04]. origins [HDH^{+}05, LCC15]. Oscillating [KEWDA18, CEM^{+}11]. oscillation [FRZVHM^{+}11, Dom23, PS16, TCC^{+}98, WTR04]. oscillations [BCR20, MMBC07]. Oshika [TWK13, TKW^{+}17]. osmoregulation [ZZ93]. Ossabaw [WKB^{+}05]. Osteichthyes [LLB^{+}20]. ostracods [LTL^{+}22]. other [JCCB15]. Otolith [BMHW13, ACT^{+}10, APGL03, BHV^{+}06, BASS11, FKUY16, GNP^{+}19, HBC07, HVHC10, Jes22, KTH^{+}15, KNO^{+}04, RSZ^{+}03, SPG^{+}16, YOY00]. otolith-based [GNP^{+}19]. Otsuchi [MWN^{+}23]. outbursts [HA07]. overexploited [ERR^{+}21]. overlap [EBFF17, KSAF13, Neu02, WP93, ZYT^{+}22]. overview [OCH99]. overwintering [GMH^{+}99, HTE^{+}03, Hea99b, HJ99, Jón99]. ovigerous [LA05]. Oxygen [JHC^{+}15, Bri94, CKA^{+}17, D’A93, JCCB15, KKK^{+}17, KCW^{+}15, Neu02, SBY^{+}15]. Oxygen-depleted [JHC^{+}15]. oxyrinchus [MCHSNEO13, RHP^{+}15]. Oyashio [KSYT97, KKNY04, MIY^{+}09, STI^{+}09, TCO^{+}05, TWKW01, TW05, TMS^{+}08, YW94]. oyster [KSM^{+}20, PKHF98, YIT^{+}22]. Ozernaya [BWS^{+}01].

Pacific [AMDM12, AI05, BB02, BF07, CLW^{+}19, FYC22, Kae23, KOWM16, KT93, KYU^{+}06, KNS97, MCM^{+}17, MVK^{+}20, NTM^{+}15, Oda94, PSM00, PFAM96, PBH^{+}04, RZM^{+}03, SAH^{+}18, Sim96, SDHB07, Woo93, YWM^{+}00, YOIW21, ZSS08, APMRH17, APMVOGMR19, ASM^{+}15, AGK^{+}08, BB03, BMH^{+}21, BBH99, BHM02, BG01, BWKM15, BBY08, BML11, BW92, CKA^{+}17, CC03, CLW^{+}19, CSK11, CH16, CAB^{+}01, Cur04, DLTI95, DL94, DPM^{+}11, Dom23, DP01, DSHL18, FCC^{+}19, FH94, FHHW98, FGGDSMF08, FKF^{+}22, FBRB12, FKSA21, FFF^{+}18, FYK^{+}13, FRHMAM^{+}06, Gar97, Gla11, GSNFL99, GAH^{+}19, GiIW^{+}20, HYW04, HKWL17, HBLC22, HJ10, Hea93, HKT^{+}03, HMS^{+}23, HKM^{+}19, HFF^{+}19, HAS^{+}19, HONH04, HLG^{+}11, HLWL12, HXC^{+}17, HHH^{+}18, HCS^{+}09, HLH^{+}17, IMS^{+}04, INM^{+}18, IST^{+}23, IFF^{+}18, IST^{+}04, IKK^{+}04, IYN^{+}09, III^{+}06, JTYB18, KNE^{+}04, Kae23, KTPM17, KSM^{+}20, KHN^{+}22]. Pacific [KPHG14, KOKM15, KSYT97, KTS15, KL01, KKNY04, KBF^{+}07, KNO^{+}04, KY17, LRS^{+}23, LAB^{+}98, LCCdS^{+}19, LYT^{+}20, LS01, Lyn03, MCM^{+}17, MESMM18, Mat06, MKK13, MSST16, MW92, MAH12, MIK07, MVK^{+}20, MTH^{+}04, MMRH^{+}16, MWP02, MSNK10, MTK^{+}07, MIY^{+}09, NTIO18, NSH^{+}17, NFKY21, NHS^{+}07, NBF^{+}01, OM10, OIA^{+}12, OWK^{+}03, OWK04, OTO^{+}09, OBA01, OUKH04, PFB^{+}16, PJO99, PMG^{+}23, PAS^{+}18, PMG^{+}94, Pol96, PKP^{+}00, RCB08, REG^{+}13, RSC96, RWI^{+}16, RBBG12, RMH^{+}19, RKZHC19, SME^{+}14, SGW^{+}21, SRR07, SMK^{+}13, SKKW02, SKHI04, SKM04, SF22, Sco95, SVEW^{+}13, SAT^{+}18, Sim92b, SC05, SMS^{+}21, Spr92, SEM^{+}14, SMDM98, SSPY08, SSP^{+}11, SK03, ST97, ST98, SK04, SP15, SPT^{+}17, TID^{+}96, TCO^{+}05, TSK^{+}92, TMS^{+}08, TKO^{+}14, TNK^{+}16, TNM^{+}02, TTI^{+}20, Tan99, Tan17a, TSK^{+}95, TAS04, TSK04, VMT^{+}23, WMD^{+}00, War95]. Pacific [WZK97, WMK^{+}99, WP93, WQI00, WQ00, WL21, YAM^{+}18, YOK^{+}17, YW94, YK96, YW07, YOIW21, YIT^{+}22, YKB08, YCS^{+}15, ZLTM11, ZHT14, ZHX^{+}20, ZWC^{+}21]. pacifica [MAS^{+}98, Tak04]. pacificus [KYU^{+}06, Mul94, Mul97]. Pagellus [GEGHPCC17, NSGL^{+}22, SFGE21]. Pagrus [Fra93, YOYK20]. palaeoenvironment [CCL^{+}05]. pallasi [BG01, CAB^{+}01, FBRB12, Tan17a, WQI00, WQ00]. pallasii [BWKM15, LYT^{+}20, REG^{+}13]. Palmyra [HK06]. Pandalus [FYKSP07, Han11, KFYP07, OA06, PBF00]. Panhandle [CMMK^{+}15]. Panulirus [Cap08, EF10]. Panulius [CB93]. Papers [BD93]. paradoxus [IMO^{+}12, KvdPBW17]. paralarvae [CG18, DBRSC16, MRL^{+}14, NII^{+}14, áRÁSG^{+}16, SCAG^{+}21]. paralarval [PS16]. Paralichthys [KUO^{+}17, SSW^{+}17]. Paralithodes [LA05]. parameter [GiIW^{+}20]. parameters [BLH98, CDG^{+}19, CH95, JGS93, MHS^{+}21]. parametric [HKWL17]. parent [EF10]. Parsons [Per23]. part [FRZVHM^{+}11, SFK^{+}20, IXW^{+}10]. particle [EvST^{+}17, IYN^{+}09, NYI^{+}13, YAM^{+}18]. particle-tracking [IYN^{+}09, NYI^{+}13, YAM^{+}18]. particles [BSS94, MFP^{+}03]. Pass [BHJ^{+}04, ZP21a]. Passes [LJH^{+}05, SKKS05, ZP21b, Coy05]. passive [BWK^{+}99, DST11, HP02]. past [LYT^{+}20, Sim92a]. PAT [BFF15]. Patagonia [LPCG23]. Patagonian [ABI^{+}21, AAI16, BBR^{+}05, HMM01, LSD^{+}21]. patagonica [BBR^{+}05]. patch [DPL02]. Patchiness [MOE06, OTO^{+}09, Bez00, FCL93]. pathway [Dom04]. pathways [GQPGA04, MMI^{+}22, SCDA10, SJB^{+}22]. pattern [BB03, Cur04, ESA09, LLCJ16, LJM^{+}10, QCM^{+}16, SNV^{+}12]. Patterns [FODCN00, SC97, ACT^{+}10, BJV^{+}17, BCBDA10, BBS99, BDAMD14, BRR05, CSB94, CG18, DTO^{+}23, DPL02, DDZ09, DABM^{+}06, FMV03, FFF^{+}18, GSBB07, HGS^{+}21, HJ10, HL07, HSH^{+}22, ICB^{+}08, JMLG06, KFS22, KMB00, KVR^{+}18, KNO^{+}04, LPCG23, LPH^{+}19, LéEPW^{+}12, LAB^{+}05, MESMM18, MBY^{+}18, MBE^{+}15, MWN^{+}23, MTH^{+}04, MWP02, NSGL^{+}22, NFKY21, QBMW99, SME^{+}14, SRR07, SPS^{+}20, SFGE21, SMK02, SKHI04, SHG12, SAG^{+}09, SWAAB20, SWS^{+}19, SMF^{+}05, TMS^{+}08, WJP^{+}01]. paucispinis [ZLTM11]. pCO [KTO^{+}11]. pealeii [DHC^{+}07]. pearlside [SSR13]. Pecten [HRS^{+}21]. Pelagic [WMD^{+}06, ARL93, BBMY93, BHS^{+}15, Buc92, CHPT20, HHK^{+}17, HRB^{+}18, HPG^{+}20, HALO00, KTPM17, KB08, LDAWM10, LPG^{+}06, LSD^{+}21, MTL^{+}22, MSR20, MTSH15, MLRS07, ODMRM98, OFS^{+}16, PM95, PAS^{+}18, PG06, RSF13, REG^{+}13, RSC96, RHG^{+}13, RG97, RD96, Shi98, TAN^{+}17b, VSÅO07, WKR^{+}18, WM06]. pelamis [And03, GCF^{+}21, LPS19, LMBL03, MSST16, MSNK10, NPLS22]. penaeid [YMB99, dBdOJdO^{+}22]. Penaeidae [MHS^{+}21]. Penaeus [BYM16]. Peninsula [GPCGdlT^{+}22, HT99, REM02, TWK13, TKW^{+}17, AGSSL^{+}22, KK00, LPCA15, PVBV19, áRÁSG^{+}16, SMF^{+}05]. perceived [SNV^{+}12]. perch [KPHG14, NH06, RBBG12, Sco95]. Perciformes [CBdSF^{+}23, RRF^{+}21]. Perfect [FC04]. performance [Dom09]. period [HMS16, Nis19, NHS^{+}07, RCG^{+}15, SMA14, ZP21a]. Periodic [REB^{+}03]. periodicity [SPM^{+}24]. periods [KEWDA18, ROH16]. permeability [HBG^{+}16]. persistence [BHH98]. personatus [KKNY92, TY04]. perspective [DL94, MSNK10, Ric96, TMMM20, WJ93]. pertinent [Bri94]. perturbations [FYC22]. Peru [AS08, CDG^{+}19, DLCQ22, GSBB07]. Peruvian [JCH04]. petrale [HTLJ20]. phase [QBMW99]. phases [GHG^{+}19]. phenology [SCTB19, SMS^{+}19, éSMB20]. phenomena [KNS97, LGM^{+}02]. phenomenal [PW12]. Phocoenoides [OM10].

Phoebastria [MJH14]. Phosichthyidae [LLB^{+}20]. Photosynthesis [PSJF93]. phyllosoma [GBAD^{+}17]. Physical [Har92, LSD^{+}21, PMG^{+}94, SCKJ^{+}18, VMG01, WBQL99, AB02, BHJ^{+}04, CCM^{+}08, CCK^{+}22, CMM06, ECM^{+}01, GQPGA04, HG98, HBG^{+}16, ITH23, LLCJ16, MEK^{+}09, MLM^{+}98, MWR^{+}98, MMB93, NKS00, ODMRM98, SBK^{+}01, ST97, ST98, WHT92]. physical/biogeochemical [MEK^{+}09]. physically [DST11, HNHP09]. physics [Bau98, FvPH^{+}16]. physiographic [KEJK00]. physiological [RPG^{+}22, DHM^{+}15, HKM^{+}21]. physiology [FDT^{+}99]. phytoplankton [FYKSP07, KWO^{+}18, OK17, PSJF93, RFM^{+}21, RP93, SWZ^{+}01, TSK^{+}92]. PICES [Woo93, Kas97, Kas98, Kas99, Liv00, Woo95, Woo97]. pieces [DBS^{+}19]. pilchardus [BJV^{+}17, BPP07, BRC^{+}03, áCGNGC19, GPL^{+}11, GVRC04, HBG^{+}16, LPSS04, MEK^{+}09, MHvD^{+}24, PBL07, SGS^{+}06, VYGT^{+}20]. pilot [KOKM15]. Pink [BRO18, BWS^{+}01, CAB^{+}01, FYA^{+}21, MAH12, PHWM96, PBF00, RZM^{+}03, TID^{+}96, Wat17, WCP^{+}01, Wil01]. pinnatifida [KNK^{+}18]. Pisces [ASCM12]. piscivory [LMB^{+}19]. Placopecten [TCS^{+}09, ZJH^{+}22]. plaice [FODCN00, FMYN06]. planktivorous [RG97, CH92]. Plankton [BF07, BMPC16, BM99a, BM99b, ECM^{+}01, HDF^{+}99, HMS^{+}23, Oda94, RPE98, YCS^{+}15, AW92, ASK99, BRO18, GTB10, GR98, LVF12, OEV^{+}10, PST03, Rob94, Rog94, SDRL96, Sko05, ST98, TKH08, UIU^{+}99]. planktonic [HL07, LTL^{+}22, Mul94, NLS^{+}24]. Plant [MFS^{+}17]. Plasticity [BGH09, HRB^{+}18]. Plata [ASCM12]. platessa [FODCN00, FMYN06]. Platichthys [YOY00]. platypterus [HLG^{+}11, MHB^{+}14, RCPS09]. Pleuragramma [BCA^{+}18]. Pleurogrammus [MFH05]. Pleuronectes [FODCN00, FMYN06]. plumchrus [BWJ03, LP10]. plume [EBO04, REM02, SMB^{+}03a, SMH^{+}92]. Point [ARL93]. points [FMM^{+}20]. Polar [KT93, WTR04, CÅP^{+}13]. pollock [AYMK01, BCBDA10, BBMY93, BBS99, CEM^{+}11, Fun07, Fun11, FYK^{+}13, HYW04, HWSS07, HONH04, IST^{+}04, KNS^{+}22, KTH^{+}15, KEWDA18, LK21, LDAWM10, MTH^{+}04, NKS00, NHS^{+}07, OTIK20, RWDA^{+}21, SS94, SB94, SADA^{+}23, Spr92, UMK20, VIS92, WSC05, Yam04, YCH^{+}15, RWDA^{+}21]. pollution [RS92]. polyactis [HGS^{+}21, KJZ97, XWL^{+}23]. polynya [MFRR96]. polyxystra [CRW20, LDH14]. Pomatomus [CTWS08, VHLM15]. pomfret [QCM^{+}16]. Pool [GAH^{+}19, FMG^{+}22]. pop [AMD^{+}16, APR^{+}08, CÅP^{+}13, DPM^{+}11, GJR18, HLG^{+}11, HKLG07, LPG^{+}06, PECG08, RHG^{+}13]. pop-up [AMD^{+}16, APR^{+}08, CÅP^{+}13, DPM^{+}11, GJR18, HLG^{+}11, HKLG07, LPG^{+}06, PECG08, RHG^{+}13]. Population [Esc98, HMTG^{+}05, LJBR20, Ric96, SPS^{+}20, BB03, BLH98, BHH98, BRR05, CAB12, CPM^{+}15, DSHL18, FPBDC11, FKUY16, GPS22, HA07, HRS^{+}21, JCH04, KPHG14, KO95, KKCL06, KKNY92, LBC23, LPH^{+}19, MLC^{+}98, NSH^{+}17, NdLOO23, PHH13, PEKL14, RCS98, RF04, RMM02, RWI^{+}16, SSW^{+}17, SGN^{+}05, SMF^{+}05, SC97, SK03, SP15, Tan02, WGW07, WSC05, YKI98, ZLTM11]. Populations [Nak98, AH97, BH97, BF07, BRO18, Buc92, CWCM14, CAB^{+}01, DAW^{+}23, IXW^{+}10, Kae23, LCH03, LBW^{+}05, MBY^{+}17, MAHG94, MPM^{+}13, PLP^{+}11, PQH16, RAK^{+}17, SGFR^{+}21, SPM^{+}19, SPLS15, TWK13, WWSE00, WQI00]. porbeagle [CJ04]. porpoises [OM10]. Port [MW92, MWGK92]. Portugal [Erz05, SBBB03]. Portuguese [TSG^{+}20]. portunid [HSH^{+}22]. Portunus [YTIS95]. position [WP93]. positively [CHPT20]. Possibility [TTI^{+}20]. Possible [KO95, BMO^{+}99, DHMT96, GEGHPCC17, LBC23, SGN^{+}05, ZHL^{+}03]. post [BPS^{+}14, MSS12, MSC^{+}17, PSS^{+}21, REL07, WKB^{+}05]. post-fertilised [PSS^{+}21]. post-larvae [WKB^{+}05]. post-larval [MSC^{+}17, REL07]. post-smolt [BPS^{+}14]. postflexion [SRR99]. postlarvae [IN00, PTS^{+}24]. postlarval [NFN00]. postsmolts [FHD98]. pot [BLG^{+}16]. Potential [AMDM12, HPL13, LPHM21, LMBL03, ASM^{+}15, ASK99, AI04, CAR^{+}10, DPK^{+}08, Dom04, DBRSC16, GIT^{+}13, HFHW19, HBO^{+}01, HMS16, ISI^{+}18, ITH23, KY17, LA05, PBL07, QCR22, SMS^{+}21, SQW^{+}99, TNK^{+}16, TTY^{+}23, óT10]. potentially [AB02]. poutassou [BC97, HEG08, MMRS16, MP18]. Power [MFS^{+}17]. pp [Gre99]. practices [Sai22]. prawn [MDR^{+}16, EvST^{+}17]. prawns [BYM16]. precision [PSC05, WSP^{+}07]. Predation [BBMY93, BG01, BLG^{+}16, CH92, Gla11, SSR13, UMK20, VFS^{+}24, Wil01, Zam01]. predator [HJ10, KSAF13, Neu02, PP01, PDD03, VCKH05, ZYT^{+}22]. predators [HRB^{+}18, HKA^{+}06, MLRS07, SBY^{+}15]. predatory [SES^{+}20]. predict [MSS12, WM06]. predictability [HP02]. predicted [ISS02]. Predicting [BK94a, CCP07, EBFF17, KTPM17, KSM^{+}20, LPCA15, LAB^{+}98, MLR10, OÅL00, SLL19, SP15, GHG^{+}19, LML^{+}03, SLZ^{+}23, VN97, XMH^{+}18]. Prediction [ITH23, WJT97, ASM^{+}15, APLG07, SL09]. predictions [BBA^{+}21, MM94b, RQN^{+}99]. predictor [WJW20]. predictors [SNL19]. preference [SSP^{+}11]. preferences [APMRH17, MYHvdL15, PLT09, RDE^{+}07, SZX^{+}08, Swa99, YMB99]. Preferential [BRFRJRLC18]. preferred [DGB^{+}16, Jan16]. preliminary [LMBL03, Ols01, SMH^{+}92]. Preparation [ZZ93]. prerecruits [HPG^{+}20]. present [Sim92a]. Presentation [KYY00]. Press [Gre99]. pressure [BAL^{+}99, Gla11]. Prey [MTK^{+}07, SMB^{+}03a, APM^{+}12, CC03, CP92, FBRB12, HL07, HNHP09, HMS16, Jan16, KSAF13, LH96, MWGK92, MWN^{+}23, Neu02, PDD03, PA14, Pol96, Rog94, SL09, SMH^{+}92, Tan99, VCKH05, WSC05, ZYT^{+}22]. prey-fish [Rog94]. Pribilof [WSC05]. primary [AYK03, GFG98, MSL^{+}05, MPM^{+}13, TKM^{+}22]. Prince [BMPC16, BG01, BWKM15, CAB^{+}01, CCSS01, ECM^{+}01, GV01, NBF^{+}01, VMG01, WJP^{+}01, WCP^{+}01]. principles [Bow11]. Prionace [GPCGdlT^{+}22, HRB^{+}18]. probability [CSS^{+}21]. probable [HDH^{+}05]. procedures [AMD^{+}16]. Process [AMK08, APGL03, PST03, WPN12]. Process-based [AMK08]. Processes [CRC11, APL01, BBB^{+}16, BHJ^{+}04, CMM06, FIDC00, LVC^{+}05, LRL^{+}06, LC95, LML^{+}03, MMRS16, NKS00, QBMW99, SHS^{+}23, SMF96, SHM05, SOTM^{+}18, VIS92, WHT92, WCP^{+}01, WJ93]. producing [GYS14]. product [HHB^{+}15]. Production [Ric96, AYK03, Col00, CP03, DMF^{+}17, DB03, ERR^{+}21, GFG98, GFO14, Kae23, KL01, KHB02, MSL^{+}05, RJHC99, Rob94, RBBG12, RWP11, Sko05, SGS^{+}06, SMF96, SMDM98, TYO21, Tan02, TKM^{+}22, WMD^{+}06, WJ93]. productive [CGMM10]. Productivity [LVM^{+}18, APL^{+}96, AMDM12, BLD^{+}03, DAW^{+}23, DB03, KMK^{+}18, LDAWM10, MCM^{+}17, Mal20, MPM^{+}13, RFM^{+}21, SHM05, TJW^{+}03, TKM^{+}22, YWI^{+}05, ZHT14]. productus [CC03, SRR07, Tan99, VMT^{+}23]. profiles [CCP07]. program [IST^{+}04, WHT92]. Projections [KNK^{+}18, NFO^{+}23]. prominent [SJB^{+}22]. promote [SES^{+}20]. promotes [SFK^{+}20]. properties [GBAD^{+}17, KSAF13, WTK^{+}16]. protected [BJCS12, BCJ^{+}13, CLKP19, CAR^{+}10, Dom04, NSH^{+}17]. protection [PVBV19]. protists [FBRB12]. provide [SWS^{+}19]. provides [YKH^{+}21]. providing [ZWL21]. Province [GAH^{+}19]. proxies [SPV96]. Proximate [PJO99]. Pseudocalanus [MLM^{+}98, MKF^{+}03]. Pseudosciaena [KJZ97]. puerulus [CB93, Cap08]. Puffins [SPT^{+}17]. Puffinus [VCKH05]. punctatus [RRF^{+}21]. purse [EPG^{+}16, GAH^{+}19, MMRH^{+}16]. purse-seine [EPG^{+}16, GAH^{+}19, MMRH^{+}16]. putative [RSZ^{+}03]. Putting [DBS^{+}19].

quality [GCW17, KUO^{+}17]. quantification [LRL^{+}06, MWN^{+}23, óóSV18]. Quantifying [BvDSDC18, PJB05, WGW07, PFSL09]. quantitative [LPG^{+}06]. Queen [JTYB18]. Queensland [SBD^{+}19]. quinqueradiata [KSMY00, UTMS06].

R [Per23]. Rachycentridae [CBdSF^{+}23]. Rachycentron [CBdSF^{+}23]. radar [HP02]. radiata [GHM21, SB06]. radioactive [Kae17]. Radiocesium [SAO^{+}17]. rainfall [GHG^{+}19, dBdOJdO^{+}22]. range [HGS^{+}21, HGG^{+}17]. ranging [HKA^{+}06]. Ranina [SBD^{+}19]. rapid [AGK^{+}08]. rate [AM18, DTC06, HK06, HMT07, KNS^{+}22, OWK^{+}03, SKT21, SPG^{+}16, ST95, TW05, TCL^{+}12, Tan17a, XDP^{+}20]. rates [BBH99, BML11, CIS20, FML^{+}14, FUA^{+}98, GHBM99, HBC07, IHHH99, KvdPBW17, LLCV18, MMBC07, MHB^{+}14, MWP02, NGGJ09, SKHN11, SF22, TA06, WSF^{+}14, ZKT07]. ratio [MWGK92]. ratios [FKUY16, MCHSNEO13, MFS^{+}17, OM10, OKT^{+}23, YOY00]. rays [CGMM10]. Re [HBPC15]. Re-evaluating [HBPC15]. reactions [VN97]. ready [SMS^{+}23]. real [ZWL21]. really [Spr92]. reappraisal [HSS19]. reared [ZZ93]. recapture [MFH05]. recognition [BB03]. recommendations [YWI^{+}05]. Reconstructing [NHNA07]. record [FPFL13]. recorded [RDE^{+}07]. Recorder [BM99a, BM99b, RPE98, YCS^{+}15]. recording [KSMY00]. records [MIK07]. Recovery [Kaw93, HMT07, LBC23, LVM^{+}18, MMMS14, ONK17, PH11, SP15]. recreational [BHS^{+}15, CCHL23, HBLC22, HKLG07, WMKR09]. recruit [CSS^{+}21, OHF12]. recruited [YMB99]. Recruitment [LOS^{+}14, MP94, Nak98, OCH99, YTIS95, ZHL^{+}03, ABI^{+}21, ACG^{+}16, APL01, APLG07, AAKMG06, ASK99, BRC04, BCJ^{+}13, BCGB14, BCR08, BUE^{+}98, BUE02, BFSV08, BDVS^{+}19, BB07, BDTR23, CCM^{+}08, CLPC18, CCC^{+}23, CRC11, CH95, Col99, CRW20, DPL^{+}20, DBGW04, Dom04, DBRSC16, EF10, FIDC00, Fun07, Fun11, FYK^{+}13, GPA^{+}21, GI13, GPS22, GMH^{+}12, GFO14, HTLJ20, HBPC15, HKWL17, HMP92, HSS19, INM^{+}18, ICB^{+}08, IFF^{+}18, KOS^{+}19, KSP^{+}22, KD98, KM93, KVR^{+}18, KWO^{+}18, LRS^{+}23, SL95, Lou10, MEK^{+}09, MMSL19, MDR^{+}16, MM94b, MWB^{+}00, NYI11, Nis19, OIA^{+}12, OS95, OHS06, OH23, PHH^{+}98, PJB05, PGL^{+}15, PCR^{+}18, RKD^{+}20, RCD^{+}99, RWLP12, SHG^{+}22, SC06, Sha13, SC05, SB07, SEM^{+}14, SQW^{+}99, SB04, SOTM^{+}18, SCS05, SP13, TSK^{+}22, TKW^{+}17, Tan17a, TD02, THL^{+}18, Tyl92, UYF92]. recruitment [VMT^{+}23, VGPL^{+}11, WPN12, WQI00, WQ00, WL21, WJW20, XMH^{+}18, YOY00, YWI^{+}05, ZLTM11, dBdOJdO^{+}22, óT10]. recruits [GGQF22, Han11]. Red [DPL^{+}20, BASS11, CP92, GPS22, KSP^{+}22, LA05, MWB^{+}00, POA^{+}17, SPM^{+}19, YOYK20]. redfish [DH11, RD96]. redfish-Calanus-microplankton [RD96]. Reduced [KHN^{+}22, JPMH20, VSÅO07]. reduction [LK21]. Reef [MSVY^{+}13, JMP^{+}14, KVR^{+}18, LéEPW^{+}12, SPM^{+}24]. reef-fish [LéEPW^{+}12]. reference [BSS94, KEJK00, SKM06]. refined [ZWC^{+}21]. reflect [SMF^{+}05]. refugium [APL^{+}08]. Regime [KYA^{+}15, SP13, AS08, BNM^{+}00, áCGNGC19, FH94, GI13, IMS^{+}04, LBSS^{+}92, SB05, Ste98]. regimes [CHF^{+}04]. region [BT99, CCM^{+}08, CC03, HFC01, HK06, ICB^{+}08, KSYT97, KKNY04, KYS15, MBE^{+}15, MLM^{+}98, MLC^{+}98, MIK07, MMI^{+}22, NZI95, NY08, PHH^{+}98, PECG08, SKKW02, SMK02, STI^{+}09, TWKW01, TW05, TMS^{+}08, TKM^{+}22, TTH15, WK03, YMK^{+}15]. regional [ERE^{+}10, FvPH^{+}16, KWB^{+}16, KPW19, LAG^{+}11, PWML12, UTMS06]. regions [CÅP^{+}13, GFG98, LBSS^{+}92, MHS^{+}21, SQW^{+}99, TCS^{+}09]. regression [MCB^{+}16]. regulated [KR10]. regulating [ETB^{+}17]. Regulation [Ric96, AI05, BCJ^{+}13, JCH04]. Reinhardtius [ÅGN^{+}04, SCDA10, YLA13]. relate [HBLC22]. related [BB07, HT99, IMS^{+}04, JGS93, LCH03, MHvD^{+}24, NHM94, RF04, SSR13]. Relating [SVEW^{+}13]. relation [AG99, BBP^{+}13, BRN^{+}95, BNM^{+}00, Bea03, BGH09, BDSM07, CDG^{+}19, CMMK^{+}15, CLT05, CG18, D’A93, DDB17, FRS^{+}05, FYC22, FHK^{+}12, GBAD^{+}17, HFC01, HBR^{+}15, HBG^{+}16, JCCB15, KNE^{+}04, KOWM16, KSC^{+}10, KCW^{+}15, KNS97, KKNY04, KBF^{+}07, KNO^{+}04, LYT^{+}20, LC95, MSM^{+}13, MHB^{+}14, MKH^{+}13, MMB93, NFN00, NY08, OE17, PSN^{+}99, RPT^{+}00, RWP11, SME^{+}14, SKHN11, SKT21, SGL04, SC06, SOTM^{+}18, SK04, SCF^{+}20, SKNT14, TKO^{+}14, TSK^{+}95, YLA13, ZSS08]. Relationship [CLPC18, HMM01, Nis19, QM01, WTK^{+}16, ZKT07, AAI16, And03, AS08, CRC11, CRVL^{+}17, LA05, SA10, SC05, SCTB19, YW94]. Relationships [BUE^{+}98, CSB94, ERR^{+}21, RAT^{+}02, Rog94, RS92, WGS^{+}08, BBY08, FCC^{+}19, HCC^{+}09, KSAF13, KGW13, Mal20, MTP07, OHF12, SPT^{+}17, WQI00, WQ00, ZHT14]. Relative [AOVAG22, DBGW04, MJH14, YOY00, BHM02, CSFC05, CP92, Coy05, DAW^{+}23, DHC^{+}07, Erz05, FFF^{+}18, HALO00, KMB00, LCCdS^{+}19, LH96, Mul94]. release [MFH05, SKHN11, Sai22]. released [HKLG07, SF22]. relevance [WL21]. religiosa [YKI98]. Remote [Sim92a, BGM^{+}18, FMV03, Hor00, MPM19, MSNK10]. Remotely [RCB08, WRTP01, ZSS08]. reorganization [SDRL96]. replenishment [CAR^{+}10]. replicated [RMM02]. Reply [Coy94]. Report [Kas98, LBSS^{+}92]. Reproduction [HSS19, VHLM15, BGP^{+}06, HYW04, Mul94, NIIS04, dBdOJdO^{+}22]. Reproductive [MSM^{+}13, AMD^{+}16, AB02, HSLP19, MMSL19, MPM19, SPM^{+}24, óT10]. requirement [GYS14]. Research [KYY00, BDBP93, CH99]. Reserve [SFA14]. reserves [PEKL14]. residence [LMB^{+}19, PSN^{+}99]. resident [SGL22]. resilience [HGG^{+}17]. resolution [LSD^{+}21, LRBJ21, MHM^{+}20]. resolutions [SLZ^{+}23]. resolved [HVHC10, ODMRM98]. resource [FCJ^{+}15, PVBV19, VPRG13]. resources [HNHP09, PFB^{+}16]. respect [HDF^{+}99]. respiration [IHHH99]. Response [DLCQ22, RPC^{+}19, Sim92b, AGK^{+}08, ABS^{+}11, FYK^{+}21, HPG^{+}20, KTS15, MHG^{+}11, RPE98, SBT20, SLM13, SB06]. Responses [SGFR^{+}21, TCL^{+}12, ECM^{+}01, RS15, WWSE00]. Restricted [CBdSF^{+}23]. result [Gla11, LAPL21]. resulting [AW92, BCR08]. results [EHW08, GTB10, HQW^{+}99, LH96, LMBL03, RPE98]. Retention [MHM^{+}20, NLS^{+}24, NH06, BSG^{+}13, BSF01b, CRC11, ETB05, FIDC00, GQPGA04, HLMS03, KR10, LRL^{+}06, MBE^{+}15, MRBBHL14, MGHS14, RPT^{+}00, SSSB03, SPLS15]. retention-dispersal [KR10]. retention/dispersion [HLMS03]. Retrospective [BSF^{+}20, MMMS14, OK17]. return [HQH^{+}06, McK13, PW14, SKHN11, TR11, WTK^{+}16, WSF^{+}14]. returning [HTL^{+}00, MFG99, TIH^{+}92]. reveal [BEF^{+}12, HCC^{+}09, LMB^{+}19, SPLS15, SPT^{+}17, UMK20, YAM^{+}18]. revealed [CPM^{+}15, FKUY16, GJR18, MWN^{+}23, YW07]. reveals [BBB^{+}19, GPS22, HKWL17, MESMM18]. reversed [KOS^{+}19]. Review [Gra98, Gre99, Par99, Bri94, DLD^{+}23, Hor00, Kae17, Spr92]. reviewers [Ano07, Ano10, BZ21]. reviewing [MRL^{+}14]. Reviews [Ano94]. Revillagigedo [SFA14]. Revision [CAB12]. revisited [Sha13]. Revisiting [TTY^{+}23]. Reyes [ARL93]. reynaudi [DBRSC16]. reynaudii [MRL^{+}14]. Rhincodon [WSP^{+}07]. Rhinoptera [CGMM10]. rhombus [OKT^{+}23, OHM^{+}10]. rhythm [SK04]. rhythms [XMW^{+}23]. rich [OUKH04, YKH^{+}21]. Ridge [SPM^{+}19]. ridley [MMRH^{+}16, PBH^{+}04]. ring [AI92]. ringens [RPG^{+}22, CRVL^{+}17, GNP^{+}19, GSBB07, HSLP19, SLL19]. Río [ASCM12]. risk [SSR13, Wil01]. Risso [BC97]. River [APL^{+}96, MFG99, Sim96, SKNT14, XDP^{+}20, XWL^{+}23, RDF^{+}11, BWS^{+}01, DLTI95, EBO04, HMT07, IK97, JMP^{+}14, LPSS04, McK13, MCG^{+}14, PW12, PW14, RFM^{+}21, SMB^{+}03a, SW05, SAO^{+}17, SMH^{+}92, SOTM^{+}18, TIH^{+}92, TH11, WSF^{+}14]. riverine [BBB^{+}16]. RNA [MWGK92]. rock [BLG^{+}16, CB93, Cap08, CRW20, FML^{+}14, HGG^{+}17, LDH14, LJM^{+}10]. rockfish [BBY08, DP01, MLRS07, PDER10, RSF13, SRR07, ZLTM11]. rocky [SPM^{+}24]. rocky-reef [SPM^{+}24]. Role [GHG^{+}19, BBB^{+}16, CHM^{+}94, DH11, DDB^{+}20, MLVO05, OEV^{+}10, TTY^{+}23, TH11]. roles [RKZHC19, UTMS06]. rookeries [CL05]. rose [GGQF22]. rosefish [SPS^{+}20]. Ross [BCA^{+}18, MKH^{+}13]. round [VCB^{+}98]. route [YAM^{+}18]. routes [OR12]. ruberrimus [BBY08]. run [TH11]. Running [JPMH20]. rupestris [CLH^{+}22].

S [Gre99, CG18]. sablefish [GJR18, KMB00, SC06, SE19, THL^{+}18]. SABRE [CH99, HQW^{+}99]. sac [BBMY93]. Sagami [TKH08]. sagax [Cur04, CCP07, GSBB07, HMS^{+}23, KYSM11, LS01, Lyn03, MYHvdL15, MGHS14, RCB08, SWS^{+}19, SSSB03, VCB^{+}98, WMD^{+}06]. Sagitta [BT99, TSK^{+}95]. Saharan [MEK^{+}09]. saida [MFRR96]. sailfish [HLG^{+}11, MHB^{+}14, RCPS09]. saira [FKSA21, INM^{+}18, IST^{+}04, III^{+}06, KHN^{+}22, KNO^{+}04, MVK^{+}20, OWK^{+}03, OWK04, OTO^{+}09, SK04, TKO^{+}14, TNK^{+}16, YW07, YOIW21]. Salangichthys [AHAM03]. salar [BPS^{+}14, DDS^{+}17, FHD98, MMMS14, RFD^{+}04, RDF^{+}11]. Salinity [MM94a, BAL^{+}99, KJZ97, KIS01, SKKS05, YMB99]. Salish [RAK^{+}17]. Salmo [BPS^{+}14, DDS^{+}17, FHD98, MMMS14, RFD^{+}04, RDF^{+}11, RDE^{+}07]. Salmon [BRO18, HTT^{+}16, XDP^{+}20, APL^{+}96, AI04, AI05, BRN^{+}95, BRPC08, BPLC11, BA12, BDSM07, BWS^{+}01, BPS^{+}14, Col00, CP92, CHF^{+}04, DDB17, DLTI95, DAW^{+}23, DDS^{+}17, DHM^{+}15, EBO04, FHD98, FYA^{+}21, Gar97, HHH^{+}16, HTL^{+}00, HFHW19, HQH^{+}06, HMT07, JPMH20, JTYB18, KNE^{+}04, Kae23, KHB02, LMB^{+}19, MRRN05, MCM^{+}17, Mal20, MSS12, McK13, MCG^{+}14, MFG99, MAH12, MMMS14, MWN^{+}23, Mor11, MWP02, PW12, PW14, PHWM96, PMT^{+}94, PMFC10, RFD^{+}04, RDF^{+}11, RFM^{+}21, RAK^{+}17, RZM^{+}03, RWLP12, SKHN11, Sai22, SMB^{+}03a, SW05, SVEW^{+}13, Sim96, SMH^{+}92, TID^{+}96, TR11, TIH^{+}92, TH11, VFS^{+}24, WTK^{+}16, Wat17, WS08, WP93, WGFR06, WGW07, WGS^{+}08, WCP^{+}01, Wil01, WSF^{+}14, YCH^{+}15, Zam01]. salmonids [Hea93, WMD^{+}00]. saltatrix [CTWS08, VHLM15]. Salvelinus [RDE^{+}07]. same [MWN^{+}23]. Samoa [Dom09, DSPH07]. sampler [COSC97, LVF12, PSC05, VCB^{+}98]. samples [MWN^{+}23, OK17, YCS^{+}15]. Sampling [AW92, BDBP93, FCJ^{+}15, LVF12, SB94]. San [Gre99, TMMM20, Zam01]. Sanctuaries [MJH14, HKA^{+}06]. sand [KKNY92, MW92, MWGK92, NNOU20, SJB^{+}22]. sandeel [BSS94, LVM^{+}18, TY04]. sandeels [PWE98]. sandy [OKU17, XMW^{+}23]. Sanriku [KNK^{+}18, MAS^{+}98, MTT^{+}17, WTK^{+}16]. sapidus [CWCM14, ERE^{+}10, OHF12, REL07, TDE09]. Sardina [BJV^{+}17, BPP07, BRC^{+}03, áCGNGC19, GPL^{+}11, GVRC04, HBG^{+}16, LPSS04, MEK^{+}09, MHvD^{+}24, PBL07, SGS^{+}06, VYGT^{+}20]. Sardine [LBSS^{+}92, AH97, BH97, BPP07, BBB^{+}16, BRC^{+}03, áCGNGC19, Cur04, CCP07, GPL^{+}11, GMH^{+}12, GVRC04, GSBB07, HZW^{+}98, HMS^{+}23, HBN^{+}21, HBG^{+}16, IYN^{+}09, ISN^{+}11, KWB^{+}16, KKS92, Kaw93, KO95, KKCL06, KM93, KWO^{+}18, KYSM11, LPSS04, LS01, Lyn03, MHM^{+}20, MYHvdL15, MGHS14, MHvD^{+}24, NHM94, NZI95, NFN00, NTIO18, NFKY21, NY08, NYI11, NYI^{+}13, Nis19, NY03, OTH09, OIA^{+}12, PBL07, PCR^{+}18, RCB08, SGFR^{+}21, SWS^{+}19, SMS^{+}21, SHK^{+}19, SGS^{+}06, SSSB03, SBBB03, SK03, SS98, TF08, VCB^{+}98, VYGT^{+}20, WMD^{+}06, WZK97, WZK^{+}98, XB09, YSW^{+}99, YKH^{+}21, YWI^{+}05, ZNI96]. sardinella [BJV^{+}17, DBB^{+}18, TAN^{+}17b, HBN^{+}21, MBE^{+}15, XB09]. sardines [KWB^{+}16, MEK^{+}09, SPG^{+}16]. Sardinian [POA^{+}17]. Sardinops [Cur04, CCP07, GSBB07, HZW^{+}98, HMS^{+}23, IYN^{+}09, ISN^{+}11, KKCL06, KYSM11, LS01, Lyn03, MYHvdL15, MGHS14, NY08, NYI11, NYI^{+}13, Nis19, NY03, OTH09, OIA^{+}12, RCB08, SWS^{+}19, SSSB03, SK03, TF08, VCB^{+}98, WMD^{+}06, WZK^{+}98, YWI^{+}05]. Sargasso [ARM16, AM18, BCR08]. Sargassum [KM94]. satellite [AMD^{+}16, ABG19, APR^{+}08, BDBP93, CÅP^{+}13, Col99, DPM^{+}11, GJR18, HLG^{+}11, HKLG07, KYY00, LVC^{+}05, LPG^{+}06, MPM19, PECG08, PH11, ROH16, RHG^{+}13, ZSS08]. satellite-tagged [ABG19]. saturation [Neu02]. saury [CLW^{+}19, FKSA21, INM^{+}18, IST^{+}04, IKK^{+}04, III^{+}06, KHN^{+}22, KNO^{+}04, KY17, MVK^{+}20, OWK^{+}03, OWK04, OTO^{+}09, SK04, TKO^{+}14, TNK^{+}16, TAS04, YW94, YK96, YW07, YOIW21]. saxatilis [NASTF10, NH06]. Scale [HHF09, PO03, BJCS12, BHH98, CHHS05, Cur04, FH94, HBLC22, HL07, HP02, KOWM16, KMB00, KVR^{+}18, Mar01, MM03, MTSH15, NH03, PWML12, PTS^{+}24, PLG^{+}10, QCR22, RHG^{+}13, SKNLD10, SGHW05, SHB^{+}11, TJW^{+}03, VPRG13, ZHT14]. scales [FCC^{+}19, LH96, RKD^{+}20, SVEW^{+}13, VYGT^{+}20]. Scaling [Par95]. scallop [BBR^{+}05, LCCS15, TCS^{+}09, ZJH^{+}22]. scallops [MMF95, TKM^{+}22]. scattering [AI92, HJ10, MTH^{+}04]. scavenger [SFL16]. scenario [LMBL03]. scenarios [POA^{+}17, RR18]. Schn [SHB^{+}11]. school [VN97]. Schooling [HALO00, Wil04]. schools [AI92, RPG^{+}22, Zam01]. Sciaenidae [ASCM12]. Science [Nak98, Woo93, FH94, HHK^{+}17, Par95, Sha95]. Scomber [AGSSL^{+}22, BC04, BRC04, BUE02, GiIW^{+}20, Jan16, KOS^{+}19, MHRC18, NK08, PVHT01, PGL^{+}15, RBPCR^{+}22, VGPL^{+}11, YWI^{+}05]. Scomberomorus [WMKR09, ZYT^{+}22]. scombrus [BC04, BRC04, BUE02, Jan16, MHRC18, PGL^{+}15, RBPCR^{+}22, VGPL^{+}11]. scope [ZJH^{+}22]. SCOR [MM03]. Scotian [MATL98, RAT^{+}02, RPC^{+}19, RPT^{+}00, SPM02]. Scottish [GMH^{+}99]. Scyphomedusae [CH92]. Sea [FYK^{+}21, HSS19, JCA^{+}16, KKNY92, LSW^{+}03, MFRR96, MKH^{+}13, OUKH04, PLT09, Ric96, SMF96, SPV96, SKNT14, SFK^{+}20, ZKT07, AG99, CL05, CAGPC21, Col00, EHW08, FRS^{+}05, FCJ^{+}15, HMM01, HMTG^{+}05, HHB^{+}15, III^{+}06, LH96, LRBJ21, MMRS16, MWB^{+}00, NBMS06, NY03, PLSO98, PBH^{+}04, RDF^{+}11, RDE^{+}07, SKHN11, SGL04, SMF^{+}05, TWK13, TCS^{+}09, TMM^{+}07, VHLM15, VPRG13, WGW07, YOYK20, ZJH^{+}22, ZZ93, VMG01, AMD^{+}16, AB02, AHKP16, AYMK01, ARM16, AM18, BCBDA10, BHC^{+}01, BH18, BRO18, BCGB14, BSF^{+}20, BSS94, BO05, BCR08, BGM^{+}18, BMO^{+}99, BCA^{+}18, BDAMD14, CLM^{+}21, CMB^{+}15, CHHS05, CCK^{+}22, CRW20, CEM^{+}11, CSS^{+}21, DPK^{+}08, DABM^{+}06, ESTJ03, ESA09, ETB05, FPBDC11, FGS95, FODCN00, FMYN06, FHD98, Fun11, GTB10, GMH^{+}99, GHBM99, GFG98, GGQF22, GPL^{+}11, GIT^{+}13, GQPGA04, GöEIOS16, GFO14, HTE^{+}03]. Sea [HGS^{+}21, HJR^{+}03, HG98, HBR^{+}99, HEG08, HLMS03, HH99, HGH93, HMS16, HCFP20, IK97, ISS02, JMP^{+}14, KKH^{+}20, KMD^{+}09, KSAF13, KJZ97, KEWDA18, KKNY04, KMK^{+}18, KWO^{+}18, LAFF15, LDDC06, LVF12, LYT^{+}20, LJBR20, LTL^{+}22, LVM^{+}18, LLSF01, LVPK11, MBH^{+}99, MLVO05, MBJ^{+}07, MSS12, MW92, MMF95, MTLL^{+}16, MKF^{+}03, Mor11, MFB^{+}09, MHvD^{+}24, NKS00, NH01, NDC05, Neu02, NHNA07, ODMRM98, OTH09, OHM^{+}10, OÅL00, OH23, PA14, PSJF93, Por22, PQH16, Ree95, REB^{+}03, RJHC99, RBPCR^{+}22, RKD^{+}20, RAK^{+}17, SGW^{+}21, SKM06, SKT21, SHG12, SS94, SAG^{+}09, Shi98, SPLY23, SCDA10, SADA^{+}23, Spe08, SBK^{+}01, SHB^{+}11, SWZ^{+}01, ST97, SP13, TID^{+}96, TSK^{+}22, TJW^{+}03, TD02, TKM^{+}22, TDT03, TLS98, TTC^{+}12, TTH15, UMK20, VYGT^{+}20, VZP98, VHJ99, VDHF08, WLWZ98, Wat17, WPN12, WQI00, WQ00, WSC05, WEW98]. Sea [YCH^{+}15, Zai92, ZYY^{+}21, ZYT^{+}22, ZHL^{+}03, ZVKŠ13]. seabass [DWHdP21, FKUY16, IUY10]. Seabird [JCH05, LAG^{+}11, APM^{+}12, LJH^{+}05, PLSO98, SRCV09]. Seabird-trawling [LAG^{+}11]. seabirds [BRR05, JCH04]. seabob [MHS^{+}21]. seabream [GEGHPCC17, SFGE21]. seafloor [FMG^{+}22]. seagrass [FKH^{+}17]. Seal [Zam01, HMS16, YKB08]. seals [TB92]. seamount [DP01, DP01]. seamounts [MBB^{+}03, SPM^{+}19, LRS^{+}23]. Seao [NFN00]. Searching [QC99]. seas [Ano99, POA^{+}17, PML06, SDRL96, SZX^{+}08, KEJK00, NFO^{+}23, SYT^{+}09]. season [FKSA21, KSM^{+}20, KKNY04, MPM19, MRBBHL14, TKO^{+}14]. Seasonal [ASM^{+}15, BMH^{+}21, GV01, HSEH16, IHHH99, KR10, KB08, LP10, MHG^{+}11, NIIS04, PQH16, SHG12, SPG^{+}16, TMMM20, VZP98, XWL^{+}23, dBdOJdO^{+}22, AAG11, And03, AGS^{+}04, FGGDSMF08, HKM^{+}19, III^{+}06, KJZ97, LCCQ^{+}22, LJM^{+}10, SYT^{+}09, SSW^{+}17, SK04, SS98, TD02, TSK^{+}95, TAN^{+}17b, VYGT^{+}20, VJ99, WJP^{+}01, YOYK20]. Seasonality [CCSS01, KL01, MESMM18]. seasonally [SGL22]. seasons [KBF^{+}07, MBE^{+}15, WBQL99]. seawater [KJZ97]. seaweed [KNK^{+}18, UTMS06]. Sebastes [BBY08, DH11, KPHG14, MLRS07, PDER10, RSF13, RBBG12, SRR99, SRR07, Sco95, ZLTM11]. second [SB04]. sediment [Lou10]. seedling [KSM^{+}20]. segmentation [Mar01]. segregation [APMVOGMR19]. sei [SMK^{+}13]. seine [EPG^{+}16, GAH^{+}19, MMRH^{+}16, TSG^{+}20, BBA^{+}21]. selected [PLT09]. selection [APGL03, BBP^{+}13, HTE^{+}03, KYSM11, LDAWM10, MTK^{+}07, SB06]. selective [VFS^{+}24]. selectivity [SMB^{+}03a]. self [BCJ^{+}13]. self-recruitment [BCJ^{+}13]. Seminar [SR93]. semisulcatus [BYM16]. Sendai [KUO^{+}17, OKU17]. Senegal [TFB^{+}17]. Senegalese [DBB^{+}18, MBE^{+}15, TAN^{+}17b]. sensed [RCB08, WRTP01, ZSS08]. sensing [BGM^{+}18, MPM19, MSNK10, Sim92a]. Sensitivity [TCS^{+}09, QC99]. sensory [FDT^{+}99]. separation [MHvD^{+}24]. Sergia [TKMS11]. series [CDG^{+}19, DLCQ22, FPBDC11, HCC^{+}09, KO95, MMBC07]. Seriola [KSMY00, TNC^{+}22, UTMS06]. services [aTCK05]. setiferus [WKB^{+}05]. Seto [FYK^{+}21, KKNY92, YOYK20, ZKT07]. Setting [FPFL13, JPHA^{+}16]. Settlement [BMOT17, CB93, Cap08, ERE^{+}10, FMYN06, HGG^{+}17, LDH14, LJM^{+}10, OHF12, PWML12, YTY96]. settling [IK97]. seven [IIS^{+}07]. Seventh [Kas99]. sex [BMH^{+}21, CLH^{+}22, SPM02, SSP^{+}11]. sex-dependent [CLH^{+}22]. sex-specific [BMH^{+}21, SPM02, SSP^{+}11]. shad [BDE^{+}19, GHG^{+}19, LAFF15]. shallow [BKvdP^{+}22, BWK^{+}99, KvdPBW17, OKU17]. shallow- [KvdPBW17]. shallow-water [BKvdP^{+}22]. Shape [HHF09, AGSSL^{+}22]. shapes [MP18]. shark [BBH99, CJ04, CIS20, CSFC05, GPCGdlT^{+}22, LCCdS^{+}19, OFS^{+}16, RHG^{+}13, SR02, WSP^{+}07]. sharks [KTPM17, MCHSNEO13, Wil04]. sharp [ZHL^{+}03]. shearwater [VCKH05]. shearwaters [BHC^{+}01]. Shedding [LS21]. Shelf [AAI16, FMM^{+}20, MCS^{+}06, SHS^{+}23, Ano99, AGS^{+}04, BO05, BT99, BDAMD14, áCGNGC19, CP03, CGMM10, CMM06, DBGW04, DABM^{+}06, EHW08, ETB^{+}17, GMH^{+}99, GI13, GP94, GHM21, HB99, HTE^{+}03, HZTS12, HWSS07, HHK^{+}10, HCWF21, KN08, LC95, LPSS04, MEK^{+}09, MSM^{+}13, MPW^{+}99, MTZG23, MMB^{+}11, PML06, PWE98, QLB^{+}05, RCG^{+}15, RHP^{+}15, SME^{+}14, SFL16, SMF96, SOTM^{+}18, WFRS93, WBQL99, WJM15, WKN^{+}95, Dd95, FYKSP07, GHV95, MFMG20, MHRC18, MATL98, PHH13, RPC^{+}19, RPT^{+}00, RAT^{+}02, SPM02, ZJH^{+}22]. shelf-edge [SMF96]. Shelikof [VIS92]. shellfish [HPL13]. shelves [LPHM21, MPM^{+}13]. Shetland [Hea99b, HJ99, Jón99, LJR^{+}22, RJHC99, Mar01]. shift [GI13, IMO^{+}12, SB05, Ste98, YCS^{+}15]. shifting [SMS^{+}21]. Shifts [LCCQ^{+}22, AS08, BKvdP^{+}22, BRN^{+}95, áCGNGC19, CCHL23, FH94, GPS22, HGS^{+}21, HFF^{+}19, JCA^{+}16, KYA^{+}15, MHRC18, OTIK20, PSM00, SP13, WWSE00]. ship [BDBP93]. shirauo [AHAM03]. Shizugawa [KU95]. shore [OKU17]. shoreline [ZP21a]. short [BHC^{+}01, DHC^{+}07, HP02, HKLG07, KOKM15, LMB^{+}19, VCKH05]. short-finned [DHC^{+}07, KOKM15]. short-tailed [BHC^{+}01, VCKH05]. short-term [LMB^{+}19]. shortfin [MCHSNEO13, OFS^{+}16, RHP^{+}15, SHS^{+}23]. should [Tyl92]. Shrimp [FYKSP07, DST11, Han11, HTP14, HSS19, KFYP07, LPH^{+}19, LPHM21, MHS^{+}21, MCB^{+}16, OA06, POA^{+}17, PBF00, ROH16, SGN^{+}05, WKB^{+}05]. shrimps [GGQF22, YMB99, dBdOJdO^{+}22]. shrinking [FMG^{+}22]. Sicilian [LGM^{+}02, LVC^{+}05]. Sicily [BGP^{+}06, BBP^{+}13, CPM^{+}15]. side [MLP22, NZI95, YAM^{+}18, ZNI96]. signal [TD02]. signatures [LCC15]. Significance [LTL^{+}22]. Significant [ZP21a]. silky [LCCdS^{+}19]. Sillaginodes [RRF^{+}21]. silver [RPC^{+}19]. silverfish [BCA^{+}18, LPCA15]. simple [WPL^{+}93]. simulate [DPK^{+}08, RHRL12]. Simulated [KTS15, VAFG95, WB93]. Simulating [BK94b, BHJ^{+}04, BC04]. Simulation [BRC04, HNHP09, LRL^{+}06, PMT^{+}94, TMN^{+}15, AI04, BLH98, GGQF22, OHM^{+}10, PKHF98, TD02]. Simulations [ODMRM98, APGL03, APL07, DLTI95, HTL^{+}00, Sim96, WJP^{+}01, Yam04]. simultaneous [MWN^{+}23]. since [MBJ^{+}07]. single [SRR07]. sinicus [LSW^{+}03, MTLL^{+}16]. sink [KSP^{+}22]. Sinking [ST95]. site [BBP^{+}13, CLH^{+}22, KMM^{+}06, PHH13]. sites [BBB^{+}19]. Sitka [HTL^{+}00]. situ [FMG^{+}22, RAT^{+}02]. six [SF22]. Sixth [Kas98]. Size [CH92, HBG^{+}16, KBS^{+}16, VFS^{+}24, APMVOGMR19, AOVAG22, BMHW13, GR98, HMS^{+}23, HKM^{+}19, HKM^{+}21, HAS^{+}19, IMS^{+}04, KvdPBW17, KNS^{+}22, MCHSNEO13, Mor11, MPM^{+}13, Oda94, OR13, PP01, PA14, REG^{+}13, RMM02, SKHN11, Sai22, TSK^{+}92, VCKH05, Wil01, XDP^{+}20]. size-at-age [HAS^{+}19, XDP^{+}20]. size-based [MCHSNEO13]. Size-dependent [CH92, Wil01]. Size-selective [VFS^{+}24]. sized [LTL^{+}22]. sizes [KFYP07]. Skagerrak [JCA^{+}16, FCJ^{+}15]. Skagerrak/Kattegat [FCJ^{+}15]. skate [GHM21, SB06]. Skipjack [AG99, GS96, VOB^{+}19, And03, GCF^{+}21, LPS19, LAB^{+}98, LMBL03, MSST16, MSNK10, NPLS22, Rog94]. slope [HFC01, LP10, MIK07, SSP^{+}07]. Small [KVR^{+}18, FFF^{+}18, HGS^{+}21, HPG^{+}20, KOWM16, KJZ97, KMM^{+}06, LH96, MTL^{+}22, MSR20, PTS^{+}24, RSC96, TAN^{+}17b, XWL^{+}23]. small-scale [PTS^{+}24]. smolt [BPS^{+}14]. smooth [PBF00]. Snake [SW05]. snapper [BASS11, CLKP19, Fra93, KSP^{+}22]. snow [SP13]. so-called [GSNFL99]. sockeye [APL^{+}96, BWS^{+}01, CHF^{+}04, DLTI95, DHM^{+}15, HTL^{+}00, HQH^{+}06, Mal20, McK13, MCG^{+}14, MFG99, PW12, PW14, PMT^{+}94, RFM^{+}21, RZM^{+}03, Sim96, TR11, TIH^{+}92, TH11, WSF^{+}14]. Solar [SAT^{+}18]. sole [BMHW13, CRW20, FODCN00, HTLJ20, LDH14, Por22]. Solea [FODCN00]. solidissima [MPM^{+}13]. Somatic [CHF^{+}04, ERR^{+}21, HBC07]. some [GP94, PJO99, SMH^{+}92, ST95, WEW98]. Sound [BG01, BWKM15, CAB^{+}01, CCSS01, ECM^{+}01, GV01, NBF^{+}01, WJP^{+}01, WKB^{+}05, WCP^{+}01, BMPC16, VMG01]. Source [KSP^{+}22, KPW19, PSM00]. sources [JCA^{+}16, KKH^{+}20, MWR^{+}98]. South [SNL19, VCB^{+}98, And03, AOVAG22, BHC^{+}01, BGH09, CHHS05, CCK^{+}22, GS99, JPHA^{+}16, KN08, LPS19, LRBJ21, MDKS93, MBKP08, NH01, NK08, QM01, SAT^{+}18, SBK^{+}01, SWZ^{+}01, Tan99, Tan02, TCC^{+}98, CIS20, CG18, DBGW04, DBRSC16, DSHL18, FML^{+}14, JHC^{+}15, LTL^{+}22, LC95, LJM^{+}10, LRBJ21, MRL^{+}14, MHM^{+}20, MWR^{+}98, OCH99, PFB^{+}16, PS06, RRF^{+}21, SBD^{+}19, TBB^{+}03, WRTP01, XTC^{+}04, ZHX^{+}20]. South-East [SBD^{+}19]. south-eastern [AOVAG22, BHC^{+}01, JPHA^{+}16, KN08, MDKS93, NH01, NK08, SBK^{+}01, SWZ^{+}01]. south-west [Tan99, Tan02, LC95]. south-western [And03, LPS19, MBKP08, SAT^{+}18, TCC^{+}98]. southeast [CP92, EHW08, SHG^{+}22, FYC22, WS08]. southeastern [CCC^{+}23, CEM^{+}11, HRB^{+}18, KY17, MW92, MHS^{+}21, SADA^{+}23]. Southern [Dom23, PS16, RHP^{+}15, TCC^{+}98, ABI^{+}21, AG99, AM18, BRO18, BGH09, CM10, CSB94, CMM06, Erz05, FML^{+}14, FHK^{+}10, FHK^{+}12, FRZVHM^{+}11, GMH^{+}12, HGG^{+}17, HHTF10, HHK^{+}10, Kae23, KOKM15, KK00, KL01, LPCG23, LVF12, LRL^{+}06, LC95, LJM^{+}10, Lyn03, MDR^{+}16, MYHvdL15, MTSH15, Mul94, MFP^{+}03, NPY^{+}15, OCCF^{+}18, PHH^{+}98, PVMP03, PECG08, RCD^{+}99, SME^{+}14, SCKJ^{+}18, SQW^{+}99, Swa99, SB06, VHCN14, VGPL^{+}11, WTR04, WMD^{+}06, Wil04, HKWL17, KGW13, RHG^{+}13, SKNLD10, Sim92b, WTR04, XMH^{+}18, FRZVHM^{+}11]. southern-central [GMH^{+}12]. southwest [BML11, CSFC05, HHK^{+}10, HCC^{+}09, OHM^{+}10, ADPC21, DBS^{+}19, SDHB07]. southwestern [BBR^{+}05, HFF^{+}19, MMSL19, MHB^{+}14, SKT21, MSM^{+}13]. space [BRFRJRLC18, HP02]. Spain [LCCQ^{+}22]. Spanish [BCR20, ZYT^{+}22]. spanner [SBD^{+}19]. spanning [PKP^{+}00]. Sparidae [Fra93]. Spatial [APMVOGMR19, BCBDA10, BKvdP^{+}22, BH97, BRPC08, BBR^{+}05, BPC^{+}16, BLG^{+}16, BDAMD14, DABM^{+}06, GHV95, HFHW19, HMS16, KYU^{+}06, KSAF13, LLCJ16, LCC15, MWB^{+}00, MWP02, MHvD^{+}24, MKH^{+}13, PHWM96, ROB05, SBY^{+}15, SGL22, TNM^{+}02, YOK^{+}17, YLA13, ZWC^{+}21, BLH98, BPLC11, Cap08, CMMK^{+}15, Cur04, DSHL18, ESA09, FCC^{+}19, GSBB07, HMM01, HHTF10, HS05, ICB^{+}08, JMLG06, KPHG14, KM94, LAFF15, LS21, LDDC06, LH96, LSD^{+}21, Mar01, MMRH^{+}16, PLP^{+}11, RKD^{+}20, RMM02, SRCV09, SVEW^{+}13, SLZ^{+}23, Spe08, SPLS15, SSPY08, SK03, TSK^{+}92, TMMM20, WWSE00]. spatially [FGS95, GYS14, HVHC10, MLVO05, PDD03, SPM^{+}24, SSP^{+}11]. Spatio [LPCG23, MHB^{+}14, SA10, YMK^{+}15, BJV^{+}17, BBA^{+}21, CAB12, DWHdP21, DLD^{+}23, EPG^{+}16, FCL93, FRBB14, GCW17, KTPM17, MWN^{+}23, PQH16, áRÁSG^{+}16]. Spatio-temporal [LPCG23, MHB^{+}14, SA10, YMK^{+}15, BJV^{+}17, BBA^{+}21, CAB12, DWHdP21, DLD^{+}23, EPG^{+}16, FCL93, FRBB14, GCW17, KTPM17, MWN^{+}23, PQH16, áRÁSG^{+}16]. Spatiotemporal [HPG^{+}20, IWK^{+}21, KMD^{+}09, NFKY21, SCTB19, DTO^{+}23, WKR^{+}18]. spawn [BG01]. spawned [CRC11, DCLC15, DBRSC16, FM93]. spawners [LC95]. Spawning [BBP^{+}13, COW^{+}99, KYS15, MFB^{+}09, RQN^{+}99, SWS^{+}19, SGS^{+}06, WJM15, ZYY^{+}21, ZYT^{+}22, ABI^{+}21, APL07, APL^{+}08, AM18, BCBDA10, BDE^{+}19, BSG^{+}13, BPP07, BDVS^{+}19, BvDSDC18, BBB^{+}19, CLKP19, CAB12, DWHdP21, DLD^{+}23, DBGW04, DBRSC16, FODCN00, FKSA21, GPA^{+}21, GGF17, GSNFL99, GöEIOS16, HONH04, IK97, III^{+}06, JGS93, KHN^{+}22, KL01, KMM^{+}06, KR10, KYSM11, LSK^{+}18, LVPK11, Lyn03, MSR20, MBE^{+}15, MDVB^{+}20, MM94b, MYHvdL15, MP18, MHB^{+}14, MMB93, NK08, NFKY21, OE17, OR12, OR13, OS95, OHS06, OH23, PSN^{+}99, PVMP03, PBL07, QBMW99, RCB08, RCPS09, RRF^{+}21, RF07, SES^{+}20, SAT^{+}18, SHK^{+}19, SQW^{+}99, SNV^{+}12, TKO^{+}14, TTI^{+}20, TH11, TLS98, TNC^{+}22, TTC^{+}12, VOB^{+}19, VDHF08, WZK97, WZK^{+}98, YIT^{+}22, ZVKŠ13, óóSV18]. speakers [Bow11]. spearfish [ABG19]. Special [Ano03a, CHPA98, KEJK00, SKM06]. Species [ARM16, MWN^{+}23, PFAM96, AOVAG22, BHS^{+}15, CIS20, CCHL23, DH11, DBB^{+}18, FYC22, HHTF10, HRS^{+}21, Hor00, HCFP20, IIS^{+}07, IMO^{+}12, JJBCW09, KT93, KMD^{+}09, KPW19, LPCA15, LVF12, LJBR20, LTL^{+}22, LéEPW^{+}12, LAG^{+}11, MDKS93, NSGL^{+}22, NTIO18, PLSO98, PJO99, PSC05, PLT09, PL03, SGL22, SLM13, SSM^{+}10, TTY^{+}23, TSG^{+}20, VPRG13, WKR^{+}18, YMK^{+}15]. Species-specific [MWN^{+}23, LAG^{+}11]. specific [BMH^{+}21, FYA^{+}21, LAG^{+}11, MCHSNEO13, MWN^{+}23, SPM02, SSP^{+}11, VMT^{+}23]. spectra [HMS^{+}23]. speed [TIH^{+}92]. Sperm [WFRS93]. spiny [EF10, SPM02, YOK^{+}17]. Spisula [MPM^{+}13]. splitting [CTWS08]. spots [MESMM18]. spp [EBO04, KNE^{+}04, MBH^{+}99, MESMM18, MLM^{+}98, MHS^{+}21, MLRS07, PDER10, RSF13, SRR99, SRR07, WP93]. spp. [BASS11]. sprat [ADPC21, BK94a, BK94b, BHV^{+}06, DPK^{+}08, HVHC10, LDDC06, MHvD^{+}24, SHB^{+}11, VDHF08]. sprattus [SHB^{+}11, ADPC21, BK94a, BK94b, BHV^{+}06, DPK^{+}08, HVHC10, LDDC06, MHvD^{+}24, SHB^{+}11, VDHF08]. spread [HDJ15]. spring [BSG^{+}13, CRC11, CP92, DTC06, ETB05, FM93, FYKSP07, GMH^{+}99, GPA^{+}21, GMH^{+}12, GDM^{+}17, HMM01, HBR^{+}99, IMS^{+}04, KOS^{+}19, KSYT97, KWO^{+}18, KNO^{+}04, LOGLD^{+}15, MBH^{+}99, Mul94, Mul97, MIY^{+}09, NKM01, NYI11, NII^{+}14, NTM^{+}15, PSJF93, SW05, SNV^{+}12, REM02]. spring-and [FM93]. spring-spawned [CRC11]. spring-spawning [BSG^{+}13, GPA^{+}21]. spring/early [MBH^{+}99]. spring/summer [SW05]. Springtime [BT99, HFC01]. Sproat [TR11]. Squalus [SPM02, YOK^{+}17]. squid [ASM^{+}15, AGS^{+}04, CG18, DHC^{+}07, DBRSC16, IMS^{+}04, ISI^{+}18, ITH23, KYU^{+}06, LCC15, MRL^{+}14, NII^{+}14, NTM^{+}15, OKT^{+}23, OHM^{+}10, PS16, SHS^{+}23, WTR04, YAM^{+}18, YWM^{+}00]. Sr [FKUY16, MFS^{+}17, YOY00]. SST [AI04, KYY00, OBA01]. St [BDVS^{+}19, CM10, D’A93, PGL^{+}15, RD96, RCD^{+}99, Swa99, SB06, VHCN14, YLA13, éSMB20]. Stability [SL95, Gar97, MAH12, MP94, PFSL09, SPLY23]. stable [DDS^{+}17, IMO^{+}12, MCHSNEO13, OM10, OKT^{+}23]. Stage [VMT^{+}23, BM99a, BSF01b, IHHH99, KSY^{+}23, KR10, LCCdS^{+}19, LS01, MCS^{+}06, SGW^{+}21, TW05]. Stage-specific [VMT^{+}23]. Staged [OR13]. stages [APGL03, ADPC21, GIT^{+}13, HG98, HBO^{+}01, LGM^{+}02, LLB^{+}20, MFP^{+}03, NH06, OA06, RS15, REG^{+}13, RWDA^{+}21, SS19, SB94, SCDA10, SK03, WPL^{+}93, XWL^{+}23]. standard [LVF12]. standing [KMK^{+}18]. State [FC04, DHM^{+}15, LRBJ21]. States [HFC01, SAH^{+}18, FPFL13, Col00, HA07, KD98, MPW^{+}99]. static [MJH14]. station [SRR07]. Statistical [KM93, BM99a]. statolith [LCC15, YAM^{+}18]. status [FKSA21, LJBR20]. steelhead [AMDM12, WWSE00]. Steller [CL05, FRS^{+}05, SMF^{+}05, TMM^{+}07]. stenolepis [HAS^{+}19, SME^{+}14, SGW^{+}21]. stepping [KPW19]. stepping-stone [KPW19]. still [Tyl92]. stimulating [WHT92]. Stock [AAKMG06, JGS93, ABI^{+}21, BML^{+}14, Bri94, EF10, Fun07, Fun11, HMM01, HBPC15, HDJ15, KMK^{+}18, KYSM11, LBC23, MRD^{+}19, NFKY21, Nis92, NII^{+}14, NTM^{+}15, Nis19, OTIK20, OHF12, OTH09, OR13, OS95, OHS06, OH23, Par96, SWAAB20, SP93, SC05, SB07, WPN12, WJM15, YWM^{+}00, ZHL^{+}03]. stock-dependent [SB07]. stock-recruitment [Fun07]. Stock-related [JGS93]. stocks [BSF^{+}20, BDVS^{+}19, CEM^{+}11, DB93, DB03, Gar97, HBO^{+}01, HPL13, JCA^{+}16, MCM^{+}17, MHM^{+}20, Ree95, RSZ^{+}03, SWS^{+}19, SC97, SRR05, WTR04]. stomach [TID^{+}96]. stone [KPW19, YTY96, YOY00]. storage [NHNA07, RDE^{+}07, WMD^{+}00]. storm [ERE^{+}10]. storms [MMF95]. straight [FPFL13]. Strait [VIS92, BGP^{+}06, BBP^{+}13, BRN^{+}95, CPM^{+}15, GDM^{+}17, GEGHPCC17, HLWL12, LLCJ16, NSGL^{+}22, PMT^{+}94, SFGE21, SMA14]. Straits [JTYB18, KBB^{+}20, RCPS09]. Strangomera [GMH^{+}12]. Strategic [LJR^{+}22]. strategies [áRÁSG^{+}16, RR18, SSR13]. strategy [ETB05, Mat06, YTIS95, ZYY^{+}21]. stratified [OA06]. streaked [OE17]. Stream [XMH^{+}18, AGK^{+}08, GS99, WFRS93]. strength [BLD^{+}03, Fra93, GPS22, MTH^{+}04, NDC05, YCH^{+}15]. strengthen [OHS06]. strengths [MSS12]. stress [RPG^{+}22, HLH^{+}17, LJM^{+}10, PSM00]. striata [EHW08]. strip [MAHG94]. Striped [SDHB07, APMRH17, APMVOGMR19, GSNFL99, NASTF10, NH06]. Strongylocentrotus [MWB^{+}00, TWK13]. Structure [PFSL09, Aut08, BKvdP^{+}22, BH97, BBB^{+}16, CPM^{+}15, FGGDSMF08, GR98, HT99, HKM^{+}19, KOWM16, KN08, KYSM11, MBKP08, MSVY^{+}13, Nis92, NdLOO23, Oda94, OTO^{+}09, OHS06, OH23, SPS^{+}20, Shi98, SHM05, SSM^{+}10, SPLS15, UIU^{+}99, Zai92]. structured [CH95, RAT^{+}02, SSW^{+}17]. structures [WJP^{+}01, ZHX^{+}20]. structuring [AB02, CAB12]. Studies [PFB^{+}16, Bri94, DPL02, PH11, PST03, PKHF98]. Study [áCGNGC19, BML^{+}14, BSG^{+}13, BLD^{+}03, BML11, CIS20, DWHdP21, FM93, FMYN06, GEGHPCC17, HB99, HQW^{+}99, HLMS03, HBN^{+}21, HSS19, KU95, LOS^{+}14, LVM^{+}18, PDD03, PDER10, SNV^{+}12, TKMS11, TSG^{+}20, VZP98, VGPL^{+}11, WSP^{+}07, ZHT14, ZSY^{+}21, VMG01]. sub [HZTS12, HPL13, LéEPW^{+}12, CAGPC21]. sub-Arctic [HPL13]. Sub-basin [CAGPC21]. sub-surface [HZTS12]. sub-tropical [LéEPW^{+}12]. subarctic [BW92, KSYT97, KTS15, MIY^{+}09, RSC96, SBT20, STI^{+}09, ST97, TID^{+}96, TSK^{+}92, TSK^{+}95, TSK04, YCS^{+}15]. Subject [Ano01b, Ano03c, Ano04b, Ano05b]. submarine [CCK^{+}22]. substantial [BMO^{+}99]. subsurface [Ree95, ZWL21]. subtidal [BAL^{+}99]. Subtropical [LLB^{+}20, HKT^{+}03, NTM^{+}15, SKHI04, ARM16]. subyearling [DDB17, LMB^{+}19]. success [CRC11, GI13, HMS^{+}23, JPMH20, KMB00, MSM^{+}13, MMSL19, MHM^{+}20, MFRR96, OIA^{+}12, PGL^{+}15, RAT^{+}02, RTK01]. successful [FMYN06, MMI^{+}22]. suckleyi [YOK^{+}17]. sufficient [DLTI95, Sim96]. suggest [HGG^{+}17]. suggests [LRBJ21]. suitability [AB02, CHPT20, CLW^{+}19, CAB12, GPL^{+}11, KOKM15, MDVB^{+}20, MHRC18]. Suitable [HCFP20, MFMG20, SLL19]. summary [BFF15]. summer [BHC^{+}01, BDAMD14, DABM^{+}06, ETB05, FHK^{+}10, HMM01, JGS93, KNO^{+}04, MBH^{+}99, Mor11, NASTF10, Oda94, RD96, SW05, SDHB07, WSC05, óóSV18]. superba [SRCV09, TBB^{+}03]. supply [BHJ^{+}04, JCA^{+}16, KNK^{+}18, MLP22, RHRL12]. supply-side [MLP22]. support [DMF^{+}17, HSEH16, JPMH20, KKH^{+}20, SMS^{+}23]. supporting [FvPH^{+}16]. surf [XMW^{+}23]. surface [AG99, Col00, CCSS01, FMM^{+}20, HZTS12, III^{+}06, NIIS04, NBMS06, NY03, RD96, SKHN11, SGL04, WK03, WGW07]. surface-layer [CCSS01]. surfclam [MPM^{+}13]. surrounding [LPSS04, SFA14]. Suruga [TKMS11]. survey [BPZR19, BH97, KvdPBW17, RPE98, SNV^{+}12]. surveys [AJ15, GHM21, OTIK20, RMM02, SYT^{+}09]. Survival [ZNI96, AHKP16, APLG07, APL07, BNM^{+}00, BSF01b, DPK^{+}08, ETB^{+}17, FHD98, FYK^{+}21, HFHW19, HTT^{+}16, HLMS03, IUY10, Jan16, KNS^{+}22, KKS92, KHB02, LS21, LS01, LML^{+}03, Mat06, MCG^{+}14, MAH12, MWP02, NYI^{+}13, PDD03, PJD14, ROH16, RAK^{+}17, RWP11, Sai22, SW05, SVEW^{+}13, TNK^{+}16, TTY^{+}23, Tan17a, VFS^{+}24, WS08, ZKT07]. survivorship [MFRR96]. sustainability [TDT03]. sustainable [aTCK05, PFB^{+}16, ZHL^{+}03]. Sv [MTH^{+}04]. swimming [KSY^{+}23, YTIS95]. swordfish [BBH99, DPM^{+}11, HBR^{+}15, SKNLD10, SAH^{+}18, SWAAB20, SRM^{+}18]. swordtip [ITH23, YAM^{+}18]. sympatric [DMF^{+}17]. Synchronicity [WTR04]. synchronization [CWCM14]. synchronous [FMV03]. synchrony [SFL16, SEM^{+}14]. Synthesis [JHK^{+}15, HL07, NKS00, NBF^{+}01, Ols01, SMS^{+}19, Tan02, Tan17a]. System [CMB^{+}15, AW92, BK94b, Col99, DL94, Gla11, HLWL12, IST^{+}23, JCH04, JMLG06, KB08, LCCQ^{+}22, MRBBHL14, OCCF^{+}18, RCG^{+}15, áRÁSG^{+}16, RMH^{+}19, Sim92b, Sko05, TKO^{+}14, TYO21, WMD^{+}06, WKB^{+}05, BRFRJRLC18, CCP07, HKA^{+}06, JCCB15, LLB^{+}20, MLRS07, MSVY^{+}13, SC06, SCKJ^{+}18]. systems [BBR^{+}05, BLD^{+}03, Sch23, Sim92a].

T. [GCF^{+}21, MSST16]. tactics [BDE^{+}19]. tag [HLG^{+}11, LPG^{+}06, MKK13, MFH05, MMMS14, WSP^{+}07]. tag-recovery [MMMS14]. tagged [ABG19, SF22]. tagging [AMD^{+}16, MBB^{+}03, PH11, SMB03b, WSP^{+}07]. tags [AMD^{+}16, APR^{+}08, CÅP^{+}13, DPM^{+}11, GJR18, HKLG07, KSMY00, NHNA07, PECG08, RDE^{+}07, RHG^{+}13, SF22, WMD^{+}00]. tailed [BHC^{+}01, VCKH05]. Taiwan [HCC^{+}09, HLWL12, TNC^{+}22, TCC^{+}98, TTC^{+}12]. Taiwanese [MTSH15]. taken [CP92]. Taking [LBC23, Par96]. Tango [SFK^{+}20]. Tanner [KBS^{+}16, RTK01]. tanneri [KBS^{+}16]. target [TSG^{+}20]. targeting [DMH16]. Tasman [MMB^{+}11, MGHS14]. Tasmanian [BLG^{+}16]. taxa [CAR^{+}10, MWN^{+}23]. taxonomic [HKT^{+}03, KMK^{+}18]. technologies [Sch23]. telemetry [GCF^{+}21]. Teleostei [MDKS93]. temperate [CAR^{+}10, FKUY16, FHK^{+}12, IUY10, PSC05, SPM^{+}24, SFK^{+}20, THH12].

Temperature [CJ04, DHMT96, DMH16, Fun07, MRD^{+}19, PVHT01, AHKP16, AG99, And03, AI05, BMOT17, BFF15, BRFRJRLC18, BAL^{+}99, BSF01a, Bri94, BCL04, CKA^{+}17, CSB94, Col00, D’A93, DTC06, Fra93, HCS^{+}09, IYN^{+}09, KJZ97, LOS^{+}14, LDH14, LA05, MSS12, MBY^{+}17, MWGK92, NBMS06, NY03, OE17, OR12, OR13, OS95, OÅL00, PMG^{+}23, PA14, QCR22, RDE^{+}07, RKD^{+}20, RTK01, SKHN11, SGL04, SPG^{+}16, SPLY23, SKKS05, TW05, TD02, TY04, THH12, UMK20, WGW07, WJT97, YOYK20, YCH^{+}15, YMB99, YKI98]. Temperature-based [DMH16]. Temperature-dependent [Fun07, QCR22]. temperatures [FHD98, FMG^{+}22, HMM01, III^{+}06, Jes22, KOS^{+}19, SCTB19]. Temporal [CMMK^{+}15, FYK^{+}21, KL01, LDDC06, WWSE00, BJV^{+}17, BCBDA10, BBA^{+}21, BDAMD14, CAB12, DWHdP21, DLD^{+}23, DABM^{+}06, EPG^{+}16, FCL93, FRBB14, GHV95, GCW17, HS05, JMLG06, KTPM17, KVR^{+}18, LAFF15, LPCG23, LLB^{+}20, MWN^{+}23, MHB^{+}14, PQH16, ROB05, áRÁSG^{+}16, SA10, SVEW^{+}13, Spe08, SBK^{+}01, TTH15, YMK^{+}15]. temporally [HVHC10]. Tenualosa [GHG^{+}19]. tenuirostris [VCKH05]. term [AH97, AS08, Bea03, BW92, BB07, Buc92, DLCQ22, IFF^{+}18, LYT^{+}20, LMB^{+}19, MLP22, OTH09, OH23, RF04, RPE98, RHRL12, RS92, SGN^{+}05, SR02, VYGT^{+}20, YW07, éSMB20]. terns [SAG^{+}09]. terrain [BPZR19]. territorial [DLD^{+}23]. test [IUY10, SB05]. Testing [CÅP^{+}13, MRL^{+}14, PJD14]. Tests [SPLS15]. Tetrapturus [ABG19, GSNFL99, HKLG07, SDHB07]. Texas [MTZG23]. Their [Sim92a, Buc92, CRC11, CRVL^{+}17, FvPH^{+}16, HDH^{+}05, HFF^{+}19, HBG^{+}16, JMLG06, KB08, LH96, MWN^{+}23, Rog94, SS94, SMH^{+}92, SHB^{+}11, SK04, VYGT^{+}20]. them [Tyl92]. Theragra [AYMK01, BCBDA10, BBMY93, Fun07, Fun11, FYK^{+}13, HYW04, HWSS07, HONH04, IST^{+}04, LDAWM10, MTH^{+}04, NKS00, NHS^{+}07, SB94, WSC05, Yam04]. there [CHM^{+}94]. Thermal [NASTF10, RFD^{+}04, ABG19, FMM^{+}20, HKM^{+}19, MTL^{+}22, Mor11, SA10, VOB^{+}19, WMD^{+}00]. thermally [OA06]. thermohaline [VSÅO07, WJP^{+}01]. thermoregulatory [HKM^{+}19, HKM^{+}21]. Third [Woo95, TTI^{+}20]. thorny [GHM21, SB06]. Thread [CMMK^{+}15]. three [APL01, CCL^{+}05, HQW^{+}99, HNHP09, KPW19, KK00, KU95, NIIS04, PSC05, SJB^{+}22, TTY^{+}23]. three-dimensional [APL01, HQW^{+}99, HNHP09, KU95]. Thunnus [AAKMG06, APR^{+}08, BCR20, BGH09, BHM02, BML11, CLT05, CSK11, DWH11, DGB^{+}16, DSPH07, Dom09, FRBB14, FHK^{+}10, FHK^{+}12, FFF^{+}18, GCQ^{+}13, GCF^{+}21, HKM^{+}19, HKM^{+}21, HFF^{+}19, HHTF10, HHK^{+}10, HK06, IFF^{+}18, KNS97, KKNY04, KBF^{+}07, LLCV18, Mat06, MKK13, MSST16, MLR10, MBB^{+}03, NPS^{+}23, Nis92, NdLOO23, PECG08, Pol96, RF04, RSZ^{+}03, RF07, RMH^{+}19, SA10, SFA14, SF22, SGL04, SL09, SAT^{+}18, SMB03b, SZX^{+}08, TTI^{+}20, VHCN14, WMD^{+}06, ZSS08]. thynnus [KKNY04, DGB^{+}16, FRBB14, GCQ^{+}13, MLR10, Pol96, RF04, RSZ^{+}03, RF07, SGL04, SL09, VHCN14]. Thysanoteuthidae [OKT^{+}23].

Thysanoteuthis [OKT^{+}23, OHM^{+}10]. Tidal [BAL^{+}99, HJR^{+}03, LHF^{+}99, VCKH05, BWK^{+}99, FRP^{+}99, GV01, GRT^{+}07, HTP14, LSW^{+}03, UYF92, XMW^{+}23, Zam01]. tidal-influenced [HTP14]. tidally [CFL^{+}99, JPMH20, JR07, SPLS15]. tiger [BYM16]. Tight [TKM^{+}22]. tilefish [MPW^{+}99, NLN^{+}21]. Time [BRFRJRLC18, HCC^{+}09, Mal20, CDG^{+}19, FPBDC11, GS96, HP02, HKM^{+}19, KO95, MMBC07, NGGJ09, Sha95, SMS^{+}19, VYGT^{+}20, VDHF08, ZWL21, ZP21a]. Time-varying [Mal20, NGGJ09]. times [PSN^{+}99]. timescale [Wat17]. timing [FYKSP07, HQH^{+}06, KSYT97, PKHF98, Sai22, TD02, TH11, TDT03]. Timothy [Per23]. Tintinnid [KT93, ST95]. Tissue [MCHSNEO13]. Tissue-specific [MCHSNEO13]. Todarodes [KYU^{+}06]. together [DBS^{+}19]. Tohoku [KKK^{+}17, TTI^{+}20]. tolerance [Bri94]. tool [BM99b, CL05, MPM19]. toothfish [MMI^{+}22]. top [Gla11, HJ10, MLRS07, VCKH05]. top-down [Gla11]. tracers [HZTS12]. Trachurus [IST^{+}23, IWK^{+}21, ISS02, KYS15, NPY^{+}15, SKM06, SYT^{+}09, SKT21, TSK^{+}22]. track [CÅP^{+}13]. Tracking [AMD^{+}16, LML^{+}03, MHRC18, EvST^{+}17, IYN^{+}09, NYI^{+}13, YAM^{+}18, YW07]. trade [MDR^{+}16]. trade-off [MDR^{+}16]. Training [BZ21]. traits [HPG^{+}20, KO95]. trajectories [SPG^{+}16, TCS^{+}09]. Trans [GR98, WTR04, Pol96]. Trans-oceanic [GR98]. trans-Pacific [Pol96]. transboundary [DTO^{+}23, NSGL^{+}22]. transect [UIU^{+}99]. transfer [Bau95]. transient [HCC^{+}09]. transition [KKNY04, Ols01, TWKW01, TW05]. transitional [SKKW02, TMS^{+}08, WMK^{+}99]. transitions [PFSL09]. transmission [BFF15]. transplant [PKHF98]. Transport [BS94, FRP^{+}99, IYN^{+}09, LHF^{+}99, NYI^{+}13, STI^{+}09, YTY96, BBS99, BK94b, BC04, BSS94, BEF^{+}12, BWK^{+}99, CCM^{+}08, CM10, CFL^{+}99, DST11, DCLC15, DBRSC16, FDT^{+}99, GP94, HT18, HQW^{+}99, HFC01, IN00, ISS02, KSM^{+}20, KKS92, KBB^{+}20, KIS01, MRL^{+}14, MHM^{+}20, MMI^{+}22, PVMP03, PWE98, RPT^{+}00, RQN^{+}99, RKD^{+}20, SKM06, SKKS05, SJB^{+}22, SFK^{+}20, TMN^{+}15, TKMS11, VSÅO07, WHT92, WKB^{+}05, WBQL99]. transportation [IST^{+}23]. trawl [AAI16, AJ15, FCJ^{+}15, GHM21, SYT^{+}09]. trawlable [BPZR19]. trawlers [WKN^{+}95]. trawling [LAG^{+}11]. trees [MCB^{+}16]. trend [MBY^{+}17, OUKH04, SMF^{+}05]. Trends [BNM^{+}00, Erz05, FMM^{+}20, JCH04, BB07, BRR05, Mor11, MHB^{+}14, RAK^{+}17, SR02, YMK^{+}15, éSMB20]. triad [AB02]. triangle [HTP14, VOB^{+}19]. Trichiurus [SCF^{+}20]. triggered [TKW^{+}17]. trituberculatus [YTIS95]. Trivial [NFO^{+}23]. Trophic [KKH^{+}20, BRO18, BBA^{+}21, CHHS05, GQPGA04, IMO^{+}12, IKK^{+}04, JPMH20, KNE^{+}04, LCCQ^{+}22, MCHSNEO13, NFO^{+}23, UIU^{+}99, WP93, Zai92]. trophic-level [CHHS05]. Trophodynamic [Yam04, AMDM12]. Tropical [ERE^{+}10, HBLC22, HHH^{+}18, MMRH^{+}16, AOVAG22, BLH98, BYM16, GSNFL99, HKT^{+}03, LéEPW^{+}12, PL03, PG06, Rog94, SKHI04, SZX^{+}08, SLZ^{+}23, SRM^{+}18, dBdOJdO^{+}22]. trout [RDE^{+}07]. trutta [RDE^{+}07]. tshawytscha [BRPC08, HTT^{+}16, HMT07, MRRN05, PMFC10, SMB^{+}03a, SW05, SVEW^{+}13, WGFR06, WGW07, WGS^{+}08, XDP^{+}20]. tsunami [KKK^{+}17, KUO^{+}17, MTT^{+}17, OKU17, OK17, ONK17, TWK13, TKW^{+}17]. Tsushima [IST^{+}23, TTH15]. tube [VPRG13]. Tuna [HBR^{+}15, RF07, VOB^{+}19, AUOGMM19, AMD^{+}16, AG99, And03, APR^{+}08, BCR20, BGH09, BHM02, CH16, DWH11, DGB^{+}16, Dom23, FRBB14, FHK^{+}10, FHK^{+}12, FFF^{+}18, GAH^{+}19, HKWL17, HKM^{+}19, HKM^{+}21, HFF^{+}19, HHTF10, HHK^{+}10, HK06, HHH^{+}18, IFF^{+}18, KKNY04, KBF^{+}07, LLCV18, LPS19, LAB^{+}98, LCH03, LMBL03, Mat06, MKK13, MSST16, MMBC07, MSNK10, MLR10, MBB^{+}03, NPLS22, NPS^{+}23, Nis92, NdLOO23, PECG08, Pol96, RF04, Rog94, RSZ^{+}03, RWI^{+}16, RBB^{+}21, RMH^{+}19, SFA14, SF22, SGL04, SL09, SAT^{+}18, SMB03b, SZX^{+}08, SLZ^{+}23, TTI^{+}20, VHCN14, WMD^{+}06, WJW20, ZHT14, ZSY^{+}21, ZHX^{+}20, ZWC^{+}21]. tunas [BLH98, Bri94, GCF^{+}21, KFH00]. turbidity [NH06, SKNT14]. turbine [WJ93]. Turbinidae [SPM^{+}24]. turbulence [FUA^{+}98, PA14, RAT^{+}02]. Turbulent [MCS^{+}06]. turtle [EBFF17]. Turtles [PKP^{+}00, HHB^{+}15, PBH^{+}04]. TurtleWatch [HHB^{+}15]. twaite [LAFF15]. twentieth [REB^{+}03]. two [AMD^{+}16, AW92, AOVAG22, BHS^{+}15, CCK^{+}22, FYC22, HRB^{+}18, IMO^{+}12, KO95, LSD^{+}21, MDKS93, MSR20, MHM^{+}20, MTT^{+}17, WSC05, óóSV18]. two-layered [AW92]. type [Lou10]. typus [WSP^{+}07]. tyrannus [FDT^{+}99, HT18, QBMW99]. Tyrrhenian [CCK^{+}22].

U.S [MFMG20, MHRC18, ZJH^{+}22]. U.S. [EHW08, HBR^{+}15, KBS^{+}16, RS92, SMS^{+}21, SNL19]. U.S.A. [LCCS15]. Ubiquitous [SS94]. Undaria [KNK^{+}18]. Underestimation [MPM^{+}13, Jes22]. underlying [Gar97, KO95]. understand [FvPH^{+}16]. Understanding [BDTR23, HGS^{+}21, LVPK11, PLP^{+}11, DST11, LML^{+}03, MTL^{+}22]. underway [COSC97, ESTJ03, LVF12, PSC05, VCB^{+}98]. undulatus [HT18, HA07]. unfished [MRD^{+}19]. unit [Dom23, NLN^{+}21, VHCN14, Wat17]. United [Col00, HFC01, HA07, KD98, MPW^{+}99, SAH^{+}18]. units [GNP^{+}19, LPH^{+}19]. Unveiling [QCM^{+}16]. updated [CC03]. upriver [CHF^{+}04, HMT07]. upwelling [AOVAG22, BDE^{+}19, BLD^{+}03, CCM^{+}08, Col99, FIDC00, GMH^{+}12, Han11, HHH^{+}16, HB92, IHS97, ICB^{+}08, JCH04, KYS15, KB08, LCCQ^{+}22, LRL^{+}06, MBE^{+}15, MHG^{+}11, MAH12, MRBBHL14, OCCF^{+}18, PHH^{+}98, PS06, RCG^{+}15, REM02, áRÁSG^{+}16, Sko05, SBD^{+}19, TCL^{+}12, TFB^{+}17, WMD^{+}06, WGW07, XH95]. urchin [MWB^{+}00, TWK13]. Uroteuthis [ITH23, YAM^{+}18]. ursinus [HMS16, YKB08]. Uruguay [ASCM12]. USA [HAS^{+}19, BRPC08, BPLC11, BASS11, GP94, GS99, NASTF10, SGL22, WFRS93, WGS^{+}08, WKB^{+}05, Zam01]. Use [HKA^{+}06, MSS12, ROH16, Bez00, BDBP93, CGMM10, FCJ^{+}15, FFF^{+}18, HLG^{+}11, KMD^{+}09, LSK^{+}18, LPG^{+}06, MJH14, PFB^{+}16, QCR22, RHP^{+}15, Sim92a, YAM^{+}18]. used [JR07, VCKH05]. Using [BGM^{+}18, DPL02, HRS^{+}21, RRF^{+}21, SRR05, TTC^{+}12, AMD^{+}16, AYK03, BPZR19, BLH98, BBY08, BSF01b, BHS^{+}15, CC03, CÅP^{+}13, CH95, DWHdP21, DPM^{+}11, DDS^{+}17, ESTJ03, FKSA21, GR98, GHM21, GRT^{+}07, HBPC15, HHF09, HBN^{+}21, IMO^{+}12, JYH^{+}18, KWB^{+}16, LAB^{+}98, LVPK11, MHM^{+}20, MIK07, MCB^{+}16, NHNA07, OTIK20, OK17, OCCF^{+}18, PP01, PECG08, PLP^{+}11, QBMW99, RHG^{+}13, RMM02, RSZ^{+}03, SSW^{+}17, SLL19, TAS04, VCB^{+}98, YOY00, YOK^{+}17, ZSS08, HBLC22]. utilisation [LAG^{+}11]. utility [XMH^{+}18]. utilization [FHK^{+}10, HKLG07, LPG^{+}06, SFA14, SF22, SGL22, SDHB07, Tan99]. utilize [VPRG13]. utilizing [WKR^{+}18].

V [BM99a, IHHH99]. values [WP93]. Vancouver [TR11, LH96, PHWM96, PBF00, Tan99, Tan02]. Variability [DGB^{+}16, GHBM99, HT99, HVHC10, HXC^{+}17, KSYT97, Lyn03, VYGT^{+}20, AYMK01, AGSSL^{+}22, And03, AM18, BH18, BMPC16, BSG^{+}13, BCGB14, BML11, BW92, BDSM07, BDVS^{+}19, Buc92, CSFC05, DPK^{+}08, DDB17, DPL^{+}20, DLCQ22, DLD^{+}23, DPL02, ETB^{+}17, FCJ^{+}15, FHHW98, FGGDSMF08, FYK^{+}13, GCQ^{+}13, GMH^{+}12, GCW17, GDM^{+}17, GFO14, HP02, HFHW19, HNHP09, HK06, HMT07, INM^{+}18, IH03, IWK^{+}21, IYN^{+}09, ISN^{+}11, KMK^{+}18, KHB02, LVC^{+}05, LDDC06, LCH03, LAPL21, MTL^{+}22, MSM^{+}13, MLP22, MAHG94, MM94b, MP18, MKF^{+}03, MWB^{+}00, MWR^{+}98, MP94, NH03, NYI11, NHS^{+}07, OWK^{+}03, OH23, OUKH04, PL03, PJB05, PGL^{+}15, PCR^{+}18, PQH16, RCS98, RPC^{+}19, RPE98, ROB05, SRCV09, SMS^{+}23, SVEW^{+}13, SGN^{+}05, SC97, SBK^{+}01, SHM05, SEM^{+}14, SCTB19, TSK^{+}92, TNK^{+}16, Tan02, TR11, Tan17a, TMMM20, TSK^{+}95, TAN^{+}17b]. variability [TBB^{+}03, VMG01, YWM^{+}00, YOIW21, ZWC^{+}21]. variable [BBB^{+}16, MCG^{+}14, Ree95]. variables [BPZR19, Erz05, HVHC10, HCC^{+}09, JYH^{+}18, JCCB15, KvdPBW17, KEJK00, MMRH^{+}16, MWP02, NdLOO23, RS92, SME^{+}14, SPM02, SCF^{+}20, WQ00]. Variation [Han11, PM95, Por22, SSR13, óT10, BMH^{+}21, BLG^{+}16, CLW^{+}19, DLCQ22, DTO^{+}23, FCL93, FKSA21, FHK^{+}10, GEGHPCC17, HFF^{+}19, HQH^{+}06, HS05, KPHG14, KOS^{+}19, KTH^{+}15, KL01, KKCL06, LLCV18, LYT^{+}20, MM03, MVK^{+}20, MWR^{+}98, NKM01, NTM^{+}15, OE17, Pol96, RSF13, Sha13, SGN^{+}05, SB04, SK04, TCO^{+}05, Tan99, TTH15, WMD^{+}00, WL21, YW07]. Variations [KNO^{+}04, NHM94, VHJ99, WZK97, CHHS05, FYK^{+}21, IST^{+}23, JMLG06, KK00, KB08, LSK^{+}18, LP10, MBY^{+}17, MTLL^{+}16, PHWM96, SKT21, ST97, ST98, SS98, TJW^{+}03, UIU^{+}99, WEW98, YSW^{+}99]. varies [NPS^{+}23]. vary [SBY^{+}15]. Varying [PKHF98, Mal20, NGGJ09]. velocity [GS96, SAG^{+}09]. VENFISH [IST^{+}04]. versus [ADPC21, BBB^{+}19, MM94b, TNK^{+}16]. Vertical [HT18, HLG^{+}11, LTL^{+}22, LPG^{+}06, MKK13, MBB^{+}03, OA06, RCG^{+}15, TF08, AYK03, APR^{+}08, BM99a, BGH09, BRC^{+}03, BT99, CCM^{+}08, CCP07, DST11, EHW08, EBO04, ETB05, GJR18, GP94, HQW^{+}99, HRB^{+}18, HHF09, HKM^{+}19, HKLG07, HCS^{+}09, KBF^{+}07, LVF12, MTH^{+}04, Mor11, PML06, SRR99, SKKW02, SMK02, SHG12, SE19, SADA^{+}23, SSSB03, SHB^{+}11, VJ99, WMK^{+}99, WJT97]. vertically [BK94b, ODMRM98, RAT^{+}02]. VI [BM99a]. via [IYN^{+}09, NYI^{+}13]. Vicinity [LHF^{+}99, HDH^{+}05]. view [Gre13]. views [GP94, WJ93]. villosus [APL^{+}08, HWSS07, LDAWM10, OR12, OR13, WPN12]. Vinciguerria [LLB^{+}20]. volcanic [KTS15, PW12]. volcano [McK13, PW14]. Volume [Ano01a, Ano01b, Ano03d, Ano03b, Ano03c, Ano04a, Ano04b, Ano05a, Ano05b, Ano06, MTH^{+}04]. vs [RMM02]. vulgaris [FIDC00]. vulnerability [FvPH^{+}16, VOB^{+}19]. vulpinus [HRB^{+}18].

W [KEJK00, KEJK00]. Wadden [SAG^{+}09]. wakame [KNK^{+}18]. wake [JR07]. Walbaum [VYGT^{+}20]. Walleye [Spr92, AYMK01, BCBDA10, BBMY93, BBS99, Fun07, Fun11, FYK^{+}13, HYW04, HWSS07, HONH04, IST^{+}04, KNS^{+}22, KTH^{+}15, KEWDA18, LK21, LDAWM10, MTH^{+}04, NKS00, NHS^{+}07, OTIK20, RWDA^{+}21, SB94, SADA^{+}23, Yam04, YCH^{+}15]. wandering [XTC^{+}04]. Warm [GAH^{+}19, IST^{+}23, TTH15, AI92, KEWDA18, Por22, SPM^{+}24, SADA^{+}23, YCS^{+}15]. warm-core [AI92]. warm-temperate [SPM^{+}24]. warm-water [YCS^{+}15]. Warming [Kae23, AGK^{+}08, CCC^{+}23, CH16, FKF^{+}22, FMG^{+}22, LéEPW^{+}12, LMBL03, OUKH04, Sim92b, SSM^{+}10, VHLM15]. Washington [WGS^{+}08, Zam01, BRPC08, BPLC11, DDB17, KBS^{+}16, MAHG94]. waste [LAG^{+}11]. Water [MFRR96, APL^{+}08, BKvdP^{+}22, Coy05, ESA09, Fra93, GTB10, GNP^{+}19, GGQF22, GJR18, GP94, HQH^{+}06, ISS02, Jes22, JMP^{+}14, KvdPBW17, KKK^{+}17, KT93, KN08, KIS01, KM94, LLB^{+}20, SL95, MHG^{+}11, MATL98, MIK07, MWN^{+}23, MBKP08, OE17, OA06, OUKH04, PSN^{+}99, QLB^{+}05, RTK01, SPM^{+}19, SAG^{+}09, STI^{+}09, WTK^{+}16, YCS^{+}15]. water-masses [MATL98]. waters [BS94, áCGNGC19, DCLC15, DWH11, DBB^{+}18, DBGW04, FHK^{+}12, HTE^{+}03, HL98, HCC^{+}09, IIS^{+}07, IHS97, JPMH20, JHC^{+}15, KL01, LLCJ16, LTL^{+}22, LP10, LPSS04, MPW^{+}99, MIY^{+}09, NZI95, NASTF10, NBH99, OKT^{+}23, OWK^{+}03, QCM^{+}16, RFD^{+}04, RD96, SME^{+}14, SKHI04, SKM04, SFA14, SLZ^{+}23, SSM^{+}10, TCO^{+}05, TA06, TCC^{+}98, WZK97, WMK^{+}99, WS08, YMB99]. wave [LRBJ21]. wavelet [MMBC07]. waves [RSC96]. weak [MRBBHL14]. weaker [RAK^{+}17]. weather [BO05, NH01, RCS98]. webs [DMF^{+}17, PAS^{+}18, SPV96, SP15]. weight [KHN^{+}22]. West [BJV^{+}17, KSC^{+}10, SME^{+}14, SMS^{+}21, Ano99, Col00, GHV95, HB99, HT99, JHC^{+}15, KBS^{+}16, MDVB^{+}20, PS06, PWE98, SPM^{+}19, SR02, Tan99, Tan02, WMKR09, DDS^{+}17, SL95, TAN^{+}17b, LC95, MM94a, MMB93]. west-central [WMKR09]. Western [MCS^{+}06, ASM^{+}15, AGSSL^{+}22, And03, AGK^{+}08, BBMY93, BGM^{+}18, CB93, Cap08, CRC11, CWCM14, EvST^{+}17, FYA^{+}21, GPCGdlT^{+}22, GS99, GAH^{+}19, HMTG^{+}05, HKT^{+}03, HBO^{+}01, HKLG07, HLWL12, INM^{+}18, IYN^{+}09, III^{+}06, KTPM17, KSYT97, KT93, KYU^{+}06, LPS19, LYT^{+}20, LLSF01, LPSS04, MEK^{+}09, MIK07, MSNK10, MBKP08, MMB^{+}11, MGHS14, MTK^{+}07, MIY^{+}09, NTIO18, NFKY21, NNOU20, OIA^{+}12, Ols01, RSC96, Rog94, SMK^{+}13, SKKW02, SKHI04, SKM04, SES^{+}20, SAT^{+}18, SPV96, SOTM^{+}18, SRM^{+}18, TCO^{+}05, TSK^{+}92, TMS^{+}08, TMN^{+}15, TSK^{+}95, TCC^{+}98, TSK04, VCB^{+}98, WMK^{+}99, YW94, YOIW21, YCS^{+}15, ZSS08, BSG^{+}13, CB93, Cap08, CAGPC21, CBdSF^{+}23, FHK^{+}10, FHK^{+}12, Mat06, MATL98, NdLOO23, PQH16, RAT^{+}02]. Westward [MRL^{+}14]. whale [KOKM15, MIK07, SP15, WSP^{+}07]. whales [KEJK00, MTK^{+}07, MKH^{+}13, SMK^{+}13, WFRS93]. Where [GGQF22, HBLC22]. Which [TSG^{+}20]. white [HKLG07, MCHSNEO13, NH06, OE17, WKB^{+}05]. white-streaked [OE17]. whiteheadi [VCB^{+}98]. whitemouth [ASCM12]. whitespotted [LJBR20]. whiting [BC97, HEG08, LVPK11, MMRS16, MP18, RRF^{+}21]. wide [KOWM16]. wide-scale [KOWM16]. width [SPG^{+}16]. widths [KTH^{+}15, KNO^{+}04]. wild [KNS^{+}22]. William [BMPC16, BG01, BWKM15, CAB^{+}01, CCSS01, ECM^{+}01, GV01, NBF^{+}01, VMG01, WJP^{+}01, WCP^{+}01]. Wind [BSG^{+}13, BWK^{+}99, BLD^{+}03, LHF^{+}99, NTM^{+}15, ASK99, DTC06, GHG^{+}19, HBPC15, JCH04, KR10, LJM^{+}10, LPSS04, OS95, PSM00, REL07, RTK01, SBD^{+}19, TF08, WGW07, WL21, XH95]. Wind-driven [LHF^{+}99, ASK99, REL07, SBD^{+}19]. wind-forced [TF08]. Wind-generated [BWK^{+}99]. Wind-induced [BSG^{+}13, NTM^{+}15, XH95]. wind-regulated [KR10]. windfield [DB93]. window [Gar97]. windows [DBB^{+}18]. winds [GPS22, SHG^{+}22]. Winter [SFK^{+}20, WKN^{+}95, BWKM15, BAL^{+}99, CRVL^{+}17, DCLC15, GTB10, GP94, GS99, HQW^{+}99, IMS^{+}04, ISI^{+}18, KOS^{+}19, MTLL^{+}16, MRHL09, Mul94, Mul97, NH03, NY08, NYI11, NII^{+}14, NTM^{+}15, RCG^{+}15, SKM04, SBT20, TKO^{+}14, WBQL99, WB93, YOYK20, óóSV18]. winter-spring [Mul94]. winter-to-spring [NYI11]. wintering [HMS16]. within [CFL^{+}99, FCJ^{+}15, JCCB15, KCW^{+}15, LCC15, MAH12, RCB08, REL07, RKZHC19, SC06, SLZ^{+}23]. workshop [LBSS^{+}92, BB02, War92]. world [LBSS^{+}92]. wrasse [CLH^{+}22]. write [Bow11].

X [Gre99]. Xiphias [SKNLD10, SAH^{+}18]. Xiphopenaeus [MHS^{+}21].

Year [HMT07, WEW98, ASCM12, AHKP16, BMPC16, BHV^{+}06, CDG^{+}19, Fra93, GPS22, Jan16, KPHG14, KOS^{+}19, KMB00, LK21, NDC05, RTK01, VGPL^{+}11, YCH^{+}15]. year-class [ASCM12, RTK01]. Year-to-year [HMT07, WEW98, KOS^{+}19]. yearling [BRPC08, PMFC10]. years [BEiI^{+}23, LYT^{+}20, MYHvdL15, Por22, SADA^{+}23, SSM^{+}10, WSC05]. Yellow [SYT^{+}09, HGS^{+}21, KJZ97, XWL^{+}23, HGS^{+}21, HJR^{+}03, KJZ97, LYT^{+}20, LJBR20, LSW^{+}03, ZYY^{+}21, ZYT^{+}22, ZHL^{+}03]. yelloweye [BBY08]. yellowfin [BCR20, BMHW13, DWH11, Dom23, GCF^{+}21, MSST16, Nis92, NdLOO23, Por22, Rog94, RWI^{+}16, SFA14, SF22, SZX^{+}08]. yellowtail [SCS05, UTMS06, XMH^{+}18]. yellowtails [KSMY00]. yield [ZHL^{+}03]. yields [KJZ97]. yolk [BBMY93]. yolk-sac [BBMY93]. young [BHV^{+}06, KPHG14, SPG^{+}16]. young-of-the-year [BHV^{+}06, KPHG14].

Zealand [CMS16, Fra93]. zone [RPG^{+}22, KSC^{+}10, LML^{+}03, MIY^{+}09, QM01, XMW^{+}23, ARM16, Dom09]. zones [BEF^{+}12, Ols01]. zoning [HHTF10]. Zooplankton [Coy05, AYK03, AS08, BW92, BMO^{+}99, CCM^{+}08, CCSS01, CSFC05, CP92, CP03, ESTJ03, ETB^{+}17, GR98, GBAD^{+}17, HH99, IST^{+}04, MM03, MWN^{+}23, MTH^{+}04, REB^{+}03, áRÁSG^{+}16, RCD^{+}99, RWP11, SR02, SWZ^{+}01, ST97, WLWZ98]. Zygochlamys [BBR^{+}05].

References

Anderson:2011:MMS

[AAG11] R. Charles Anderson, M. Shiham Adam, and Joaquim I. Goes. From monsoons to mantas: seasonal distribution of Manta alfredi in the Maldives. Fisheries Oceanography, 20 (2):104–113, March 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Alemany:2016:DIB

[AAI16] Daniela Alemany, Eduardo M. Acha, and Oscar O. Iribarne. Distribution and intensity of bottom trawl fisheries in the Patagonian Shelf Large Marine Ecosystem and its relationship with marine fronts. Fisheries Oceanography, 25(2): 183–192, March 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arregui:2006:SER

[AAKMG06] Igor Arregui, Haritz Arrizabalaga, David S. Kirby, and Juan Manuel Martín-González. Stock–environment–recruitment models for North Atlantic albacore (Thunnus alalunga). Fisheries Oceanography, 15(5):402–412, September 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Agostini:2002:OTM

[AB02] Vers N. Agostini and Andrew Bakun. ‘Ocean triads’ in the Mediterranean Sea: physical mechanisms potentially structuring reproductive habitat suitability (with example application to European anchovy, Engraulis encrasicolus). Fisheries Oceanography, 11(3):129–142, May 2002. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arostegui:2019:MTN

[ABG19] Martin C. Arostegui, Camrin D. Braun, and Peter Gaube. Movement and thermal niche of the first satellite-tagged Mediterranean spearfish (Tetrapturus belone). Fisheries Oceanography, 28(3):327–333, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Alberto:2021:DCM

[ABI^{+}21] María Luz Torres Alberto, Nicolás Bodnariuk, Marcela Ivanovic, Martín Saraceno, and Eduardo Marcelo Acha. Dynamics of the confluence of Malvinas and Brazil currents, and a southern Patagonian spawning ground, explain recruitment fluctuations of the main stock of Illex argentinus. Fisheries Oceanography, 30(2):127–141, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Auth:2011:RFL

[ABS^{+}11] Toby D. Auth, Richard D. Brodeur, Heather L. Soulen, Lorenzo Ciannelli, and William T. Peterson. The response of fish larvae to decadal changes in environmental forcing factors off the Oregon coast. Fisheries Oceanography, 20(4): 314–328, July 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Aldanondo:2016:IDG

[ACG^{+}16] Naroa Aldanondo, Unai Cotano, Nerea Goikoetxea, Guillermo Boyra, Leire Ibaibarriaga, and Xabier Irigoien. Interannual differences in growth and hatch-date distributions of early juvenile European anchovy in the Bay of Biscay: implications for recruitment. Fisheries Oceanography, 25(2): 147–163, March 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cabrero:2019:SSS

[áCGNGC19] águeda Cabrero, Gonzalo González-Nuevo, Jesús Gago, and Jose Manuel Cabanas. Study of sardine (Sardina pilchardus) regime shifts in the Iberian Atlantic shelf waters. Fisheries Oceanography, 28(3):305–316, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Aldanondo:2010:GMP

[ACT^{+}10] Naroa Aldanondo, Unai Cotano, Massimo Tiepolo, Guillermo Boyra, and Xabier Irigoien. Growth and movement patterns of early juvenile European anchovy (Engraulis encrasicolus L.) in the Bay of Biscay based on otolith microstructure and chemistry. Fisheries Oceanography, 19(3):196–208, May 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Atwood:2010:IME

[ADAHL10] Elizabeth Atwood, Janet T. Duffy-Anderson, John K. Horne, and Carol Ladd. Influence of mesoscale eddies on ichthyoplankton assemblages in the Gulf of Alaska. Fisheries Oceanography, 19(6):493–507, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Alonso:2021:OVE

[ADPC21] Virginia Andrea García Alonso, Marina Vera Diaz, Marcelo Pájaro, and Fabiana Lia Capitanio. Ontogeny versus environmental forcing off the Southwest Atlantic Ocean: Nutritional condition of Fuegian sprat (Sprattus fuegensis) early stages. Fisheries Oceanography, 30(6):653–665, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Andrade:1999:STF

[AG99] Humber Agrelli Andrade and Carlos Alberto Eiras Garcia. Skipjack tuna fishery in relation to sea surface temperature off the southern Brazilian coast. Fisheries Oceanography, 8 (4):245–254, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Aoyama:2008:MBR

[AGK^{+}08] Michio Aoyama, Hirohito Goto, Hitomi Kamiya, Ikuo Kaneko, Satoshi Kawae, Hiroki Kodama, Yasushi Konishi, Kenn-Ichi Kusumoto, Hisato Miura, Eiichi Moriyama, Kiyoshi Murakami, Toshiya Nakano, Futoshi Nozaki, Daisuke Sasano, Takao Shimizu, Hiroki Suzuki, Yasushi Takatsuki, and Akito Toriyama. Marine biogeochemical response to a rapid warming in the main stream of the Kuroshio in the western North Pacific. Fisheries Oceanography, 17(3):206–218, May 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Aadlandsvik:2004:MAD

[ÅGN^{+}04] Bjørn Ådlandsvik, Agnes C. Gundersen, Kjell H. Nedreaas, Anne Stene, and Ole T. Albert. Modelling the advection and diffusion of eggs and larvae of Greenland halibut (Reinhardtius hippoglossoides) in the north-east Arctic. Fisheries Oceanography, 13(6):403–415, November 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arkhipkin:2004:ISE

[AGS^{+}04] A. I. Arkhipkin, R. Grzebielec, A. M. Sirota, A. V. Remeslo, I. A. Polishchuk, and D. A. J. Middleton. The influence of seasonal environmental changes on ontogenetic migrations of the squid Loligo gahi on the Falkland shelf. Fisheries Oceanography, 13(1):1–9, January 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anaya-Godinez:2022:IOC

[AGSSL^{+}22] Eduardo Anaya-Godínez, Claudia A. Silva-Segundo, Mauricio F. Landaeta, René Funes-Rodríguez, Oscar Trujillo-Millán, Anidia Blanco-Jarvio, Carolina Galván-Tirado, Fausto Valenzuela-Quiñonez, and Emigdio Marín-Enríquez. Influence of oceanographic conditions on the body shape variability of Scomber japonicus larvae from the western coast of the Baja California Peninsula. Fisheries Oceanography, 31(3):225–237, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Alheit:1997:LTC

[AH97] Jürgen Alheit and Eberhard Hagen. Long-term climate forcing of European herring and sardine populations. Fisheries Oceanography, 6(2):130–139, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arai:2003:CAL

[AHAM03] T. Arai, H. Hayano, H. Asami, and N. Miyazaki. Coexistence of anadromous and lacustrine life histories of the shirauo, Salangichthys microdon. Fisheries Oceanography, 12 (2):134–139, March 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Akimova:2016:MET

[AHKP16] A. Akimova, M. Hufnagl, M. Kreus, and M. A. Peck. Modeling the effects of temperature on the survival and growth of North Sea cod (Gadus morhua) through the first year of life. Fisheries Oceanography, 25(3):193–209, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Aoki:1992:AOF

[AI92] Ichiro Aoki and Tadashi Inagaki. Acoustic observations of fish schools and scattering layers in a Kuroshio warm-core ring and its environs. Fisheries Oceanography, 1(2):137–142, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Azumaya:2004:EPI

[AI04] Tomonori Azumaya and Yukimasa Ishida. An evaluation of the potential influence of SST and currents on the oceanic migration of juvenile and immature chum salmon (Oncorhynchus keta) by a simulation model. Fisheries Oceanography, 13(1):10–23, January 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Azumaya:2005:MBC

[AI05] Tomonori Azumaya and Yukimasa Ishida. Mechanism of body cavity temperature regulation of chum salmon (Oncorhynchus keta) during homing migration in the North Pacific Ocean. Fisheries Oceanography, 14(2):81–96, March 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Axelsen:2015:EBT

[AJ15] Bjørn Erik Axelsen and Espen Johnsen. An evaluation of the bottom trawl surveys in the Benguela Current Large Marine Ecosystem. Fisheries Oceanography, 24(S1):74–87, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ayala:2018:GRV

[AM18] Daniel Jiro Ayala and Peter Munk. Growth rate variability of larval European eels (Anguilla anguilla) across the extensive eel spawning area in the southern Sargasso Sea. Fisheries Oceanography, 27(6):525–535, November 2018. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Abascal:2016:TBT

[AMD^{+}16] Francisco J. Abascal, Antonio Medina, José M. De La Serna, Dolores Godoy, and Guillermo Aranda. Tracking bluefin tuna reproductive migration into the Mediterranean Sea with electronic pop-up satellite archival tags using two tagging procedures. Fisheries Oceanography, 25(1):54–66, January 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Atcheson:2012:PTE

[AMDM12] Margaret E. Atcheson, Katherine W. Myers, Nancy D. Davis, and Nathan J. Mantua. Potential trophodynamic and environmental drivers of steelhead (Oncorhynchus mykiss) productivity in the North Pacific Ocean. Fisheries Oceanography, 21(5):321–335, September 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Andersen:2008:PBM

[AMK08] K. H. Andersen, C. Möllmann, and F. W. Köster. Processbased model for direct and indirect effects of hydrographic conditions on Central Baltic cod (Gadus morhua) egg mortality. Fisheries Oceanography, 17(2):84–88, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Andrade:2003:RBS

[And03] H. A. Andrade. The relationship between the skipjack tuna (Katsuwonus pelamis) fishery and seasonal temperature variability in the south-western Atlantic. Fisheries Oceanography, 12(1):10–18, January 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:1994:BR

[Ano94] Anonymous. Book reviews. Fisheries Oceanography, 3(3): 217–220, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:1995:A

[Ano95a] Anonymous. Acknowledgements. Fisheries Oceanography, 4(1):83, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:1995:O

[Ano95b] Anonymous. Obituary. Fisheries Oceanography, 4(1):81–82, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:1999:ICF

[Ano99] Anonymous. Investigation of Calanus finmarchicus migrations between oceanic and shelf seas off north-west Europe (ICOS). Fisheries Oceanography, 8(S1):i, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2000:Eb

[Ano00a] Anonymous. Errata. Fisheries Oceanography, 9(4):377–378, December 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2000:Ea

[Ano00b] Anonymous. Erratum. Fisheries Oceanography, 9(1):120, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2001:IVA

[Ano01a] Anonymous. Index to volume 10: Author. Fisheries Oceanography, 10(4):380–381, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2001:IVS

[Ano01b] Anonymous. Index to volume 10: Subject. Fisheries Oceanography, 10(4):376–380, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2001:I

[Ano01c] Anonymous. Introduction. Fisheries Oceanography, 10(S1): v, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2001:MMM

[Ano01d] Anonymous. Michael M. Mullin 1937–2000. Fisheries Oceanography, 10(1):131, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2002:E

[Ano02] Anonymous. Errata. Fisheries Oceanography, 11(3):191, May 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2003:GSI

[Ano03a] Anonymous. GLOBEC special issue: Foreword. Fisheries Oceanography, 12(4–5):221–222, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2003:IVA

[Ano03b] Anonymous. Index to volume 12: Author. Fisheries Oceanography, 12(6):593–594, November 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2003:IVS

[Ano03c] Anonymous. Index to volume 12: Subject. Fisheries Oceanography, 12(6):584–593, November 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2003:IV

[Ano03d] Anonymous. Indexes to volume 12. Fisheries Oceanography, 12(6):595–597, November 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2004:IVA

[Ano04a] Anonymous. Index to volume 13: Author. Fisheries Oceanography, 13(6):422–423, November 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2004:IVS

[Ano04b] Anonymous. Index to volume 13: Subject. Fisheries Oceanography, 13(6):416–422, November 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2005:IVA

[Ano05a] Anonymous. Index to volume 14: Author. Fisheries Oceanography, 14(6):474–475, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2005:IVS

[Ano05b] Anonymous. Index to volume 14: Subject. Fisheries Oceanography, 14(6):468–474, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2006:VC

[Ano06] Anonymous. Volume contents. Fisheries Oceanography, 15 (6):i–iii, November 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2007:LR

[Ano07] Anonymous. List of reviewers. Fisheries Oceanography, 16 (6):578, November 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2010:LR

[Ano10] Anonymous. List of reviewers. Fisheries Oceanography, 19 (3):254–255, May 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2011:Ca

[Ano11a] Anonymous. Corrigendum. Fisheries Oceanography, 20(1): 92–93, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See [PLG^{+}10].

Anonymous:2011:Cb

[Ano11b] Anonymous. Corrigendum. Fisheries Oceanography, 20(6): 583, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2012:C

[Ano12] Anonymous. Corrigendum. Fisheries Oceanography, 21(2– 3):226–227, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2014:E

[Ano14] Anonymous. Erratum. Fisheries Oceanography, 23(6):459, November 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2015:C

[Ano15] Anonymous. Corrigendum. Fisheries Oceanography, 24(4): 393–394, July 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2016:N

[Ano16] Anonymous. Note. Fisheries Oceanography, 25(S1):5, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2017:C

[Ano17] Anonymous. Corrigendum. Fisheries Oceanography, 26(6): 705, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See [JJBCW09].

Anonymous:2018:IIa

[Ano18a] Anonymous. Issue information. Fisheries Oceanography, 27 (1):i–iv, January 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2018:IIb

[Ano18b] Anonymous. Issue information. Fisheries Oceanography, 27 (2):i–iv, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2018:IIc

[Ano18c] Anonymous. Issue information. Fisheries Oceanography, 27(3):i–iv, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2018:IId

[Ano18d] Anonymous. Issue information. Fisheries Oceanography, 27(4):i–iv, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2018:IIe

[Ano18e] Anonymous. Issue information. Fisheries Oceanography, 27 (5):i–iv, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2018:IIf

[Ano18f] Anonymous. Issue information. Fisheries Oceanography, 27 (6):i–iv, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:C

[Ano19a] Anonymous. Corrigendum. Fisheries Oceanography, 28(1): 113–115, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:IIa

[Ano19b] Anonymous. Issue information. Fisheries Oceanography, 28 (1):i–iv, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:IIb

[Ano19c] Anonymous. Issue information. Fisheries Oceanography, 28 (2):i–iv, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:IIc

[Ano19d] Anonymous. Issue information. Fisheries Oceanography, 28(3):i–iv, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:IId

[Ano19e] Anonymous. Issue information. Fisheries Oceanography, 28(4):i–iv, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:IIe

[Ano19f] Anonymous. Issue information. Fisheries Oceanography, 28 (5):i–iv, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2019:IIf

[Ano19g] Anonymous. Issue information. Fisheries Oceanography, 28 (6):i–iv, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Fisheries Oceanography, 29 (1):i–iv, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Fisheries Oceanography, 29 (2):i–iv, March 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Fisheries Oceanography, 29(3):i–iv, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Fisheries Oceanography, 29(4):i–iv, July 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Fisheries Oceanography, 29 (5):i–iv, September 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Fisheries Oceanography, 29 (6):i–iv, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2021:IIa

[Ano21a] Anonymous. Issue information. Fisheries Oceanography, 30 (1):i–iv, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2021:IIb

[Ano21b] Anonymous. Issue information. Fisheries Oceanography, 30 (2):i–iv, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2021:IIc

[Ano21c] Anonymous. Issue information. Fisheries Oceanography, 30(3):i–iv, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2021:IId

[Ano21d] Anonymous. Issue information. Fisheries Oceanography, 30(4):i–iv, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2021:IIe

[Ano21e] Anonymous. Issue information. Fisheries Oceanography, 30 (5):i–iv, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2021:IIf

[Ano21f] Anonymous. Issue information. Fisheries Oceanography, 30 (6):i–iv, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2022:IIa

[Ano22a] Anonymous. Issue information. Fisheries Oceanography, 31 (1):i–iv, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2022:IIb

[Ano22b] Anonymous. Issue information. Fisheries Oceanography, 31 (2):i–iv, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2022:IIc

[Ano22c] Anonymous. Issue information. Fisheries Oceanography, 31(3):i–iv, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2022:IId

[Ano22d] Anonymous. Issue information. Fisheries Oceanography, 31(4):i–iv, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anonymous:2022:IIe

[Ano22e] Anonymous. Issue information. Fisheries Oceanography, 31 (5):i–iv, September 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Avendano:2022:RAD

[AOVAG22] Otilio Avendaño, Jaime Otero, Iván Velázquez-Abunader, and Ángel Guerra. Relative abundance distribution and body size changes of two co-occurring octopus species, Octopus americanus and Octopus maya, in a tropical upwelling area (south-eastern Gulf of Mexico). Fisheries Oceanography, 31(4):402–415, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Allain:2003:SPL

[APGL03] Gwenhael Allain, Pierre Petitgas, Patrick Grellier, and Pascal Lazure. The selection process from larval to juvenile stages of anchovy (Engraulis encrasicolus) in the Bay of Biscay investigated by Lagrangian simulations and comparative otolith growth. Fisheries Oceanography, 12(4–5):407–418, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Adkison:1996:AMC

[APL^{+}96] Milo D. Adkison, Randall M. Peterman, Michael F. Lapointe, Darren M. Gillis, and Josh Korman. Alternative models of climatic effects on sockeye salmon, Oncorhynchus nerka, productivity in Bristol Bay, Alaska, and the Fraser River, British Columbia. Fisheries Oceanography, 5(3–4): 137–152, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Allain:2001:IMO

[APL01] G. Allain, P. Petitgas, and P. Lazure. The influence of mesoscale ocean processes on anchovy (Engraulis encrasicolus) recruitment in the Bay of Biscay estimated with a three-dimensional hydrodynamic mode. Fisheries Oceanography, 10(2):151–163, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Allain:2007:IES

[APL07] Gwenhael Allain, Pierre Petitgas, and Pascal Lazure. The influence of environment and spawning distribution on the survival of anchovy (Engraulis encrasicolus) larvae in the Bay of Biscay (NE Atlantic) investigated by biophysical simulations. Fisheries Oceanography, 16(6):506–514, November 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arimitsu:2008:DSD

[APL^{+}08] Mayumi L. Arimitsu, John F. Piatt, Michael A. Litzow, Alisa A. Abookire, Marc D. Romano, and Martin D. Robards. Distribution and spawning dynamics of capelin (Mallotus villosus) in Glacier Bay, Alaska: a cold water refugium. Fisheries Oceanography, 17(2):137–146, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Allain:2007:BML

[APLG07] Gwenhael Allain, Pierre Petitgas, Pascal Lazure, and Patrick Grellier. Biophysical modelling of larval drift, growth and survival for the prediction of anchovy (Engraulis encrasicolus) recruitment in the Bay of Biscay (NE Atlantic). Fisheries Oceanography, 16(6):489–505, November 2007. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arimitsu:2012:OGS

[APM^{+}12] Mayumi L. Arimitsu, John F. Piatt, Erica N. Madison, Jeffrey S. Conaway, and Nicola Hillgruber. Oceanographic gradients and seabird prey community dynamics in glacial fjords. Fisheries Oceanography, 21(2–3):148–169, March– May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Acosta-Pachon:2017:HPS

[APMRH17] Tatiana A. Acosta-Pachón, Raúl O. Martínez-Rincón, and Michael G. Hinton. Habitat preferences of striped marlin (Kajikia audax) in the eastern Pacific Ocean. Fisheries Oceanography, 26(6):615–624, November 2017. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Acosta-Pachon:2019:SSS

[APMVOGMR19] Tatiana A. Acosta-Pachón, Mariana L. Muzquiz-Villalobos, Sofia Ortega-García, and Raúl O. Martínez-Rincón. Spatial segregation of striped marlin (Kajikia audax) by size in the eastern Pacific Ocean. Fisheries Oceanography, 28(2): 203–211, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Arrizabalaga:2008:BTT

[APR^{+}08] H. Arrizabalaga, J. G. Pereira, F. Royer, B. Galuardi, N. Goñi, I. Artetxe, I. Arregi, and M. Lutcavage. Bigeye tuna (Thunnus obesus) vertical movements in the Azores Islands determined with pop-up satellite archival tags. Fisheries Oceanography, 17(2):74–83, March 2008. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Roura:2016:LSC

[áRÁSG^{+}16] álvaro Roura, X. Antón Álvarez-Salgado, Ángel F. González, María Gregori, Gabriel Rosón, Jaime Otero, and Ángel Guerra. Life strategies of cephalopod paralarvae in a coastal upwelling system (NW Iberian Peninsula): insights from zooplankton community and spatio-temporal analyses. Fisheries Oceanography, 25(3):241–258, May 2016. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Adams:1993:OEC

[ARL93] Peter B. Adams, Stephen Ralston, and Thomas E. Laidig. Occurrence of an exceptional catch of pelagic juvenile lingcod (Ophiodon elongatus) off Point Reyes, California. Fisheries Oceanography, 2(2):97–100, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ayala:2016:SCD

[ARM16] Daniel Ayala, Lasse Riemann, and Peter Munk. Species composition and diversity of fish larvae in the Subtropical Convergence Zone of the Sargasso Sea from morphology and DNA barcoding. Fisheries Oceanography, 25(1):85–104, January 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ayon:2008:CLT

[AS08] Patricia Ayón and Gordon Swartzman. Changes in the longterm distribution of zooplankton in the Humboldt Current ecosystem off Peru, 1961–2005, and its relationship to regime shifts and environmental factors. Fisheries Oceanography, 17 (6):421–431, November 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Acha:2012:CIY

[ASCM12] E. M. Acha, C. G. Simionato, C. Carozza, and H. Mianzan. Climate-induced year-class fluctuations of whitemouth croaker Micropogonias furnieri (Pisces, Sciaenidae) in the Río de la Plata estuary, Argentina–Uruguay. Fisheries Oceanography, 21(2–3):58–77, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Asplin:1999:NWD

[ASK99] Lars Asplin, Anne Gro Vea Salvanes, and Jon Bent Kristoffersen. Nonlocal wind-driven fjord–coast advection and its potential effect on plankton and fish recruitment. Fisheries Oceanography, 8(4):255–263, December 1999. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Alabia:2015:SPF

[ASM^{+}15] Irene D. Alabia, Sei-Ichi Saitoh, Robinson Mugo, Hiromichi Igarashi, Yoichi Ishikawa, Norihisa Usui, Masafumi Kamachi, Toshiyuki Awaji, and Masaki Seito. Seasonal potential fishing ground prediction of neon flying squid (Ommastrephes bartramii) in the western and central North Pacific. Fisheries Oceanography, 24(2):190–203, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Crow:2005:TSE

[aTCK05] J. D. Schumacher a.k.a. Two Crow and Gordon H. Kruse. Toward sustainable ecosystem services from the Aleutian Archipelago. Fisheries Oceanography, 14(S1):277–291, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Abad-Uribarren:2019:MEI

[AUOGMM19] Alberto Abad-Uribarren, Sofía Ortega-García, David March, and Antonio Medina. Modeling environmental influence on Atlantic bluefin tuna bycatch by Mexican longliners in the Gulf of Mexico. Fisheries Oceanography, 28(6):672–685, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Auth:2008:DCS

[Aut08] Toby D. Auth. Distribution and community structure of ichthyoplankton from the northern and central California Current in May 2004–06. Fisheries Oceanography, 17(4): 316–331, July 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Anderson:1992:SER

[AW92] John T. Anderson and Ian T. Webster. Sampling error resulting from depth-dependent plankton distributions and current differentials in a two-layered system. Fisheries Oceanography, 1(3):208–215, September 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Aita:2003:EOV

[AYK03] Maki N. Aita, Yasuhiro Yamanaka, and Michio J. Kishi. Effects of ontogenetic vertical migration of zooplankton on annual primary production — using NEMURO embedded in a general circulation model. Fisheries Oceanography, 12(4–5): 284–290, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Akira:2001:IVG

[AYMK01] N. Akira, T. Yanagimoto, K. Mito, and S. Katakura. Interannual variability in growth of walleye pollock, Theragra chalcogramma, in the central Bering Sea. Fisheries Oceanography, 10(4):367–375, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bracis:2012:IGI

[BA12] Chloe Bracis and James J. Anderson. An investigation of the geomagnetic imprinting hypothesis for salmon. Fisheries Oceanography, 21(2–3):170–181, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Becognee:2006:ACC

[BAB^{+}06] P. Bécognée, C. Almeida, A. Barrera, A. Hernández-Guerra, and S. Hernández-León. Annual cycle of clupeiform larvae around Gran Canaria Island, Canary Islands. Fisheries Oceanography, 15(4):293–300, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Blanton:1999:TSF

[BAL^{+}99] Jackson O. Blanton, Julie Amft, Richard A. Luettich, Jr., James L. Hench, and James H. Churchill. Tidal and subtidal fluctuations in temperature, salinity and pressure for the winter 1996 larval ingress experiment — Beaufort Inlet, NC. Fisheries Oceanography, 8(S2):134–152, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Black:2011:MOG

[BASS11] Bryan A. Black, Robert J. Allman, Isaac D. Schroeder, and Michael J. Schirripa. Multidecadal otolith growth histories for red and gray snapper (Lutjanus spp.) in the northern Gulf of Mexico, USA. Fisheries Oceanography, 20(5):347– 356, September 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Baumann:1995:CTE

[Bau95] Michael Baumann. A comment on transfer efficiencies. Fisheries Oceanography, 4(3):264–266, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Baumann:1998:FMM

[Bau98] Michael Baumann. The fallacy of the missing middle: physics → … → fisheries. Fisheries Oceanography, 7(1):63– 65, April 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bakun:2002:IIP

[BB02] Andrew Bakun and Kenneth Broad. The IRI–IPRC Pacific Climate-Fisheries Workshop. Fisheries Oceanography, 11(3): 189–190, May 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bakun:2003:ELF

[BB03] Andrew Bakun and Kenneth Broad. Environmental ‘loopholes’ and fish population dynamics: comparative pattern recognition with focus on El Niño effects in the Pacific. Fisheries Oceanography, 12(4–5):458–473, September 2003. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brunel:2007:LTT

[BB07] Thomas Brunel and Jean Boucher. Long-term trends in fish recruitment in the north-east Atlantic related to climate change. Fisheries Oceanography, 16(4):336–349, July 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bourdaud:2021:ICC

[BBA^{+}21] Pierre Bourdaud, Frida Ben Rais Lasram, Emma Araignous, Juliette Champagnat, Samantha Grusd, Ghassen Halouani, Tarek Hattab, Boris Leroy, Quentin Noguès, Aurore Raoux, Georges Safi, and Nathalie Niquil. Impacts of climate change on the Bay of Seine ecosystem: Forcing a spatio-temporal trophic model with predictions from an ecological niche model. Fisheries Oceanography, 30(5):471–489, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bonanno:2016:EPD

[BBB^{+}16] A. Bonanno, M. Barra, G. Basilone, S. Genovese, P. Rumolo, S. Goncharov, S. Popov, B. Buongiorno Nardelli, D. Iudicone, G. Procaccini, S. Aronica, B. Patti, G. Giacalone, R. Ferreri, I. Fontana, G. Tranchida, S. Mangano, M. Pulizzi, A. Gargano, A. Di Maria, and S. Mazzola. Environmental processes driving anchovy and sardine distribution in a highly variable environment: the role of the coastal structure and riverine input. Fisheries Oceanography, 25(5):471–490, September 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Burrows:2019:DBR

[BBB^{+}19] Makenzie Burrows, Jeremy S. Browning, Mya Breitbart, Steven A. Murawski, and Ernst B. Peebles. DNA barcoding reveals clear delineation between spawning sites for neritic versus oceanic fishes in the Gulf of Mexico. Fisheries Oceanography, 28(2):228–239, March 2019. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bigelow:1999:EES

[BBH99] Keith A. Bigelow, Christofer H. Boggs, and Xi He. Environmental effects on swordfish and blue shark catch rates in the US North Pacific longline fishery. Fisheries Oceanography, 8(3):178–198, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bailey:1993:PWP

[BBMY93] Kevin M. Bailey, Richard D. Brodeur, Nazila Merati, and Mary M. Yoklavich. Predation on walleye pollock (Theragra chakograrnma) eggs and yolk-sac larvae by pelagic crustacean invertebrates in the western Gulf of Alaska. Fisheries Oceanography, 2(1):30–39, March 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Basilone:2013:SSS

[BBP^{+}13] G. Basilone, A. Bonanno, B. Patti, S. Mazzola, M. Barra, A. Cuttitta, and R. McBride. Spawning site selection by European anchovy (Engraulis encrasicolus) in relation to oceanographic conditions in the Strait of Sicily. Fisheries Oceanography, 22(4):309–323, July 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bogazzi:2005:SCB

[BBR^{+}05] Eugenia Bogazzi, Ana Baldoni, Andrés Rivas, Patricia Martos, Raúl Reta, J. M. (Lobo) Orensanz, Mario Lasta, Patricia Dell’Arciprete, and Francisco Werner. Spatial correspondence between areas of concentration of Patagonian scallop (Zygochlamys patagonica) and frontal systems in the southwestern Atlantic. Fisheries Oceanography, 14(5):359– 376, September 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bailey:1999:ATW

[BBS99] Kevin M. Bailey, Nicholas A. Bond, and Phyllis J. Stabeno. Anomalous transport of walleye pollock larvae linked to ocean and atmospheric patterns in May 1996. Fisheries Oceanography, 8(4):264–273, December 1999. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bonhommeau:2009:HFC

[BBT^{+}09] Sylvain Bonhommeau, Bruno Blanke, Anne-Marie Tréguier, Nicolas Grima, Etienne Rivot, Youen Vermard, Eric Greiner, and Olivier Le Pape. How fast can the European eel (Anguilla anguilla) larvae cross the Atlantic Ocean? Fisheries Oceanography, 18(6):371–385, November 2009. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Black:2008:ECG

[BBY08] Bryan A. Black, George W. Boehlert, and Mary M. Yoklavich. Establishing climate–growth relationships for yelloweye rockfish (Sebastes ruberrimus) in the northeast Pacific using a dendrochronological approach. Fisheries Oceanography, 17(5):368–379, September 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bartsch:1997:NMD

[BC97] J. Bartsch and S. Coombs. A numerical model of the dispersion of blue whiting larvae, Micromesistius poutassou (Risso), in the eastern North Atlantic. Fisheries Oceanography, 6(3):141–154, October 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bartsch:2004:IBM

[BC04] J. Bartsch and S. H. Coombs. An individual-based model of the early life history of mackerel (Scomber scombrus) in the eastern North Atlantic, simulating transport, growth and mortality. Fisheries Oceanography, 13(6):365–379, November 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brooks:2018:ELH

[BCA^{+}18] Cassandra M. Brooks, Jilda Alicia Caccavo, Julian Ashford, Robert Dunbar, Kimberly Goetz, Mario La Mesa, and Lorenzo Zane. Early life history connectivity of Antarctic silverfish (Pleuragramma antarctica) in the Ross Sea. Fisheries Oceanography, 27(3):274–287, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bacheler:2010:STP

[BCBDA10] Nathan M. Bacheler, Lorenzo Ciannelli, Kevin M. Bailey, and Janet T. Duffy-Anderson. Spatial and temporal patterns of walleye pollock (Theragra chalcogramma) spawning in the eastern Bering Sea inferred from egg and larval distributions. Fisheries Oceanography, 19(2):107–120, March 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beggs:2014:LCG

[BCGB14] Steven E. Beggs, Massimiliano Cardinale, Richard J. Gowen, and Valerio Bartolino. Linking cod (Gadus morhua) and climate: investigating variability in Irish Sea cod recruitment. Fisheries Oceanography, 23(1):54–64, January 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Basterretxea:2013:DRL

[BCJ^{+}13] Gotzon Basterretxea, Ignacio A. Catalán, Antoni Jordi, Itziar Álvarez, Miquel Palmer, and Ana Sabatés. Dynamic regulation of larval fish self-recruitment in a marine protected area. Fisheries Oceanography, 22(6):477–495, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Buckley:2004:OTF

[BCL04] Lawrence J. Buckley, Elaine M. Caldarone, and R. Gregory Lough. Optimum temperature and food-limited growth of larval Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) on Georges Bank. Fisheries Oceanography, 13(2):134–140, March 2004. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bonhommeau:2008:FEE

[BCR08] Sylvain Bonhommeau, Emmanuel Chassot, and Etienne Rivot. Fluctuations in European eel (Anguilla anguilla) recruitment resulting from environmental changes in the Sargasso Sea. Fisheries Oceanography, 17(1):32–44, January 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Baez:2020:COE

[BCR20] José Carlos Báez, Ivone A. Czerwinski, and María Lourdes Ramos. Climatic oscillations effect on the yellowfin tuna (Thunnus albacares) Spanish captures in the Indian Ocean. Fisheries Oceanography, 29(6):572–583, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brander:1993:PAD

[BD93] Keith M. Brander and Robert R. Dickson. Papers in appreciation of David Henry Cushing FRS. Fisheries Oceanography, 2(3–4):109–113, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Busby:2014:STP

[BDAMD14] M. S. Busby, J. T. Duffy-Anderson, K. L. Mier, and L. G. De Forest. Spatial and temporal patterns in summer ichthyoplankton assemblages on the eastern Bering Sea shelf 1996– 2007. Fisheries Oceanography, 23(3):270–287, May 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brown:1993:USI

[BDBP93] R. M. Brown, K. L. Denman, G. A. Borstad, and J. R. Parks. The use of satellite imagery to direct research ship sampling operations. Fisheries Oceanography, 2(3–4):184–190, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Balde:2019:BSE

[BDE^{+}19] Bocar Sabaly Baldé, Julian Döring, Werner Ekau, Malick Diouf, and Patrice Brehmer. Bonga shad (Ethmalosa fimbriata) spawning tactics in an upwelling environment. Fisheries Oceanography, 28(6):686–697, November 2019. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brodeur:2007:IIV

[BDSM07] Richard D. Brodeur, Elizabeth A. Daly, Robert A. Schabetsberger, and Kathryn L. Mier. Interannual and interdecadal variability in juvenile Coho salmon (Oncorhynchus kisutch) diets in relation to environmental changes in the northern California Current. Fisheries Oceanography, 16(5):395–408, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Burbank:2023:UFI

[BDTR23] Jacob Burbank, Rachel A. DeJong, François Turcotte, and Nicolas Rolland. Understanding factors influencing Atlantic herring (Clupea harengus) recruitment: From egg deposition to juveniles. Fisheries Oceanography, 32(2):147–159, March 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brosset:2019:EVC

[BDVS^{+}19] Pablo Brosset, Thomas Doniol-Valcroze, Douglas P. Swain, Caroline Lehoux, Elisabeth Van Beveren, Baye C. Mbaye, Kim Emond, and Stéphane Plourde. Environmental variability controls recruitment but with different drivers among spawning components in Gulf of St. Lawrence herring stocks. Fisheries Oceanography, 28(1):1–17, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beaugrand:2003:LTC

[Bea03] Gregory Beaugrand. Long-term changes in copepod abundance and diversity in the north-east Atlantic in relation to fluctuations in the hydroclimatic environment. Fisheries Oceanography, 12(4–5):270–283, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Berry:2012:MDA

[BEF^{+}12] Oliver Berry, Phillip England, David Fairclough, Gary Jackson, and Jim Greenwood. Microsatellite DNA analysis and hydrodynamic modelling reveal the extent of larval transport and gene flow between management zones in an exploited marine fish (Glaucosoma hebraicum). Fisheries Oceanography, 21(4):243–254, July 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bograd:2023:FOF

[BEiI^{+}23] Steven J. Bograd, Martin Edwards, Shin ichi Ito, Janet Nye, and Emily Chappell. Fisheries Oceanography: The first 30 years and new challenges in the 21st century. Fisheries Oceanography, 32(1):3–9, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bez:2000:ULI

[Bez00] N. Bez. On the use of Lloyd’s index of patchiness. Fisheries Oceanography, 9(4):372–376, December 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Batten:2007:PPB

[BF07] Sonia D. Batten and Howard J. Freeland. Plankton populations at the bifurcation of the North Pacific Current. Fisheries Oceanography, 16(6):536–546, November 2007. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bauer:2015:OPD

[BFF15] Robert Klaus Bauer, Fabien Forget, and Jean-Marc Fromentin. Optimizing PAT data transmission: assessing the accuracy of temperature summary data to estimate environmental conditions. Fisheries Oceanography, 24(6):533–539, November 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Borja:2008:COR

[BFSV08] Angel Borja, Almudena Fontán, Jon Sáenz, and Victoriano Valencia. Climate, oceanography, and recruitment: the case of the Bay of Biscay anchovy (Engraulis encrasicolus). Fisheries Oceanography, 17(6):477–493, November 2008. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bishop:2001:PPH

[BG01] Mary A. Bishop and S. Patrick Green. Predation on Pacific herring (Clupea pallasi) spawn by birds in Prince William Sound, Alaska. Fisheries Oceanography, 10(S1):149–158, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bestley:2009:PVB

[BGH09] Sophie Bestley, John S. Gunn, and Mark A. Hindell. Plasticity in vertical behaviour of migrating juvenile southern bluefin tuna (Thunnus maccoyii) in relation to oceanography of the south Indian Ocean. Fisheries Oceanography, 18 (4):237–254, July 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brigolin:2018:URS

[BGM^{+}18] D. Brigolin, P. Girardi, P. I. Miller, W. Xu, D. Nachite, M. Zucchetta, and F. Pranovi. Using remote sensing indicators to investigate the association of landings with fronts: Application to the Alboran Sea (western Mediterranean Sea). Fisheries Oceanography, 27(5):408–416, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Basilone:2006:EHC

[BGP^{+}06] G. Basilone, C. Guisande, B. Patti, S. Mazzola, A. Cuttitta, A. Bonanno, A. R. Vergara, and I. Maneiro. Effect of habitat conditions on reproduction of the European anchovy (Engraulis encrasicolus) in the Strait of Sicily. Fisheries Oceanography, 15(4):271–280, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Barange:1997:SSC

[BH97] Manuel Barange and Ian Hampton. Spatial structure of co-occurring anchovy and sardine populations from acoustic data: implications for survey design. Fisheries Oceanography, 6(2):94–108, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Barbeaux:2018:OMC

[BH18] Steven J. Barbeaux and Anne B. Hollowed. Ontogeny matters: Climate variability and effects on fish distribution in the eastern Bering Sea. Fisheries Oceanography, 27(1):1–15, January 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Baduini:2001:MMS

[BHC^{+}01] C. L. Baduini, K. D. Hyrenbach, K. O. Coyle, A. Pinchuk, V. Mendenhall, and G. L. Hunt, Jr. Mass mortality of short-tailed shearwaters in the south-eastern Bering Sea during summer 1997. Fisheries Oceanography, 10(1):117–130, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bryant:1998:BSA

[BHH98] Andrew D. Bryant, Dagmar Hainbucher, and Michael Heath. Basin-scale advection and population persistence of Calanus finmarchicus. Fisheries Oceanography, 7(3–4):235–244, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brown:2004:SLS

[BHJ^{+}04] C. A. Brown, S. A. Holt, G. A. Jackson, D. A. Brooks, and G. J. Holt. Simulating larval supply to estuarine nursery areas: how important are physical processes to the supply of larvae to the Aransas Pass Inlet? Fisheries Oceanography, 13 (3):181–196, May 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bigelow:2002:AHB

[BHM02] Keith A. Bigelow, John Hampton, and Naozumi Miyabe. Application of a habitat-based model to estimate effective longline fishing effort and relative abundance of Pacific bigeye tuna (Thunnus obesus). Fisheries Oceanography, 11(3): 143–155, May 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brodie:2015:MOH

[BHS^{+}15] Stephanie Brodie, Alistair J. Hobday, James A. Smith, Jason D. Everett, Matt D. Taylor, Charles A. Gray, and Iain M. Suthers. Modelling the oceanic habitats of two pelagic species using recreational fisheries data. Fisheries Oceanography, 24(5):463–477, September 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Baumann:2006:LGE

[BHV^{+}06] Hannes Baumann, Hans-Harald Hinrichsen, Rudi Voss, Daniel Stepputtis, Wlodzimierz Grygiel, Lotte W. Clausen, and Axel Temming. Linking growth to environmental histories in central Baltic young-of-the-year sprat, Sprattus sprattus: an approach based on otolith microstructure analysis and hydrodynamic modelling. Fisheries Oceanography, 15 (6):465–476, November 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Basterretxea:2012:MBA

[BJCS12] Gotzon Basterretxea, Antoni Jordi, Ignacio A. Catalán, and Ana Sabatés. Model-based assessment of local-scale fish larval connectivity in a network of marine protected areas. Fisheries Oceanography, 21(4):291–306, July 2012. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bacha:2017:EES

[BJV^{+}17] Mahmoud Bacha, Mohamed Ahmed Jeyid, Vincent Vantrepotte, David Dessailly, and Rachid Amara. Environmental effects on the spatio-temporal patterns of abundance and distribution of Sardina pilchardus and sardinella off the Mauritanian coast (North-West Africa). Fisheries Oceanography, 26(3):282–298, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bartsch:1994:PDS

[BK94a] Joachim Bartsch and Rainer Knust. Predicting the dispersion of sprat larvae (Sprattus sprattus (L.)) in the German Bight. Fisheries Oceanography, 3(4):292–296, December 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bartsch:1994:SDV

[BK94b] Joachim Bartsch and Rainer Knust. Simulating the dispersion of vertically migrating sprat larvae (Sprattus sprattus (L.)) in the German Bight with a circulation and transport model system. Fisheries Oceanography, 3(2):92–105, June 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bahamon:2022:SBS

[BKvdP^{+}22] Nixon Bahamon, Johannes N. Kathena, Anja K. van der Plas, Paulus Kainge, Jorge Paramo, and Ana Gordoa. Spatial and biomass structure of shallow-water cape hake (Merluccius capensis) in the light of episodic environmental shifts. Fisheries Oceanography, 31(5):524–538, September 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Botsford:2003:WSB

[BLD^{+}03] Louis W. Botsford, Cathryn A. Lawrence, Edward P. Dever, Alan Hastings, and John Largier. Wind strength and biological productivity in upwelling systems: an idealized study. Fisheries Oceanography, 12(4–5):245–259, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Briceno:2016:SVM

[BLG^{+}16] Felipe Briceño, Rafael León, Caleb Gardner, Alistair J. Hobday, Jessica André, Stewart D. Frusher, and Gretta T. Pecl. Spatial variation in mortality by in-pot predation in the Tasmanian rock lobster fishery. Fisheries Oceanography, 25(S1): 6–18, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bertignac:1998:SPD

[BLH98] M. Bertignac, P. Lehodey, and J. Hampton. A spatial population dynamics simulation model of tropical tunas using a habitat index based on environmental parameters. Fisheries Oceanography, 7(3–4):326–334, December 1998. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beare:1999:CPR

[BM99a] D. Beare and E. McKenzie. Continuous plankton recorder data and diel vertical migration in stage V and VI Calanus finmarchicus: a statistical analysis. Fisheries Oceanography, 8(S1):126–137, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beare:1999:MLM

[BM99b] D. Beare and E. McKenzie. The multinomial logit model: a new tool for exploring continuous plankton recorder data. Fisheries Oceanography, 8(S1):25–39, September 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beck:2021:SVP

[BMH^{+}21] Jessie Beck, Pamela E. Michael, Michelle Hester, Hannahrose M. Nevins, Erica Donnelly-Greenan, Corinne Gibble, Elizabeth M. Phillips, Colleen Young, and Shannon Fitzgerald. Seasonal variation of Pacific northern fulmar bycatch: Implications for age and sex-specific mortality. Fisheries Oceanography, 30(3):253–263, May 2021. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Black:2013:OBM

[BMHW13] Bryan A. Black, Mary E. Matta, Thomas E. Helser, and Thomas K. Wilderbuer. Otolith biochronologies as multidecadal indicators of body size anomalies in yellowfin sole (Limanda aspera). Fisheries Oceanography, 22(6):523–532, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Briand:2011:SVA

[BML11] Karine Briand, Brett Molony, and Patrick Lehodey. A study on the variability of albacore (Thunnus alalunga) longline catch rates in the southwest Pacific Ocean. Fisheries Oceanography, 20(6):517–529, November 2011. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bartolino:2014:FFS

[BML^{+}14] Valerio Bartolino, Piotr Margonski, Martin Lindegren, Hans. W. Linderholm, Massimiliano Cardinale, David Rayner, Håkan Wennhage, and Michele Casini. Forecasting fish stock dynamics under climate change: Baltic herring (Clupea harengus) as a case study. Fisheries Oceanography, 23(3):258–269, May 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brodeur:1999:ESI

[BMO^{+}99] Richard D. Brodeur, Claudia E. Mills, James E. Overland, Gary E. Walters, and James D. Schumacher. Evidence for a substantial increase in gelatinous zooplankton in the Bering Sea, with possible links to climate change. Fisheries Oceanography, 8(4):296–306, December 1999. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Barret:2017:SBA

[BMOT17] Léo Barret, Gilles Miron, Patrick Ouellet, and Réjean Tremblay. Settlement behavior of American lobster (Homarus americanus): effect of female origin and developmental temperature. Fisheries Oceanography, 26(1):69–82, January 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Batten:2016:PIE

[BMPC16] S. D. Batten, S. Moffitt, W. S. Pegau, and R. Campbell. Plankton indices explain interannual variability in Prince William Sound herring first year growth. Fisheries Oceanography, 25(4):420–432, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beamish:2000:TCM

[BNM^{+}00] Richard J. Beamish, D. J. Noakes, G. A. Mcfarlane, W. Pinnix, R. Sweeting, and J. King. Trends in Coho marine survival in relation to the regime concept. Fisheries Oceanography, 9(1):114–119, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bond:2005:IEW

[BO05] Nicholas A. Bond and James E. Overland. The importance of episodic weather events to the ecosystem of the Bering Sea shelf. Fisheries Oceanography, 14(2):97–111, March 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bower:2011:FPH

[Bow11] John R. Bower. Four principles to help non-native speakers of English write clearly. Fisheries Oceanography, 20(1):89–91, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Boyra:2016:SDJ

[BPC^{+}16] Guillermo Boyra, Marian Peña, Unai Cotano, Xabier Irigoien, Anna Rubio, and Enrique Nogueira. Spatial dynamics of juvenile anchovy in the Bay of Biscay. Fisheries Oceanography, 25(5):529–543, September 2016. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bi:2011:CSC

[BPLC11] Hongsheng Bi, William T. Peterson, Jesse Lamb, and Edmundo Casillas. Copepods and salmon: characterizing the spatial distribution of juvenile salmon along the Washington and Oregon coast, USA. Fisheries Oceanography, 20(2): 125–138, March 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bellier:2007:HFS

[BPP07] Edwige Bellier, Benjamin Planque, and Pierre Petitgas. Historical fluctuations in spawning location of anchovy (Engraulis encrasicolus) and sardine (Sardina pilchardus) in the Bay of Biscay during 1967–73 and 2000–2004. Fisheries Oceanography, 16(1):1–15, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Byron:2014:MMP

[BPS^{+}14] Carrie J. Byron, Andrew J. Pershing, Jason D. Stockwell, Huijie Xue, and John Kocik. Migration model of post-smolt Atlantic salmon (Salmo salar) in the Gulf of Maine. Fisheries Oceanography, 23(2):172–189, March 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Baker:2019:MTA

[BPZR19] Matthew R. Baker, Wayne Palsson, Mark Zimmermann, and Christopher N. Rooper. Model of trawlable area using benthic terrain and oceanographic variables — informing survey design and habitat maps in the Gulf of Alaska. Fisheries Oceanography, 28(6):629–657, November 2019. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Boyra:2003:MVD

[BRC^{+}03] G. Boyra, L. Rueda, S. H. Coombs, S. Sundby, B. Ådlandsvik, M. Santos, and A. Uriarte. Modelling the vertical distribution of eggs of anchovy (Engraulis encrasicolus) and sardine (Sardina pilchardus). Fisheries Oceanography, 12(4–5):381– 395, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bartsch:2004:SMS

[BRC04] J. Bartsch, D. Reid, and S. H. Coombs. Simulation of mackerel (Scomber scombrus) recruitment with an individualbased model and comparison with field data. Fisheries Oceanography, 13(6):380–391, November 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bautista-Romero:2018:PDF

[BRFRJRLC18] J. Jesús Bautista-Romero, Rene Funes-Rodríguez, Sylvia P. A. Jiménez-Rosenberg, and Daniel B. Lluch-Cota. Preferential distribution of fish larvae in the California Current System: Time, space, and temperature. Fisheries Oceanography, 27(3):259–273, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brill:1994:RTO

[Bri94] Richard W. Brill. A review of temperature and oxygen tolerance studies of tunas pertinent to fisheries oceanography, movement models and stock assessments. Fisheries Oceanography, 3(3):204–216, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Beamish:1995:DCS

[BRN^{+}95] Richard J. Beamish, Brian E. Riddell, Chrys-Ellen M. Neville, Barbara L. Thomson, and Ziyang Zhang. Declines in Chinook salmon catches in the Strait of Georgia in relation to shifts in the marine environment. Fisheries Oceanography, 4(3):243–256, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Batten:2018:PSI

[BRO18] Sonia Dawn Batten, Gregory T. Ruggerone, and Ivonne Ortiz. Pink salmon induce a trophic cascade in plankton populations in the southern Bering Sea and around the Aleutian Islands. Fisheries Oceanography, 27(6):548–559, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bi:2008:SDO

[BRPC08] Hongsheng Bi, Rachel E. Ruppel, William T. Peterson, and Edmundo Casillas. Spatial distribution of ocean habitat of yearling Chinook (Oncorhynchus tshawytscha) and Coho (Oncorhynchus kisutch) salmon off Washington and Oregon, USA. Fisheries Oceanography, 17(6):463–476, November 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Byrd:2005:DPP

[BRR05] G. Vernon Byrd, Heather M. Renner, and Martin Renner. Distribution patterns and population trends of breeding seabirds in the Aleutian Islands. Fisheries Oceanography, 14(S1):139–159, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

BjORke:1994:TLJ

[BS94] Herman BjØRke and Roald Saetre. Transport of larvae and juvenile fish into central and northern Norwegian waters.

Fisheries Oceanography, 3(2):106–119, June 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bradbury:2001:ITA

[BSF01a] Ian R. Bradbury, Paul V. R. Snelgrove, and Sandra Fraser. The influence of temperature on advective loss of Atlantic cod (Gadus morhua) eggs from the inshore environment. Fisheries Oceanography, 10(4):342–352, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brickman:2001:MRS

[BSF01b] D. Brickman, N. L. Shackell, and K. T. Frank. Modelling the retention and survival of Browns Bank haddock larvae using an early life stage model. Fisheries Oceanography, 10(3):284– 296, September 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bentley:2020:RAI

[BSF^{+}20] Jacob W. Bentley, Natalia Serpetti, Clive J. Fox, Johanna J. Heymans, and David G. Reid. Retrospective analysis of the influence of environmental drivers on commercial stocks and fishing opportunities in the Irish Sea. Fisheries Oceanography, 29(5):415–435, September 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bauer:2013:WIV

[BSG^{+}13] Robert Klaus Bauer, Daniel Stepputtis, Ulf Gräwe, Christopher Zimmermann, and Cornelius Hammer. Wind-induced variability in coastal larval retention areas: a case study on Western Baltic spring-spawning herring. Fisheries Oceanography, 22(5):388–399, September 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Berntsen:1994:MTP

[BSS94] Jarle Berntsen, Dankert W. Skagen, and Einar Svendsen. Modelling the transport of particles in the North Sea with reference to sandeel larvae. Fisheries Oceanography, 3(2):81– 91, June 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brodeur:1999:SAC

[BT99] Richard D. Brodeur and Makoto Terazaki. Springtime abundance of chaetognaths in the shelf region of the northern Gulf of Alaska, with observations on the vertical distribution and feeding of Sagitta elegans. Fisheries Oceanography, 8(2):93– 103, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brickman:2007:OBM

[BTGM07] David Brickman, Lorna Taylor, Ásta Gudmundsdóttir, and Gudrun Marteinsdóttir. Optimized biophysical model for Icelandic cod (Gadus morhua) larvae. Fisheries Oceanography, 16(5):448–458, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Buckley:1992:LTV

[Buc92] L. J. Buckley. Long-term variability of pelagic fish populations and their environment. Fisheries Oceanography, 1(2): 190–191, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Borja:1998:RBA

[BUE^{+}98] Angel Borja, Andre’s Uriarte, Joseba Egaña, Lorenzo Motos, and Victor Valencia. Relationships between anchovy (Engraulis encrasicolus) recruitment and environment in the Bay of Biscay (1967–1996). Fisheries Oceanography, 7(3–4): 375–380, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Borja:2002:EFR

[BUE02] A. Borja, A. Uriarte, and J. Egaña. Environmental factors and recruitment of mackerel, Scomber scombrus L. 1758, along the north-east Atlantic coasts of Europe. Fisheries Oceanography, 11(2):116–127, March 2002. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brunel:2018:QIG

[BvDSDC18] Thomas Brunel, Cindy J. G. van Damme, Melvin Samson, and Mark Dickey-Collas. Quantifying the influence of geography and environment on the northeast Atlantic mackerel spawning distribution. Fisheries Oceanography, 27(2):

159–173, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Brodeur:1992:LTV

[BW92] Richard D. Brodeur and Daniel M. Ware. Long-term variability in zooplankton biomass in the subarctic Pacific Ocean. Fisheries Oceanography, 1(1):32–38, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Batten:2003:LDD

[BWJ03] Sonia D. Batten, David W. Welch, and Tanya Jonas. Latitudinal differences in the duration of development of Neocalanus plumchrus copepodites. Fisheries Oceanography, 12 (3):201–208, May 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Blanton:1999:WGT

[BWK^{+}99] Jackson O. Blanton, Francisco E. Werner, Andras Kapolnai, Brian O. Blanton, David Knott, and Elizabeth L. Wenner. Wind-generated transport of fictitious passive larvae into shallow tidal estuaries. Fisheries Oceanography, 8(S2): 210–223, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bishop:2015:PHC

[BWKM15] Mary Anne Bishop, Jordan T. Watson, Kathy Kuletz, and Tawna Morgan. Pacific herring (Clupea pallasii) consumption by marine birds during winter in Prince William Sound, Alaska. Fisheries Oceanography, 24(1):1–13, January 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bugaev:2001:IMA

[BWS^{+}01] V. F. Bugaev, D. W. Welch, M. M. Selifonov, L. E. Grachev, and J. P. Eveson. Influence of the marine abundance of pink (Oncorhynchus gorbuscha) and sockeye salmon (O. nerka) on growth of Ozernaya River sockeye. Fisheries Oceanography, 10(1):26–32, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bishop:2016:WEC

[BYM16] Janet Bishop, Yimin Ye, and David A. Milton. What are the effects of climate on the offshore and onshore movements of adult grooved tiger prawns Penaeus semisulcatus in tropical northern Australia? Fisheries Oceanography, 25(4):349–361, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Bograd:2021:TGA

[BZ21] Steven J. Bograd and Gayle Barbin Zydlewski. Training a generation of avid journal manuscript reviewers. Fisheries Oceanography, 30(1):1–2, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cooney:2001:ECJ

[CAB^{+}01] Robert T. Cooney, J. R. Allen, M. A. Bishop, D. L. Eslinger, T. Kline, B. L. Norcross, C. P. Mcroy, J. Milton, J. Olsen, V. Patrick, A. J. Paul, D. Salmon, D. Scheel, G. L. Thomas, S. L. Vaughan, and T. M. Willette. Ecosystem controls of juvenile pink salmon (Onchorynchus gorbuscha) and Pacific herring (Clupea pallasi) populations in Prince William Sound, Alaska. Fisheries Oceanography, 10(S1):1– 13, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Crespi-Abril:2012:RPS

[CAB12] Augusto C. Crespi-Abril and Pedro J. Barón. Revision of the population structuring of Illex argentinus (Castellanos, 1960) and a new interpretation based on modelling the spatio-temporal environmental suitability for spawning and nursery. Fisheries Oceanography, 21(2–3):199–214, March– May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Carbonell:2021:CDL

[CAGPC21] Aina Carbonell, Alberto Aparicio-González, Vanesa Papiol, and Joan Enric Cartes. Composition and distribution of the larval decapod community in the deep sea of the Western Mediterranean Sea Balearic Sub-basin. Fisheries Oceanography, 30(2):205–218, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Caputi:2008:ILC

[Cap08] Nick Caputi. Impact of the Leeuwin Current on the spatial distribution of the puerulus settlement of the western rock lobster (Panulirus cygnus) and implications for the fishery of Western Australia. Fisheries Oceanography, 17(2):147–152, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chittenden:2013:TMT

[CÅP^{+}13] Cedar M. Chittenden, BjØRn Ådlandsvik, Ole-Petter Pedersen, David Righton, and Audun H. Rikardsen. Testing a model to track fish migrations in polar regions using pop-up satellite archival tags. Fisheries Oceanography, 22(1):1–13, January 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

CrecHriou:2010:FRE

[CAR^{+}10] R. Crec’Hriou, F. Alemany, E. Roussel, A. Chassanite, J. Y. Marinaro, J. Mader, E. Rochel, and S. Planes. Fisheries replenishment of early life taxa: potential export of fish eggs and larvae from a temperate marine protected area. Fisheries Oceanography, 19(2):135–150, March 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Caputi:1993:EEP

[CB93] Nicolavito Caputi and Rhys S. Brown. The effect of environment on puerulus settlement of the western rock lobster (Panulius cygnus) in Western Australia. Fisheries Oceanography, 2(1):1–10, March 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coimbra:2023:RCC

[CBdSF^{+}23] Maria Raquel M. Coimbra, Emilly Benevides, Renata da Silva Farias, Bruno C. N. R. da Silva, Sara Cloux, Vicente Pérez-Muñuzuri, Manuel Vera, and Rodrigo Torres. Restricted connectivity for cobia Rachycentron canadum (Perciformes: Rachycentridae) in the Western Atlantic Ocean. Fisheries Oceanography, 32(6):495–508, November 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cass-Calay:2003:FEL

[CC03] Shannon L. Cass-Calay. The feeding ecology of larval Pacific hake (Merluccius productus) in the California Current region: an updated approach using a combined OPC/ MOCNESS to estimate prey biovolume. Fisheries Oceanography, 12(1):34–48, January 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Casey:2023:IOW

[CCC^{+}23] Flynn Casey, James H. Churchill, Geoffrey W. Cowles, Tracy L. Pugh, Richard A. Wahle, Kevin D. E. Stokesbury, and Robert P. Glenn. The impact of ocean warming on juvenile American lobster recruitment off southeastern Massachusetts. Fisheries Oceanography, 32(2):229–244, March 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Crear:2023:CIS

[CCHL23] Daniel P. Crear, Tobey H. Curtis, Cliff P. Hutt, and Yong-Woo Lee. Climate-influenced shifts in a highly migratory species recreational fishery. Fisheries Oceanography, 32(4): 327–340, July 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ciannelli:2022:IAP

[CCK^{+}22] Lorenzo Ciannelli, Antonio Cannavacciuolo, Peter Konstandinidis, Alice Mirasole, Jennifer A. T. K. Wong-Ala, Maria Teresa Guerra, Isabella D’Ambra, Emilio Riginella, and Daniela Cianelli. Ichthyoplankton assemblages and physical characteristics of two submarine canyons in the south central Tyrrhenian Sea. Fisheries Oceanography, 31(5):480– 496, September 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Causey:2005:PHM

[CCL^{+}05] Douglas Causey, Debra G. Corbett, Christine Lefèvre, Dixie L. West, Arkady B. Savinetsky, Nina K. Kiseleva, and Bulat F. Khassanov. The palaeoenvironment of humans and marine birds of the Aleutian Islands: three millennia of change. Fisheries Oceanography, 14(S1):259–276, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Carr:2008:IDV

[CCM^{+}08] Sarah D. Carr, Xavier J. Capet, James C. Mcwilliams, J. Timothy Pennington, and Francisco P. Chavez. The influence of diel vertical migration on zooplankton transport and recruitment in an upwelling region: estimates from a coupled behavioral-physical model. Fisheries Oceanography, 17 (1):1–15, January 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Curtis:2007:PVP

[CCP07] K. Alexandra Curtis, David M. Checkley, Jr., and Pierre Pepin. Predicting the vertical profiles of anchovy (Engraulis mordax) and sardine (Sardinops sagax) eggs in the California Current System. Fisheries Oceanography, 16(1):68–84, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cooney:2001:SSL

[CCSS01] R. T. Cooney, K. O. Coyle, E. Stockmar, and C. Stark. Seasonality in surface-layer net zooplankton communities in Prince William Sound, Alaska. Fisheries Oceanography, 10 (S1):97–109, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Castillo:2019:ADP

[CDG^{+}19] Ramiro Castillo, Luciano Dalla Rosa, Walter García Diaz, Lauro Madureira, Mariano Gutierrez, Luís Vásquez, and Rolf Koppelmann. Anchovy distribution off Peru in relation to abiotic parameters: a 32-year time series from 1985 to 2017. Fisheries Oceanography, 28(4):389–401, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coyle:2011:CCS

[CEM^{+}11] K. O. Coyle, L. B. Eisner, F. J. Mueter, A. I. Pinchuk, M. A. Janout, K. D. Cieciel, E. V. Farley, and A. G. Andrews. Climate change in the southeastern Bering Sea: impacts on pollock stocks and implications for the oscillating control hypothesis. Fisheries Oceanography, 20(2):139–156, March 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Churchill:1999:CLF

[CFL^{+}99] James H. Churchill, Richard B. Forward, Jr., Richard A. Luettich, Jr., James L. Hench, William F. Hettler, Larry B. Crowder, and Jackson O. Blanton. Circulation and larval fish transport within a tidally dominated estuary. Fisheries Oceanography, 8(S2):173–189, December 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

CostadeAraujo:2018:DPL

[CG18] Carolina Costa de Araujo and Maria A. Gasalla. Distribution patterns of loliginid squid paralarvae in relation to the oceanographic features off the South Brazil Bight (22^{o}– 25^{o}S). Fisheries Oceanography, 27(1):63–75, January 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chust:2019:EMD

[CGI^{+}19] Guillem Chust, Nerea Goikoetxea, Leire Ibaibarriaga, Yolanda Sagarminaga, Igor Arregui, Almudena Fontán, Xabier Irigoien, and Haritz Arrizabalaga. Earlier migration and distribution changes of albacore in the Northeast Atlantic. Fisheries Oceanography, 28(5):505–516, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Craig:2010:HUC

[CGMM10] J. K. Craig, P. C. Gillikin, M. A. Magelnicki, and L. N. May. Habitat use of cownose rays (Rhinoptera bonasus) in a highly productive, hypoxic continental shelf ecosystem. Fisheries Oceanography, 19(4):301–317, July 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cowan:1992:SDP

[CH92] James H. Cowan, Jr. and Edward D. Houde. Size-dependent predation on marine fish larvae by Ctenophores, Scyphomedusae, and Planktivorous fish. Fisheries Oceanography, 1 (2):113–126, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cochrane:1995:SAU

[CH95] K. L. Cochrane and L. Hutchings. A structured approach to using biological and environmental parameters to forecast anchovy recruitment. Fisheries Oceanography, 4(2):102–127, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Crowder:1999:ASR

[CH99] Larry B. Crowder and Donald E. Hoss. Applications of SABRE to research and management. Fisheries Oceanography, 8(S2):247–252, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Christian:2016:CAT

[CH16] James R. Christian and John Holmes. Changes in albacore tuna habitat in the northeast Pacific Ocean under anthropogenic warming. Fisheries Oceanography, 25(5):544–554, September 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Crossin:2004:SES

[CHF^{+}04] Glenn T. Crossin, Scott G. Hinch, Anthony P. Farrell, David A. Higgs, and Michael C. Healey. Somatic energy of sockeye salmon Oncorhynchus nerka at the onset of upriver migration: a comparison among ocean climate regimes. Fisheries Oceanography, 13(5):345–349, September 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chiba:2005:NSC

[CHHS05] S. Chiba, Y. Hirota, S. Hasegawa, and T. Saino. North– south contrasts in decadal scale variations in lower trophiclevel ecosystems in the Japan Sea. Fisheries Oceanography, 14(6):401–412, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Castonguay:1994:TRO

[CHM^{+}94] Martin Castonguay, Peter V. Hodson, Christopher Moriarty, Kenneth F. Drinkwater, and Brian M. Jessop. Is there a role of ocean environment in American and European eel decline? Fisheries Oceanography, 3(3):197–203, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coombs:1998:GSI

[CHPA98] Steve Coombs, Roger Harris, Ian Perry, and Juergen Alheit. GLOBEC special issue: Foreword. Fisheries Oceanography, 7(3–4):175, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Champion:2020:OHS

[CHPT20] Curtis Champion, Alistair J. Hobday, Gretta T. Pecl, and Sean R. Tracey. Oceanographic habitat suitability is positively correlated with the body condition of a coastal-pelagic fish. Fisheries Oceanography, 29(1):100–110, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coelho:2020:CGG

[CIS20] Rui Coelho, Paulo Infante, and Miguel N. Santos. Comparing GLM, GLMM, and GEE modeling approaches for catch rates of bycatch species: a case study of blue shark fisheries in the South Atlantic. Fisheries Oceanography, 29(2): 169–184, March 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Campana:2004:TDA

[CJ04] S. E. Campana and W. N. Joyce. Temperature and depth associations of porbeagle shark (Lamna nasus) in the northwest Atlantic. Fisheries Oceanography, 13(1):52–64, January 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Carlisle:2017:ITO

[CKA^{+}17] Aaron B. Carlisle, Randall E. Kochevar, Martin C. Arostegui, James E. Ganong, Michael Castleton, Jason Schratwieser, and Barbara A. Block. Influence of temperature and oxygen on the distribution of blue marlin (Makaira nigricans) in the Central Pacific. Fisheries Oceanography, 26 (1):34–48, January 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Call:2005:ECA

[CL05] Katherine A. Call and Thomas R. Loughlin. An ecological classification of Alaskan Steller sea lion (Eumetopias jubatus) rookeries: a tool for conservation/management. Fisheries Oceanography, 14(S1):212–222, November 2005. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cresci:2022:GWC

[CLH^{+}22] Alessandro Cresci, Torkel Larsen, Kim T. Halvorsen, Caroline M. F. Durif, Reidun Bjelland, Howard I. Browman, and Anne Berit Skiftesvik. Goldsinny wrasse (Ctenolabrus rupestris) have a sex-dependent magnetic compass for maintaining site fidelity. Fisheries Oceanography, 31(2):164–171, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Claro:2019:BCS

[CLKP19] Rodolfo Claro, Kenyon C. Lindeman, Andrew S. Kough, and Claire B. Paris. Biophysical connectivity of snapper spawning aggregations and marine protected area management alternatives in Cuba. Fisheries Oceanography, 28(1):33–42, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Canu:2021:NNA

[CLM^{+}21] Donata Melaku Canu, Célia Laurent, Elisabetta B. Morello, Stefano Querin, Giuseppe Scarcella, Nedo Vrgoc, Carlo Froglia, Silvia Angelini, and Cosimo Solidoro. Nephrops norvegicus in the Adriatic Sea: Connectivity modeling, essential fish habitats, and management area network. Fisheries Oceanography, 30(4):349–365, July 2021. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cartes:2018:CRA

[CLPC18] J. E. Cartes, C. López-Pérez, and A. Carbonell. Condition and recruitment of Aristeus antennatus at great depths (to 2,300 m) in the Mediterranean: Relationship with environmental factors. Fisheries Oceanography, 27(2):114–126, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chen:2005:DAT

[CLT05] I-Ching Chen, Pei-Fen Lee, and Wann-Nian Tzeng. Distribution of albacore (Thunnus alalunga) in the Indian Ocean and its relation to environmental factors. Fisheries Oceanography, 14(1):71–80, January 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chang:2019:MIE

[CLW^{+}19] Yi-Jay Chang, Kuo-Wei Lan, William A. Walsh, Jhen Hsu, and Chih hao Hsieh. Modelling the impacts of environmental variation on habitat suitability for Pacific saury in the Northwestern Pacific Ocean. Fisheries Oceanography, 28(3): 291–304, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chasse:2010:LLT

[CM10] Joël Chassé and Robert J. Miller. Lobster larval transport in the southern Gulf of St. Lawrence. Fisheries Oceanography, 19(5):319–338, September 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Carpi:2015:COM

[CMB^{+}15] Piera Carpi, Michela Martinelli, Andrea Belardinelli, Aniello Russo, Enrico Arneri, Alessandro Coluccelli, and Alberto Santojanni. Coupling an oceanographic model to a fishery observing system through mixed models: the importance of fronts for anchovy in the Adriatic Sea. Fisheries Oceanography, 24(6):521–532, November 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

CristieFranco:2006:MPP

[CMM06] Bárbara Cristie Franco, José Henrique Muelbert, and Mauricio Magalhães Mata. Mesoscale physical processes and the distribution and composition of ichthyoplankton on the southern Brazilian shelf break. Fisheries Oceanography, 15 (1):37–43, January 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chagaris:2015:TSA

[CMMK^{+}15] David Chagaris, Behzad Mahmoudi, Frank Muller-Karger, Wade Cooper, and Keith Fischer. Temporal and spatial availability of Atlantic thread herring, Opisthonema oglinum, in relation to oceanographic drivers and fishery landings on the Florida panhandle. Fisheries Oceanography, 24(3):257–273, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cryer:2016:NZE

[CMS16] Martin Cryer, Pamela M. Mace, and Kevin J. Sullivan. New Zealand’s ecosystem approach to fisheries management. Fisheries Oceanography, 25(S1):57–70, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cole:1999:ECS

[Col99] James Cole. Environmental conditions, satellite imagery, and clupeoid recruitment in the northern Benguela upwelling system. Fisheries Oceanography, 8(1):25–38, March 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cole:2000:CSS

[Col00] James Cole. Coastal sea surface temperature and Coho salmon production off the north-west United States. Fisheries Oceanography, 9(1):1–16, March 2000. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Checkley:1997:CUF

[COSC97] David M. Checkley, Jr., Peter B. Ortner, Lawrence R. Settle, and Shailer R. Cummings. A continuous, underway fish egg sampler. Fisheries Oceanography, 6(2):58–73, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Checkley:1999:SHA

[COW^{+}99] David M. Checkley, Jr., Peter B. Ortner, Francisco E. Werner, Lawrence R. Settle, and Shailer R. Cummings. Spawning habitat of the Atlantic menhaden in Onslow Bay, North Carolina. Fisheries Oceanography, 8(S2):22–36, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coyle:1994:R

[Coy94] K. O. Coyle. Reply. Fisheries Oceanography, 3(1):79, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coyle:2005:ZDA

[Coy05] Kenneth O. Coyle. Zooplankton distribution, abundance and biomass relative to water masses in eastern and central Aleutian Island passes. Fisheries Oceanography, 14(S1):77–92, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coyle:1992:IDP

[CP92] K. O. Coyle and A. J. Paul. Interannual differences in prey taken by capelin, herring, and red salmon relative to zooplankton abundance during the spring bloom in a southeast Alaskan embayment. Fisheries Oceanography, 1(4):294–305, December 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Coyle:2003:ACZ

[CP03] Kenneth O. Coyle and Alexei I. Pinchuk. Annual cycle of zooplankton abundance, biomass and production on the northern Gulf of Alaska shelf, October 1997 through October 2000. Fisheries Oceanography, 12(4–5):327–338, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cuttitta:2015:LPS

[CPM^{+}15] Angela Cuttitta, Bernardo Patti, Teresa Maggio, Enza Maria Quinci, Anna Maria Pappalardo, Venera Ferrito, Vito De Pinto, Marco Torri, Francesca Falco, Aldo Nicosia, Marianna Musco, Grazia Maria Armeri, Francesco Placenti, Giorgio Tranchida, Roberta Mifsud, Angelo Bonanno, and Salvatore Mazzola. Larval population structure of Engraulis encrasicolus in the Strait of Sicily as revealed by morphometric and genetic analysis. Fisheries Oceanography, 24(2): 135–149, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Churchill:2011:PCR

[CRC11] James H. Churchill, Jeffrey Runge, and Changsheng Chen. Processes controlling retention of spring-spawned Atlantic cod (Gadus morhua) in the western Gulf of Maine and their relationship to an index of recruitment success. Fisheries Oceanography, 20(1):32–46, January 2011. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Contreras:2017:GML

[CRVL^{+}17] Jorge E. Contreras, Camilo Rodriguez-Valentino, Mauricio F. Landaeta, Guido Plaza, Manuel I. Castillo, and Mónica Alvarado-Niño. Growth and mortality of larval anchoveta Engraulis ringens, in northern Chile during winter and their relationship with coastal hydrographic conditions. Fisheries Oceanography, 26(6):603–614, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cooper:2020:EDF

[CRW20] Daniel Cooper, Lauren A. Rogers, and Thomas Wilderbuer. Environmentally driven forecasts of northern rock sole (Lepidopsetta polyxystra) recruitment in the eastern Bering Sea. Fisheries Oceanography, 29(2):111–121, March 2020. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Chen:1994:RBT

[CSB94] Xiao Hong Chen, Peter A. Shelton, and Alistair J. Bath. Relationships between temperature and landing patterns in the southern Labrador and eastern Newfoundland inshore cod fishery. Fisheries Oceanography, 3(3):191–196, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cotton:2005:ECV

[CSFC05] Peter A. Cotton, David W. Sims, Sam Fanshawe, and Mark Chadwick. The effects of climate variability on zooplankton and basking shark (Cetorhinus maximus) relative abundance off southwest Britain. Fisheries Oceanography, 14(2): 151–155, March 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Childers:2011:MBJ

[CSK11] John Childers, Stephanie Snyder, and Suzanne Kohin. Migration and behavior of juvenile North Pacific albacore (Thunnus alalunga). Fisheries Oceanography, 20(3):157–173, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Cresci:2021:LCE

[CSS^{+}21] Alessandro Cresci, Anne D. Sandvik, Pål N. Sævik, Bjørn Ådlandsvik, Maria Josefina Olascoaga, Philippe Miron, Caroline M. F. Durif, Anne Berit Skiftesvik, Howard I. Browman, and Frode Vikebø. The lunar compass of European glass eels (Anguilla anguilla) increases the probability that they recruit to North Sea coasts. Fisheries Oceanography, 30 (3):315–330, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Callihan:2008:CSB

[CTWS08] Jody L. Callihan, Lynn T. Takata, Ryan J. Woodland, and David H. Secor. Cohort splitting in bluefish, Pomatomus saltatrix, in the US mid-Atlantic Bight. Fisheries Oceanography, 17(3):191–205, May 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Curtis:2004:FSS

[Cur04] K. Alexandra Curtis. Fine scale spatial pattern of Pacific sardine (Sardinops sagax) and northern anchovy (Engraulis mordax) eggs. Fisheries Oceanography, 13(4):239–254, July 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Carlson:1994:LE

[CW94] H. R. Carlson and A. C. Wertheimer. Letter to the editor. Fisheries Oceanography, 3(1):78–79, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Carlotti:1998:LEM

[CW98] FranÇOis Carlotti and Karl-Ulrich Wolf. A Lagrangian ensemble model of Calanus finmarchicus coupled with a 1D ecosystem model. Fisheries Oceanography, 7(3–4):191–204, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Colton:2014:ESD

[CWCM14] A. R. Colton, M. J. Wilberg, V. J. Coles, and T. J. Miller. An evaluation of the synchronization in the dynamics of blue crab (Callinectes sapidus) populations in the western Atlantic. Fisheries Oceanography, 23(2):132–146, March 2014.

CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

DAmours:1993:DCQ

[D’A93] Denis D’Amours. The distribution of cod (Gadus morhua) in relation to temperature and oxygen level in the Gulf of St. Lawrence. Fisheries Oceanography, 2(1):24–29, March 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Duffy-Anderson:2006:STP

[DABM^{+}06] J. T. Duffy-Anderson, M. S. Busby, K. L. Mier, C. M. Deliyanides, and P. J. Stabeno. Spatial and temporal patterns in summer ichthyoplankton assemblages on the eastern Bering Sea shelf 1996–2000. Fisheries Oceanography, 15(1): 80–94, January 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Davis:2023:CAO

[DAW^{+}23] Melanie J. Davis, James Anthony, Eric J. Ward, Julie Firman, and Christopher Lorion. Coherence among Oregon Coast Coho salmon populations highlights increasing relative importance of marine conditions for productivity. Fisheries Oceanography, 32(3):293–310, May 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dickson:1993:ECW

[DB93] R. R. Dickson and K. M. Brander. Effects of a changing windfield on cod stocks of the North Atlantic. Fisheries Oceanography, 2(3–4):124–153, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dutil:2003:CPN

[DB03] Jean-Denis Dutil and Keith Brander. Comparing productivity of North Atlantic cod (Gadus morhua) stocks and limits to growth production. Fisheries Oceanography, 12(4–5):502– 512, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Diankha:2018:COE

[DBB^{+}18] Ousmane Diankha, Aliou Ba, Patrice Brehmer, Timothée Brochier, Bamol Ali Sow, Modou Thiaw, Amadou Thierno Gaye, Fambaye Ngom, and Hervé Demarcq. Contrasted optimal environmental windows for both sardinella species in Senegalese waters. Fisheries Oceanography, 27(4):351–365, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

deBarros:2022:SRI

[dBdOJdO^{+}22] Matheus Souza Ferreira de Barros, José Gilmar Cavalcante de Oliveira Júnior, Cícero Diogo Lins de Oliveira, Taciana Kramer Pinto, Vandick da Silva Batista, and Nidia Noemi Fabré. Seasonal rainfall influences reproduction and recruitment of tropical penaeid shrimps: Implications to fisheries management. Fisheries Oceanography, 31(2): 191–204, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dagorn:2013:HMD

[DBFW13] Laurent Dagorn, Nicolas Bez, Thomas Fauvel, and Emily Walker. How much do fish aggregating devices (FADs) modify the floating object environment in the ocean? Fisheries Oceanography, 22(3):147–153, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dimmlich:2004:RIG

[DBGW04] W. F. Dimmlich, W. G. Breed, M. Geddes, and T. M. Ward. Relative importance of gulf and shelf waters for spawning and recruitment of Australian anchovy, Engraulis australis, in South Australia. Fisheries Oceanography, 13(5):310–323, September 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Downey-Breedt:2016:MTI

[DBRSC16] N. J. Downey-Breedt, M. J. Roberts, W. H. H. Sauer, and N. Chang. Modelling transport of inshore and deep-spawned chokka squid (Loligo reynaudi) paralarvae off South Africa: the potential contribution of deep spawning to recruitment. Fisheries Oceanography, 25(1):28–43, January 2016. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

DoSouto:2019:PPT

[DBS^{+}19] Marina Do Souto, Daniel R. Brown, Valeria Segura, Rubén Negri, Brenda Temperoni, Georgina Cepeda, Maria D.

Viñas, Fabiana L. Capitanio, and Marina V. Diaz. Putting the pieces together: Recent growth, nutritional condition, and mortality of Engraulis anchoita larvae in the Southwest Atlantic. Fisheries Oceanography, 28(5):597–611, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

DeCelles:2015:MTW

[DCLC15] Gregory DeCelles, Geoffrey Cowles, Chang Liu, and Steven Cadrin. Modeled transport of winter flounder larvae spawned in coastal waters of the Gulf of Maine. Fisheries Oceanography, 24(5):430–444, September 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Davidson:1995:MAC

[Dd95] Fraser J. M. Davidson and Brad deYoung. Modelling advection of cod eggs and larvae on the Newfoundland Shelf. Fisheries Oceanography, 4(1):33–51, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dale:2017:IVF

[DDB17] Katherine E. Dale, Elizabeth A. Daly, and Richard D. Brodeur. Interannual variability in the feeding and condition of subyearling Chinook salmon off Oregon and Washington in relation to fluctuating ocean conditions. Fisheries Oceanography, 26(1):1–16, January 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Diaz:2020:ERE

[DDB^{+}20] Marina V. Diaz, Marina Do Souto, Paola Betti, Brenda Temperoni, Agustín Schiariti, Laura Machinandiarena, Daniel R. Brown, and Gustavo J. Macchi. Evaluating the role of endogenous and exogenous features on larval hake nutritional condition. Fisheries Oceanography, 29(6):584–596, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dixon:2017:ADN

[DDS^{+}17] Heather J. Dixon, J. Brian Dempson, Timothy F. Sheehan, Mark D. Renkawitz, and Michael Power. Assessing the diet of North American Atlantic salmon (Salmo salar L.) off the West Greenland coast using gut content and stable isotope analyses. Fisheries Oceanography, 26(5):555–568, September 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dransfeld:2009:DPI

[DDZ09] Leonie Dransfeld, Oonagh Dwane, and Alain F. Zuur. Distribution patterns of ichthyoplankton communities in different ecosystems of the Northeast Atlantic. Fisheries Oceanography, 18(6):470–475, November 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dulcic:2000:CCA

[DG00] Dulčić and Grbec. Climate change and Adriatic ichthyofauna. Fisheries Oceanography, 9(2):187–191, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Domingues:2016:VPE

[DGB^{+}16] Ricardo Domingues, Gustavo Goni, Francis Bringas, Barbara Muhling, David Lindo-Atichati, and John Walter. Variability of preferred environmental conditions for Atlantic bluefin tuna (Thunnus thynnus) larvae in the Gulf of Mexico during 1993–2011. Fisheries Oceanography, 25(3):320– 336, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Devine:2011:REC

[DH11] Jennifer A. Devine and Richard L. Haedrich. The role of environmental conditions and exploitation in determining dynamics of redfish (Sebastes species) in the Northwest Atlantic. Fisheries Oceanography, 20(1):66–81, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dawe:2007:OCE

[DHC^{+}07] E. G. Dawe, L. C. Hendrickson, E. B. Colbourne, K. F. Drinkwater, and M. A. Showell. Ocean climate effects on the relative abundance of short-finned (Illex illecebrosus) and long-finned (Loligo pealeii) squid in the northwest Atlantic Ocean. Fisheries Oceanography, 16(4):303–316, July 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Drenner:2015:ECP

[DHM^{+}15] S. Matthew Drenner, Scott G. Hinch, Eduardo G. Martins, Nathan B. Furey, Timothy D. Clark, Steven J. Cooke, David A. Patterson, David Robichaud, David W. Welch, Anthony P. Farrell, and Richard E. Thomson. Environmental conditions and physiological state influence estuarine movements of homing sockeye salmon. Fisheries Oceanography, 24 (4):307–324, July 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Drinkwater:1996:TPF

[DHMT96] K. F. Drinkwater, G. C. Harding, K. H. Mann, and N. Tanner. Temperature as a possible factor in the increased abundance of American lobster, homarus americanust during the 1980s and early 1990s. Fisheries Oceanography, 5(3–4):176– 193, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Deimling:1994:FDE

[DL94] Elisabeth A. Deimling and William J. Liss. Fishery development in the eastern North Pacific: a natural-cultural system perspective, 1888–1976. Fisheries Oceanography, 3(1):60– 77, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dan:2022:LTS

[DLCQ22] Ya Ting Dan, Bi Lin Liu, Xin Jun Chen, and Wei Guo Qian. Long-term series variation in fishery biology of Dosidicus gigas off Peru: Response to climate variability from 2008 to 2020. Fisheries Oceanography, 31(6):629–640, November 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

DiStefano:2023:IST

[DLD^{+}23] Marine Di Stefano, Térence Legrand, Antonio Di Franco, David Nerini, and Vincent Rossi. Insights into the spatiotemporal variability of spawning in a territorial coastal fish by combining observations, modelling and literature review. Fisheries Oceanography, 32(1):70–90, January 2023. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dat:1995:CSH

[DLTI95] Claire G. Dat, Paul H. Leblond, Keith A. Thomson, and W. James Ingraham, Jr. Computer simulations of homeward-migrating Fraser River sockeye salmon: is compass orientation a sufficient direction-finding mechanism in the north-east Pacific Ocean? Fisheries Oceanography, 4(3): 209–216, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See comments [Sim96].

Dias:2017:BFW

[DMF^{+}17] Ester Dias, Pedro Morais, Ana M. Faria, Carlos Antunes, and Joel C. Hoffman. Benthic food webs support the production of sympatric flatfish larvae in estuarine nursery habitat. Fisheries Oceanography, 26(4):507–512, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dunn:2016:TBT

[DMH16] Daniel C. Dunn, Jerry H. Moxley, and Patrick N. Halpin. Temperature-based targeting in a multispecies fishery under climate change. Fisheries Oceanography, 25(2):105–118, March 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Domeier:2004:PLR

[Dom04] Michael L. Domeier. A potential larval recruitment pathway originating from a Florida marine protected area. Fisheries Oceanography, 13(5):287–294, September 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Domokos:2009:EEF

[Dom09] Réka Domokos. Environmental effects on forage and longline fishery performance for albacore (Thunnus alalunga) in the American Samoa Exclusive Economic Zone. Fisheries Oceanography, 18(6):419–438, November 2009. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Domokos:2023:INN

[Dom23] Réka Domokos. Influence of El Niño–Southern Oscillation on bigeye and yellowfin tuna longline catch per unit effort in the equatorial Pacific. Fisheries Oceanography, 32(6):527– 540, November 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dower:2001:HAL

[DP01] John F. Dower and R. Ian Perry. High abundance of larval rockfish over Cobb Seamount, an isolated seamount in the Northeast Pacific. Fisheries Oceanography, 10(3):268–274, September 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Daewel:2008:CEI

[DPK^{+}08] Ute Daewel, Myron A. Peck, Wilfried Kühn, Michael A. St. John, Irina Alekseeva, and Corinna Schrum. Coupling ecosystem and individual-based models to simulate the influence of environmental variability on potential growth and survival of larval sprat (Sprattus sprattus L.) in the North Sea. Fisheries Oceanography, 17(5):333–351, September 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dower:2002:UPS

[DPL02] John F. Dower, Pierre Pepin, and William C. Leggett. Using patch studies to link mesoscale patterns of feeding and growth in larval fish to environmental variability. Fisheries Oceanography, 11(4):219–232, July 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Daly:2020:RKC

[DPL^{+}20] Benjamin Daly, Carolina Parada, Timothy Loher, Sarah Hinckley, Albert J. Hermann, and David Armstrong. Red king crab larval advection in Bristol Bay: Implications for recruitment variability. Fisheries Oceanography, 29(6):505– 525, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dewar:2011:MBS

[DPM^{+}11] Heidi Dewar, Eric D. Prince, Michael K. Musyl, Richard W. Brill, Chugey Sepulveda, Jiangang Luo, David Foley, Eric S. Orbesen, Michael L. Domeier, Nicole Nasby-Lucas, Derke Snodgrass, R. Michael Laurs, John P. Hoolihan, Barbara A. Block, and Lianne M. Mcnaughton. Movements and behaviors of swordfish in the Atlantic and Pacific Oceans examined using pop-up satellite archival tags. Fisheries Oceanography, 20(3):219–241, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dragon:2018:MSP

[DSHL18] A.-C. Dragon, I. Senina, N. T. Hintzen, and P. Lehodey. Modelling South Pacific jack mackerel spatial population dynamics and fisheries. Fisheries Oceanography, 27(2):97–113, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Domokos:2007:OIA

[DSPH07] Réka Domokos, Michael P. Seki, Jeffrey J. Polovina, and Donald R. Hawn. Oceanographic investigation of the American Samoa albacore (Thunnus alalunga) habitat and longline fishing grounds. Fisheries Oceanography, 16(6):555–572, November 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Daewel:2011:TMC

[DST11] Ute Daewel, Corinna Schrum, and Axel Temming. Towards a more complete understanding of the life cycle of brown shrimp (Crangon crangon): modelling passive larvae and juvenile transport in combination with physically forced vertical juvenile migration. Fisheries Oceanography, 20(6):479– 496, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Drinkwater:2006:IWT

[DTC06] K. F. Drinkwater, M. J. Tremblay, and M. Comeau. The influence of wind and temperature on the catch rate of the American lobster (Homarus americanus) during spring fisheries off eastern Canada. Fisheries Oceanography, 15(2): 150–165, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

DeFilippo:2023:CDP

[DTO^{+}23] Lukas B. DeFilippo, James T. Thorson, Cecilia A. O’Leary, Stan Kotwicki, Jerry Hoff, James N. Ianelli, Vladimir V. Kulik, and Andre E. Punt. Characterizing dominant patterns of spatiotemporal variation for a transboundary groundfish assemblage. Fisheries Oceanography, 32(6):541–558, November 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dell:2011:EYT

[DWH11] James Dell, Chris Wilcox, and Alistair J. Hobday. Estimation of yellowfin tuna (Thunnus albacares) habitat in waters adjacent to Australia’s East Coast: making the most of commercial catch data. Fisheries Oceanography, 20(5):383–396, September 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Dambrine:2021:CEF

[DWHdP21] Chloé Dambrine, Mathieu Woillez, Martin Huret, and Hélène de Pontual. Characterising essential fish habitat using spatio-temporal analysis of fishery data: a case study of the European seabass spawning areas. Fisheries Oceanography, 30(4):413–428, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Eguchi:2017:POB

[EBFF17] Tomoharu Eguchi, Scott R. Benson, David G. Foley, and Karin A. Forney. Predicting overlap between drift gillnet fishing and leatherback turtle habitat in the California Current ecosystem. Fisheries Oceanography, 26(1):17–33, January 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Emmett:2004:VDJ

[EBO04] Robert L. Emmett, Richard D. Brodeur, and Philip M. Orton. The vertical distribution of juvenile salmon (Oncorhynchus spp.) and associated fishes in the Columbia River plume. Fisheries Oceanography, 13(6):392–402, November 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Eslinger:2001:PDO

[ECM^{+}01] David L. Eslinger, R. Ted Cooney, C. Peter Mcroy, Alison Ward, Thomas C. Kline, Jr., E. Paul Simpson, Jia Wang, and Jennifer R. Allen. Plankton dynamics: observed and modelled responses to physical conditions in Prince William Sound, Alaska. Fisheries Oceanography, 10(S1):81–96, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ehrhardt:2010:DRP

[EF10] Nelson M. Ehrhardt and Mark D. Fitchett. Dependence of recruitment on parent stock of the spiny lobster, Panulirus argus, in Florida. Fisheries Oceanography, 19(6):434–447, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Edwards:2008:DBS

[EHW08] K. P. Edwards, J. A. Hare, and F. E. Werner. Dispersal of black sea bass (Centropristis striata) larvae on the southeast U.S. continental shelf: results of a coupled vertical larval behavior — 3D circulation model. Fisheries Oceanography, 17(4):299–315, July 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Escalle:2016:EFM

[EPG^{+}16] Lauriane Escalle, Maria Grazia Pennino, Daniel Gaertner, Pierre Chavance, Alicia Delgado de Molina, Hervé Demarcq, Evgeny Romanov, and Bastien Merigot. Environmental factors and megafauna spatio-temporal co-occurrence with purse-seine fisheries. Fisheries Oceanography, 25(4):433–447, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Eggleston:2010:TSE

[ERE^{+}10] David B. Eggleston, Nathalie B. Reyns, Lisa L. Etherington, Gayle R. Plaia, and Lian Xie. Tropical storm and environmental forcing on regional blue crab (Callinectes sapidus) settlement. Fisheries Oceanography, 19(2):89–106, March 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Erisman:2021:RAS

[ERR^{+}21] Brad E. Erisman, Erin M. Reed, Martha J. Román, Ismael Mascareñas-Osorio, Peter van der Sleen, Catalina López-Sagástegui, Octavio Aburto-Oropeza, Kirsten Rowell, and Bryan A. Black. Relationships among somatic growth, climate, and fisheries production in an overexploited marine fish from the Gulf of California, Mexico. Fisheries Oceanography, 30(5):556–568, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Erzini:2005:TAL

[Erz05] Karim Erzini. Trends in NE Atlantic landings (southern Portugal): identifying the relative importance of fisheries and environmental variables. Fisheries Oceanography, 14(3):195– 209, May 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ehrich:2009:LSP

[ESA09] Siegfried Ehrich, Vanessa Stelzenmüller, and Sara Adlerstein. Linking spatial pattern of bottom fish assemblages with water masses in the North Sea. Fisheries Oceanography, 18(1):36–50, January 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Elliott:2016:BCD

[ESA^{+}16] Meredith L. Elliott, Annie E. Schmidt, Sara Acosta, Russell Bradley, Pete Warzybok, Keith M. Sakuma, John C. Field, and Jaime Jahncke. Brandt’s cormorant diet (1994–2012) indicates the importance of fall ocean conditions for northern anchovy in central California. Fisheries Oceanography, 25(5): 515–528, September 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Escribano:1998:PDC

[Esc98] Ruben Escribano. Population dynamics of Calanus chilensis in the Chilean Eastern Boundary Humboldt Current. Fisheries Oceanography, 7(3–4):245–251, December 1998. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

emond:2020:LTT

[éSMB20] Kim émond, Bernard Sainte-Marie, and Joël Bêty. Longterm trends and drivers of larval phenology and abundance of dominant brachyuran crabs in the Gulf of St. Lawrence (Canada). Fisheries Oceanography, 29(2):185–200, March 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Edvardsen:2003:AZA

[ESTJ03] A. Edvardsen, D. Slagstad, K. S. Tande, and P. Jaccard. Assessing zooplankton advection in the Barents Sea using underway measurements and modelling. Fisheries Oceanography, 12(2):61–74, March 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Emsley:2005:EVM

[ETB05] S. M. Emsley, G. A. Tarling, and M. T. Burrows. The effect of vertical migration strategy on retention and dispersion in the Irish Sea during spring–summer. Fisheries Oceanography, 14(3):161–174, May 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Espinasse:2017:MRI

[ETB^{+}17] Boris Espinasse, Vigdis Tverberg, Sünnje L. Basedow, Tore Hattermann, Ole Anders Nøst, Jon Albretsen, Jofrid Skardhamar, and Ketil Eiane. Mechanisms regulating inter-annual variability in zooplankton advection over the Lofoten shelf, implications for cod larvae survival. Fisheries Oceanography, 26(3):299–315, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Everett:2017:DEK

[EvST^{+}17] Jason D. Everett, Erik van Sebille, Matthew D. Taylor, Iain M. Suthers, Christopher Setio, Paulina Cetina-Heredia, and James A. Smith. Dispersal of Eastern King Prawn larvae in a western boundary current: New insights from particle tracking. Fisheries Oceanography, 26(5):513–525, September 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Friedenberg:2012:FDL

[FBRB12] Laura E. Friedenberg, Stephen M. Bollens, and Gretchen Rollwagen-Bollens. Feeding dynamics of larval Pacific herring (Clupea pallasi) on natural prey assemblages: the importance of protists. Fisheries Oceanography, 21(2–3):95– 108, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Frank:2004:POS

[FC04] Kenneth T. Frank and Jae S. Choi. In a perfect ocean: the state of fisheries and ecosystems in the North Atlantic Ocean. Fisheries Oceanography, 13(2):141–143, March 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Feng:2019:ICS

[FCC^{+}19] Yongjiu Feng, Li Cui, Xinjun Chen, Lijuan Chen, and Qianqian Yang. Impacts of changing spatial scales on CPUE– factor relationships of Ommastrephes bartramii in the northwest Pacific. Fisheries Oceanography, 28(2):143–158, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Feekings:2015:USS

[FCJ^{+}15] Jordan Feekings, Asbjørn Christensen, Patrik Jonsson, Rikke Frandsen, Mats Ulmestrand, Sten Munch-Petersen, and Bo Andersen. The use of at-sea-sampling data to dissociate environmental variability in Norway lobster (Nephrops norvegicus) catches to improve resource exploitation efficiency within the Skagerrak/Kattegat trawl fishery. Fisheries Oceanography, 24(4):383–392, July 2015. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Frank:1993:CST

[FCL93] Kenneth T. Frank, James E. Carscadden, and William C. Leggett. Causes of spatio-temporal variation in the patchiness of larval fish distributions: differential mortality or behaviour? Fisheries Oceanography, 2(3–4):114–123, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Forward:1999:BSP

[FDT^{+}99] Richard B. Forward, Jr., M. C. De Vries, R. A. Tankersley, D.. Rittschof, William F. Hettler, J. S. Burke, J. M. Welch, and Donald E. Hoss. Behaviour and sensory physiology of Atlantic menhaden larvae, Brevoortia tyrannus, during horizontal transport. Fisheries Oceanography, 8(S2):37–56, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fujioka:2018:HUM

[FFF^{+}18] Ko Fujioka, Hiromu Fukuda, Seishiro Furukawa, Yaoki Tei, Suguru Okamoto, and Seiji Ohshimo. Habitat use and movement patterns of small (age-0) juvenile Pacific bluefin tuna (Thunnus orientalis) relative to the Kuroshio. Fisheries Oceanography, 27(3):185–198, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Franco-Gordo:2008:ISV

[FGGDSMF08] C. Franco-Gordo, E. Godínez-Domínguez, E. Suárez-Morales, and J. Freire. Interannual and seasonal variability of the diversity and structure of ichthyoplankton assemblages in the central Mexican Pacific. Fisheries Oceanography, 17 (3):178–190, May 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fiksen:1995:SEF

[FGS95] Ø. Fiksen, J. Giske, and D. Slagstad. A spatially explicit fitness-based model of capelin migrations the Barents Sea. Fisheries Oceanography, 4(3):193–208, September 1995. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Francis:1994:DSR

[FH94] Robert C. Francis and Steven R. Hare. Decadal-scale regime shifts in the large marine ecosystems of the North-east Pacific: a case for historical science. Fisheries Oceanography, 3 (4):279–291, December 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Friedland:1998:MTE

[FHD98] Kevin D. Friedland, Lars P. Hansen, and David A. Dunkley. Marine temperatures experienced by postsmolts and the survival of Atlantic salmon, Salmo salar L., in the North Sea area. Fisheries Oceanography, 7(1):22–34, April 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Francis:1998:EIC

[FHHW98] Robert C. Francis, Steven R. Hare, Anne B. Hollowed, and Warren S. Wooster. Effects of interdecadal climate variability on the oceanic ecosystems of the NE Pacific. Fisheries Oceanography, 7(1):1–21, April 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fujioka:2010:IVS

[FHK^{+}10] Ko Fujioka, Alistair J. Hobday, Ryo Kawabe, Kazushi Miyashita, Kentaro Honda, Tomoyuki Itoh, and Yoshimi Takao. Interannual variation in summer habitat utilization by juvenile southern bluefin tuna (Thunnus maccoyii) in southern Western Australia. Fisheries Oceanography, 19(3): 183–195, May 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fujioka:2012:DBJ

[FHK^{+}12] Ko Fujioka, Alistair J. Hobday, Ryo Kawabe, Kazushi Miyashita, Yoshimi Takao, Osamu Sakai, and Tomoyuki Itoh. Departure behaviour of juvenile southern bluefin tuna (Thunnus maccoyii) from southern Western Australia temperate waters in relation to the Leeuwin Current. Fisheries Oceanography, 21(4):269–280, July 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Faure:2000:IRP

[FIDC00] Valerie Faure, Cheikh Abdellahi Inejih, Herve Demarcq, and Philippe Cury. The importance of retention processes in upwelling areas for recruitment of Octopus vulgaris: the example of the Arguin Bank (Mauritania). Fisheries Oceanography, 9(4):343–355, December 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Franco:2022:IWD

[FKF^{+}22] Ana C. Franco, Hongsik Kim, Hartmut Frenzel, Curtis Deutsch, Debby Ianson, U. Rashid Sumaila, and Philippe D. Tortell. Impact of warming and deoxygenation on the habitat distribution of Pacific halibut in the Northeast Pacific. Fisheries Oceanography, 31(6):601–614, November 2022. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fukuta:2017:OCE

[FKH^{+}17] Atsushi Fukuta, Yasuhiro Kamimura, Masakazu Hori, Masahiro Nakaoka, Tsutomu Noda, Yoh Yamashita, Tsuguo Otake, and Jun Shoji. Offshore currents explain the discontinuity of a fish community in the seagrass bed along the Japanese archipelago. Fisheries Oceanography, 26(1):65–68, January 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fuji:2021:ESG

[FKSA21] Taiki Fuji, Yutaka Kurita, Satoshi Suyama, and Daisuke Ambe. Estimating the spawning ground of Pacific saury Cololabis saira by using the distribution and geographical variation in maturation status of adult fish during the main spawning season. Fisheries Oceanography, 30(4):382–396, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fuji:2016:IEN

[FKUY16] Taiki Fuji, Akihide Kasai, Masahiro Ueno, and Yoh Yamashita. Importance of estuarine nursery areas for the adult population of the temperate seabass Lateolabrax japonicus, as revealed by otolith Sr : Ca ratios. Fisheries Oceanography, 25(4):448–456, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fossum:1993:SSA

[FM93] P. Fossum and E. Moksness. A study of spring-and autumnspawned herring (Clupea harengus L.) larvae in the Norwegian Coastal Current during spring 1990. Fisheries Oceanography, 2(2):73–81, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Friedland:2022:MAB

[FMG^{+}22] Kevin D. Friedland, Travis Miles, Andrew G. Goode, Eric N. Powell, and Damian C. Brady. The Middle Atlantic Bight Cold Pool is warming and shrinking: Indices from in situ autumn seafloor temperatures. Fisheries Oceanography, 31 (2):217–223, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Feenstra:2014:EID

[FML^{+}14] John Feenstra, Richard McGarvey, Adrian Linnane, André E. Punt, and Nigel Bean. Environmental influences on daily commercial catch rates of South Australia’s southern rock lobster (Jasus edwardsii). Fisheries Oceanography, 23(4): 362–374, July 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Friedland:2020:TCP

[FMM^{+}20] Kevin D. Friedland, Ryan E. Morse, James P. Manning, Donald Christopher Melrose, Travis Miles, Andrew G. Goode, Damian C. Brady, Josh T. Kohut, and Eric N. Powell. Trends and change points in surface and bottom thermal environments of the US Northeast Continental Shelf Ecosystem. Fisheries Oceanography, 29(5):396–414, September 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Freon:2003:IRS

[FMV03] P. Fréon, C. Mullon, and B. Voisin. Investigating remote synchronous patterns in fisheries. Fisheries Oceanography, 12(4–5):443–457, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fox:2006:IIB

[FMYN06] C. J. Fox, P. Mccloghrie, E. F. Young, and R. D. M. Nash. The importance of individual behaviour for successful settlement of juvenile plaice (Pleuronectes platessa L.): a modelling and field study in the eastern Irish Sea. Fisheries Oceanography, 15(4):301–313, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fox:2000:PSC

[FODCN00] Fox, O’Brien, Dickey-Collas, and Nash. Patterns in the spawning of cod (Gadus morhua L.), sole (Solea solea L.) and plaice (Pleuronectes platessa L.) in the Irish Sea as determined by generalized additive modelling. Fisheries Oceanography, 9(1):33–49, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fassler:2011:DLM

[FPBDC11] Sascha M. M. Fässler, Mark R. Payne, Thomas Brunel, and Mark Dickey-Collas. Does larval mortality influence population dynamics? An analysis of North Sea herring (Clupea harengus) time series. Fisheries Oceanography, 20(6):530– 543, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Frank:2013:SRS

[FPFL13] Kenneth T. Frank, Brian Petrie, Jonathan A. D. Fisher, and William C. Leggett. Setting the record straight on drivers of changing ecosystem states. Fisheries Oceanography, 22(2): 143–146, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Francis:1993:DWT

[Fra93] Malcolm P. Francis. Does water temperature determine year class strength in New Zealand snapper (Pagrus auratus, Sparidae)? Fisheries Oceanography, 2(2):65–72, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fromentin:2014:OCE

[FRBB14] Jean-Marc Fromentin, Gabriel Reygondeau, Sylvain Bonhommeau, and Gregory Beaugrand. Oceanographic changes and exploitation drive the spatio-temporal dynamics of Atlantic bluefin tuna (Thunnus thynnus). Fisheries Oceanography, 23(2):147–156, March 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Funes-Rodriguez:2006:INN

[FRHMAM^{+}06] René Funes-Rodríguez, Alejandro Hinojosa-Medina, Gerardo Aceves-Medina, Sylvia P. A. Jiménez-Rosenberg, and J. Jesús Bautista-Romero. Influences of El Niño on assemblages of mesopelagic fish larvae along the Pacific coast of Baja California Sur. Fisheries Oceanography, 15(3):244–255, May 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Forward:1999:TFL

[FRP^{+}99] Richard B. Forward, Jr., K. A. Reinsel, D. S. Peters, R. A. Tankersley, James H. Churchill, Larry B. Crowder, William F. Hettler, Stanley M. Warlen, and M. D. Green. Transport of fish larvae through a tidal inlet. Fisheries Oceanography, 8(S2):153–172, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fadely:2005:ISS

[FRS^{+}05] Brian S. Fadely, Bruce W. Robson, Jeremy T. Sterling, Angie Greig, and Katherine A. Call. Immature Steller sea lion (Eumetopias jubatus) dive activity in relation to habitat features of the eastern Aleutian Islands. Fisheries Oceanography, 14(S1):243–258, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Funes-Rodriguez:2011:MFL

[FRZVHM^{+}11] René Funes-Rodríguez, Alejandro Zárate-Villafranco, Alejandro Hinojosa-Medina, Rogelio González-Armas, and Sergio Hernández-Trujillo. Mesopelagic fish larval assemblages during El Niño-southern oscillation (1997–2001) in the southern part of the California Current. Fisheries Oceanography, 20(4):329–346, July 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fiksen:1998:MIL

[FUA^{+}98] Ø. Fiksen, A. C. W. Utne, D. L. Aksnes, K. Eiane, J. V. Helvik, and S. Sundby. Modelling the influence of light, turbulence and ontogeny on ingestion rates in larval cod and herring. Fisheries Oceanography, 7(3–4):355–363, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Funamoto:2007:TDS

[Fun07] Tetsuichiro Funamoto. Temperature-dependent stockrecruitment model for walleye pollock (Theragra chalcogramma) around northern Japan. Fisheries Oceanography, 16(6):515– 525, November 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Funamoto:2011:CWP

[Fun11] Tetsuichiro Funamoto. Causes of walleye pollock (Theragra chalcogramma) recruitment decline in the northern Sea of Japan: implications for stock management. Fisheries Oceanography, 20(2):95–103, March 2011. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Frusher:2016:PFF

[FvPH^{+}16] Stewart Frusher, Ingrid van Putten, Marcus Haward, Alistair J. Hobday, Neil J. Holbrook, Sarah Jennings, Nadine Marshall, Sarah Metcalf, Gretta T. Pecl, and Malcolm Tull. From physics to fish to folk: supporting coastal regional communities to understand their vulnerability to climate change in Australia. Fisheries Oceanography, 25(S1):19–28, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Frost:2021:DDE

[FYA^{+}21] Tessa J. Frost, Ellen M. Yasumiishi, Beverly A. Agler, Milo D. Adkison, and Megan V. McPhee. Density-dependent effects of eastern Kamchatka pink salmon (Oncorhynchus gorbuscha) and Japanese chum salmon (O. keta) on agespecific growth of western Alaska chum salmon. Fisheries Oceanography, 30(1):99–109, January 2021. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Feng:2022:CHF

[FYC22] Zhiping Feng, Wei Yu, and Xinjun Chen. Concurrent habitat fluctuations of two economically important marine species in the Southeast Pacific Ocean off Chile in relation to ENSO perturbations. Fisheries Oceanography, 31(1):123–134, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Funamoto:2013:ABF

[FYK^{+}13] Tetsuichiro Funamoto, Orio Yamamura, Tokihiro Kono, Tomonori Hamatsu, and Akira Nishimura. Abiotic and biotic factors affecting recruitment variability of walleye pollock (Theragra chalcogramma) off the Pacific coast of Hokkaido, Japan. Fisheries Oceanography, 22(3):193–206, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fujita:2021:TVH

[FYK^{+}21] Tatsunori Fujita, Masayuki Yamamoto, Naoaki Kono, Takeshi Tomiyama, Koichi Sugimatsu, and Michio Yoneda. Temporal variations in hatch date and early survival of Japanese anchovy (Engraulis japonicus) in response to environmental factors in the central Seto Inland Sea, Japan. Fisheries Oceanography, 30(5):527–541, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Fuentes-Yaco:2007:SPB

[FYKSP07] C. Fuentes-Yaco, P. A. Koeller, S. Sathyendranath, and T. Platt. Shrimp (Pandalus borealis) growth and timing of the spring phytoplankton bloom on the Newfoundland– Labrador Shelf. Fisheries Oceanography, 16(2):116–129, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Griffiths:2019:JFE

[GAH^{+}19] Shane P. Griffiths, Valerie Allain, Simon D. Hoyle, Tim A. Lawson, and Simon J. Nicol. Just a FAD? Ecosystem impacts of tuna purse-seine fishing associated with fish aggregating devices in the western Pacific Warm Pool Province. Fisheries Oceanography, 28(1):94–112, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gargett:1997:OSW

[Gar97] Ann E. Gargett. The optimal stability ‘window’: a mechanism underlying decadal fluctuations in North Pacific salmon stocks? Fisheries Oceanography, 6(2):109–117, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Greer:2017:ABL

[GBAD^{+}17] Adam T. Greer, Christian Briseño-Avena, Alison L. Deary, Robert K. Cowen, Frank J. Hernandez, and William M. Graham. Associations between lobster phyllosoma and gelatinous zooplankton in relation to oceanographic properties in the northern Gulf of Mexico. Fisheries Oceanography, 26(6): 693–704, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Govinden:2021:BSK

[GCF^{+}21] Rodney Govinden, Manuela Capello, Fabien Forget, John D. Filmalter, and Laurent Dagorn. Behavior of skipjack (Katsuwonus pelamis), yellowfin (Thunnus albacares), and bigeye (T. obsesus) tunas associated with drifting fish aggregating devices (dFADs) in the Indian Ocean, assessed through acoustic telemetry. Fisheries Oceanography, 30(5):542–555, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Garcia:2013:CIE

[GCQ^{+}13] A. García, D. Cortés, J. Quintanilla, T. Rámirez, L. Quintanilla, J. M. Rodríguez, and F. Alemany. Climate-induced environmental conditions influencing interannual variability of Mediterranean bluefin (Thunnus thynnus) larval growth. Fisheries Oceanography, 22(4):273–287, July 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Guan:2017:EST

[GCW17] Lisha Guan, Yong Chen, and James A. Wilson. Evaluating spatio-temporal variability in the habitat quality of Atlantic cod (Gadus morhua) in the Gulf of Maine. Fisheries Oceanography, 26(1):83–96, January 2017. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Guan:2017:IVA

[GDM^{+}17] Lu Guan, John F. Dower, Skip M. McKinnell, Pierre Pepin, Evgeny A. Pakhomov, and Brian P. V. Hunt. Interannual variability in the abundance and composition of spring larval fish assemblages in the Strait of Georgia (British Columbia, Canada) from 2007 to 2010. Fisheries Oceanography, 26(6): 638–654, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gutierrez-Estrada:2017:IPD

[GEGHPCC17] Juan Carlos Gutiérrez-Estrada, Juan Gil-Herrera, Inmaculada Pulido-Calvo, and Ivone Alejandra Czerwinski. Is it possible to differentiate between environmental and fishery effects on abundance-biomass variation? A case study of blackspot seabream (Pagellus bogaraveo) in the Strait of Gibraltar. Fisheries Oceanography, 26(4):455–475, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gao:1998:MAC

[GFG98] Huiwang Gao, Shizuo Feng, and Yuping Guan. Modelling annual cycles of primary production in different regions of the Bohai Sea. Fisheries Oceanography, 7(3–4):258–264, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Guraslan:2014:MIC

[GFO14] Ceren Güraslan, Bettina A. Fach, and Temel Oguz. Modeling the impact of climate variability on Black Sea anchovy recruitment and production. Fisheries Oceanography, 23(5): 436–457, September 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gargano:2017:ECB

[GGF17] Francesco Gargano, Germana Garofalo, and Fabio Fiorentino. Exploring connectivity between spawning and nursery areas of Mullus barbatus (L., 1758) in the Mediterranean through a dispersal model. Fisheries Oceanography, 26(4):476–497, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gargano:2022:WDR

[GGQF22] Francesco Gargano, Germana Garofalo, Federico Quattrocchi, and Fabio Fiorentino. Where do recruits come from? Backward Lagrangian simulation for the deep water rose shrimps in the Central Mediterranean Sea. Fisheries Oceanography, 31(4):369–383, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gallego:1999:VGR

[GHBM99] A. Gallego, M. R. Heath, D. J. Basford, and Brian R. Mackenzie. Variability in growth rates of larval haddock in the northern North Sea. Fisheries Oceanography, 8(2):77– 92, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Giri:2019:RLP

[GHG^{+}19] Sandip Giri, Sugata Hazra, Pramit Ghosh, Amit Ghosh, Sourav Das, Abhra Chanda, Isha Das, Kunal Chakraborty, Anirban Mukhopadhyay, and Sourav Maity. Role of lunar phases, rainfall, and wind in predicting Hilsa shad (Tenualosa ilisha) catch in the northern Bay of Bengal. Fisheries Oceanography, 28(5):567–575, September 2019. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Grieve:2021:MIC

[GHM21] Brian D. Grieve, Jonathan A. Hare, and W. David McElroy. Modeling the impacts of climate change on thorny skate (Amblyraja radiata) on the Northeast US shelf using trawl and longline surveys. Fisheries Oceanography, 30(3):300– 314, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gomes:1995:STC

[GHV95] Manuel C. Gomes, Richard L. Haedrich, and M. Guadalupe Villagarcia. Spatial and temporal changes in the groundfish assemblages on the north-east Newfoundland/Labrador Shelf, north-west Atlantic, 1978–1991. Fisheries Oceanography, 4(2):85–101, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Goikoetxea:2013:LBR

[GI13] Nerea Goikoetxea and Xabier Irigoien. Links between the recruitment success of northern European hake (Merluccius merluccius L.) and a regime shift on the NE Atlantic continental shelf. Fisheries Oceanography, 22(6):459–476, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Guo:2020:MMP

[GiIW^{+}20] Chenying Guo, Shin ichi Ito, Nicholas C. Wegner, Laura N. Frank, Emmanis Dorval, Kathryn A. Dickson, and Dane H. Klinger. Metabolic measurements and parameter estimations for bioenergetics modelling of Pacific chub mackerel Scomber japonicus. Fisheries Oceanography, 29(3):215–226, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Giannoulaki:2013:CPH

[GIT^{+}13] M. Giannoulaki, M. Iglesias, M. P. Tugores, A. Bonanno, B. Patti, A. De Felice, I. Leonori, J. L. Bigot, V. TiČIna, M. M. Pyrounaki, K. Tsagarakis, A. Machias, S. Somarakis, E. Schismenou, E. Quinci, G. Basilone, A. Cuttitta, F. Campanella, J. Miquel, D. Oñate, D. Roos, and V. Valavanis. Characterizing the potential habitat of European anchovy Engraulis encrasicolus in the Mediterranean Sea, at different life stages. Fisheries Oceanography, 22(2):69–89, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Goetz:2018:WGM

[GJR18] Frederick W. Goetz, Andrew J. Jasonowicz, and Steven B. Roberts. What goes up must come down: Diel vertical migration in the deep-water sablefish (Anoplopoma fimbria) revealed by pop-up satellite archival tags. Fisheries Oceanography, 27(2):127–142, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Glaser:2011:DAE

[Gla11] Sarah M. Glaser. Do albacore exert top-down pressure on northern anchovy? Estimating anchovy mortality as a result of predation by juvenile North Pacific albacore in the California current system. Fisheries Oceanography, 20(3):

242–257, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gallego:1999:MSM

[GMH^{+}99] A. Gallego, J. Mardaljevic, M. R. Heath, D. Hainbucher, and D. Slagstad. A model of the spring migration into the North Sea by Calanus finmarchicus overwintering off the Scottish continental shelf. Fisheries Oceanography, 8(S1):107–125, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gomez:2012:ISU

[GMH^{+}12] Fabian Gomez, Aldo Montecinos, Samuel Hormazabal, Luis A. Cubillos, Marco Correa-Ramirez, and Francisco P. Chavez. Impact of spring upwelling variability off southerncentral Chile on common sardine (Strangomera bentincki) recruitment. Fisheries Oceanography, 21(6):405–414, November 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Garces:2019:AER

[GNP^{+}19] Cristóbal Garcés, Edwin J. Niklitschek, Guido Plaza, Francisco Cerna, Mathieu Leisen, Pamela Toledo, and Fernando Barra. Anchoveta Engraulis ringens along the Chilean coast: Management units, demographic units and water masses: Insights from multiple otolith-based approaches. Fisheries Oceanography, 28(6):735–750, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gucu:2016:RCS

[GöEIOS16] Ali Cemal Gucu, özgür Emek Inanmaz, Meltem Ok, and Serdar Sakinan. Recent changes in the spawning grounds of Black Sea anchovy, Engraulis encrasicolus. Fisheries Oceanography, 25(1):67–84, January 2016. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Govoni:1994:EVL

[GP94] John J. Govoni and Leonard J. Pietrafesa. Eulerian views of layered water currents, vertical distribution of some larval fishes, and inferred advective transport over the continental shelf off North Carolina, USA, in winter. Fisheries Oceanography, 3(2):120–132, June 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Garcia:2021:ADN

[GPA^{+}21] Théo Garcia, Benjamin Planque, Per Arneberg, Bjarte Bogstad, Øystein Skagseth, and Maik Tiedemann. An appraisal of the drivers of Norwegian spring-spawning herring (Clupea harengus) recruitment. Fisheries Oceanography, 30 (2):159–173, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Godinez-Padilla:2022:MCI

[GPCGdlT^{+}22] Carlos Javier Godínez-Padilla, José Leonardo Castillo-Géniz, Benigno Hernández de la Torre, Luis Vicente González-Ania, and Marlon H. Román-Verdesoto. Marineclimate interactions with the blue shark (Prionace glauca) catches in the western coast of Baja California Peninsula, Mexico. Fisheries Oceanography, 31(3):291–318, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Giannoulaki:2011:HSM

[GPL^{+}11] Marianna Giannoulaki, Maria M. Pyrounaki, Bernard Liorzou, Iole Leonori, Vasilis D. Valavanis, Konstantinos Tsagarakis, Jean L. Bigot, David Roos, Andrea De Felice, Fabio Campanella, Stylianos Somarakis, Enrico Arneri, and Athanassios Machias. Habitat suitability modelling for sardine juveniles (Sardina pilchardus) in the Mediterranean Sea. Fisheries Oceanography, 20(5):367–382, September 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Goldberg:2022:AER

[GPS22] Danielle A. Goldberg, Lee M. Paramore, and Frederick S. Scharf. Analysis of environment-recruitment associations for a coastal red drum population reveals consistent link between year class strength and early shifts in nearshore winds. Fisheries Oceanography, 31(1):56–69, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gonzalez-Quiros:2004:IMP

[GQPGA04] Rafael González-Quirós, Ananda Pascual, Damiá Gomis, and Ricardo Anadón. Influence of mesoscale physical forcing on trophic pathways and fish larvae retention in the central Cantabrian Sea. Fisheries Oceanography, 13(6):351–364, November 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gallienne:1998:TOC

[GR98] Christopher P. Gallienne and David B. Robins. Transoceanic characterization of zooplankton community size structure using an optical plankton counter. Fisheries Oceanography, 7(2):147–158, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Graham:1998:BR

[Gra98] Jeffrey B. Graham. Book review. Fisheries Oceanography, 7 (2):171–173, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Greene:1999:BRF

[Gre99] Charles H. Greene. Book review: Fundamentals of Acoustical Oceanography, by Herman Medwin and Clarence S. Clay, Academic Press, San Diego. 1994. 712 pp. ISBN 0-12-487570-X. $US 75 (hard cover). Fisheries Oceanography, 8(2):157– 158, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Greene:2013:TMB

[Gre13] Charles H. Greene. Towards a more balanced view of marine ecosystems. Fisheries Oceanography, 22(2):140–142, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Groger:2007:GAC

[GRT^{+}07] J. P. Gröger, R. A. Rountree, U. H. Thygesen, D. Jones, D. Martins, Q. Xu, and B. J. Rothschild. Geolocation of Atlantic cod (Gadus morhua) movements in the Gulf of Maine using tidal information. Fisheries Oceanography, 16(4):317– 335, July 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gauldie:1996:SVD

[GS96] R. W. Gauldie and G. D. Sharp. Skipjack velocity, dwell time and migration. Fisheries Oceanography, 5(2):100–113, June 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Govoni:1999:EFL

[GS99] Govoni and Spach. Exchange and flux of larval fishes across the western Gulf Stream front south of Cape Hatteras, USA, in winter. Fisheries Oceanography, 8(S2):77–92, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gutierrez:2007:AER

[GSBB07] Mariano Gutiérrez, Gordon Swartzman, Arnaud Bertrand, and Sophie Bertrand. Anchovy (Engraulis ringens) and sardine (Sardinops sagax) spatial dynamics and aggregation patterns in the Humboldt Current ecosystem, Peru, from 1983–2003. Fisheries Oceanography, 16(2):155–168, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

GonzalezArmas:1999:CSA

[GSNFL99] Rogelio González Armas, Oscar Sosa-Nishizaki, René Funes Rodríguez, and Victor Andrés Levy Pérez. Confirmation of the spawning area of the striped marlin, Tetrapturus audax, in the so-called core area of the eastern tropical Pacific off Mexico. Fisheries Oceanography, 8(3):238–242, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gaardsted:2010:MCA

[GTB10] Frank Gaardsted, Kurt S. Tande, and Sünnje L. Basedow. Measuring copepod abundance in deep-water winter habitats in the NE Norwegian Sea: intercomparison of results from laser optical plankton counter and multinet. Fisheries Oceanography, 19(6):480–492, November 2010. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Gay:2001:SHT

[GV01] Shelton. M. Gay III and S. L. Vaughan. Seasonal hydrography and tidal currents of bays and fjords in Prince William Sound, Alaska. Fisheries Oceanography, 10(S1):159–193, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Guisande:2004:CCA

[GVRC04] C. Guisande, A. R. Vergara, I. Riveiro, and J. M. Cabanas. Climate change and abundance of the Atlantic–Iberian sardine (Sardina pilchardus). Fisheries Oceanography, 13(2): 91–101, March 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Glaser:2014:NLD

[GYS14] Sarah M. Glaser, Hao Ye, and George Sugihara. A nonlinear, low data requirement model for producing spatially explicit fishery forecasts. Fisheries Oceanography, 23(1):45–53, January 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hare:2007:MLB

[HA07] Jonathan A. Hare and Kenneth W. Able. Mechanistic links between climate and fisheries along the east coast of the United States: explaining population outbursts of Atlantic croaker (Micropogonias undulatus). Fisheries Oceanography, 16(1):31–45, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Humston:2000:SML

[HALO00] Robert Humston, Jerald S. Ault, Molly Lutcavage, and Donald B. Olson. Schooling and migration of large pelagic fishes relative to environmental cues. Fisheries Oceanography, 9 (2):136–146, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hannah:2011:VDO

[Han11] Robert W. Hannah. Variation in the distribution of ocean shrimp (Pandalus jordani) recruits: links with coastal upwelling and climate change. Fisheries Oceanography, 20(4): 305–313, July 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Harrison:1992:DME

[Har92] P. J. Harrison. Dynamics of marine ecosystems: Biological– physical interactions in the oceans. Fisheries Oceanography, 1(1):111–112, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Holsman:2019:CEB

[HAS^{+}19] Kirstin K. Holsman, Kerim Aydin, Jane Sullivan, Tom Hurst, and Gordon H. Kruse. Climate effects and bottom-up controls on growth and size-at-age of Pacific halibut (Hippoglossus stenolepis) in Alaska (USA). Fisheries Oceanography, 28(3):345–358, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hsieh:1992:GCC

[HB92] William W. Hsieh and George J. Boer. Global climate change and ocean upwelling. Fisheries Oceanography, 1(4):333–338, December 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hainbucher:1999:CEN

[HB99] D. Hainbucher and J. O. Backhaus. Circulation of the eastern North Atlantic and north-west European continental shelf — a hydrodynamic modelling study. Fisheries Oceanography, 8 (S1):1–12, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinrichsen:2007:IBM

[HBC07] H.-H. Hinrichsen, V. Buehler, and C. Clemmesen. An individual-based model for the direct conversion of otolith into somatic growth rates. Fisheries Oceanography, 16(3): 207–215, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Huret:2016:SPB

[HBG^{+}16] M. Huret, P. Bourriau, P. Gatti, F. Dumas, and P. Petitgas. Size, permeability and buoyancy of anchovy (Engraulis encrasicolus) and sardine (Sardina pilchardus) eggs in relation to their physical environment in the Bay of Biscay. Fisheries Oceanography, 25(6):582–597, November 2016. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Haulsee:2022:WDB

[HBLC22] Danielle E. Haulsee, Hannah E. Blondin, Ryan K. Logan, and Larry B. Crowder. Where do the billfish go? Using recreational catch data to relate local and basin scale environmental conditions to billfish occurrence in the Eastern Tropical Pacific. Fisheries Oceanography, 31(2):135–148, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Holmes:2021:ILF

[HBN^{+}21] Elizabeth Eli Holmes, Smitha Br, Kumar Nimit, Sourav Maity, David M. Checkley, Jr., Mark L. Wells, and Vera L. Trainer. Improving landings forecasts using environmental covariates: a case study on the Indian oil sardine (Sardinella longiceps). Fisheries Oceanography, 30(6):623–642, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinrichsen:2001:PAE

[HBO^{+}01] H.-H. Hinrichsen, U. Böttcher, R. Oeberst, R. Voss, and A. Lehmann. The potential for advective exchange of the early life stages between the western and eastern Baltic cod (Gadus morhua L.) stocks. Fisheries Oceanography, 10(3): 249–258, September 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hare:2015:REE

[HBPC15] Jonathan A. Hare, Elizabeth N. Brooks, Michael C. Palmer, and James H. Churchill. Re-evaluating the effect of wind on recruitment in Gulf of Maine Atlantic cod (Gadus morhua) using an environmentally-explicit stock recruitment model. Fisheries Oceanography, 24(1):90–105, January 2015. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1999:CFS

[HBR^{+}99] M. R. Heath, J. O. Backhaus, K. Richardson, E. McKenzie, D. Slagstad, D. Beare, J. Dunn, J. G. Fraser, A. Gallego, D. Hainbucher, S. Hay, S. Jónasdóttir, H. Madden, J. Mardaljevic, and A. Schacht. Climate fluctuations and the spring invasion of the North Sea by Calanus finmarchicus. Fisheries Oceanography, 8(S1):163–176, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hsu:2015:TSC

[HBR^{+}15] Ango C. Hsu, André M. Boustany, Jason J. Roberts, Jui-Han Chang, and Patrick N. Halpin. Tuna and swordfish catch in the U.S. northwest Atlantic longline fishery in relation to mesoscale eddies. Fisheries Oceanography, 24(6):508–520, November 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hsieh:2009:TSA

[HCC^{+}09] Chih-Hao Hsieh, Chih-Shin Chen, Tai-Sheng Chiu, Kuo-Tien Lee, Feng-Jen Shieh, Jia-Yi Pan, and Ming-An Lee. Time series analyses reveal transient relationships between abundance of larval anchovy and environmental variables in the coastal waters southwest of Taiwan. Fisheries Oceanography, 18(2):102–117, March 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Husson:2020:SHF

[HCFP20] Bérengère Husson, Gregoire Certain, Anatoly Filin, and Benjamin Planque. Suitable habitats of fish species in the Barents Sea. Fisheries Oceanography, 29(6):526–540, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hurst:2009:EOT

[HCS^{+}09] Thomas P. Hurst, Daniel W. Cooper, Joel S. Scheingross, Erin M. Seale, Benjamin J. Laurel, and Mara L. Spencer. Effects of ontogeny, temperature, and light on vertical movements of larval Pacific cod (Gadus macrocephalus). Fisheries Oceanography, 18(5):301–311, September 2009. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Howard:2021:ECO

[HCWF21] Rebecca A. Howard, Lorenzo Ciannelli, W. Waldo Wakefield, and Melanie R. Fewings. The effects of climate, oceanography, and habitat on the distribution and abundance of northern California Current continental shelf groundfishes. Fisheries Oceanography, 30(6):707–725, November 2021. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1999:FCO

[HDF^{+}99] M. R. Heath, J. Dunn, J. G. Fraser, S. J. Hay, and H. Madden. Field calibration of the optical plankton counter with respect to Calanus finmarchicus. Fisheries Oceanography, 8 (S1):13–24, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Harding:2005:LLH

[HDH^{+}05] G. C. Harding, K. F. Drinkwater, C. G. Hannah, J. D. Pringle, J. Prena, J. W. Loder, S. Pearre, Jr., and W. P. Vass. Larval lobster (Homarus americanus) distribution and drift in the vicinity of the Gulf of Maine offshore banks and their probable origins. Fisheries Oceanography, 14(2): 112–137, March 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hughes:2015:CSI

[HDJ15] K. M. Hughes, L. Dransfeld, and M. P. Johnson. Climate and stock influences on the spread and locations of catches in the northeast Atlantic mackerel fishery. Fisheries Oceanography, 24(6):540–552, November 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Healey:1993:OEN

[Hea93] M. C. Healey. Ocean ecology of North Pacific salmonids. Fisheries Oceanography, 2(1):41–42, 1993, March 1993. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1999:MHM

[Hea99a] M. Heath. In memoriam — Heather Madden 1971–97. Fisheries Oceanography, 8(S1):v, September 1999. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1999:AMC

[Hea99b] M. R. Heath. The ascent migration of Calanus finmarchicus from overwintering depths in the Faroe–Shetland Channel. Fisheries Oceanography, 8(S1):84–99, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1999:I

[Hea99c] M. R. Heath. Introduction. Fisheries Oceanography, 8(S1): vii–viii, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heino:2008:MHD

[HEG08] Mikko Heino, Georg H. Engelhard, and Olav Rune GodØ. Migrations and hydrography determine the abundance fluctuations of blue whiting (Micromesistius poutassou) in the Barents Sea. Fisheries Oceanography, 17(2):153–163, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hare:2001:SIS

[HFC01] Jonathan A. Hare, Michael P. Fahay, and Robert K. Cowen. Springtime ichthyoplankton of the slope region off the northeastern United States of America: larval assemblages, relation to hydrography and implications for larval transport. Fisheries Oceanography, 10(2):164–192, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hiraoka:2019:IVD

[HFF^{+}19] Yuko Hiraoka, Ko Fujioka, Hiromu Fukuda, Mikio Watai, and Seiji Ohshimo. Interannual variation of the diet shifts and their effects on the fatness and growth of age-0 Pacific bluefin tuna (Thunnus orientalis) off the southwestern Pacific coast of Japan. Fisheries Oceanography, 28(4):419–433, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Henderson:2019:SVO

[HFHW19] Mark Henderson, Jerome Fiechter, David D. Huff, and Brian K. Wells. Spatial variability in ocean-mediated growth potential is linked to Chinook salmon survival. Fisheries Oceanography, 28(3):334–344, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1998:BPM

[HG98] Heath and A. Gallego. Bio-physical modelling of the early life stages of haddock, Melanogrammus aeglefinus, in the North Sea. Fisheries Oceanography, 7(2):110–125, July 1998. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinojosa:2017:DED

[HGG^{+}17] Ivan A. Hinojosa, Caleb Gardner, Bridget S. Green, Andrew Jeffs, Rafael Leon, and Adrian Linnane. Differing environmental drivers of settlement across the range of southern rock lobster (Jasus edwardsii) suggest resilience of the fishery to climate change. Fisheries Oceanography, 26(1):49–64, January 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Holligan:1993:WCD

[HGH93] Patrick M. Holligan, Stephen B. Groom, and Derek S. Harbour. What controls the distribution of the coccolithophore, Emiliania huxleyi, in the North Sea? Fisheries Oceanography, 2(3–4):175–183, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Han:2021:UPD

[HGS^{+}21] Qingpeng Han, Arnaud Grüss, Xiujuan Shan, Xianshi Jin, and James T. Thorson. Understanding patterns of distribution shifts and range expansion/contraction for small yellow croaker (Larimichthys polyactis) in the Yellow Sea. Fisheries Oceanography, 30(1):69–84, January 2021. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hirota:1999:ZBS

[HH99] Yuichi Hirota and Seizo Hasegawa. The zooplankton biomass in the Sea of Japan. Fisheries Oceanography, 8(4):274–283, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Howell:2015:ETP

[HHB^{+}15] Evan A. Howell, Aimee Hoover, Scott R. Benson, Helen Bailey, Jeffrey J. Polovina, Jeffrey A. Seminoff, and Peter H. Dutton. Enhancing the TurtleWatch product for leatherback sea turtles, a dynamic habitat model for ecosystem-based management. Fisheries Oceanography, 24(1):57–68, January 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hernandez:2009:EDO

[HHF09] Frank J. Hernandez, Jonathan A. Hare, and Dariusz P. Fey. Evaluating diel, ontogenetic and environmental effects on larval fish vertical distribution using generalized additive models for location, scale and shape. Fisheries Oceanography, 18 (4):224–236, July 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hassrick:2016:EOD

[HHH^{+}16] Jason L. Hassrick, Mark J. Henderson, David D. Huff, William J. Sydeman, Megan C. Sabal, Jeffrey A. Harding, Arnold J. Ammann, Eric D. Crandall, Eric P. Bjorkstedt, John Carlos Garza, and Sean A. Hayes. Early ocean distribution of juvenile Chinook salmon in an upwelling ecosystem. Fisheries Oceanography, 25(2):133–146, March 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hu:2018:HAC

[HHH^{+}18] Chen Hu, Daniel P. Harrison, Michael G. Hinton, Zachary C. Siegrist, and Dale A. Kiefer. Habitat analysis of the commercial tuna of the Eastern tropical Pacific Ocean. Fisheries Oceanography, 27(5):417–434, September 2018. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Honda:2010:ADD

[HHK^{+}10] Kentaro Honda, Alistair J. Hobday, Ryo Kawabe, Naoki Tojo, Ko Fujioka, Yoshimi Takao, and Kazushi Miyashita. Age-dependent distribution of juvenile southern bluefin tuna (Thunnus maccoyii) on the continental shelf off southwest Australia determined by acoustic monitoring. Fisheries Oceanography, 19(2):151–158, March 2010. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Harrison:2017:PHA

[HHK^{+}17] Daniel P. Harrison, Michael G. Hinton, Suzanne Kohin, Edward M. Armstrong, Stephanie Snyder, Frank O’Brien, and Dale K. Kiefer. The pelagic habitat analysis module for ecosystem-based fisheries science and management. Fisheries Oceanography, 26(3):316–335, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hobday:2010:DSZ

[HHTF10] Alistair J. Hobday, Jason R. Hartog, Trent Timmiss, and Josh Fielding. Dynamic spatial zoning to manage southern bluefin tuna (Thunnus maccoyii) capture in a multi-species longline fishery. Fisheries Oceanography, 19(3):243–253, May 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1999:DAO

[HJ99] M. R. Heath and S. H. Jónasdóttir. Distribution and abundance of overwintering Calanus finmarchicus in the Faroe– Shetland Channel. Fisheries Oceanography, 8(S1):40–60, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hazen:2010:MPD

[HJ10] Elliott L. Hazen and David W. Johnston. Meridional patterns in the deep scattering layers and top predator distribution in the central equatorial Pacific. Fisheries Oceanography, 19(6):427–433, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hao:2003:TFC

[HJR^{+}03] Wei Hao, Su Jian, Wan Ruijing, Wang Lei, and Lin Yi’an. Tidal front and the convergence of anchovy (Engraulis japonicus) eggs in the Yellow Sea. Fisheries Oceanography, 12 (4–5):434–442, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Howell:2006:NNE

[HK06] Evan A. Howell and Donald R. Kobayashi. El Niño effects in the Palmyra Atoll region: oceanographic changes and bigeye tuna (Thunnus obesus) catch rate variability. Fisheries Oceanography, 15(6):477–489, November 2006. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hyrenbach:2006:UMS

[HKA^{+}06] K. D. Hyrenbach, C. Keiper, S. G. Allen, D. G. Ainley, and D. J. Anderson. Use of marine sanctuaries by far-ranging predators: commuting flights to the California Current System by breeding Hawaiian albatrosses. Fisheries Oceanography, 15(2):95–103, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Horodysky:2007:HUV

[HKLG07] Andrij Z. Horodysky, David W. Kerstetter, Robert J. Latour, and John E. Graves. Habitat utilization and vertical movements of white marlin (Tetrapturus albidus) released from commercial and recreational fishing gears in the western North Atlantic Ocean: inferences from short duration pop-up archival satellite tags. Fisheries Oceanography, 16 (3):240–256, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hino:2019:CVT

[HKM^{+}19] Haruhiko Hino, Takashi Kitagawa, Takayuki Matsumoto, Yoshinori Aoki, and Shingo Kimura. Changes to vertical thermoregulatory movements of juvenile bigeye tuna (Thunnus obesus) in the northwestern Pacific Ocean with time of day, seasonal ocean vertical thermal structure, and body size. Fisheries Oceanography, 28(4):359–371, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hino:2021:DBP

[HKM^{+}21] Haruhiko Hino, Takashi Kitagawa, Takayuki Matsumoto, Yoshinori Aoki, and Shingo Kimura. Development of behavioral and physiological thermoregulatory mechanisms with body size in juvenile bigeye tuna Thunnus obesus. Fisheries Oceanography, 30(3):219–231, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hidaka:2003:BTC

[HKT^{+}03] Kiyotaka Hidaka, Kouichi Kawaguchi, Toshiyuki Tanabe, Mio Takahashi, and Tsunemi Kubodera. Biomass and taxonomic composition of micronekton in the western tropical– subtropical Pacific. Fisheries Oceanography, 12(2):112–125, March 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Harford:2017:NPM

[HKWL17] William J. Harford, Mandy Karnauskas, John F. Walter, and Hui Liu. Non-parametric modeling reveals environmental effects on bluefin tuna recruitment in Atlantic, Pacific, and Southern Oceans. Fisheries Oceanography, 26(4):396–412, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hernandez-Leon:1998:ACE

[HL98] Santiago Hernández-León. Annual cycle of epiplanktonic copepods in Canary Island waters. Fisheries Oceanography, 7(3–4):252–257, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:2007:SLS

[HL07] M. R. Heath and R. G. Lough. A synthesis of large-scale patterns in the planktonic prey of larval and juvenile cod (Gadus morhua). Fisheries Oceanography, 16(2):169–185, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hoolihan:2011:VHU

[HLG^{+}11] John P. Hoolihan, Jiangang Luo, C. P. Goodyear, Eric S. Orbesen, and Eric D. Prince. Vertical habitat use of sailfish (Istiophorus platypterus) in the Atlantic and eastern Pacific, derived from pop-up satellite archival tag data. Fisheries Oceanography, 20(3):192–205, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hurst:2017:ECD

[HLH^{+}17] Thomas P. Hurst, Benjamin J. Laurel, Eric Hanneman, Scott A. Haines, and Michele L. Ottmar. Elevated CO_{2} does not exacerbate nutritional stress in larvae of a Pacific flatfish. Fisheries Oceanography, 26(3):336–349, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinrichsen:2003:DLF

[HLMS03] H.-H. Hinrichsen, A. Lehmann, C. Möllmann, and J. O. Schmidt. Dependency of larval fish survival on retention/ dispersion in food limited environments: the Baltic Sea as a case study. Fisheries Oceanography, 12(4–5):425–433, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hsieh:2012:LFA

[HLWL12] Hung-Yen Hsieh, Wen-Tseng Lo, Long-Jing Wu, and Dong-Chung Liu. Larval fish assemblages in the Taiwan Strait, western North Pacific: linking with monsoon-driven mesoscale current system. Fisheries Oceanography, 21(2–3): 125–147, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hansen:2001:RBS

[HMM01] J. E. Hansen, P. Martos, and A. Madirolas. Relationship between spatial distribution of the Patagonian stock of Argentine anchovy, Engraulis anchoita, and sea temperatures during late spring to early summer. Fisheries Oceanography, 10(2):193–206, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Helbig:1992:EIR

[HMP92] J. Helbig, G. Mertz, and P. Pepin. Environmental influences on the recruitment of Newfoundland/Labrador cod. Fisheries Oceanography, 1(1):39–56, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Horimoto:2016:SAB

[HMS16] Takanori Horimoto, Yoko Mitani, and Yasunori Sakurai. Spatial association between northern fur seal (Callorhinus ursinus) and potential prey distribution during the wintering period in the northern Sea of Japan. Fisheries Oceanography, 25(1):44–53, January 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinchliffe:2023:PSS

[HMS^{+}23] Charles Hinchliffe, Paloma A. Matis, Hayden T. Schilling, Jason D. Everett, Anthony G. Miskiewicz, Pierre Pepin, Daniel S. Falster, and Iain M. Suthers. Plankton size spectra as an indicator of larval success in Pacific sardine (Sardinops sagax). Fisheries Oceanography, 32(2):196–212, March 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hyun:2007:YYV

[HMT07] Saang-Yoon Hyun, Katherine W. Myers, and André Talbot. Year-to-year variability in ocean recovery rate of Columbia River Upriver Bright fall Chinook salmon (Oncorhynchus tshawytscha). Fisheries Oceanography, 16(4):350–362, July 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hearn:2005:PDE

[HMTG^{+}05] Alex Hearn, Priscilla Martínez, M. Veronica Toral-Granda, Juan Carlos Murillo, and Jeffrey Polovina. Population dynamics of the exploited sea cucumber Isostichopus fuscus in the western Galápagos Islands, Ecuador. Fisheries Oceanography, 14(5):377–385, September 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinckley:2009:SPM

[HNHP09] S. Hinckley, J. M. Napp, A. J. Hermann, and C. Parada. Simulation of physically mediated variability in prey resources of a larval fish: a three-dimensional NPZ model. Fisheries Oceanography, 18(4):201–223, July 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Honda:2004:MJW

[HONH04] Satoshi Honda, Tatsuki Oshima, Akira Nishimura, and Tsutomu Hattori. Movement of juvenile walleye pollock, Theragra chalcogramma, from a spawning ground to a nursery ground along the Pacific coast of Hokkaido, Japan. Fisheries Oceanography, 13(S1):84–98, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Horne:2000:AAR

[Hor00] John K. Horne. Acoustic approaches to remote species identification: a review. Fisheries Oceanography, 9(4):356–371, December 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Helbig:2002:ESS

[HP02] J. A. Helbig and P. Pepin. The effects of short space and time scale current variability on the predictability of passive ichthyoplankton distributions: an analysis based on HF radar observations. Fisheries Oceanography, 11(3):175–188, May 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hernandez:2020:SAD

[HPG^{+}20] Alejandra Hernández, Guido Plaza, Jairo Gutiérrez, Francisco Cerna, and Edwin J. Niklitschek. Spatiotemporal analysis of the daily growth traits of the prerecruits of a small pelagic fish in response to environmental drivers. Fisheries Oceanography, 29(6):457–469, November 2020. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hollowed:2013:PMF

[HPL13] Anne Babcock Hollowed, Benjamin Planque, and Harald Loeng. Potential movement of fish and shellfish stocks from the sub-Arctic to the Arctic Ocean. Fisheries Oceanography, 22(5):355–370, September 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hodgson:2006:MFC

[HQH^{+}06] Sayre Hodgson, Thomas P. Quinn, Ray Hilborn, Robert C. Francis, and Donald E. Rogers. Marine and freshwater climatic factors affecting interannual variation in the timing of return migration to fresh water of sockeye salmon (Oncorhynchus nerka). Fisheries Oceanography, 15(1):1–24, January 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hare:1999:LTD

[HQW^{+}99] Jonathan A. Hare, John A. Quinlan, Francisco E. Werner, Brian O. Blanton, John J. Govoni, Richard B. Forward, Jr., Lawrence R. Settle, and Donald E. Hoss. Larval transport during winter in the SABRE study area: results of a coupled vertical larval behaviour–three-dimensional circulation model. Fisheries Oceanography, 8(S2):57–76, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heard:2018:PDV

[HRB^{+}18] Matthew Heard, Paul J. Rogers, Barry D. Bruce, Nicolas E. Humphries, and Charlie Huveneers. Plasticity in the diel vertical movement of two pelagic predators (Prionace glauca and Alopias vulpinus) in the southeastern Indian Ocean. Fisheries Oceanography, 27(3):199–211, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hold:2021:UBM

[HRS^{+}21] Natalie Hold, Peter Robins, Claire L. Szostek, Gwladys Lambert, Harriet Lincoln, Lewis Le Vay, Ewen Bell, and Michel J. Kaiser. Using biophysical modelling and population genetics for conservation and management of an exploited species, Pecten maximus L. Fisheries Oceanography, 30(6):740–756, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hunt:2005:OEA

[HS05] George L. Hunt, Jr. and Phyllis J. Stabeno. Oceanography and ecology of the Aleutian Archipelago: spatial and temporal variation. Fisheries Oceanography, 14(S1):292–306, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hobday:2016:SFD

[HSEH16] Alistair J. Hobday, Claire M. Spillman, J. Paige Eveson, and Jason R. Hartog. Seasonal forecasting for decision support in marine fisheries and aquaculture. Fisheries Oceanography, 25 (S1):45–56, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hewitt:2022:MOF

[HSH^{+}22] Daniel E. Hewitt, Hayden T. Schilling, Roshan Hanamseth, Jason D. Everett, Junde Li, Moninya Roughan, Daniel D. Johnson, Iain M. Suthers, and Matthew D. Taylor. Mesoscale oceanographic features drive divergent patterns in connectivity for co-occurring estuarine portunid crabs. Fisheries Oceanography, 31(6):587–600, November 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hernandez-Santoro:2019:EED

[HSLP19] Carola Hernández-Santoro, Mauricio F. Landaeta, and Jorge Castillo Pizarro. Effect of ENSO on the distribution and concentration of catches and reproductive activity of anchovy Engraulis ringens in northern Chile. Fisheries Oceanography, 28(3):241–255, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hunerlage:2019:RRB

[HSS19] Kim Hünerlage, Volker Siegel, and Reinhard Saborowski. Reproduction and recruitment of the brown shrimp Crangon crangon in the inner German Bight (North Sea): an interannual study and critical reappraisal. Fisheries Oceanography, 28(6):708–722, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hernandez-Trujillo:1999:VCS

[HT99] Sergio Hernandez-Trujillo. Variability of community structure of Copepoda related to El Niño 1982–83 and 1987–88 along the west coast of Baja California Peninsula, Mexico. Fisheries Oceanography, 8(4):284–295, December 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hale:2018:VDL

[HT18] Edward A. Hale and Timothy E. Targett. Vertical distribution of larval Atlantic menhaden (Brevoortia tyrannus) and Atlantic croaker (Micropogonias undulatus): Implications for vertical migratory behaviour and transport. Fisheries Oceanography, 27(3):222–231, May 2018. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Halvorsen:2003:HSO

[HTE^{+}03] Elisabeth Halvorsen, Kurt S. Tande, Are Edvardsen, Dag Slagstad, and Ole Petter Pedersen. Habitat selection of overwintering Calanus finmarchicus in the NE Norwegian Sea and shelf waters off Northern Norway in 2000–02. Fisheries Oceanography, 12(4–5):339–351, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Healey:2000:CSE

[HTL^{+}00] Michael C. Healey, Keith A. Thomson, Paul H. Leblond, Leonardo Huato, Scott G. Hinch, and Carl J. Walters. Computer simulations of the effects of the Sitka eddy on the migration of sockeye salmon returning to British Columbia. Fisheries Oceanography, 9(3):271–281, September 2000. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Haltuch:2020:ODP

[HTLJ20] Melissa A. Haltuch, Nick Tolimieri, Qi Lee, and Michael G. Jacox. Oceanographic drivers of petrale sole recruitment in the California Current ecosystem. Fisheries Oceanography, 29(2):122–136, March 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hufnagl:2014:MLT

[HTP14] Marc Hufnagl, Axel Temming, and Thomas Pohlmann. The missing link: tidal-influenced activity a likely candidate to close the migration triangle in brown shrimp Crangon crangon (Crustacea, Decapoda). Fisheries Oceanography, 23(3):

242–257, May 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hertz:2016:IOC

[HTT^{+}16] E. Hertz, M. Trudel, S. Tucker, T. D. Beacham, C. Parken, D. Mackas, and A. Mazumder. Influences of ocean conditions and feeding ecology on the survival of juvenile Chinook salmon (Oncorhynchus tshawytscha). Fisheries Oceanography, 25(4):407–419, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hinrichsen:2010:VLB

[HVHC10] Hans-Harald Hinrichsen, Rudi Voss, Bastian Huwer, and Catriona Clemmesen. Variability of larval Baltic sprat (Sprattus sprattus L.) otolith growth: a modeling approach combining spatially and temporally resolved biotic and abiotic environmental key variables. Fisheries Oceanography, 19(6):463–479, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heifetz:2005:CAI

[HWS^{+}05] Jonathan Heifetz, Bruce L. Wing, Robert P. Stone, Patrick W. Malecha, and Dean L. Courtney. Corals of the Aleutian Islands. Fisheries Oceanography, 14(S1):131–138, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hollowed:2007:EOC

[HWSS07] Anne Babcock Hollowed, Christopher D. Wilson, Phyllis J. Stabeno, and Sigrid A. Salo. Effect of ocean conditions on the cross-shelf distribution of walleye pollock (Theragra chalcogramma) and capelin (Mallotus villosus). Fisheries Oceanography, 16(2):142–154, March 2007. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hsu:2017:VPN

[HXC^{+}17] Ango C. Hsu, Huijie Xue, Fei Chai, Peng Xiu, and Yu-San Han. Variability of the Pacific North Equatorial Current and its implications on Japanese eel (Anguilla japonica) larval migration. Fisheries Oceanography, 26(3):251–267, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Hamatsu:2004:DCR

[HYW04] Tomonori Hamatsu, Keizou Yabuki, and Kazutoshi Watanabe. Decadal changes in reproduction of walleye pollock (Theragra chalcogramma) off the Pacific coast of northern Japan. Fisheries Oceanography, 13(S1):74–83, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Han:2012:LJE

[HZTS12] Yu-San Han, Heng Zhang, Yu-Heng Tseng, and Mao-Lin Shen. Larval Japanese eel (Anguilla japonica) as sub-surface current bio-tracers on the East Asia continental shelf. Fisheries Oceanography, 21(4):281–290, July 2012. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Heath:1998:MDL

[HZW^{+}98] M. Heath, H. Zenitani, Y. Watanabe, R. Kimura, and M. Ishida. Modelling the dispersal of larval Japanese sardine, Sardinops melanostictus, by the Kuroshio Current in 1993 and 1994. Fisheries Oceanography, 7(3–4):335–346, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Irigoien:2008:EJB

[ICB^{+}08] Xabier Irigoien, Unai Cotano, Guillermo Boyra, Maria Santos, Paula Alvarez, Pantxika Otheguy, Egoitz Etxebeste, Andres Uriarte, Luis Ferrer, and Leire Ibaibarriaga. From egg to juvenile in the Bay of Biscay: spatial patterns of anchovy (Engraulis encrasicolus) recruitment in a non-upwelling region. Fisheries Oceanography, 17(6):446–462, November 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ishida:2018:LTC

[IFF^{+}18] Yukimasa Ishida, Hiromu Fukuda, Ko Fujioka, Osamu Sakai, Yuko Hiraoka, Kazuhiro Oshima, Shuya Nakatsuka, Nobuaki Suzuki, and Hiroyuki Shimada. Long-term changes in recruitment of age-0 Pacific bluefin tuna (Thunnus orientalis) and environmental conditions around Japan. Fisheries Oceanography, 27(1):41–48, January 2018. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Irigoien:2003:IVC

[IH03] Xabier Irigoien and Roger P. Harris. Interannual variability of Calanus helgolandicus in the English Channel. Fisheries Oceanography, 12(4–5):317–326, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ingvarsdottir:1999:SCR

[IHHH99] A. Ingvarsdóttir, D. F. Houlihan, M. R. Heath, and S. J. Hay. Seasonal changes in respiration rates of copepodite stage V Calanus finmarchicus (Gunnerus). Fisheries Oceanography, 8(S1):73–83, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ings:1997:IEU

[IHS97] D. W. Ings, J. K. Horne, and D. C. Schneider. Influence of episodic upwelling on capelin, Mallotusvillosus, and Atlantic cod, Gadusmorhua, catches in Newfoundland coastal waters. Fisheries Oceanography, 6(1):41–48, March 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Iwahashi:2006:ESS

[III^{+}06] Masayuki Iwahashi, Yutaka Isoda, Shin-Ichi Ito, Yoshioki Oozeki, and Satoshi Suyama. Estimation of seasonal spawning ground locations and ambient sea surface temperatures for eggs and larvae of Pacific saury (Cololabis saira) in the western North Pacific. Fisheries Oceanography, 15(2): 125–138, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ibaibarriaga:2007:ELD

[IIS^{+}07] L. Ibaibarriaga, X. Irigoien, M. Santos, L. Motos, J. M. Fives, C. Franco, A. Lago De Lanzós, S. Acevedo, M. Bernal, N. Bez, G. Eltink, A. Farinha, C. Hammer, S. A. Iversen, S. P. Milligan, and D. G. Reid. Egg and larval distributions of seven fish species in north-east Atlantic waters. Fisheries Oceanography, 16(3):284–293, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Iseki:1997:DSJ

[IK97] Kazuo Iseki and Yoko Kiyomoto. Distribution and settling of Japanese anchovy (Engraulis japonicus) eggs at the spawning ground off Changjiang River in the East China Sea. Fisheries Oceanography, 6(3):205–210, October 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ito:2004:IDF

[IKK^{+}04] Shin-Ichi Ito, Michio J. Kishi, Yutaka Kurita, Yoshioki Oozeki, Yasuhiro Yamanaka, Bernard A. Megrey, and Francisco E. Werner. Initial design for a fish bioenergetics model of Pacific saury coupled to a lower trophic ecosystem model. Fisheries Oceanography, 13(S1):111–124, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Iitembu:2012:COT

[IMO^{+}12] Johannes A. Iitembu, Todd W. Miller, Koji Ohmori, Abraham Kanime, and Sean Wells. Comparison of ontogenetic trophic shift in two hake species, Merluccius capensis and Merluccius paradoxus, from the Northern Benguela Current ecosystem (Namibia) using stable isotope analysis. Fisheries Oceanography, 21(2–3):215–225, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ichii:2004:DBS

[IMS^{+}04] Taro Ichii, Kedarnath Mahapatra, Mitsuo Sakai, Denzo Inagake, and Yoshihiro Okada. Differing body size between the autumn and the winter–spring cohorts of neon flying squid (Ommastrephes bartramii) related to the oceanographic regime in the North Pacific: a hypothesis. Fisheries Oceanography, 13(5):295–309, September 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Incze:2000:MTL

[IN00] Lewis S. Incze and Christopher E. Naimie. Modelling the transport of lobster (Homarus americanus) larvae and postlarvae in the Gulf of Maine. Fisheries Oceanography, 9(1): 99–113, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ichii:2018:OFA

[INM^{+}18] Taro Ichii, Haruka Nishikawa, Kedarnath Mahapatra, Hiroshi Okamura, Hiromichi Igarashi, Mitsuo Sakai, Satoshi Suyama, Masayasu Nakagami, Miyako Naya, Norihisa Usui, and Yoshihiro Okada. Oceanographic factors affecting interannual recruitment variability of Pacific saury (Cololabis saira) in the central and western North Pacific. Fisheries Oceanography, 27(5):445–457, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Igarashi:2018:IPH

[ISI^{+}18] Hiromichi Igarashi, Sei-Ichi Saitoh, Yoichi Ishikawa, Masafumi Kamachi, Norihisa Usui, Mitsuo Sakai, and Yutaka Imamura. Identifying potential habitat distribution of the neon flying squid (Ommastrephes bartramii) off the eastern coast of Japan in winter. Fisheries Oceanography, 27(1): 16–27, January 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Itoh:2011:EVG

[ISN^{+}11] Sachihiko Itoh, Toshiro Saruwatari, Haruka Nishikawa, Ichiro Yasuda, Kosei Komatsu, Atsushi Tsuda, Takashi Setou, and Manabu Shimizu. Environmental variability and growth histories of larval Japanese sardine (Sardinops melanostictus) and Japanese anchovy (Engraulis japonicus) near the frontal area of the Kuroshio. Fisheries Oceanography, 20(2):114–124, March 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Iversen:2002:AHM

[ISS02] Svein A. Iversen, Morten D. Skogen, and Einar Svendsen. Availability of horse mackerel (Trachurus trachurus) in the north-eastern North Sea, predicted by the transport of Atlantic water. Fisheries Oceanography, 11(4):245–250, July 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ito:2004:CVP

[IST^{+}04] Shin-Ichi Ito, Hiroya Sugisaki, Atsushi Tsuda, Orio Yamamura, and Kuniaki Okuda. Contributions of the VENFISH program: meso-zooplankton, Pacific saury (Cololabis saira) and walleye pollock (Theragra chalcogramma) in the northwestern Pacific. Fisheries Oceanography, 13(S1):1–9, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Igeta:2023:EIV

[IST^{+}23] Yosuke Igeta, Chiyuki Sassa, Motomitsu Takahashi, Mizuki Kuga, Satoshi Kitajima, Taku Wagawa, Shoko Abe, Chikako Watanabe, Takashi Setou, Hirohiko Nakamura, and Naoki Hirose. Effect of interannual variations of Kuroshio– Tsushima Warm Current system on the transportation of juvenile Japanese jack mackerel (Trachurus japonicus) to the Pacific coast of Japan. Fisheries Oceanography, 32(1): 133–146, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ito:2023:PPF

[ITH23] Takeshi Ito, Katsumi Takayama, and Naoki Hirose. Prediction of potential fishing grounds of swordtip squid (Uroteuthis edulis) based on a physical–biochemical coupled model. Fisheries Oceanography, 32(6):559–570, November 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Islam:2010:GDS

[IUY10] Md. Shahidul Islam, Masahiro Ueno, and Yoh Yamashita. Growth-dependent survival mechanisms during the early life of a temperate seabass (Lateolabrax japonicus): field test of the ‘growth–mortality’ hypothesis. Fisheries Oceanography, 19(3):230–242, May 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ishikawa:2021:SVO

[IWK^{+}21] Kazuo Ishikawa, Chikako Watanabe, Takahiko Kameda, Tsutomu Tokeshi, Hikari Horie, Daisuke Hashida, Toshiyuki Ookawa, Takashi Takeda, Masahiro Kuno, Yuki Suzuki, Shozo Takamura, Ryousuke Fukumoto, and Sachihiko Itoh. Spatiotemporal variability in the occurrence of juvenile Japanese jack mackerel Trachurus japonicus along coastal areas of the Kuroshio Current. Fisheries Oceanography, 30 (5):569–583, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Incze:2010:CLH

[IXW^{+}10] Lewis Incze, Huijie Xue, Nicholas Wolff, Danya Xu, Carl Wilson, Robert Steneck, Richard Wahle, Peter Lawton, Neal Pettigrew, and Yong Chen. Connectivity of lobster (Homarus americanus) populations in the coastal Gulf of Maine: Part II. Coupled biophysical dynamics. Fisheries Oceanography, 19(1):1–20, January 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Itoh:2009:TET

[IYN^{+}09] Sachihiko Itoh, Ichiro Yasuda, Haruka Nishikawa, Hideharu Sasaki, and Yoshikazu Sasai. Transport and environmental temperature variability of eggs and larvae of the Japanese anchovy (Engraulis japonicus) and Japanese sardine (Sardinops melanostictus) in the western North Pacific estimated via numerical particle-tracking experiments. Fisheries Oceanography, 18(2):118–133, March 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jansen:2016:FYS

[Jan16] Teunis Jansen. First-year survival of North East Atlantic mackerel (Scomber scombrus) from 1998 to 2012 appears to be driven by availability of Calanus, a preferred copepod prey. Fisheries Oceanography, 25(4):457–469, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jonsson:2016:RDC

[JCA^{+}16] Per R. Jonsson, Hanna Corell, Carl André, Henrik Svedäng, and Per-Olav Moksnes. Recent decline in cod stocks in the North Sea–Skagerrak–Kattegat shifts the sources of larval supply. Fisheries Oceanography, 25(3):210–228, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Johnson-Colegrove:2015:IDA

[JCCB15] Angela Johnson-Colegrove, Lorenzo Ciannelli, and Richard D. Brodeur. Ichthyoplankton distribution and abundance in relation to nearshore dissolved oxygen levels and other environmental variables within the Northern California Current System. Fisheries Oceanography, 24(6):495–507, November 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jahncke:2004:TCF

[JCH04] J. Jahncke, D. M. Checkley, Jr., and G. L. Hunt, Jr. Trends in carbon flux to seabirds in the Peruvian upwelling system: effects of wind and fisheries on population regulation. Fisheries Oceanography, 13(3):208–223, May 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jahncke:2005:SDA

[JCH05] J. Jahncke, K. O. Coyle, and George L. Hunt, Jr. Seabird distribution, abundance and diets in the eastern and central Aleutian Islands. Fisheries Oceanography, 14(S1):160–177, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jessop:2022:OWT

[Jes22] Brian M. Jessop. Oceanic water temperatures less than 20^{∘} C may largely adjust for underestimation of European elver otolith ages. Fisheries Oceanography, 31(3):353–367, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jakobsson:1993:SRC

[JGS93] J. Jakobsson, A. Gudmundsdóttir, and G. Stefánsson. Stockrelated changes in biological parameters of the Icelandic summer spawning herring. Fisheries Oceanography, 2(3–4):260– 277, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jarre:2015:ODB

[JHC^{+}15] A. Jarre, L. Hutchings, M. Crichton, K. Wieland, T. Lamont, L. K. Blamey, C. Illert, E. Hill, and M. van den Berg. Oxygen-depleted bottom waters along the west coast of South Africa, 1950–2011. Fisheries Oceanography, 24(S1): 56–73, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jarre:2015:SCE

[JHK^{+}15] Astrid Jarre, Laurence Hutchings, Stephen P. Kirkman, Anja Kreiner, Pedro C. M. Tchipalanga, Paulus Kainge, Uatjavi Uanivi, Anja K. van der Plas, Laura K. Blamey, Janet C. Coetzee, Tarron Lamont, Toufiek Samaai, Hans M. Verheye, Dawit G. Yemane, Bjoern E. Axelsen, Marek Ostrowski, Erling K. Stenevik, and Harald Loeng. Synthesis: climate effects on biodiversity, abundance and distribution of marine organisms in the Benguela. Fisheries Oceanography, 24(S1):

122–149, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Juan-Jorda:2009:GSA

[JJBCW09] Maria José Juan-Jordá, John A. Barth, M. E. Clarke, and W. W. Wakefield. Groundfish species associations with distinct oceanographic habitats in the Northern California Current. Fisheries Oceanography, 18(1):1–19, January 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See correction [Ano17].

Juan-Jorda:2017:C

[JJBCW17] Maria José Juan-Jordá, John A. Barth, M. E. Clarke, and W. W. Wakefield. Corrigendum. Fisheries Oceanography, 26 (6):706, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See [SPT^{+}17].

Jaureguizar:2006:FAN

[JMLG06] Andrés J. Jaureguizar, Roberto Menni, Carlos Lasta, and Raúl Guerrero. Fish assemblages of the northern Argentine coastal system: spatial patterns and their temporal variations. Fisheries Oceanography, 15(4):326–344, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Johns:2014:ARW

[JMP^{+}14] E. M. Johns, B. A. Muhling, R. C. Perez, F. E. Müller-Karger, N. Melo, R. H. Smith, J. T. Lamkin, T. L. Gerard, and E. Malca. Amazon River water in the northeastern Caribbean Sea and its effect on larval reef fish assemblages during April 2009. Fisheries Oceanography, 23(6):472–494, November 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jonasdottir:1999:LCC

[Jón99] S. H. Jónasdóttir. Lipid content of Calanus finmarchicus during overwintering in the Faroe–Shetland Channel. Fisheries Oceanography, 8(S1):61–72, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jennings:2016:SOE

[JPHA^{+}16] Sarah Jennings, Sean Pascoe, Sophie Hall-Aspland, Bastien Le Bouhellec, Ana Norman-Lopez, Andrew Sullivan, and Gretta Pecl. Setting objectives for evaluating management adaptation actions to address climate change impacts in south-eastern Australian fisheries. Fisheries Oceanography, 25(S1):29–44, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

James:2020:RTG

[JPMH20] Samantha E. James, Evgeny A. Pakhomov, Natalie Mahara, and Brian P. V. Hunt. Running the trophic gauntlet: Empirical support for reduced foraging success in juvenile salmon in tidally mixed coastal waters. Fisheries Oceanography, 29 (3):290–295, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Johnston:2007:FFO

[JR07] D. W. Johnston and A. J. Read. Flow-field observations of a tidally driven island wake used by marine mammals in the Bay of Fundy, Canada. Fisheries Oceanography, 16(5):422– 435, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Journey:2018:EDG

[JTYB18] M. L. Journey, M. Trudel, G. Young, and B. R. Beckman. Evidence for depressed growth of juvenile Pacific salmon (Oncorhynchus) in Johnstone and Queen Charlotte Straits, British Columbia. Fisheries Oceanography, 27(2):174–183, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Jalali:2018:MCF

[JYH^{+}18] Ali Jalali, Mary Young, Zhi Huang, Harry Gorfine, and Daniel Ierodiaconou. Modelling current and future abundances of benthic invertebrates using bathymetric LiDAR and oceanographic variables. Fisheries Oceanography, 27(6): 587–601, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kaeriyama:2017:ODF

[Kae17] Hideki Kaeriyama. Oceanic dispersion of Fukushima-derived radioactive cesium: a review. Fisheries Oceanography, 26(2): 99–113, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kaeriyama:2023:WCI

[Kae23] Masahide Kaeriyama. Warming climate impacts on production dynamics of southern populations of Pacific salmon in the North Pacific Ocean. Fisheries Oceanography, 32(1): 121–132, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kashiwai:1997:PFA

[Kas97] Makoto Kashiwai. PICES Fifth Annual Meeting. Fisheries Oceanography, 6(3):211–212, October 1997. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kashiwai:1998:PSA

[Kas98] Makoto Kashiwai. PICES Sixth Annual Meeting (Report). Fisheries Oceanography, 7(1):66–67, April 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kashiwai:1999:HPS

[Kas99] Makoto Kashiwai. Highlights of the PICES Seventh Annual Meeting. Fisheries Oceanography, 8(3):243–244, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kawasaki:1993:RCF

[Kaw93] T. Kawasaki. Recovery and collapse of the Far Eastern sardine. Fisheries Oceanography, 2(3–4):244–253, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Krishnakumar:2008:SIV

[KB08] P. K. Krishnakumar and G. S. Bhat. Seasonal and interannual variations of oceanographic conditions off Mangalore coast (Karnataka, India) in the Malabar upwelling system during 1995–2004 and their influences on the pelagic fishery. Fisheries Oceanography, 17(1):45–60, January 2008. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kerr:2020:DBF

[KBB^{+}20] Makenzie Kerr, Jeremy Browning, Eva-Maria Bønnelycke, Yingjun Zhang, Chuanmin Hu, Maickel Armenteros, Steven Murawski, Ernst Peebles, and Mya Breitbart. DNA barcoding of fish eggs collected off northwestern Cuba and across the Florida Straits demonstrates egg transport by mesoscale eddies. Fisheries Oceanography, 29(4):340–348, July 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kitagawa:2007:HVM

[KBF^{+}07] Takashi Kitagawa, André M. Boustany, Charles J. Farwell, Thomas D. Williams, Michael R. Castleton, and Barbara A. Block. Horizontal and vertical movements of juvenile bluefin tuna (Thunnus orientalis) in relation to seasons and oceanographic conditions in the eastern Pacific Ocean. Fisheries Oceanography, 16(5):409–421, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Keller:2016:SMG

[KBS^{+}16] Aimee A. Keller, John C. Buchanan, Erin Steiner, Doug Draper, Aaron Chappell, Peter Frey, and Melissa A. Head. Size at maturity for grooved Tanner crab (Chionoecetes tanneri) along the U.S. west coast (Washington to California). Fisheries Oceanography, 25(3):292–305, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Keller:2015:ODF

[KCW^{+}15] Aimee A. Keller, Lorenzo Ciannelli, W. Waldo Wakefield, Victor Simon, John A. Barth, and Stephen D. Pierce. Occurrence of demersal fishes in relation to near-bottom oxygen levels within the California Current large marine ecosystem. Fisheries Oceanography, 24(2):162–176, March 2015. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kendall:1998:DRF

[KD98] Arthur W. Kendall and Gary J. Duker. The development of recruitment fisheries oceanography in the United States. Fisheries Oceanography, 7(2):69–88, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kasamatsu:2000:DMW

[KEJK00] Fujio Kasamatsu, Paul Ensor, Gerald G. Joyce, and Noritugu Kimura. Distribution of minke whales in the Bellingshausen and Amundsen Seas (60^{o}W–120^{o}W), with special reference to environmental/physiographic variables. Fisheries Oceanography, 9(3):214–223, September 2000. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimmel:2018:CDA

[KEWDA18] David G. Kimmel, Lisa B. Eisner, Matthew T. Wilson, and Janet T. Duffy-Anderson. Copepod dynamics across warm and cold periods in the eastern Bering Sea: Implications for walleye pollock (Gadus chalcogrammus) and the oscillating control hypothesis. Fisheries Oceanography, 27(2):143–158, March 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kirby:2000:DOM

[KFH00] David S. Kirby, Øyvind Fiksen, and Paul J. B. Hart. A dynamic optimisation model for the behaviour of tunas at ocean fronts. Fisheries Oceanography, 9(4):328–342, December 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kanaji:2022:MMP

[KFS22] Yu Kanaji, Noriko Funasaka, and Hiroko Sasaki. Mesoscale movement patterns of common bottlenose dolphins along Kuroshio Current and its bifurcation. Fisheries Oceanography, 31(4):429–442, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Koeller:2007:DSP

[KFYP07] P. A. Koeller, C. Fuentes-Yaco, and T. Platt. Decreasing shrimp (Pandalus borealis) sizes off Newfoundland and Labrador — environment or fishing? Fisheries Oceanography, 16(2):105–115, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Koslow:2013:FAS

[KGW13] J. Anthony Koslow, Ralf Goericke, and William Watson. Fish assemblages in the Southern California Current: relationships with climate, 1951–2008. Fisheries Oceanography, 22(3):207–219, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Koslow:2002:CVM

[KHB02] J. Anthony Koslow, Alistair J. Hobday, and George W. Boehlert. Climate variability and marine survival of Coho salmon (Oncorhynchus kisutch) in the Oregon production area. Fisheries Oceanography, 11(2):65–77, March 2002. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kakehi:2022:RBW

[KHN^{+}22] Shigeho Kakehi, Midori Hashimoto, Miyako Naya, Shin ichi Ito, Hiroomi Miyamoto, and Satoshi Suyama. Reduced body weight of Pacific saury (Cololabis saira) causes delayed initiation of spawning migration. Fisheries Oceanography, 31 (3):319–332, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimura:2023:OAF

[Kim23] Shingo Kimura. On the occasion of the 30th anniversary of Fisheries Oceanography. Fisheries Oceanography, 32(1):1–2, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimura:2001:FDL

[KIS01] Shingo Kimura, Takashi Inoue, and Takashige Sugimoto. Fluctuation in the distribution of low-salinity water in the North Equatorial Current and its effect on the larval transport of the Japanese eel. Fisheries Oceanography, 10(1):51– 60, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kim:1997:ESA

[KJZ97] Suam Kim, Sukgeun Jung, and Chang Ik Zhang. The effect of seasonal anomalies of seawater temperature and salinity on the fluctuation in yields of small yellow croaker, Pseudosciaena polyactis, in the Yellow Sea. Fisheries Oceanography, 6(1):1–9, March 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kim:2000:EVN

[KK00] Suam Kim and Sukyung Kang. Ecological variations and El Niño effects off the southern coast of the Korean Peninsula during the last three decades. Fisheries Oceanography, 9(3):

239–247, September 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kim:2006:EDD

[KKCL06] Jin Yeong Kim, Suam Kim, Young Min Choi, and Jae Bong Lee. Evidence of density-dependent effects on population variation of Japanese sardine (Sardinops melanosticta) off Korea. Fisheries Oceanography, 15(4):345–349, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Karu:2020:TSL

[KKH^{+}20] Fukutaro Karu, Toru Kobari, Taiga Honma, Takeru Kanayama, Koji Suzuki, Naoki Yoshie, and Gen Kume. Trophic sources and linkages to support mesozooplankton community in the Kuroshio of the East China Sea. Fisheries Oceanography, 29(5):442–456, September 2020. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kakehi:2017:IDO

[KKK^{+}17] Shigeho Kakehi, Takashi Kamiyama, Yoshimasa Kaga, Kimiaki Naiki, and Shinnosuke Kaga. Improvement in the dissolved oxygen concentration and water exchange in Ofunato Bay, Japan, after the collapse of the bay-mouth breakwater by the 2011 Tohoku Earthquake and tsunami. Fisheries Oceanography, 26(2):114–127, March 2017. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimura:1992:NAP

[KKNY92] Shingo Kimura, Michio J. Kishi, Hideaki Nakata, and Yoh Yamashita. A numerical analysis of population dynamics of the sand lance (Ammodytes personatus) in the eastern Seto Inland Sea, Japan. Fisheries Oceanography, 1(4):321–332, December 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kitagawa:2004:DBI

[KKNY04] Takashi Kitagawa, Shingo Kimura, Hideaki Nakata, and Harumi Yamada. Diving behavior of immature, feeding Pacific bluefin tuna (Thunnus thynnus orientalis) in relation to season and area: the East China Sea and the Kuroshio– Oyashio transition region. Fisheries Oceanography, 13(3):

161–180, May 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kasai:1992:MTS

[KKS92] Akihide Kasai, Michio J. Kishi, and Takashige Sugimoto. Modeling the transport and survival of Japanese sardine larvae in and around the Kuroshio Current. Fisheries Oceanography, 1(1):1–10, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kim:2001:TVS

[KL01] Jinyeong Kim and Nancy C. H. Lo. Temporal variation of seasonality of egg production and the spawning biomass of Pacific anchovy, Engraulis japonicus, in the southern waters of Korea in 1983–1994. Fisheries Oceanography, 10(3):297– 310, September 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kishida:1993:SAI

[KM93] Tatsu Kishida and Hiroyuki Matsuda. Statistical analyses of intra- and interspecific density effects on recruitment of chub mackerel and sardine in Japan. Fisheries Oceanography, 2 (3–4):278–287, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Komatsu:1994:ISF

[KM94] Teruhisa Komatsu and Shin-Ichi Murakami. Influence of a Sargassum forest on the spatial distribution of water flow. Fisheries Oceanography, 3(4):256–266, December 1994. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

King:2000:DSP

[KMB00] J. R. King, G. A. Mcfarlane, and Richard J. Beamish. Decadal-scale patterns in the relative year class success of sablefish (Anoplopoma fimbria). Fisheries Oceanography, 9 (1):62–70, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Katsanevakis:2009:SDH

[KMD^{+}09] Stelios Katsanevakis, Christos D. Maravelias, Dimitrios Damalas, Aristomenis P. Karageorgis, Efthymia V. Tsitsika, Christos Anagnostou, and Costas Papaconstantinou.

Spatiotemporal distribution and habitat use of commercial demersal species in the eastern Mediterranean Sea. Fisheries Oceanography, 18(6):439–457, November 2009. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kobari:2018:GVT

[KMK^{+}18] Toru Kobari, Wataru Makihara, Takahiro Kawafuchi, Kie Sato, and Gen Kume. Geographic variability in taxonomic composition, standing stock, and productivity of the mesozooplankton community around the Kuroshio Current in the East China Sea. Fisheries Oceanography, 27(4):336–350, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimura:2006:ELS

[KMM^{+}06] Yobuo Kimura, Michael J. Miller, Gen Minagawa, Shun Watanabe, Akira Shinoda, Jun Aoyama, Tadashi Inagaki, and Katsumi Tsukamoto. Evidence of a local spawning site of marine eels along northeastern Japan, based on the distribution of small leptocephali. Fisheries Oceanography, 15(2): 183–190, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Keane:2008:LFA

[KN08] John P. Keane and Francisco J. Neira. Larval fish assemblages along the south-eastern Australian shelf: linking mesoscale non-depth-discriminate structure and water masses. Fisheries Oceanography, 17(4):263–280, July 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kaeriyama:2004:CFE

[KNE^{+}04] M. Kaeriyama, M. Nakamura, R. Edpalina, J. R. Bower, H. Yamaguchi, R. V. Walker, and K. W. Myers. Change in feeding ecology and trophic dynamics of Pacific salmon (Oncorhynchus spp.) in the central Gulf of Alaska in relation to climate events. Fisheries Oceanography, 13(3):197–207, May 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kakehi:2018:PNS

[KNK^{+}18] Shigeho Kakehi, Kimiaki Naiki, Takuya Kodama, Taku Wagawa, Hiroshi Kuroda, and Shin-Ichi Ito. Projections of nutrient supply to a wakame (Undaria pinnatifida) seaweed farm on the Sanriku Coast of Japan. Fisheries Oceanography, 27(4):323–335, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kurita:2004:VPD

[KNO^{+}04] Yutaka Kurita, Yutaka Nemoto, Yoshioki Oozeki, Ken-Ichi Hayashizaki, and Hitoshi Ida. Variations in patterns of daily changes in otolith increment widths of 0+ Pacific saury, Cololabis saira, off Japan by hatch date in relation to the northward feeding migration during spring and summer. Fisheries Oceanography, 13(S1):54–62, December 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimura:1997:MAT

[KNS97] Shingo Kimura, Munenori Nakai, and Takashige Sugimoto. Migration of albacore, Thunnus alalunga, in the North Pacific Ocean in relation to large oceanic phenomena. Fisheries Oceanography, 6(2):51–57, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kajiwara:2022:EES

[KNS^{+}22] Keitaro Kajiwara, Mitsuhiro Nakaya, Kota Suzuki, Yota Kano, and Tetsuya Takatsu. Effect of egg size on the growth rate and survival of wild walleye pollock Gadus chalcogrammus larvae. Fisheries Oceanography, 31(3):238–254, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kawasaki:1995:PMU

[KO95] Tsuyoshi Kawasaki and Michio Omori. Possible mechanisms underlying fluctuations in the Far Eastern sardine population inferred from time series of two biological traits. Fisheries Oceanography, 4(3):238–242, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kanaji:2015:EHS

[KOKM15] Yu Kanaji, Makoto Okazaki, Toshiya Kishiro, and Tomio Miyashita. Estimation of habitat suitability for the southern form of the short-finned pilot whale (Globicephala macrorhynchus) in the North Pacific. Fisheries Oceanography, 24(1):14–25, January 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kaneko:2019:DER

[KOS^{+}19] Hitoshi Kaneko, Takeshi Okunishi, Takashi Seto, Hiroshi Kuroda, Sachihiko Itoh, Shinya Kouketsu, and Daisuke Hasegawa. Dual effects of reversed winter–spring temperatures on year-to-year variation in the recruitment of chub mackerel (Scomber japonicus). Fisheries Oceanography, 28 (2):212–227, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kanaji:2016:BSO

[KOWM16] Yu Kanaji, Makoto Okazaki, Hikaru Watanabe, and Tomio Miyashita. Biogeography of small odontocetes in relation to wide-scale oceanographic structure in the North Pacific Ocean. Fisheries Oceanography, 25(2):119–132, March 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kamin:2014:ISV

[KPHG14] Lisa M. Kamin, Katie J. Palof, Jonathan Heifetz, and Anthony J. Gharrett. Interannual and spatial variation in the population genetic composition of young-of-the-year Pacific Ocean perch (Sebastes alutus) in the Gulf of Alaska. Fisheries Oceanography, 23(1):1–17, January 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kendall:2019:GRS

[KPW19] Matthew S. Kendall, Matthew Poti, and Arliss Winship. Is Guam a regional source, destination, or stepping-stone for larvae of three fisheries species? Fisheries Oceanography, 28 (2):159–170, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Knickle:2010:SSW

[KR10] D. Craig Knickle and George A. Rose. Seasonal spawning and wind-regulated retention-dispersal of early life stage Atlantic cod (Gadus morhua) in a Newfoundland fjord. Fisheries Oceanography, 19(5):397–411, September 2010. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kitchens:2014:HAD

[KR14] Larissa L. Kitchens and Jay R. Rooker. Habitat associations of dolphinfish larvae in the Gulf of Mexico. Fisheries Oceanography, 23(6):460–471, November 2014. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kempf:2013:SAP

[KSAF13] Alexander Kempf, V. Stelzenmüller, A. Akimova, and J. Floeter. Spatial assessment of predator–prey relationships in the North Sea: the influence of abiotic habitat properties on the spatial overlap between 0-group cod and grey gurnard. Fisheries Oceanography, 22(3):174–192, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Keller:2010:DFI

[KSC^{+}10] Aimee A. Keller, Victor Simon, Francis Chan, W. W. Wakefield, M. E. Clarke, John A. Barth, Dan Kamikawa, and Erica L. Fruh. Demersal fish and invertebrate biomass in relation to an offshore hypoxic zone along the US West Coast. Fisheries Oceanography, 19(1):76–87, January 2010. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kakehi:2020:PLT

[KSM^{+}20] Shigeho Kakehi, Hidekazu Shirai, Shinya Magome, Takamasa Takagi, Katsuaki Okabe, Kazufumi Takayanagi, Masami Hamaguchi, Hiroshi Ito, and Takashi Kamiyama. Predicting the larval transport of Pacific oyster Crassostrea gigas during the seedling collection season. Fisheries Oceanography, 29(6):484–504, November 2020. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kasai:2000:BIY

[KSMY00] Akihide Kasai, Wataru Sakamoto, Yasushi Mitsunaga, and Shotaro Yamamoto. Behaviour of immature yellowtails (Seriola quinqueradiata) observed by electronic data-recording tags. Fisheries Oceanography, 9(3):259–270, September 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Karnauskas:2022:SSR

[KSP^{+}22] Mandy Karnauskas, Kyle W. Shertzer, Claire B. Paris, Nicholas A. Farmer, Theodore S. Switzer, Susan K. Lowerre-Barbieri, G. Todd Kellison, Ruoying He, and Ana C. Vaz. Source–sink recruitment of red snapper: Connectivity between the Gulf of Mexico and Atlantic Ocean. Fisheries Oceanography, 31(6):571–586, November 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Karaki:2023:IMJ

[KSY^{+}23] Tatsuro Karaki, Kei Sakamoto, Goro Yamanaka, Shingo Kimura, and Akihide Kasai. Inshore migration of Japanese eel Anguilla japonica encouraged by active horizontal swimming during the glass eel stage. Fisheries Oceanography, 32(5):419–430, September 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kasai:1997:VTM

[KSYT97] Hiromi Kasai, Hiroaki Saito, Akira Yoshimori, and Satoru Taguchi. Variability in timing and magnitude of spring bloom in the Oyashio region, the western subarctic Pacific off Hokkaido, Japan. Fisheries Oceanography, 6(2):118–129, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kato:1993:TCI

[KT93] Satoshi Kato and Akira Taniguchi. Tintinnid ciliates as indicator species of different water masses in the western North Pacific Polar Front. Fisheries Oceanography, 2(3–4):166–174, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kano:2015:AVO

[KTH^{+}15] Yota Kano, Tetsuya Takatsu, Yutaro Hashimoto, Yuta Inagaki, and Toshikuni Nakatani. Annual variation in otolith increment widths of walleye pollock (Gadus chalcogrammus) larvae in Funka Bay, Hokkaido, Japan. Fisheries Oceanography, 24(4):325–334, July 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kimura:2011:EEP

[KTO^{+}11] Ryo Kimura, Hideki Takami, Tsuneo Ono, Toshihiro Onitsuka, and Yukihiro Nojiri. Effects of elevated pCO_{2} on the early development of the commercially important gastropod, Ezo abalone Haliotis discus hannai. Fisheries Oceanography, 20(5):357–366, September 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kai:2017:PST

[KTPM17] Mikihiko Kai, James T. Thorson, Kevin R. Piner, and Mark N. Maunder. Predicting the spatio-temporal distributions of pelagic sharks in the western and central North Pacific. Fisheries Oceanography, 26(5):569–582, September 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kearney:2015:SER

[KTS15] Kelly A. Kearney, Desiree Tommasi, and Charles Stock. Simulated ecosystem response to volcanic iron fertilization in the subarctic Pacific Ocean. Fisheries Oceanography, 24(5):395– 413, September 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kishi:1995:TDN

[KU95] Michio J. Kishi and Masato Uchiyama. A three-dimensional numerical model for a mariculture nitrogen cycle: case study in Shizugawa Bay, Japan. Fisheries Oceanography, 4(4):303– 316, December 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kurita:2017:IGT

[KUO^{+}17] Yutaka Kurita, Shinji Uehara, Yuji Okazaki, Tomoko Sakami, Ryogen Nambu, and Takeshi Tomiyama. Impact of the great tsunami in 2011 on the quality of nursery grounds for juvenile Japanese flounder Paralichthys olivaceus in Sendai Bay, Japan. Fisheries Oceanography, 26(2):165– 180, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kainge:2017:EEV

[KvdPBW17] P. Kainge, A. K. van der Plas, C. H. Bartholomae, and K. Wieland. Effects of environmental variables on survey catch rates and distribution by size of shallow- and deepwater Cape hakes, Merluccius capensis and Merluccius paradoxus off Namibia. Fisheries Oceanography, 26(6):680–692, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Klein:2018:SST

[KVR^{+}18] Maria Klein, Elisabeth Van Beveren, Diana Rodrigues, Ester A. Serrão, Jennifer E. Caselle, Emanuel J. Gonçalves, and Rita Borges. Small scale temporal patterns of recruitment and hatching of Atlantic horse mackerel (L.) at a nearshore reef area. Fisheries Oceanography, 27(6):505–516, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kaplan:2016:CCS

[KWB^{+}16] Isaac C. Kaplan, Gregory D. Williams, Nicholas A. Bond, Albert J. Hermann, and Samantha A. Siedlecki. Cloudy with a chance of sardines: forecasting sardine distributions using regional climate models. Fisheries Oceanography, 25(1): 15–27, January 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kodama:2018:IRJ

[KWO^{+}18] Taketoshi Kodama, Taku Wagawa, Seiji Ohshimo, Haruyuki Morimoto, Naoki Iguchi, Ken-Ichi Fukudome, Tsuneo Goto, Motomitsu Takahashi, and Tohya Yasuda. Improvement in recruitment of Japanese sardine with delays of the spring phytoplankton bloom in the Sea of Japan. Fisheries Oceanography, 27(4):289–301, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kuroda:2017:IDP

[KY17] Hiroshi Kuroda and Katsumi Yokouchi. Interdecadal decrease in potential fishing areas for Pacific saury off the southeastern coast of Hokkaido, Japan. Fisheries Oceanography, 26(4):439–454, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kirkman:2015:RSD

[KYA^{+}15] Stephen P. Kirkman, Dawit Yemane, Lara J. Atkinson, John Kathena, Silvi E. Nsiangango, Larvika Singh, Bjorn E. Axelsen, and Toufiek Samaai. Regime shifts in demersal assemblages of the Benguela Current Large Marine Ecosystem: a comparative assessment. Fisheries Oceanography, 24(S1):15– 30, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kreiner:2015:SHC

[KYS15] Anja Kreiner, Dawit Yemane, and Erling Kåre Stenevik. Spawning habitats of Cape horse mackerel (Trachurus capensis) in the northern Benguela upwelling region. Fisheries Oceanography, 24(S1):46–55, March 2015. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kreiner:2011:SSL

[KYSM11] Anja Kreiner, Dawit Yemane, Erling K. Stenevik, and Nadine E. Moroff. The selection of spawning location of sardine (Sardinops sagax) in the northern Benguela after changes in stock structure and environmental conditions. Fisheries Oceanography, 20(6):560–569, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kawabata:2006:SDJ

[KYU^{+}06] Atsushi Kawabata, Akihiko Yatsu, Yasuhiro Ueno, Satoshi Suyama, and Yutaka Kurita. Spatial distribution of the Japanese common squid, Todarodes pacificus, during its northward migration in the western North Pacific Ocean. Fisheries Oceanography, 15(2):113–124, March 2006. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Kuno:2000:PSI

[KYY00] Masahiro Kuno, Takashi Yamakawa, and Hirokatsu Yamada. Presentation of SST images from satellite NOAA on the Internet by the Fisheries Research Institute of Mie, Japan. Fisheries Oceanography, 9(3):282–284, September 2000. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Loher:2005:HCA

[LA05] T. Loher and D. A. Armstrong. Historical changes in the abundance and distribution of ovigerous red king crabs (Paralithodes camtschaticus) in Bristol Bay (Alaska), and potential relationship with bottom temperature. Fisheries Oceanography, 14(4):292–306, July 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lehodey:1998:PST

[LAB^{+}98] Patrick Lehodey, Jean-Michel André, Michel Bertignac, John Hampton, Anne Stoens, Christophe Menkes, Laurent Memery, and Nicolas Grima. Predicting skipjack tuna forage distributions in the equatorial Pacific using a coupled dynamical bio-geochemical model. Fisheries Oceanography, 7(3–4):317–325, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Logerwell:2005:GPD

[LAB^{+}05] E. A. Logerwell, K. Aydin, S. Barbeaux, E. Brown, M. E. Conners, S. Lowe, J. W. Orr, I. Ortiz, R. Reuter, and P. Spencer. Geographic patterns in the demersal ichthyofauna of the Aleutian Islands. Fisheries Oceanography, 14 (S1):93–112, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

LaMesa:2015:MET

[LAFF15] Gabriele La Mesa, Aldo Annunziatellis, Elio Filidei, Jr., and Caterina Maria Fortuna. Modeling environmental, temporal and spatial effects on twaite shad (Alosa fallax) by-catches in the central Mediterranean Sea. Fisheries Oceanography, 24(2):107–117, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Louzao:2011:STI

[LAG^{+}11] M. Louzao, J. M. Arcos, B. Guijarro, M. Valls, and D. Oro. Seabird-trawling interactions: factors affecting species-specific to regional community utilisation of fisheries waste. Fisheries Oceanography, 20(4):263–277, July 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lopes:2021:IAF

[LAPL21] Priscila F. M. Lopes, Lorena C. A. Andrade, Maria Grazia Pennino, and Tatiana S. Leite. The inter-annual fishing variability in Octopus insularis (Leite & Haimovici 2008) as a result of oceanographic factors. Fisheries Oceanography, 30 (5):515–526, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Langan:2023:TSR

[LBC23] Joseph A. Langan, Richard J. Bell, and Jeremy S. Collie. Taking stock: Is recovery of a depleted population possible in a changing climate? Fisheries Oceanography, 32(1):15–27, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lluch-Belda:2005:CMF

[LBLCLC05] Daniel Lluch-Belda, Daniel B. Lluch-Cota, and Salvador E. Lluch-Cota. Changes in marine faunal distributions and ENSO events in the California Current. Fisheries Oceanography, 14(6):458–467, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lluch-Belda:1992:SAR

[LBSS^{+}92] D. Lluch-Belda, R. A. Schwartzlose, R. Serra, R. Parrish, T. Kawasaki, D. Hedgecock, and R. J. M. Crawford. Sardine and anchovy regime fluctuations of abundance in four regions of the world oceans: a workshop report. Fisheries Oceanography, 1(4):339–347, December 1992. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lough:2005:GBM

[LBW^{+}05] R. G. Lough, L. J. Buckley, F. E. Werner, J. A. Quinlan, and K. Pehrson Edwards. A general biophysical model of larval cod (Gadus morhua) growth applied to populations on Georges Bank. Fisheries Oceanography, 14(4):241–262, July 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lima:1995:DAS

[LC95] Ivan D. Lima and Jorge P. Castello. Distribution and abundance of South-west Atlantic anchovy spawners (Engraulis anchoita) in relation to oceanographic processes in the southern Brazilian shelf. Fisheries Oceanography, 4(1):1– 16, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Liu:2015:SDE

[LCC15] Bi Lin Liu, Yong Chen, and Xin Jun Chen. Spatial difference in elemental signatures within early ontogenetic statolith for identifying Jumbo flying squid natal origins. Fisheries Oceanography, 24(4):335–346, July 2015. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lennert-Cody:2019:IEL

[LCCdS^{+}19] Cleridy E. Lennert-Cody, Shelley C. Clarke, Alexandre Aires da Silva, Mark N. Maunder, Peter J. S. Franks, Marlon Román, Arthur J. Miller, and Mihoko Minami. The importance of environment and life stage on interpretation of silky shark relative abundance indices for the equatorial Pacific Ocean. Fisheries Oceanography, 28(1):43–53, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Laiz-Carrion:2022:SST

[LCCQ^{+}22] Raúl Laiz-Carrión, Águeda Cabrero, José María Quintanilla, Alma Hernández, Amaya Uriarte, Jesús Gago, José María Rodríguez, Carmen Piñeiro, Alberto García, and Fran Saborido-Rey. Shifts in the seasonal trophic ecology of larvae and juveniles of European hake (Merluccius merluccius): From the Galician upwelling system (NW Spain). Fisheries Oceanography, 31(5):539–553, September 2022. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Liu:2015:CBS

[LCCS15] Chang Liu, Geoffrey W. Cowles, James H. Churchill, and Kevin D. E. Stokesbury. Connectivity of the bay scallop (Argopecten irradians) in Buzzards Bay, Massachusetts, U.S.A. Fisheries Oceanography, 24(4):364–382, July 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lehodey:2003:MCR

[LCH03] Patrick Lehodey, Fei Chai, and John Hampton. Modelling climate-related variability of tuna populations from a coupled ocean–biogeochemical-populations dynamics model. Fisheries Oceanography, 12(4–5):483–494, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Logerwell:2010:IPH

[LDAWM10] Elizabeth A. Logerwell, Janet Duffy-Anderson, Matthew Wilson, and Denise Mckelvey. The influence of pelagic habitat selection and interspecific competition on productivity of juvenile walleye pollock (Theragra chalcogramma) and capelin (Mallotus villosus) in the Gulf of Alaska. Fisheries Oceanography, 19(4):262–278, July 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lee:2006:TSV

[LDDC06] Orla Lee, Bret S. Danilowicz, and Mark Dickey-Collas. Temporal and spatial variability in growth and condition of dab (Limanda limanda) and sprat (Sprattus sprattus) larvae in the Irish Sea. Fisheries Oceanography, 15(6):490–507, November 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Laurel:2014:ETG

[LDH14] Benjamin J. Laurel, Courtney Danley, and Scott Haines. The effects of temperature on growth, development and settlement of northern rock sole larvae (Lepidopsetta polyxystra). Fisheries Oceanography, 23(6):495–505, November 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lloyd:2012:OWA

[LéEPW^{+}12] Penn Lloyd, éva E. Plagányi, Scarla J. Weeks, Marites Magno-Canto, and Gyula Plagányi. Ocean warming alters species abundance patterns and increases species diversity in an African sub-tropical reef-fish community. Fisheries Oceanography, 21(2–3):78–94, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lafuente:2002:HPI

[LGM^{+}02] J. García Lafuente, A. García, S. Mazzola, L. Quintanilla, J. Delgado, A. Cuttita, and B. Patti. Hydrographic phenomena influencing early life stages of the Sicilian Channel anchovy. Fisheries Oceanography, 11(1):31–44, January 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Logerwell:1996:DSB

[LH96] Elizabeth A. Logerwell and N. Brent Hargreaves. The distribution of sea birds relative to their fish prey off Vancouver Island: opposing results at large and small spatial scales. Fisheries Oceanography, 5(3–4):163–175, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Luettich:1999:BTW

[LHF^{+}99] Richard A. Luettich, Jr., James L. Hench, Crystal W. Fulcher, Francisco E. Werner, Brian O. Blanton, and James H. Churchill. Barotropic tidal and wind-driven larval transport in the vicinity of a barrier island inlet. Fisheries Oceanography, 8(S2):190–209, December 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ladd:2005:MEE

[LHM^{+}05] Carol Ladd, George L. Hunt, Jr., Calvin W. Mordy, Sigrid A. Salo, and Phyllis J. Stabeno. Marine environment of the eastern and central Aleutian Islands. Fisheries Oceanography, 14(S1):22–38, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Livingston:2000:HPE

[Liv00] Patricia Livingston. Highlights of the PICES Eighth Annual Meeting. Fisheries Oceanography, 9(2):192–193, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Li:2020:PSD

[LJBR20] Min Li, Yan Jiao, Rujia Bi, and Yiping Ren. Population status and distribution of whitespotted conger (Conger myriaster) in Yellow Sea: an important migratory species along coastal China with limited data. Fisheries Oceanography, 29 (1):32–45, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ladd:2005:HFS

[LJH^{+}05] Carol Ladd, Jaime Jahncke, George L. Hunt, Jr., Kenneth O. Coyle, and Phyllis J. Stabeno. Hydrographic features and seabird foraging in Aleutian passes. Fisheries Oceanography, 14(S1):178–195, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Linnane:2010:IWS

[LJM^{+}10] Adrian Linnane, Charles James, John Middleton, Peter Hawthorne, and Matthew Hoare. Impact of wind stress anomalies on the seasonal pattern of southern rock lobster (Jasus edwardsii) settlement in South Australia. Fisheries Oceanography, 19(4):290–300, July 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lennox:2022:SIB

[LJR^{+}22] Robert J. Lennox, Claudia Junge, Jan Reubens, Abdirahman Omar, Ingunn Skjelvan, and Knut Wiik Vollset. Strategic importance of the Bergen–Shetland Corridor to marine biology and oceanography of the Atlantic Ocean. Fisheries Oceanography, 31(5):471–479, September 2022. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lamb:2021:CDD

[LK21] Jesse F. Lamb and David G. Kimmel. The contribution of diet to the dramatic reduction of the 2013 year class of Gulf of Alaska walleye pollock (Gadus chalcogrammus). Fisheries Oceanography, 30(6):757–771, November 2021. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lemus:2020:SWI

[LLB^{+}20] Diego Lemus, Mauricio F. Landaeta, Fernando Balbontín, Juan Carlos Saavedra-Nievas, Vivian Valenzuela, and Hernán Miles. Subtropical water influences temporal fluctuations of early life stages of Vinciguerria lucetia (Osteichthyes: Phosichthyidae) in the Humboldt Current System (1998–2004). Fisheries Oceanography, 29(1):23–31, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

LaMesa:2016:SDP

[LLCJ16] Mario La Mesa, Gabriele La Mesa, Barbara Catalano, and Christopher D. Jones. Spatial distribution pattern and physical — biological interactions in the larval notothenioid fish assemblages from the Bransfield Strait and adjacent waters. Fisheries Oceanography, 25(6):624–636, November 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lan:2018:ABI

[LLCV18] Kuo-Wei Lan, Ming-An Lee, Chin-Pei Chou, and Ali Haghi Vayghan. Association between the interannual variation in the oceanic environment and catch rates of bigeye tuna (Thunnus obesus) in the Atlantic Ocean. Fisheries Oceanography, 27(5):395–407, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lloret:2001:FLE

[LLSF01] Josep Lloret, Jordi Lleonart, Ignasi Solé, and Jean-Marc Fromentin. Fluctuations of landings and environmental conditions in the north-western Mediterranean Sea. Fisheries Oceanography, 10(1):33–50, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Litz:2019:EDS

[LMB^{+}19] Marisa N. C. Litz, Jessica A. Miller, Richard D. Brodeur, Elizabeth A. Daly, Laurie A. Weitkamp, Adam G. Hansen, and Andrew M. Claiborne. Energy dynamics of subyearling Chinook salmon reveal the importance of piscivory to short-term growth during early marine residence. Fisheries Oceanography, 28(3):273–290, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Loukos:2003:PCS

[LMBL03] Harilaos Loukos, Patrick Monfray, Laurent Bopp, and Patrick Lehodey. Potential changes in skipjack tuna (Katsuwonus pelamis) habitat from a global warming scenario: modelling approach and preliminary results. Fisheries Oceanography, 12(4–5):474–482, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Logerwell:2003:TEP

[LML^{+}03] E. A. Logerwell, N. Mantua, P. W. Lawson, R. C. Francis, and V. N. Agostini. Tracking environmental processes in the coastal zone for understanding and predicting Oregon Coho (Oncorhynchus kisutch) marine survival. Fisheries Oceanography, 12(6):554–568, November 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lezama-Ochoa:2015:BCH

[LOGLD^{+}15] Ainhoa Lezama-Ochoa, Daniel Grados, Anne Lebourges-Dhaussy, Xabier Irigoien, Alexis Chaigneau, and Arnaud Bertrand. Biological characteristics of the hydrological landscapes in the Bay of Biscay in spring 2009. Fisheries Oceanography, 24(1):26–41, January 2015. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Landa:2014:RDB

[LOS^{+}14] Camilla S. Landa, Geir Ottersen, Svein Sundby, Gjert E. Dingsør, and Jan E. Stiansen. Recruitment, distribution boundary and habitat temperature of an arcto-boreal gadoid in a climatically changing environment: a case study on Northeast Arctic haddock (Melanogrammus aeglefinus). Fisheries Oceanography, 23(6):506–520, November 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lough:2010:JCG

[Lou10] R. Gregory Lough. Juvenile cod (Gadus morhua) mortality and the importance of bottom sediment type to recruitment on Georges Bank. Fisheries Oceanography, 19(2):159–181, March 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Liu:2010:SIA

[LP10] Hui Liu and William T. Peterson. Seasonal and inter-annual variations in the abundance and biomass of Neocalanus plumchrus in continental slope waters off Oregon. Fisheries Oceanography, 19(5):354–369, September 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

LaMesa:2015:PEL

[LPCA15] Mario La Mesa, Andrea Piñones, Barbara Catalano, and Julian Ashford. Predicting early life connectivity of Antarctic silverfish, an important forage species along the Antarctic Peninsula. Fisheries Oceanography, 24(2):150–161, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Landaeta:2023:STP

[LPCG23] Mauricio F. Landaeta, Lissette D. Paredes, Manuel I. Castillo, and M. Teresa González. Spatio-temporal patterns of ichthyoplankton in southern Chilean Patagonia: βdiversity and associated environmental factors. Fisheries Oceanography, 32(4):341–351, July 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Luo:2006:VHU

[LPG^{+}06] Jiangang Luo, Eric D. Prince, C. Phillip Goodyear, Brian E. Luckhurst, and Joseph E. Serafy. Vertical habitat utilization by large pelagic animals: a quantitative framework and numerical method for use with pop-up satellite tag data. Fisheries Oceanography, 15(3):208–229, May 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

LeCorre:2019:ACP

[LPH^{+}19] Nicolas Le Corre, Pierre Pepin, Guoqi Han, Zhimin Ma, and Paul V. R. Snelgrove. Assessing connectivity patterns among management units of the Newfoundland and Labrador shrimp population. Fisheries Oceanography, 28(2): 183–202, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

LeCorre:2021:PIC

[LPHM21] Nicolas Le Corre, Pierre Pepin, Guoqi Han, and Zhimin Ma. Potential impact of climate change on northern shrimp habitats and connectivity on the Newfoundland and Labrador continental shelves. Fisheries Oceanography, 30(3):331–347, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

LauserColetto:2019:OOA

[LPS19] Juliano Lauser Coletto, Marcelo Peres Pinho, and Lauro Saint Pastous Madureira. Operational oceanography applied to skipjack tuna (Katsuwonus pelamis) habitat monitoring and fishing in south-western Atlantic. Fisheries Oceanography, 28(1):82–93, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lloret:2004:IFI

[LPSS04] J. Lloret, I. Palomera, J. Salat, and I. Sole. Impact of freshwater input and wind on landings of anchovy (Engraulis encrasicolus) and sardine (Sardina pilchardus) in shelf waters surrounding the Ebre (Ebro) River delta (north-western Mediterranean). Fisheries Oceanography, 13(2):102–110, March 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lyttle:2021:AHR

[LRBJ21] Casey Tara Lyttle, Christo Rautenbach, Björn Backeberg, and Astrid Jarre. An analysis of high-resolution modelled wave heights along the South African south coast suggests recent deterioration of sea state. Fisheries Oceanography, 30(6):679–696, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lett:2006:SQE

[LRL^{+}06] Christophe Lett, Claude Roy, Anne Levasseur, Carl D. Van Der Lingen, and Christian Mullon. Simulation and quantification of enrichment and retention processes in the southern Benguela upwelling ecosystem. Fisheries Oceanography, 15(5):363–372, September 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lavery:2023:EON

[LRS^{+}23] Madeline A. K. Lavery, Christopher N. Rooper, Kota Sawada, Kari Fenske, Vladimir Kulik, and Kyum Joon Park. Effects of oceanography on North Pacific armorhead recruitment in the Emperor Seamounts. Fisheries Oceanography, 32(2):160–176, March 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Logerwell:2001:MES

[LS01] Elizabeth A. Logerwell and Paul E. Smith. Mesoscale eddies and survival of late stage Pacific sardine (Sardinops sagax) larvae. Fisheries Oceanography, 10(1):13–25, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Loeng:2015:NCE

[LS15] Harald Loeng and Erling Kåre Stenevik. NansClim — climate effects on biodiversity, abundance and distribution of marine organisms. Fisheries Oceanography, 24(S1):iii–iv, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Langangen:2021:SLL

[LS21] Øystein Langangen and Leif Christian Stige. Shedding light on the link between the spatial distribution of eggs and survival in Northeast Arctic cod. Fisheries Oceanography, 30 (4):429–436, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Luzenti:2021:PBD

[LSD^{+}21] Elvio Agustín Luzenti, Guillermo Martín Svendsen, Mariana Degrati, Nadia Soledad Curcio, Raúl Alberto González, and Silvana Laura Dans. Physical and biological drivers of pelagic fish distribution at high spatial resolution in two Patagonian Gulfs. Fisheries Oceanography, 30(4):397–412, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Langangen:2018:MDV

[LSK^{+}18] Øystein Langangen, Leif Christian Stige, Kristina Ø. Kvile, Natalia A. Yaragina, Jon Egil Skjæraasen, Frode B. Vikebø, and Geir Ottersen. Multi-decadal variations in spawning ground use in Northeast Arctic haddock (Melanogrammus aeglefinus). Fisheries Oceanography, 27(5):435–444, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Liu:2003:ACS

[LSW^{+}03] G. M. Liu, S. Sun, H. Wang, Y. Zhang, B. Yang, and P. Ji. Abundance of Calanus sinicus across the tidal front in the Yellow Sea, China. Fisheries Oceanography, 12(4–5):291– 298, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Liao:2022:VDP

[LTL^{+}22] Tongchen Liao, Yehui Tan, Hui Liu, Linghui Yu, and Kaizhi Li. Vertical distribution of planktonic ostracods (Halocyprididae) in the northeastern and central South China Sea: Significance of large-sized species in deep waters. Fisheries Oceanography, 31(5):497–509, September 2022. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lafuente:2005:AVH

[LVC^{+}05] Jesús García Lafuente, Juan Miguel Vargas, Francisco Criado, Alberto García, Javier Delgado, and Salvatore Mazzola. Assessing the variability of hydrographic processes influencing the life cycle of the Sicilian Channel anchovy, Engraulis encrasicolus, by satellite imagery. Fisheries Oceanography, 14(1):32–46, January 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lelievre:2012:ECU

[LVF12] S. Lelievre, S. Vaz, and C. Fox. The effectiveness of the continuous underway fish egg sampler compared with standard vertical plankton hauls for sampling the eggs of demersal species in the southern North Sea. Fisheries Oceanography, 21(2–3):109–124, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lindegren:2018:PRF

[LVM^{+}18] Martin Lindegren, Mikael Van Deurs, Brian R. MacKenzie, Lotte Worsoe Clausen, Asbjørn Christensen, and Anna Rindorf. Productivity and recovery of forage fish under climate change and fishing: North Sea sandeel as a case study. Fisheries Oceanography, 27(3):212–221, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Loots:2011:UWC

[LVPK11] Christophe Loots, Sandrine Vaz, Benjamin Planque, and Philippe Koubbi. Understanding what controls the spawning distribution of North Sea whiting (Merlangius merlangus) using a multi-model approach. Fisheries Oceanography, 20(1):18–31, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lynn:2003:VSH

[Lyn03] Ronald J. Lynn. Variability in the spawning habitat of Pacific sardine (Sardinops sagax) off southern and central California. Fisheries Oceanography, 12(6):541–553, November 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Li:2020:LTV

[LYT^{+}20] Haoyu Li, Shu Yang, Qisheng Tang, Xin Zhou, and Yao Sun. Long-term variation in the abundance of Pacific herring (Clupea pallasii) from the Yellow Sea in the western North Pacific and its relation to climate over the past 590 years. Fisheries Oceanography, 29(1):56–65, January 2020. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miller:2012:DSE

[MAH12] Sara E. Miller, Milo D. Adkison, and Lewis J. Haldorson. Differences in stability effects on the marine survival of hatchery pink salmon (Oncorhynchus gorbuscha) within the upwelling and downwelling domains of the northeast Pacific Ocean. Fisheries Oceanography, 21(6):430–444, November 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcconnaughey:1994:IVC

[MAHG94] Robert A. Mcconnaughey, David A. Armstrong, Barbara M. Hickey, and Donald R. Gunderson. Interannual variability in coastal Washington Dungeness crab (Cancer magister) populations: larval advection and the coastal landing strip. Fisheries Oceanography, 3(1):22–38, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Malick:2020:TVR

[Mal20] Michael J. Malick. Time-varying relationships between ocean conditions and sockeye salmon productivity. Fisheries Oceanography, 29(3):265–275, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Maravelias:2001:HAA

[Mar01] Christos D. Maravelias. Habitat associations of Atlantic herring in the Shetland area: influence of spatial scale and geographic segmentation. Fisheries Oceanography, 10(3):259– 267, September 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miyashita:1998:AII

[MAS^{+}98] Kazushi Miyashita, Ichiro Aoki, Koutaro Seno, Kenji Taki, and Takashi Ogishima. Acoustic identification of isada krill, Euphausia pacifica Hansen, off the Sanriku coast, northeastern Japan. Fisheries Oceanography, 6(4):266–271, February 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Matsukawa:2006:CCS

[Mat06] Yasuo Matsukawa. Carrying capacity and survival strategy for the Pacific bluefin tuna, Thunnus orientalis, in the Western Pacific. Fisheries Oceanography, 15(2):104–112, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mclaren:1998:FLC

[MATL98] I. A. Mclaren, Patricia Avendaño, Christopher T. Taggart, and S. E. Lochmann. Feeding by larval cod in different water-masses on Western Bank, Scotian Shelf. Fisheries Oceanography, 6(4):250–265, February 1998. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Musyl:2003:VMB

[MBB^{+}03] Michael K. Musyl, Richard W. Brill, Christofer H. Boggs, Daniel S. Curran, Thomas K. Kazama, and Michael P. Seki. Vertical movements of bigeye tuna (Thunnus obesus) associated with islands, buoys, and seamounts near the main Hawaiian Islands from archival tagging data. Fisheries Oceanography, 12(3):152–169, May 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mbaye:2015:DSA

[MBE^{+}15] Baye Cheikh Mbaye, Timothée Brochier, Vincent Echevin, Alban Lazar, Marina Lévy, Evan Mason, Amadou Thierno Gaye, and Eric Machu. Do Sardinella aurita spawning seasons match local retention patterns in the Senegalese– Mauritanian upwelling region? Fisheries Oceanography, 24 (1):69–89, January 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Madden:1999:SES

[MBH^{+}99] H. Madden, D. Beare, M. R. Heath, J. G. Fraser, and A. Gallego. The spring/early summer distribution of Calanus spp. in the northern North Sea and adjacent areas. Fisheries Oceanography, 8(S1):138–152, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mamie:2007:ADG

[MBJ^{+}07] J. C. J. Mamie, D. J. Beare, E. G. Jones, M. Kienzle, H. Dobby, M. R. Heath, and D. G. Reid. Aspects of the distribution and growth of bluemouth (Helicolenus dactylopterus, Delaroche 1809) since its invasion of the northern North Sea in 1991. Fisheries Oceanography, 16(1):85–94, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Muhling:2008:LFA

[MBKP08] B. A. Muhling, L. E. Beckley, J. A. Koslow, and A. F. Pearce. Larval fish assemblages and water mass structure off the oligotrophic south-western Australian coast. Fisheries Oceanography, 17(1):16–31, January 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mavruk:2017:TLF

[MBY^{+}17] Sinan Mavruk, Fethi Bengil, Hacer Yeldan, Meltem Manasirli, and Dursun Avsar. The trend of Lessepsian fish populations with an emphasis on temperature variations in Iskenderun Bay, the Northeastern Mediterranean. Fisheries Oceanography, 26(5):542–554, September 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mavruk:2018:IAP

[MBY^{+}18] Sinan Mavruk, Fethi Bengil, Ahsen Yüksek, Caner Enver özyurt, Volkan Barış Kiyağa, and Dursun Avşar. Intraannual patterns of coastal larval fish assemblages along environmental gradients in the northeastern Mediterranean. Fisheries Oceanography, 27(3):232–245, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Montero:2016:BSF

[MCB^{+}16] Jose T. Montero, Tanya A. Chesney, Jennifer R. Bauer, John T. Froeschke, and Jim Graham. Brown shrimp (Farfantepenaeus aztecus) density distribution in the Northern Gulf of Mexico: an approach using boosted regression trees. Fisheries Oceanography, 25(3):337–348, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

McKinnell:2014:OAE

[MCG^{+}14] Skip McKinnell, Enrique Curchitser, Kees Groot, Masahide Kaeriyama, and Marc Trudel. Oceanic and atmospheric extremes motivate a new hypothesis for variable marine survival of Fraser River sockeye salmon. Fisheries Oceanography, 23(4):322–341, July 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Malpica-Cruz:2013:TSS

[MCHSNEO13] Luis Malpica-Cruz, Sharon Z. Herzka, Oscar Sosa-Nishizaki, and Miguel A. Escobedo-Olvera. Tissue-specific stable isotope ratios of shortfin mako (Isurus oxyrinchus) and white (Carcharodon carcharias) sharks as indicators of size-based differences in foraging habitat and trophic level. Fisheries Oceanography, 22(6):429–445, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

McKinnell:2013:CKV

[McK13] Skip McKinnell. Challenges for the Kasatoshi volcano hypothesis as the cause of a large return of sockeye salmon (Oncorhynchus nerka) to the Fraser River in 2010. Fisheries Oceanography, 22(4):337–344, July 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Malick:2017:ENP

[MCM^{+}17] Michael J. Malick, Sean P. Cox, Franz J. Mueter, Brigitte Dorner, and Randall M. Peterman. Effects of the North Pacific Current on the productivity of 163 Pacific salmon stocks. Fisheries Oceanography, 26(3):268–281, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Meekan:2006:TMM

[MCS^{+}06] M. G. Meekan, J. H. Carleton, C. R. Steinberg, A. D. Mckinnon, R. Brinkman, P. J. Doherty, A. Halford, S. Duggan, and L. Mason. Turbulent mixing and mesoscale distributions of late-stage fish larvae on the NW Shelf of Western Australia. Fisheries Oceanography, 15(1):44–59, January 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Matsuura:1993:DAT

[MDKS93] Yasunobu Matsuura, Ana Claudia De Silva Garcia, Mario Katsuragawa, and Kazuko Suzuki. Distribution and abundance of two species of codlet (Teleostei, Bregmacerotidae) larvae from the south-eastern Brazilian Bight. Fisheries Oceanography, 2(2):82–90, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcleay:2016:BMA

[MDR^{+}16] Lachlan Mcleay, Mark Doubell, Shane Roberts, Cameron Dixon, Lorenzo Andreacchio, Charles James, John Luick, and John Middleton. A biophysical model to assess the trade-off between larval recruitment and catch in southern Australia’s largest prawn fishery. Fisheries Oceanography, 25(2):164–182, March 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mbaye:2020:MAM

[MDVB^{+}20] Baye Mbaye, Thomas Doniol-Valcroze, Pablo Brosset, Martin Castonguay, Elisabeth Van Beveren, Andrew Smith, Caroline Lehoux, David Brickman, Zeliang Wang, and Stéphane Plourde. Modelling Atlantic mackerel spawning habitat suitability and its future distribution in the north-west Atlantic. Fisheries Oceanography, 29(1):84–99, January 2020. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Machu:2009:ECR

[MEK^{+}09] E. Machu, O. Ettahiri, S. Kifani, A. Benazzouz, A. Makaoui, and H. Demarcq. Environmental control of the recruitment of sardines (Sardina pilchardus) over the western Saharan shelf between 1995 and 2002: a coupled physical/ biogeochemical modelling experiment. Fisheries Oceanography, 18(5):287–300, September 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Marin-Enriquez:2018:EMO

[MESMM18] Emigdio Marín-Enríquez, Javier Seoane, and Arturo Muhlia-Melo. Environmental modeling of occurrence of dolphinfish (Coryphaena spp.) in the Pacific Ocean off Mexico reveals seasonality in abundance, hot spots and migration patterns. Fisheries Oceanography, 27(1):28–40, January 2018. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Munk:2009:SNS

[MFB^{+}09] Peter Munk, Clive J. Fox, Loes J. Bolle, Cindy J. G. Van Damme, Petter Fossum, and Gerd Kraus. Spawning of North Sea fishes linked to hydrographic features. Fisheries Oceanography, 18(6):458–469, November 2009. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mckinnell:1999:AND

[MFG99] Skip Mckinnell, Howard J. Freeland, and Sarah D. Groulx. Assessing the northern diversion of sockeye salmon returning to the Fraser River, BC. Fisheries Oceanography, 8(2):104– 114, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcdermott:2005:EMA

[MFH05] Susanne F. Mcdermott, L. W. Fritz, and V. Haist. Estimating movement and abundance of Atka mackerel (Pleurogrammus monopterygius) with tag–release–recapture data. Fisheries Oceanography, 14(S1):113–130, November 2005. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mazur:2020:DCA

[MFMG20] Mackenzie D. Mazur, Kevin D. Friedland, M. Conor Mc-Manus, and Andrew G. Goode. Dynamic changes in American lobster suitable habitat distribution on the Northeast U.S. Shelf linked to oceanographic conditions. Fisheries Oceanography, 29(4):349–365, July 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mullon:2003:PIM

[MFP^{+}03] C. Mullon, P. Fréon, C. Parada, C. Van Der Lingen, and J. Huggett. From particles to individuals: modelling the early stages of anchovy (Engraulis capensis/encrasicolus) in the southern Benguela. Fisheries Oceanography, 12(4–5): 396–406, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Michaud:1996:FSS

[MFRR96] Josée Michaud, Louis Fortier, Peter Rowe, and René Ramseier. Feeding success and survivorship of Arctic cod larvae, Boreogadus saida, in the Northeast Water polynya (Greenland Sea). Fisheries Oceanography, 5(2):120–135, June 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miki:2017:CAR

[MFS^{+}17] Shizuho Miki, Ken Fujimoto, Yuya Shigenobu, Daisuke Ambe, Hideki Kaeriyama, Kaori Takagi, Tsuneo Ono, Tomowo Watanabe, Hiroya Sugisaki, and Takami Morita. Concentrations of ^{90}Sr and ^{137}Cs/^{90}Sr activity ratios in marine fishes after the Fukushima Dai-ichi Nuclear Power Plant accident. Fisheries Oceanography, 26(2):221–233, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mullaney:2014:EAL

[MGHS14] Thomas J. Mullaney, Bronwyn M. Gillanders, Elizabeth C. Heagney, and Iain M. Suthers. Entrainment and advection of larval sardine, Sardinops sagax, by the East Australian Current and retention in the western Tasman Front. Fisheries Oceanography, 23(6):554–567, November 2014. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mourato:2014:STT

[MHB^{+}14] Bruno L. Mourato, Fábio Hazin, Keith Bigelow, Michael Musyl, Felipe Carvalho, and Humberto Hazin. Spatio-temporal trends of sailfish, Istiophorus platypterus catch rates in relation to spawning ground and environmental factors in the equatorial and southwestern Atlantic Ocean. Fisheries Oceanography, 23(1):32–44, January 2014. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcilwain:2011:SCD

[MHG^{+}11] Jennifer L. Mcilwain, Euan S. Harvey, Simon Grove, Glenn Shiell, Hamed Al Oufi, and Nasr Al Jardani. Seasonal changes in a deep-water fish assemblage in response to monsoon-generated upwelling events. Fisheries Oceanography, 20(6):497–516, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

McGrath:2020:ICB

[MHM^{+}20] Alice M. McGrath, Juliet C. Hermes, Coleen L. Moloney, Claude Roy, Gildas Cambon, Steven Herbette, and Carl D. van der Lingen. Investigating connectivity between two sardine stocks off South Africa using a high-resolution IBM: Retention and transport success of sardine eggs. Fisheries Oceanography, 29(2):137–151, March 2020. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

McManus:2018:TSA

[MHRC18] M. Conor McManus, Jonathan A. Hare, David E. Richardson, and Jeremy S. Collie. Tracking shifts in Atlantic mackerel (Scomber scombrus) larval habitat suitability on the Northeast U.S. Continental Shelf. Fisheries Oceanography, 27(1):49–62, January 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miazaki:2021:GPL

[MHS^{+}21] Lizandra Fernandes Miazaki, Gisele Salgado Heckler, Ana Paula Freitas Santos, Antonio Leão Castilho, Régis Augusto Pescinelli, and Rogerio Caetano Costa. Growth parameters, longevity, and mortality of the seabob shrimp Xiphopenaeus spp. (Decapoda: Penaeidae) in four important fishing regions of southeastern Brazil. Fisheries Oceanography, 30(5):499– 514, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Munk:2024:SSL

[MHvD^{+}24] Peter Munk, Bastian Huwer, Mikael van Deurs, Matthias Kloppmann, and Anne Sell. Spatial separation of larval sprat (Sprattus sprattus) and sardine (Sardina pilchardus) related to hydrographical characteristics in the North Sea.

Fisheries Oceanography, 33(1):e12656:1–e12656:??, January 2024. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Minamikawa:2007:DBB

[MIK07] Shingo Minamikawa, Toshihide Iwasaki, and Toshiya Kishiro. Diving behaviour of a Baird’s beaked whale, Berardius bairdii, in the slope water region of the western North Pacific: first dive records using a data logger. Fisheries Oceanography, 16(6):573–577, November 2007. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Murase:2009:ACB

[MIY^{+}09] Hiroto Murase, Morio Ichihara, Hiroki Yasuma, Hikaru Watanabe, Shiroh Yonezaki, Hiroshi Nagashima, Shigeyuki Kawahara, and Kazushi Miyashita. Acoustic characterization of biological backscatterings in the Kuroshio–Oyashio inter-frontal zone and subarctic waters of the western North Pacific in spring. Fisheries Oceanography, 18(6):386–401, November 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Michael:2014:RIS

[MJH14] P. E. Michael, J. Jahncke, and K. D. Hyrenbach. Relative influence of static and dynamic features on black-footed albatross (Phoebastria nigripes) habitat use in central California sanctuaries. Fisheries Oceanography, 23(1):18–31, January 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mollmann:2003:MCP

[MKF^{+}03] C. Möllmann, G. Kornilovs, M. Fetter, F. W. Köster, and H.-H. Hinrichsen. The marine copepod, Pseudocalanus elongatus, as a mediator between climate variability and fisheries in the Central Baltic Sea. Fisheries Oceanography, 12(4–5): 360–368, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Murase:2013:SDA

[MKH^{+}13] Hiroto Murase, Toshihide Kitakado, Takashi Hakamada, Koji Matsuoka, Shigetoshi Nishiwaki, and Mikio Naganobu. Spatial distribution of Antarctic minke whales (Balaenoptera bonaerensis) in relation to spatial distributions of krill in the Ross Sea, Antarctica. Fisheries Oceanography, 22(3):154– 173, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Matsumoto:2013:VBB

[MKK13] Takayuki Matsumoto, Takashi Kitagawa, and Shingo Kimura. Vertical behavior of bigeye tuna (Thunnus obesus) in the northwestern Pacific Ocean based on archival tag data. Fisheries Oceanography, 22(3):234–246, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miller:1998:CIB

[MLC^{+}98] Charles B. Miller, Daniel R. Lynch, FranÇOis Carlotti, Wendy Gentleman, and Craig V. W. Lewis. Coupling of an individual-based population dynamic model of Calanus finmarchicus to a circulation model for the Georges Bank region. Fisheries Oceanography, 7(3–4):219–234, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcgillicuddy:1998:ADA

[MLM^{+}98] D. J. Mcgillicuddy, D. R. Lynch, A. M. Moore, W. C. Gentleman, C. S. Davis, and C. J. Meise. An adjoint data assimilation approach to diagnosis of physical and biological controls on Pseudocalanus spp. in the Gulf of Maine–Georges Bank region. Fisheries Oceanography, 7(3–4):205–218, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

McGeady:2022:LTI

[MLP22] Ryan McGeady, Colm Lordan, and Anne Marie Power. Longterm interannual variability in larval dispersal and connectivity of the Norway lobster (Nephrops norvegicus) around Ireland: When supply-side matters. Fisheries Oceanography, 31(3):255–270, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Muhling:2010:POA

[MLR10] Barbara A. Muhling, John T. Lamkin, and Mitchell A. Roffer. Predicting the occurrence of Atlantic bluefin tuna (Thunnus thynnus) larvae in the northern Gulf of Mexico: building a classification model from archival data. Fisheries Oceanography, 19(6):526–539, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mills:2007:DTP

[MLRS07] K. L. Mills, T. Laidig, S. Ralston, and W. J. Sydeman. Diets of top predators indicate pelagic juvenile rockfish (Sebastes spp.) abundance in the California Current System. Fisheries Oceanography, 16(3):273–283, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Maes:2005:SEI

[MLVO05] Joachim Maes, Karin E. Limburg, Anton Van De Putte, and Frans Ollevier. A spatially explicit, individual-based model to assess the role of estuarine nurseries in the early life history of North Sea herring, Clupea harengus. Fisheries Oceanography, 14(1):17–31, January 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mertz:1994:EIG

[MM94a] G. Mertz and R. A. Myers. The ecological impact of the Great Salinity Anomaly in the northern North-west Atlantic. Fisheries Oceanography, 3(1):1–14, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mertz:1994:MMP

[MM94b] G. Mertz and R. A. Myers. Match/mismatch predictions of spawning duration versus recruitment variability. Fisheries Oceanography, 3(4):236–245, December 1994. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mckinnell:2003:ISN

[MM03] S. M. Mckinnell and D. L. Mackas. Intercalibrating SCOR, NORPAC and bongo nets and the consequences for interpreting decadal-scale variation in zooplankton biomass in the Gulf of Alaska. Fisheries Oceanography, 12(2):126–133, March 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Myers:1993:CSR

[MMB93] R. A. Myers, G. Mertz, and C. A. Bishop. Cod spawning in relation to physical and biological cycles of the northern North-west Atlantic. Fisheries Oceanography, 2(3–4):154– 165, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mullaney:2011:ELF

[MMB^{+}11] T. J. Mullaney, A. G. Miskiewicz, M. E. Baird, P. T. P. Burns, and I. M. Suthers. Entrainment of larval fish assemblages from the inner shelf into the East Australian Current and into the western Tasman Front. Fisheries Oceanography, 20(6):434–447, September 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Menard:2007:COT

[MMBC07] F. Ménard, F. Marsac, E. Bellier, and B. Cazelles. Climatic oscillations and tuna catch rates in the Indian Ocean: a wavelet approach to time series analysis. Fisheries Oceanography, 16(1):95–104, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miyake:1995:LDS

[MMF95] Hideo Miyake, Mitsuaki Matsuoka, and Kunio Furuya. Loss of and damage to scallops due to storms in the Sea of Okhotsk. Fisheries Oceanography, 4(4):293–302, December 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mori:2022:MEL

[MMI^{+}22] Mao Mori, Kohei Mizobata, Taro Ichii, Philippe Ziegler, and Takehiro Okuda. Modeling the egg and larval transport pathways of the Antarctic toothfish (Dissostichus mawsoni) in the East Antarctic region: New insights into successful transport connections. Fisheries Oceanography, 31(1):19–39, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miller:2014:RAA

[MMMS14] Alicia S. Miller, Timothy J. Miller, Katherine E. Mills, and Timothy F. Sheehan. Retrospective analysis of Atlantic salmon (Salmo salar) marine growth and condition in the northwest Atlantic based on tag-recovery data. Fisheries Oceanography, 23(2):103–115, March 2014. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Montero:2016:CES

[MMRH^{+}16] Jose T. Montero, Raul O. Martinez-Rincon, Selina S. Heppell, Martin Hall, and Michael Ewal. Characterizing environmental and spatial variables associated with the incidental catch of olive ridley (Lepidochelys olivacea) in the Eastern Tropical Pacific purse-seine fishery. Fisheries Oceanography, 25(1):1–14, January 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Martin:2016:CFB

[MMRS16] Paloma Martin, Francesc Maynou, Laura Recasens, and Ana Sabatés. Cyclic fluctuations of blue whiting (Micromesistius poutassou) linked to open-sea convection processes in the northwestern Mediterranean. Fisheries Oceanography, 25(3): 229–240, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Marrari:2019:IEC

[MMSL19] Marina Marrari, Gustavo J. Macchi, Betina Santos, and Ezequiel Leonarduzzi. Influence of environmental conditions on the reproductive success and recruitment of the Argentine hake Merluccius hubbsi (southwestern Atlantic Ocean). Fisheries Oceanography, 28(1):66–81, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Maynou:2006:PEL

[MOE06] F. Maynou, M. P. Olivar, and M. Emelianov. Patchiness of eggs, larvae and juveniles of European hake Merluccius merluccius from the NW Mediterranean. Fisheries Oceanography, 15(5):390–401, September 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Morita:2011:BST

[Mor11] K. Morita. Body size trends along vertical and thermal gradients of chum salmon in the Bering Sea during summer. Fisheries Oceanography, 20(3):258–262, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Myers:1994:RVO

[MP94] Ransom A. Myers and Pierre Pepin. Recruitment variability and oceanographic stability. Fisheries Oceanography, 3(4):

246–255, December 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miesner:2018:OVS

[MP18] Anna K. Miesner and Mark R. Payne. Oceanographic variability shapes the spawning distribution of blue whiting (Micromesistius poutassou). Fisheries Oceanography, 27(6):623– 638, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Munroe:2013:UPP

[MPM^{+}13] D. M. Munroe, E. N. Powell, R. Mann, J. M. Klinck, and E. E. Hofmann. Underestimation of primary productivity on continental shelves: evidence from maximum size of extant surfclam (Spisula solidissima) populations. Fisheries Oceanography, 22(3):220–233, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Martinez-Pita:2019:IBF

[MPM19] Inés Martínez-Pita and Oscar Moreno. Improving bivalve fishery management: The development of a tool to forecast bivalve reproductive season based on environmental data from satellite remote sensing observations. Fisheries Oceanography, 28(6):698–707, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Marsh:1999:TKC

[MPW^{+}99] Robert Marsh, Brian Petrie, Christopher R. Weidman, Robert R. Dickson, John W. Loder, Charles G. Hannah, Kenneth Frank, and Ken Drinkwater. The 1882 tilefish kill — a cold event in shelf waters off the north-eastern United States? Fisheries Oceanography, 8(1):39–49, March 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Moyano:2014:LFD

[MRBBHL14] M. Moyano, J. M. Rodríguez, V. M. Benítez-Barrios, and S. Hernández-León. Larval fish distribution and retention in the Canary Current system during the weak upwelling season. Fisheries Oceanography, 23(3):191–209, May 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mullowney:2019:TIG

[MRD^{+}19] Darrell R. J. Mullowney, George A. Rose, Earl G. Dawe, Sherrylynn Rowe, Gary L. Maillet, and Eric J. Pedersen. Temperature influences on growth of unfished juvenile northern cod (Gadus morhua) during stock collapse. Fisheries Oceanography, 28(5):612–627, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Moyano:2009:LFA

[MRHL09] M. Moyano, J. M. Rodríguez, and S. Hernández-León. Larval fish abundance and distribution during the late winter bloom off Gran Canaria Island, Canary Islands. Fisheries Oceanography, 18(1):51–61, January 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Martins:2014:MTC

[MRL^{+}14] Rodrigo S. Martins, Michael J. Roberts, Christophe Lett, Nicolette Chang, Coleen L. Moloney, Maurício G. Camargo, and Erica A. G. Vidal. Modelling transport of chokka squid (Loligo reynaudii) paralarvae off South Africa: reviewing, testing and extending the ‘Westward Transport Hypothesis’. Fisheries Oceanography, 23(2):116–131, March 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Macfarlane:2005:JCS

[MRRN05] R. Bruce Macfarlane, Stephen Ralston, Chantell Royer, and Elizabeth C. Norton. Juvenile Chinook salmon (Oncorhynchus tshawytscha) growth on the central California coast during the 1998 El Niño and 1999 La Niña. Fisheries Oceanography, 14(5):321–332, September 2005. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mohan:2017:IOC

[MSC^{+}17] John A. Mohan, Tracey T. Sutton, April B. Cook, Kevin M. Boswell, and R. J. David Wells. Influence of oceanographic conditions on abundance and distribution of post-larval and juvenile carangid fishes in the northern Gulf of Mexico. Fisheries Oceanography, 26(5):526–541, September 2017. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mordy:2005:NPP

[MSL^{+}05] Calvin W. Mordy, Phyllis J. Stabeno, Carol Ladd, Stephan Zeeman, David P. Wisegarver, Sigrid A. Salo, and George L. Hunt, Jr. Nutrients and primary production along the eastern Aleutian Island Archipelago. Fisheries Oceanography, 14 (S1):55–76, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Midway:2020:MDA

[MSL^{+}20] Stephen R. Midway, Amy M. Schueller, Robert T. Leaf, Geneviève M. Nesslage, and Raymond M. Mroch III. Macroscale drivers of Atlantic and Gulf menhaden growth. Fisheries Oceanography, 29(3):252–264, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Marrari:2013:RSA

[MSM^{+}13] Marina Marrari, Sergio R. Signorini, Charles R. McClain, Marcelo Pájaro, Patricia Martos, María Delia Viñas, Jorge Hansen, Roxana Dimauro, Georgina Cepeda, and Claudio Buratti. Reproductive success of the Argentine anchovy, Engraulis anchoita, in relation to environmental variability at a mid-shelf front (Southwestern Atlantic Ocean). Fisheries Oceanography, 22(3):247–261, May 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mugo:2010:HCS

[MSNK10] Robinson Mugo, Sei-Ichi Saitoh, Akira Nihira, and Tadaaki Kuroyama. Habitat characteristics of skipjack tuna (Katsuwonus pelamis) in the western North Pacific: a remote sensing perspective. Fisheries Oceanography, 19(5):382–396, September 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Maynou:2020:CSH

[MSR20] Francesc Maynou, Ana Sabatés, and Vanesa Raya. Changes in the spawning habitat of two small pelagic fish in the Northwestern Mediterranean. Fisheries Oceanography, 29 (2):201–213, March 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Martinson:2012:UJS

[MSS12] Ellen C. Martinson, Houston H. Stokes, and Dennis L. Scarnecchia. Use of juvenile salmon growth and temperature change indices to predict groundfish post age-0 yr class strengths in the Gulf of Alaska and eastern Bering Sea. Fisheries Oceanography, 21(4):307–319, July 2012. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Matsumoto:2016:CBS

[MSST16] Takayuki Matsumoto, Keisuke Satoh, Yasuko Semba, and Mikio Toyonaga. Comparison of the behavior of skipjack (Katsuwonus pelamis), yellowfin (Thunnus albacares) and bigeye (T. obesus) tuna associated with drifting FADs in the equatorial central Pacific Ocean. Fisheries Oceanography, 25 (6):565–581, November 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Muhling:2013:LFA

[MSVY^{+}13] Barbara A. Muhling, Ryan H. Smith, Lourdes Vásquez-Yeomans, John T. Lamkin, Elizabeth M. Johns, Laura Carrillo, Eloy Sosa-Cordero, and Estrella Malca. Larval fish assemblages and mesoscale oceanographic structure along the Mesoamerican Barrier Reef System. Fisheries Oceanography, 22(5):409–428, September 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miyashita:2004:DCV

[MTH^{+}04] Kazushi Miyashita, Koutarou Tetsumura, Satoshi Honda, Tatsuki Oshima, Ryo Kawabe, and Kei Sasaki. Diel changes in vertical distribution patterns of zooplankton and walleye pollock (Theragra chalcogramma) off the Pacific coast of eastern Hokkaido, Japan, estimated by the volume back scattering strength (Sv) difference method. Fisheries Oceanography, 13(S1):99–110, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Murase:2007:PSC

[MTK^{+}07] Hiroto Murase, Tsutomu Tamura, Hiroshi Kiwada, Yoshihiro Fujise, Hikaru Watanabe, Hiroshi Ohizumi, Shiroh Yonezaki, Hiroshi Okamura, and Shigeyuki Kawahara. Prey selection of common minke (Balaenoptera acutorostrata) and Bryde’s (Balaenoptera edeni) whales in the western North Pacific in 2000 and 2001. Fisheries Oceanography, 16(2):186–201, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ma:2022:ITN

[MTL^{+}22] Shuyang Ma, Yongjun Tian, Jianchao Li, Peilong Ju, Peng Sun, Zhenjiang Ye, Yang Liu, and Yoshiro Watanabe. Incorporating thermal niche to benefit understanding climateinduced biological variability in small pelagic fishes in the Kuroshio ecosystem. Fisheries Oceanography, 31(2):172–190, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Molinero:2016:CDW

[MTLL^{+}16] Juan Carlos Molinero, Li-Chun Tseng, Lucia Lopez-Lopez, Ulrich Sommer, Sami Souissi, and Jiang-Shiou Hwang. Climate-driven winter variations of Calanus sinicus abundance in the East China Sea. Fisheries Oceanography, 25(6): 555–564, November 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Maravelias:2007:EMD

[MTP07] Christos D. Maravelias, Efthymia V. Tsitsika, and Costas Papaconstantinou. Evidence of Morocco dentex (Dentex maroccanus) distribution in the NE Mediterranean and relationships with environmental factors determined by generalized additive modelling. Fisheries Oceanography, 16(3): 294–302, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Michael:2015:EAB

[MTSH15] Pamela E. Michael, Geoffrey N. Tuck, Peter Strutton, and Alistair Hobday. Environmental associations with broadscale Japanese and Taiwanese pelagic longline effort in the southern Indian and Atlantic Oceans. Fisheries Oceanography, 24(5):478–493, September 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Muraoka:2017:EGE

[MTT^{+}17] Daisuke Muraoka, Hitoshi Tamaki, Hideki Takami, Yutaka Kurita, and Tomohiko Kawamura. Effects of the 2011 Great East Japan Earthquake and tsunami on two kelp bed communities on the Sanriku coast. Fisheries Oceanography, 26 (2):128–140, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

McAskill:2023:NLF

[MTZG23] Shannan McAskill, James M. Tolan, Glenn A. Zapfe, and Simon J. Geist. Neritic larval fish assemblages across the Texas shelf in the northwestern Gulf of Mexico following Hurricane Harvey. Fisheries Oceanography, 32(3):280–292, May 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mullin:1994:DRP

[Mul94] Michael M. Mullin. Distribution and reproduction of the planktonic copepod, Calanus pacificus, off southern California during winter-spring of 1992, relative to 1989–91. Fisheries Oceanography, 3(2):142–157, June 1994. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mullin:1997:DCP

[Mul97] Michael Mullin. The demography of Calanus pacificus during winter–spring Californian El Niño conditions, 1992–1993: implications for anchovy? Fisheries Oceanography, 6(1):10– 18, March 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Miyamoto:2020:GVF

[MVK^{+}20] Hiroomi Miyamoto, Dharmamony Vijai, Hideaki Kidokoro, Kazuaki Tadokoro, Tsuyoshi Watanabe, Taiki Fuji, and Satoshi Suyama. Geographic variation in feeding of Pacific saury Cololabis saira in June and July in the North Pacific Ocean. Fisheries Oceanography, 29(6):558–571, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcgurk:1992:PSL

[MW92] Michael D. Mcgurk and H. David Warburton. Pacific sand lance of the Port Moller estuary, southeastern Bering Sea: an estuarine-dependent early life history. Fisheries Oceanography, 1(4):306–320, December 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Morgan:2000:SVR

[MWB^{+}00] Lance E. Morgan, Stephen R. Wing, Louis W. Botsford, Carolyn J. Lundquist, and Jennifer M. Diehl. Spatial variability in red sea urchin (Strongylocentrotus franciscanus) recruitment in northern California. Fisheries Oceanography, 9(1): 83–98, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mcgurk:1992:RDR

[MWGK92] M. D. Mcgurk, H. D. Warburton, M. Galbraith, and W. C. Kusser. RNA–DNA ratio of herring and sand lance larvae from Port Moller, Alaska: Comparison with prey concentration and temperature. Fisheries Oceanography, 1(3):193– 207, September 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Minegishi:2023:SSP

[MWN^{+}23] Yuki Minegishi, Marty Kwok-Shing Wong, Mako Nakao, Yuichiro Nishibe, Aiko Tachibana, Yoo-Jun Kim, and Susumu Hyodo. Species-specific patterns in spatio-temporal dynamics of juvenile chum salmon and their zooplankton prey in Otsuchi Bay, Japan, revealed by simultaneous eDNA quantification of diverse taxa from the same water samples. Fisheries Oceanography, 32(3):311–326, May 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mueter:2002:SCP

[MWP02] Franz J. Mueter, Dan M. Ware, and Randall M. Peterman. Spatial correlation patterns in coastal environmental variables and survival rates of salmon in the north-east Pacific Ocean. Fisheries Oceanography, 11(4):205–218, July 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Murphy:1998:IVS

[MWR^{+}98] E. J. Murphy, J. L. Watkins, K. Reid, P. N. Trathan, I. Everson, J. P. Croxall, J. Priddle, M. A. Brandon, A. S. Brierley, and E. Hofmann. Interannual variability of the South Georgia marine ecosystem: biological and physical sources of variation in the abundance of krill. Fisheries Oceanography, 7(3–4):381–390, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Mhlongo:2015:SHP

[MYHvdL15] Nandipha Mhlongo, Dawit Yemane, Marc Hendricks, and Carl D. van der Lingen. Have the spawning habitat preferences of anchovy (Engraulis encrasicolus) and sardine (Sardinops sagax) in the southern Benguela changed in recent years? Fisheries Oceanography, 24(S1):1–14, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:1998:TSR

[Nak98] Hideaki Nakata. Towards a science of recruitment in fish populations. Fisheries Oceanography, 6(4):272–274, February 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nelson:2010:THS

[NASTF10] Gary A. Nelson, Michael P. Armstrong, Jennifer Stritzel-Thomson, and Kevin D. Friedland. Thermal habitat of striped bass (Morone saxatilis) in coastal waters of northern Massachusetts, USA, during summer. Fisheries Oceanography, 19(5):370–381, September 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Norcross:2001:SLH

[NBF^{+}01] Brenda L. Norcross, Evelyn D. Brown, Robert J. Foy, Michele Frandsen, Shelton M. Gay, Thomas C. Kline, Jr., Doran M. Mason, E. Vincent Patrick, A. J. Paul, and Kevin D. E. Stokesbury. A synthesis of the life history and ecology of juvenile Pacific herring in Prince William Sound, Alaska. Fisheries Oceanography, 10(S1):42–57, December 2001. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Norcross:1999:CMD

[NBH99] Brenda L. Norcross, Arny Blanchard, and Brenda A. Holladay. Comparison of models for defining nearshore flatfish nursery areas in Alaskan waters. Fisheries Oceanography, 8 (1):50–67, March 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nielsen:2006:ILB

[NBMS06] Anders Nielsen, Keith A. Bigelow, Michael K. Musyl, and John R. Sibert. Improving light-based geolocation by including sea surface temperature. Fisheries Oceanography, 15 (4):314–325, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nash:2005:ILH

[NDC05] R. D. M. Nash and M. Dickey-Collas. The influence of life history dynamics and environment on the determination of year class strength in North Sea herring (Clupea harengus L.). Fisheries Oceanography, 14(4):279–291, July 2005. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nobrega:2023:IBO

[NdLOO23] Marcelo Francisco Nóbrega, Marcelo Gomes de Lira, Matheus Assis Oliveira, and Jorge Eduardo Lins Oliveira. Interactions between oceanographic variables and population structure of the yellowfin tuna Thunnus albacares (Bonnaterre, 1788) in the Western Central Atlantic. Fisheries Oceanography, 32(2):213–228, March 2023. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Neuenfeldt:2002:IOS

[Neu02] Stefan Neuenfeldt. The influence of oxygen saturation on the distributional overlap of predator (cod, Gadus morhua) and prey (herring, Clupea harengus) in the Bornholm Basin of the Baltic Sea. Fisheries Oceanography, 11(1):11–17, January 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Niino:2021:SSP

[NFKY21] Yohei Niino, Sho Furuichi, Yasuhiro Kamimura, and Ryuji Yukami. Spatiotemporal spawning patterns and early growth of Japanese sardine in the western North Pacific during the recent stock increase. Fisheries Oceanography, 30(6):643– 652, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:2000:ADP

[NFN00] Hideaki Nakata, Shigeo Funakoshi, and Motohiko Nakamura. Alternating dominance of postlarval sardine and anchovy caught by coastal fishery in relation to the Kuroshio meander in the Enshu-nada Seao. Fisheries Oceanography, 9(3):248– 258, September 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nilsen:2023:TGD

[NFO^{+}23] Ina Nilsen, Filippa Fransner, Are Olsen, Jerry Tjiputra, Robinson Hordoir, and Cecilie Hansen. Trivial gain of downscaling in future projections of higher trophic levels in the Nordic and Barents Seas. Fisheries Oceanography, 32(5): 479–493, September 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Neuheimer:2009:MLC

[NGGJ09] A. B. Neuheimer, W. C. Gentleman, C. L. Galloway, and C. L. Johnson. Modeling larval Calanus finmarchicus on Georges Bank: time-varying mortality rates and a cannibalism hypothesis. Fisheries Oceanography, 18(3):147–160, May 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Napp:2001:ACS

[NH01] Jeffrey M. Napp and George L. Hunt, Jr. Anomalous conditions in the south-eastern Bering Sea 1997: linkages among climate, weather, ocean, and biology. Fisheries Oceanography, 10(1):61–68, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:2003:DSV

[NH03] Kaoru Nakata and Kiyotaka Hidaka. Decadal-scale variability in the Kuroshio marine ecosystem in winter. Fisheries Oceanography, 12(4–5):234–244, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

North:2006:RMW

[NH06] E. W. North and E. D. Houde. Retention mechanisms of white perch (Morone americana) and striped bass (Morone saxatilis) early-life stages in an estuarine turbidity maximum: an integrative fixed-location and mapping approach. Fisheries Oceanography, 15(6):429–450, November 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:1994:VFA

[NHM94] Kaoru Nakata, Akio Hada, and Yasuo Matsukawa. Variations in food abundance for Japanese sardine larvae related to the Kuroshio meander. Fisheries Oceanography, 3(1):39– 49, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Neuenfeldt:2007:RMI

[NHNA07] Stefan Neuenfeldt, Hans-Harald Hinrichsen, Anders Nielsen, and Ken H. Andersen. Reconstructing migrations of individual cod (Gadus morhua L.) in the Baltic Sea by using electronic data storage tags. Fisheries Oceanography, 16(6): 526–535, November 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishimura:2007:IVH

[NHS^{+}07] Akira Nishimura, Tomonori Hamatsu, Osamu Shida, Ikuo Mihara, and Takashi Mutoh. Interannual variability in hatching period and early growth of juvenile walleye pollock, Theragra chalcogramma, in the Pacific coastal area of Hokkaido. Fisheries Oceanography, 16(3):229–239, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2014:IPJ

[NII^{+}14] Haruka Nishikawa, Hiromichi Igarashi, Yoichi Ishikawa, Mitsuo Sakai, Yoshiki Kato, Masahito Ebina, Norihisa Usui, Masafumi Kamachi, and Toshiyuki Awaji. Impact of paralarvae and juveniles feeding environment on the neon flying squid (Ommastrephes bartramii) winter–spring cohort stock. Fisheries Oceanography, 23(4):289–303, July 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:2004:SCR

[NIIS04] Kaoru Nakata, Hiroshi Itoh, Tadafumi Ichikawa, and Katsuyuki Sasaki. Seasonal changes in the reproduction of three oncaeid copepods in the surface layer of the Kuroshio Extension. Fisheries Oceanography, 13(S1):21–33, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishida:1992:CSS

[Nis92] Tsutomu Nishida. Considerations of stock structure of yellowfin tuna (Thunnus albacares) in the Indian Ocean based on fishery data. Fisheries Oceanography, 1(2):143–152, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2019:RBR

[Nis19] Haruka Nishikawa. Relationship between recruitment of Japanese sardine (Sardinops melanostictus) and environment of larval habitat in the low-stock period (1995–2010). Fisheries Oceanography, 28(2):131–142, March 2019. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Neira:2008:IBS

[NK08] Francisco J. Neira and John P. Keane. Ichthyoplanktonbased spawning dynamics of blue mackerel (Scomber australasicus) in south-eastern Australia: links to the East Australian Current. Fisheries Oceanography, 17(4):281–298, July 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:2001:IVS

[NKM01] K. Nakata, S. Koyama, and Y. Matsukawa. Interannual variation in spring biomass and gut content composition of copepods in the Kuroshio Current, 1971–89. Fisheries Oceanography, 10(4):329–341, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Napp:2000:SBP

[NKS00] J. M. Napp, A. W. Kendall, Jr., and J. D. Schumacher. A synthesis of biological and physical processes affecting the feeding environment of larval walleye pollock (Theragra chalcogramma) in the eastern Bering Sea. Fisheries Oceanography, 9(2):147–162, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nesslage:2021:EDG

[NLN^{+}21] Geneviève Nesslage, Vyacheslav Lyubchich, Paul Nitschke, Erik Williams, Churchill Grimes, and John Wiedenmann. Environmental drivers of golden tilefish (Lopholatilus chamaeleonticeps) commercial landings and catch-per-unit-effort. Fisheries Oceanography, 30(5):608–622, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nguyen:2024:REP

[NLS^{+}24] Bich Vi Viviane Nguyen, Yonggang Liu, Christopher D. Stallings, Mya Breitbart, Steven A. Murawski, Robert H. Weisberg, Makenzie Kerr, Eva-Maria S. Bønnelycke, and Ernst B. Peebles. Retention and export of planktonic fish eggs in the northeastern Gulf of Mexico. Fisheries Oceanography, 33(1):e12655:1–e12655:??, January 2024. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2020:IDC

[NNOU20] Tetsuya Nishikawa, Yukinobu Nakamura, Shigeyoshi Okamoto, and Hiroshi Ueda. Interannual decrease in condition factor of the western sand lance Ammodytes japonicus in Japan in the last decade: Evidence for food-limited decline of the catch. Fisheries Oceanography, 29(1):52–55, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nataniel:2022:MIC

[NPLS22] Anildo Nataniel, Maria Grazia Pennino, Jon Lopez, and Maria Soto. Modelling the impacts of climate change on skipjack tuna (Katsuwonus pelamis) in the Mozambique Channel. Fisheries Oceanography, 31(2):149–163, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nickels:2023:JAT

[NPS^{+}23] Catherine F. Nickels, Elan J. Portner, Owyn Snodgrass, Barbara Muhling, and Heidi Dewar. Juvenile albacore tuna (Thunnus alalunga) foraging ecology varies with environmental conditions in the California Current Large Marine Ecosystem. Fisheries Oceanography, 32(5):431–447, September 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Naranjo:2015:FJM

[NPY^{+}15] Laura Naranjo, Francisco Plaza, Eleuterio Yáñez, María Ángela Barbieri, and Felipe Sánchez. Forecasting of jack mackerel landings (Trachurus murphyi) in central-southern Chile through neural networks. Fisheries Oceanography, 24(3): 219–228, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nadal:2022:HCD

[NSGL^{+}22] Irene Nadal, Simone Sammartino, Jesús García-Lafuente, José C. Sánchez Garrido, Juan Gil-Herrera, Manuel Hidalgo, and Pilar Hernández. Hydrodynamic connectivity and dispersal patterns of a transboundary species (Pagellus bogaraveo) in the Strait of Gibraltar and adjacent basins. Fisheries Oceanography, 31(4):384–401, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Narimatsu:2017:EMP

[NSH^{+}17] Yoji Narimatsu, Yasutoki Shibata, Tsutomu Hattori, Toshikazu Yano, and Jiro Nagao. Effects of a marineprotected area occurred incidentally after the Great East Japan Earthquake on the Pacific cod (Gadus macrocephalus) population off northeastern Honshu, Japan. Fisheries Oceanography, 26(2):181–192, March 2017. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakayama:2018:CCI

[NTIO18] Shin-Ichiro Nakayama, Akinori Takasuka, Momoko Ichinokawa, and Hiroshi Okamura. Climate change and interspecific interactions drive species alternations between anchovy and sardine in the western North Pacific: Detection of causality by convergent cross mapping. Fisheries Oceanography, 27(4):312–322, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2015:WIS

[NTM^{+}15] Haruka Nishikawa, Takahiro Toyoda, Shuhei Masuda, Yoichi Ishikawa, Yuji Sasaki, Hiromichi Igarashi, Mitsuo Sakai, Masaki Seito, and Toshiyuki Awaji. Wind-induced stock variation of the neon flying squid (Ommastrephes bartramii) winter–spring cohort in the subtropical North Pacific Ocean. Fisheries Oceanography, 24(3):229–241, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Noto:2003:EBM

[NY03] Masayuki Noto and Ichiro Yasuda. Empirical biomass model for the Japanese sardine, Sardinops melanostictus, with sea surface temperature in the Kuroshio Extension. Fisheries Oceanography, 12(1):1–9, January 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2008:JSS

[NY08] Haruka Nishikawa and Ichiro Yasuda. Japanese sardine (Sardinops melanostictus) mortality in relation to the winter mixed layer depth in the Kuroshio Extension region. Fisheries Oceanography, 17(5):411–420, September 2008. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2011:IWS

[NYI11] Haruka Nishikawa, Ichiro Yasuda, and Sachihiko Itoh. Impact of winter-to-spring environmental variability along the Kuroshio jet on the recruitment of Japanese sardine (Sardinops melanostictus). Fisheries Oceanography, 20(6): 570–582, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nishikawa:2013:TSJ

[NYI^{+}13] Haruka Nishikawa, Ichiro Yasuda, Sachihiko Itoh, Kosei Komatsu, Hideharu Sasaki, Yoshikazu Sasai, and Yoshioki Oozeki. Transport and survival of Japanese sardine (Sardinops melanostictus) eggs and larvae via particletracking experiments. Fisheries Oceanography, 22(6):509– 522, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Nakata:1995:DFA

[NZI95] K. Nakata, H. Zenitani, and D. Inagake. Differences in food availability for Japanese sardine larvae between the frontal region and the waters on the offshore side of Kuroshio. Fisheries Oceanography, 4(1):68–79, March 1995. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ouellet:2006:VDB

[OA06] Patrick Ouellet and Jean-Pierre Allard. Vertical distribution and behaviour of shrimp Pandalus borealis larval stages in thermally stratified water columns: laboratory experiment and field observations. Fisheries Oceanography, 15(5):373– 389, September 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ottersen:2000:PTB

[OÅL00] Geir Ottersen, Bjørn Ådlandsvik, and Harald Loeng. Predicting the temperature of the Barents Sea. Fisheries Oceanography, 9(2):121–135, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Overland:2001:NPA

[OBA01] James E. Overland, Nicholas A. Bond, and Jennifer Miletta Adams. North Pacific atmospheric and SST anomalies in 1997: Links to ENSO? Fisheries Oceanography, 10(1):69– 80, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ortega-Cisneros:2018:EEC

[OCCF^{+}18] Kelly Ortega-Cisneros, Kevern L. Cochrane, Elizabeth A. Fulton, Rebecca Gorton, and Ekaterina Popova. Evaluating the effects of climate change in the southern Benguela upwelling system using the Atlantis modelling framework. Fisheries Oceanography, 27(5):489–503, September 2018. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ortner:1999:SAB

[OCH99] Peter B. Ortner, Larry B. Crowder, and Donald E. Hoss. The South Atlantic Bight recruitment experiment: introduction and overview. Fisheries Oceanography, 8(S2):1–6, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Odate:1994:PAS

[Oda94] Tsuneo Odate. Plankton abundance and size structure in the northern North Pacific Ocean in early summer. Fisheries Oceanography, 3(4):267–278, December 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Oguz:1998:SBS

[ODMRM98] Temel Oguz, Hugh W. Ducklow, Paola Malanotte-Rizzoli, and James W. Murray. Simulations of the Black Sea pelagic ecosystem by 1-D, vertically resolved, physical–biochemical models. Fisheries Oceanography, 7(3–4):300–304, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ohta:2017:IAV

[OE17] Itaru Ohta and Akihiko Ebisawa. Inter-annual variation of the spawning aggregations of the white-streaked grouper Epinephelus ongus, in relation to the lunar cycle and water temperature fluctuation. Fisheries Oceanography, 26(3):350– 363, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Olivar:2010:ROC

[OEV^{+}10] M. Pilar Olivar, Mikhail Emelianov, Fernando Villate, Ibon Uriarte, Francesc Maynou, Ignacio Álvarez, and Elvira Morote. The role of oceanographic conditions and plankton availability in larval fish assemblages off the Catalan coast (NW Mediterranean). Fisheries Oceanography, 19(3):209– 229, May 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ohshimo:2016:DBL

[OFS^{+}16] Seiji Ohshimo, Yuki Fujinami, Ko Shiozaki, Mikihiko Kai, Yasuko Semba, Nobuhiro Katsumata, Daisuke Ochi, Hiroaki Matsunaga, Hiroshi Minami, Masashi Kiyota, and Kotaro Yokawa. Distribution, body length, and abundance of blue shark and shortfin mako offshore of northeastern Japan, as determined from observed pelagic longline data, 2000–2014. Fisheries Oceanography, 25(3):259–276, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ottersen:2023:LTV

[OH23] Geir Ottersen and Rebecca E. Holt. Long-term variability in spawning stock age structure influences climate–recruitment link for Barents Sea cod. Fisheries Oceanography, 32(1): 91–105, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ogburn:2012:BCC

[OHF12] Matthew B. Ogburn, Megan Hall, and Richard B. Forward, Jr. Blue crab (Callinectes sapidus) larval settlement in North Carolina: environmental forcing, recruit–stock relationships, and numerical modeling. Fisheries Oceanography, 21(2–3): 20–32, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Onitsuka:2010:NSM

[OHM^{+}10] G. Onitsuka, N. Hirose, K. Miyahara, T. Ota, J. Hatayama, Y. Mitsunaga, and T. Goto. Numerical simulation of the migration and distribution of diamond squid (Thysanoteuthis rhombus) in the southwest Sea of Japan. Fisheries Oceanography, 19(1):63–75, January 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ottersen:2006:CSS

[OHS06] Geir Ottersen, Dag Ø. Hjermann, and Nils Chr. Stenseth. Changes in spawning stock structure strengthen the link between climate and recruitment in a heavily fished cod (Gadus morhua) stock. Fisheries Oceanography, 15(3):230–243, May 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Okunishi:2012:MAE

[OIA^{+}12] Takeshi Okunishi, Shin-Ichi Ito, Daisuke Ambe, Akinori Takasuka, Takahiko Kameda, Kazuaki Tadokoro, Takashi Setou, Kosei Komatsu, Atsushi Kawabata, Hiroshi Kubota, Tadafumi Ichikawa, Hiroya Sugisaki, Taketo Hashioka, Yasuhiro Yamanaka, Naoki Yoshie, and Tomowo Watanabe. A modeling approach to evaluate growth and movement for recruitment success of Japanese sardine (Sardinops melanostictus) in the western Pacific. Fisheries Oceanography, 21 (2–3):44–57, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Okumura:2017:RAP

[OK17] Yutaka Okumura and Shinnosuke Kaga. Retrospective analysis of phytoplankton assemblages on the iwate coast before and after the 2011 tsunami using cryopreserved DNA samples. Fisheries Oceanography, 26(2):234–250, March 2017.

CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ohshimo:2023:DGD

[OKT^{+}23] Seiji Ohshimo, Taketoshi Kodama, Atsushi Tawa, Hiroshige Tanaka, Yosuke Tanaka, Go Yasuhara, Michihiro Tokuyasu, and Yoichi Minami. Distribution and geographic differences in stable isotope ratios of diamond squid Thysanoteuthis rhombus (Cephalopoda: Thysanoteuthidae) in Japanese waters. Fisheries Oceanography, 32(3):255–266, May 2023. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Okazaki:2017:LEM

[OKU17] Yuji Okazaki, Yutaka Kurita, and Shinji Uehara. Limited effect of the massive tsunami caused by the 2011 Great East Japan Earthquake on the shallow sandy shore demersal fish assemblages in Sendai Bay. Fisheries Oceanography, 26(2): 155–164, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Olson:2001:BDW

[Ols01] Donald B. Olson. Biophysical dynamics of western transition zones: a preliminary synthesis. Fisheries Oceanography, 10 (2):133–150, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ohizumi:2010:DSI

[OM10] Hiroshi Ohizumi and Nobuyuki Miyazaki. Differences in stable isotope ratios of Dall’s porpoises (Phocoenoides dalli) between coastal and oceanic areas of the North Pacific. Fisheries Oceanography, 19(4):257–261, July 2010. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Oozeki:2017:FER

[ONK17] Yoshioki Oozeki, Kaoru Nakata, and Michio J. Kishi. Fisheries effects and recovery from the earthquake and tsunami of the Great East Japan Earthquake. Fisheries Oceanography, 26(2):97–98, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

oskarsson:2018:OQT

[óóSV18] Gudmundur J. óskarsson, Sólveig R. ólafsdóttir, Torsteinn Sigurdsson, and Hédinn Valdimarsson. Observation and quantification of two incidents of mass fish kill of Icelandic summer spawning herring (Clupea harengus) in the winter 2012/2013. Fisheries Oceanography, 27(4):302–311, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Olafsdottir:2012:ITB

[OR12] Anna H. Olafsdottir and George A. Rose. Influences of temperature, bathymetry and fronts on spawning migration routes of Icelandic capelin (Mallotus villosus). Fisheries Oceanography, 21(2–3):182–198, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Olafsdottir:2013:SSM

[OR13] Anna H. Olafsdottir and George A. Rose. Staged spawning migration in Icelandic capelin (Mallotus villosus): effects of temperature, stock size and maturity. Fisheries Oceanography, 22(6):446–458, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ottersen:1995:ETW

[OS95] Geir Ottersen and Svein Sundby. Effects of temperature, wind and spawning stock biomass on recruitment of arcto-Norwegian cod. Fisheries Oceanography, 4(4):278–292, December 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

oskarsson:2010:VRP

[óT10] Gudmundur J. óskarsson and Christopher T. Taggart. Variation in reproductive potential and influence on Icelandic herring recruitment. Fisheries Oceanography, 19(5):412–426, September 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ohshimo:2009:LTS

[OTH09] Seiji Ohshimo, Hiroshige Tanaka, and Yoshiaki Hiyama. Long-term stock assessment and growth changes of the Japanese sardine (Sardinops melanostictus) in the Sea of Japan and East China Sea from 1953 to 2006. Fisheries Oceanography, 18(5):346–358, September 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

OLeary:2020:ACD

[OTIK20] Cecilia A. O’Leary, James T. Thorson, James N. Ianelli, and Stan Kotwicki. Adapting to climate-driven distribution shifts using model-based indices and age composition from multiple surveys in the walleye pollock (Gadus chalcogrammus) stock assessment. Fisheries Oceanography, 29(6):541–557, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Oozeki:2009:PSM

[OTO^{+}09] Yoshioki Oozeki, Akinori Takasuka, Hiroshi Okamura, Hiroshi Kubota, and Ryo Kimura. Patchiness structure and mortality of Pacific saury Cololabis saira larvae in the northwestern Pacific. Fisheries Oceanography, 18(5):328–345, September 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ozaki:2004:IVE

[OUKH04] Koji Ozaki, Shin-Ichi Uye, Teruichi Kusumoto, and Tetsuo Hagino. Interannual variability of the ecosystem of the Kii Channel, the Inland Sea of Japan, as influenced by bottom intrusion of cold and nutrient-rich water from the Pacific Ocean, and a recent trend of warming and oligotrophication. Fisheries Oceanography, 13(1):65–79, January 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Oozeki:2003:GRV

[OWK^{+}03] Yoshioki Oozeki, Yoshiro Watanabe, Yutaka Kurita, Kaoru Nakata, and Daiji Kitagawa. Growth rate variability of Pacific saury, Cololabis saira, larvae in the Kuroshio waters. Fisheries Oceanography, 12(4–5):419–424, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Oozeki:2004:EFA

[OWK04] Yoshioki Oozeki, Yoshiro Watanabe, and Daiji Kitagawa. Environmental factors affecting larval growth of Pacific saury, Cololabis saira, in the northwestern Pacific Ocean.

Fisheries Oceanography, 13(S1):44–53, December 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pitois:2014:GLC

[PA14] Sophie G. Pitois and Mike Armstrong. The growth of larval cod and haddock in the Irish Sea: a model with temperature, prey size and turbulence forcing. Fisheries Oceanography, 23(5):417–435, September 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parsons:1995:SFS

[Par95] Tim Parsons. Scaling fisheries: the science of measuring the effects of fishing, 1855–1955. Fisheries Oceanography, 4(2): 191–192, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parsons:1996:TSF

[Par96] Timothy R. Parsons. Taking stock of fisheries management. Fisheries Oceanography, 5(3–4):224–226, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parsons:1999:BR

[Par99] Parsons. Book review. Fisheries Oceanography, 8(4):307– 308, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Piatt:2018:BPF

[PAS^{+}18] John F. Piatt, Mayumi L. Arimitsu, William J. Sydeman, Sarah Ann Thompson, Heather Renner, Stephani Zador, David Douglas, Scott Hatch, Arthur Kettle, and Jeff Williams. Biogeography of pelagic food webs in the North Pacific. Fisheries Oceanography, 27(4):366–380, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Perry:2000:EIA

[PBF00] R. Ian Perry, James A. Boutillier, and Michael G. G. Foreman. Environmental influences on the availability of smooth pink shrimp, Pandalus jordani, to commercial fishing gear off Vancouver Island, Canada. Fisheries Oceanography, 9 (1):50–61, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Polovina:2004:FMH

[PBH^{+}04] Jeffrey J. Polovina, George H. Balazs, Evan A. Howell, Denise M. Parker, Michael P. Seki, and Peter H. Dutton. Forage and migration habitat of loggerhead (Caretta caretta) and olive ridley (Lepidochelys olivacea) sea turtles in the central North Pacific Ocean. Fisheries Oceanography, 13(1): 36–51, January 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Planque:2007:MPS

[PBL07] Benjamin Planque, Edwige Bellier, and Pascal Lazure. Modelling potential spawning habitat of sardine (Sardina pilchardus) and anchovy (Engraulis encrasicolus) in the Bay of Biscay. Fisheries Oceanography, 16(1):16–30, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Politikos:2018:CVS

[PCR^{+}18] Dimitrios V. Politikos, Enrique N. Curchitser, Kenneth A. Rose, David M. Checkley, Jr., and Jerome Fiechter. Climate variability and sardine recruitment in the California Current: a mechanistic analysis of an ecosystem model. Fisheries Oceanography, 27(6):602–622, November 2018. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pepin:2003:SES

[PDD03] P. Pepin, J. F. Dower, and F. J. M. Davidson. A spatially explicit study of prey–predator interactions in larval fish: assessing the influence of food and predator abundance on larval growth and survival. Fisheries Oceanography, 12(1): 19–33, January 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Petersen:2010:NSI

[PDER10] Christine H. Petersen, Patrick T. Drake, Christopher A. Edwards, and Stephen Ralston. A numerical study of inferred rockfish (Sebastes spp.) larval dispersal along the central California coast. Fisheries Oceanography, 19(1):21–41, January 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Patterson:2008:MBL

[PECG08] Toby A. Patterson, Karen Evans, Thor I. Carter, and John S. Gunn. Movement and behaviour of large southern bluefin tuna (Thunnus maccoyii) in the Australian region determined using pop-up satellite archival tags. Fisheries Oceanography, 17(5):352–367, September 2008. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Puckett:2014:LDP

[PEKL14] Brandon J. Puckett, David B. Eggleston, Patrick C. Kerr, and Richard A. Luettich, Jr. Larval dispersal and population connectivity among a network of marine reserves. Fisheries Oceanography, 23(4):342–361, July 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Perry:2023:MTR

[Per23] R. Ian Perry. In memoriam: Timothy R. Parsons (1932– 2022). Fisheries Oceanography, 32(1):10–14, January 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pearcy:1996:SAE

[PFAM96] William G. Pearcy, Joseph P. Fisher, Gen Anma, and Toshimi Meguro. Species associations of epipelagic nekton of the North Pacific Ocean, 1978–1993. Fisheries Oceanography, 5(1):1–20, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parada:2016:SPI

[PFB^{+}16] C. Parada, S. Frusher, R. H. Bustamante, E. Di Lorenzo, P. Bernal, M. Cryer, A. Dunn, R. Garreaud, M. Gutierrez, S. Jennings, A. Montecinos, S. Neira, R. A. Quiñones, K. Takahashi, R. Tascheri, and B. Yannicelli. South Pacific integrated ecosystem studies meeting: toward conservation and sustainable use of marine resources in the South Pacific. Fisheries Oceanography, 25(S1):1–4, April 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Petrie:2009:SSE

[PFSL09] Brian Petrie, Kenneth T. Frank, Nancy L. Shackell, and William C. Leggett. Structure and stability in exploited marine fish communities: quantifying critical transitions. Fisheries Oceanography, 18(2):83–101, March 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Prince:2006:HBH

[PG06] Eric D. Prince and C. Phillip Goodyear. Hypoxia-based habitat compression of tropical pelagic fishes. Fisheries Oceanography, 15(6):451–464, November 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Plourde:2015:EEV

[PGL^{+}15] Stéphane Plourde, François Grégoire, Caroline Lehoux, Peter S. Galbraith, Martin Castonguay, and Marc Ringuette. Effect of environmental variability on body condition and recruitment success of Atlantic mackerel (Scomber scombrus L.) in the Gulf of St. Lawrence. Fisheries Oceanography, 24 (4):347–363, July 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Patterson:2011:DST

[PH11] Toby A. Patterson and Klaas Hartmann. Designing satellite tagging studies: estimating and optimizing data recovery. Fisheries Oceanography, 20(6):449–461, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Painting:1998:EBM

[PHH^{+}98] S. J. Painting, L. Hutchings, J. A. Huggett, J. L. Korrûbel, A. J. Richardson, and H. M. Verheye. Environmental and biological monitoring for forecasting anchovy recruitment in the southern Benguela upwelling region. Fisheries Oceanography, 7(3–4):364–374, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pepin:2013:MDC

[PHH13] Pierre Pepin, Guoqi Han, and Erica J. Head. Modelling the dispersal of Calanus finmarchicus on the Newfoundland Shelf: implications for the analysis of population dynamics from a high frequency monitoring site. Fisheries Oceanography, 22(5):371–387, September 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Perry:1996:SVF

[PHWM96] R. Ian Perry, N. Brent Hargreaves, Brenda J. Waddell, and David L. Mackas. Spatial variations in feeding and condition of juvenile pink and chum salmon off Vancouver Island, British Columbia. Fisheries Oceanography, 5(2):73–88, June 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pitchford:2005:QEI

[PJB05] Jonathan W. Pitchford, Alex James, and John Brindley. Quantifying the effects of individual and environmental variability in fish recruitment. Fisheries Oceanography, 14(2): 156–160, March 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Petrik:2014:IDL

[PJD14] Colleen M. Petrik, Rubao Ji, and Cabell S. Davis. Interannual differences in larval haddock survival: hypothesis testing with a 3D biophysical model of Georges Bank. Fisheries Oceanography, 23(6):521–553, November 2014. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Payne:1999:PCS

[PJO99] Susan A. Payne, B. Alan Johnson, and Robert S. Otto. Proximate composition of some north-eastern Pacific forage fish species. Fisheries Oceanography, 8(3):159–177, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Powell:1998:VTO

[PKHF98] Eric N. Powell, John M. Klinck, Eileen E. Hofmann, and Susan Ford. Varying the timing of oyster transplant: implications for management from simulation studies. Fisheries Oceanography, 6(4):213–237, February 1998. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Polovina:2000:TEM

[PKP^{+}00] Jeffrey J. Polovina, Donald R. Kobayashi, Denise M. Parker, Michael P. Seki, and George H. Balazs. Turtles on the edge: movement of loggerhead turtles (Caretta caretta) along oceanic fronts, spanning longline fishing grounds in the central North Pacific, 1997–1998. Fisheries Oceanography, 9 (1):71–82, March 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Piontkovski:2003:CSD

[PL03] Sergey A. Piontkovski and Michael R. Landry. Copepod species diversity and climate variability in the tropical Atlantic Ocean. Fisheries Oceanography, 12(4–5):352–359, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Prince:2010:OSH

[PLG^{+}10] Eric D. Prince, Jiangang Luo, C. Phillip Goodyear, John P. Hoolihan, Derke Snodgrass, Eric S. Orbesen, Joseph E. Serafy, Mauricio Ortiz, and Michael J. Schirripa. Ocean scale hypoxia-based habitat compression of Atlantic istiophorid billfishes. Fisheries Oceanography, 19(6):448–462, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See corrigendum [Ano11a].

Planque:2011:UWC

[PLP^{+}11] Benjamin Planque, Christophe Loots, Pierre Petitgas, Ulf LindstrøM, and Sandrine Vaz. Understanding what controls the spatial distribution of fish populations using a multimodel approach. Fisheries Oceanography, 20(1):1–17, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parrish:1998:ECL

[PLSO98] Julia K. Parrish, Norm Lemberg, and Laura South-Oryshchyn. Effects of colony location and nekton abundance on the at-sea distribution of four seabird species. Fisheries Oceanography, 7(2):126–135, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Persohn:2009:HPS

[PLT09] Cecile Persohn, Pascal Lorance, and Verena M. Trenkel. Habitat preferences of selected demersal fish species in the Bay of Biscay and Celtic Sea, North-East Atlantic. Fisheries Oceanography, 18(4):268–285, July 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parrish:1995:VCF

[PM95] Richard H. Parrish and Donna L. Mallicoate. Variation in the condition factors of California pelagic fishes and associated environmental factors. Fisheries Oceanography, 4(2): 171–190, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Peterson:2010:ODH

[PMFC10] William T. Peterson, Cheryl A. Morgan, Joseph P. Fisher, and Edmundo Casillas. Ocean distribution and habitat associations of yearling Coho (Oncorhynchus kisutch) and Chinook (O. tshawytscha) salmon in the northern California Current. Fisheries Oceanography, 19(6):508–525, November 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Polovina:1994:PBC

[PMG^{+}94] Jeffrey J. Polovina, Gary T. Mitchum, Nick E. Graham, Mitchell P. Craig, Edward E. Demartini, and Elizabeth N. Flint. Physical and biological consequences of a climate event in the central North Pacific. Fisheries Oceanography, 3(1): 15–21, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Phillips:2023:ITP

[PMG^{+}23] Elizabeth M. Phillips, Michael J. Malick, Stéphane Gauthier, Melissa A. Haltuch, Mary E. Hunsicker, Sandra L. Parker-Stetter, and Rebecca E. Thomas. The influence of temperature on Pacific hake co-occurrence with euphausiids in the California Current Ecosystem. Fisheries Oceanography, 32 (3):267–279, May 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Petitgas:2006:ODB

[PML06] Pierre Petitgas, Stéphanie Magri, and Pascal Lazure. Onedimensional biophysical modelling of fish egg vertical distributions in shelf seas. Fisheries Oceanography, 15(5):413–428, September 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Peterman:1994:SJS

[PMT^{+}94] Randall M. Peterman, Silvio G. Marinone, Keith A. Thomson, Ian D. Jardine, Robert N. Crittenden, Paul H. Leblond, and Carl J. Walters. Simulation of juvenile sockeye salmon (Oncorhynchus nerka) migrations in the Strait of Georgia, British Columbia. Fisheries Oceanography, 3(4):221–235, December 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Perry:2003:SIM

[PO03] R. Ian Perry and Rosemary E. Ommer. Scale issues in marine ecosystems and human interactions. Fisheries Oceanography, 12(4–5):513–522, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Palmas:2017:MGR

[POA^{+}17] Francesco Palmas, Antonio Olita, Piero Addis, Roberto Sorgente, and Andrea Sabatini. Modelling giant red shrimp larval dispersal in the Sardinian seas: density and connectivity scenarios. Fisheries Oceanography, 26(3):364–378, May 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Polovina:1996:DVT

[Pol96] Jeffrey J. Polovina. Decadal variation in the trans-Pacific migration of northern bluefin tuna (Thunnus thynnus) coherent with climate-induced change in prey abundance. Fisheries Oceanography, 5(2):114–119, June 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Porter:2022:VDY

[Por22] Steven M. Porter. Variation in the distribution of yellowfin sole Limanda aspera larvae in warm and cold years in the eastern Bering Sea. Fisheries Oceanography, 31(1):108–122, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Paradis:2001:MCL

[PP01] A. R. Paradis and P. Pepin. Modelling changes in the length– frequency distributions of fish larvae using field estimates of predator abundance and size distributions. Fisheries Oceanography, 10(2):217–234, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Puerta:2016:SVC

[PQH16] Patricia Puerta, Antoni Quetglas, and Manuel Hidalgo. Seasonal variability of cephalopod populations: a spatiotemporal approach in the Western Mediterranean Sea. Fisheries Oceanography, 25(4):373–389, July 2016. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Petitgas:2013:ICC

[PRDC^{+}13] Pierre Petitgas, Adriaan D. Rijnsdorp, Mark Dickey-Collas, Georg H. Engelhard, Myron A. Peck, John K. Pinnegar, Ken Drinkwater, Martin Huret, and Richard D. M. Nash. Impacts of climate change on the complex life cycles of fish. Fisheries Oceanography, 22(2):121–139, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pickett:2006:EUE

[PS06] Mark H. Pickett and Franklin B. Schwing. Evaluating upwelling estimates off the west coasts of North and South America. Fisheries Oceanography, 15(3):256–269, May 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Perretti:2016:INN

[PS16] Charles T. Perretti and Mohammad Sedarat. The influence of the El Niño Southern Oscillation on paralarval market squid (Doryteuthis opalescens). Fisheries Oceanography, 25 (5):491–499, September 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pepin:2005:APC

[PSC05] P. Pepin, P. V. R. Snelgrove, and K. P. Carter. Accuracy and precision of the continuous underway fish egg sampler (CUFES) and bongo nets: a comparison of three species of temperate fish. Fisheries Oceanography, 14(6):432–447, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Platt:1993:PCP

[PSJF93] Trevor Platt, Shubha Sathyendranath, Ian Joint, and Michael J. R. Fasham. Photosynthesis characteristics of the phytoplankton in the Celtic Sea during late spring. Fisheries Oceanography, 2(3–4):191–201, December 1993. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parrish:2000:MLW

[PSM00] Richard H. Parrish, Franklin B. Schwing, and Roy Mendelssohn. Mid-latitude wind stress: the energy source for climatic shifts in the North Pacific Ocean. Fisheries Oceanography, 9(3):224–238, September 2000. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Page:1999:CHS

[PSN^{+}99] F. H. Page, M. Sinclair, C. E. Naimie, J. W. Loder, R. J. Losier, P. L. Berrien, and R. G. Lough. Cod and haddock spawning on Georges Bank in relation to water residence times. Fisheries Oceanography, 8(3):212–226, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parker:2021:BPF

[PSS^{+}21] Steven J. Parker, Svein Sundby, Darren Stevens, Davide Di Blasi, Stefano Schiaparelli, and Laura Ghigliotti. Buoyancy of post-fertilised Dissostichus mawsoni eggs and implications for early life history. Fisheries Oceanography, 30(6):697–706, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pedersen:2003:HMF

[PST03] Ole P. Pedersen, Dag Slagstad, and Kurt S. Tande. Hydrodynamic model forecasts as a guide for process studies on plankton and larval fish. Fisheries Oceanography, 12(4–5): 369–380, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pineda:2024:CCA

[PTS^{+}24] Jesús Pineda, Carolyn Tepolt, Vicke Starczak, Phil Alatalo, and Sara Shapiro. Concentration and condition of American lobster postlarvae in small-scale convergences. Fisheries Oceanography, 33(1):e12657:1–e12657:??, January 2024. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pennino:2019:BRP

[PVBV19] Maria Grazia Pennino, Raul Vilela, Jose M. Bellido, and Francisco Velasco. Balancing resource protection and fishing activity: The case of the European hake in the northern Iberian Peninsula. Fisheries Oceanography, 28(1):54–65, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Perrotta:2001:TCA

[PVHT01] Ricardo G. Perrotta, María D. Viñas, Daniel R. Hernandez, and Leonardo Tringali. Temperature conditions in the Argentine chub mackerel (Scomber japonicus) fishing ground: implications for fishery management. Fisheries Oceanography, 10(3):275–283, September 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parada:2003:MEB

[PVMP03] C. Parada, C. D. Van Der Lingen, C. Mullon, and P. Penven. Modelling the effect of buoyancy on the transport of anchovy (Engraulis capensis) eggs from spawning to nursery grounds in the southern Benguela: an IBM approach. Fisheries Oceanography, 12(3):170–184, May 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parsons:2012:DVA

[PW12] Timothy R. Parsons and Frank A. Whitney. Did volcanic ash from Mt. Kasatoshi in 2008 contribute to a phenomenal increase in Fraser River sockeye salmon (Oncorhynchus nerka) in 2010? Fisheries Oceanography, 21(5):374–377, September 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Parsons:2014:EKV

[PW14] Tim Parsons and Frank Whitney. On the effect of the Kasatoshi volcano on the large return of sockeye salmon (Oncorhynchus nerka) to the Fraser River in 2010. Fisheries Oceanography, 23(1):101–102, January 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Proctor:1998:MTL

[PWE98] Roger Proctor, Peter J. Wright, and Andrew Everitt. Modelling the transport of larval sandeels on the north-west European shelf. Fisheries Oceanography, 7(3–4):347–354, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Pershing:2012:LSC

[PWML12] Andrew J. Pershing, Richard A. Wahle, Patrick C. Meyers, and Peter Lawton. Large-scale coherence in New England lobster (Homarus americanus), settlement and associations with regional atmospheric conditions. Fisheries Oceanography, 21(5):348–362, September 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Quinlan:1999:SGE

[QBMW99] John A. Quinlan, Brian O. Blanton, Thomas J. Miller, and Francisco E. Werner. From spawning grounds to the estuary: using linked individual-based and hydrodynamic models to interpret patterns and processes in the oceanic phase of Atlantic menhaden Brevoortia tyrannus life history. Fisheries Oceanography, 8(S2):224–246, December 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Quinlan:1999:SSL

[QC99] John A. Quinlan and Larry B. Crowder. Searching for sensitivity in the life history of Atlantic menhaden: inferences from a matrix model. Fisheries Oceanography, 8(S2):124– 133, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Quinzan:2016:UIE

[QCM^{+}16] M. Quinzán, J. Castro, M. Marín, G. Costas, S. Monserrat, A. Amores, E. Massutí, and M. Hidalgo. Unveiling the influence of the environment on the migration pattern of the Atlantic pomfret (Brama brama) in North-eastern Atlantic waters. Fisheries Oceanography, 25(6):610–623, November 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Quinn:2022:IUD

[QCR22] Brady K. Quinn, Joël Chassé, and Rémy Rochette. Impact of the use of different temperature-dependent larval development functions on estimates of potential large-scale connectivity of American lobster. Fisheries Oceanography, 31(5): 554–569, September 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Quattrini:2005:DLF

[QLB^{+}05] A. M. Quattrini, D. G. Lindquist, F. M. Bingham, T. E. Lankford, and J. J. Govoni. Distribution of larval fishes among water masses in Onslow Bay, North Carolina: implications for cross-shelf exchange. Fisheries Oceanography, 14(6):413–431, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Quinones:2001:RBF

[QM01] R. A. Quiñones and R. M. Montes. Relationship between freshwater input to the coastal zone and the historical landings of the benthic/demersal fish Eleginops maclovinus in central-south Chile. Fisheries Oceanography, 10(4):311–328, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ruff:2017:SSC

[RAK^{+}17] Casey P. Ruff, Joseph H. Anderson, Iris M. Kemp, Neala W. Kendall, Peter A. Mchugh, Antonio Velez-Espino, Correigh M. Greene, Marc Trudel, Carrie A. Holt, Kristen E. Ryding, and Kit Rawson. Salish Sea Chinook salmon exhibit weaker coherence in early marine survival trends than coastal populations. Fisheries Oceanography, 26(6):625–637, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reiss:2002:RAV

[RAT^{+}02] Christian S. Reiss, Ayal Anis, Christopher T. Taggart, JohnF. Dower, and Barry Ruddick. Relationships among vertically structured in situ measures of turbulence, larval fish abundance and feeding success and copepods on Western Bank, Scotian Shelf. Fisheries Oceanography, 11(3):156– 174, May 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rouyer:2021:EDA

[RBB^{+}21] Tristan Rouyer, Sylvain Bonhommeau, Guilaume Bal, Olivier Derridj, and Jean-Marc Fromentin. The environment drives Atlantic bluefin tuna availability in the Gulf of Lions. Fisheries Oceanography, 30(5):490–498, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rooper:2012:GPP

[RBBG12] Christopher N. Rooper, Jennifer L. Boldt, Sonia Batten, and Christopher Gburski. Growth and production of Pacific Ocean perch (Sebastes alutus) in nursery habitats of the Gulf of Alaska. Fisheries Oceanography, 21(6):415–429, November 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rodriguez-Basalo:2022:FBA

[RBPCR^{+}22] Augusto Rodríguez-Basalo, Antonio Punzón, Elvira Ceballos-Roa, Gabriel Jordà, José Manuel González-Irusta, and Enric Massutí. Fisheries-based approach to disentangle mackerel (Scomber scombrus) migration in the Cantabrian Sea. Fisheries Oceanography, 31(4):443–455, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reiss:2008:RSS

[RCB08] Christian S. Reiss, David M. Checkley, Jr., and Steven J. Bograd. Remotely sensed spawning habitat of Pacific sardine (Sardinops sagax) and northern anchovy (Engraulis mordax) within the California Current. Fisheries Oceanography, 17 (2):126–136, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Runge:1999:CCZ

[RCD^{+}99] Jeffrey A. Runge, Martin Castonguay, Y. De Lafontaine, Marc Ringuette, and J.-L. Beaulieu. Covariation in climate, zooplankton biomass and mackerel recruitment in the southern Gulf of St Lawrence. Fisheries Oceanography, 8(2):139– 149, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rodriguez:2015:VDM

[RCG^{+}15] J. M. Rodriguez, A. Cabrero, J. Gago, C. Guevara-Fletcher, M. Herrero, A. Hernandez de Rojas, A. Garcia, R. Laiz-Carrion, A. R. Vergara, P. Alvarez, C. Piñeiro, and F. Saborido-Rey. Vertical distribution and migration of fish larvae in the NW Iberian upwelling system during the winter mixing period: implications for cross-shelf distribution. Fisheries Oceanography, 24(3):274–290, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Richardson:2009:ISF

[RCPS09] David E. Richardson, Robert K. Cowen, Eric D. Prince, and Su Sponaugle. Importance of the Straits of Florida spawning ground to Atlantic sailfish (Istiophorus platypterus) and blue marlin (Makaira nigricans). Fisheries Oceanography, 18(6): 402–418, November 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Radach:1998:AWV

[RCS98] Günther Radach, Francois Carlotti, and Andreas Spangenberg. Annual weather variability and its influence on the population dynamics of Calanus finmarchicus. Fisheries Oceanography, 7(3–4):272–281, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Runge:1996:CPE

[RD96] Jeffrey A. Runge and Yves De Lafontaine. Characterization of the pelagic ecosystem in surface waters of the northern Gulf of St. Lawrence in early summer: the larval redfish-Calanus-microplankton interaction. Fisheries Oceanography, 5(1):21–37, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rikardsen:2007:MTD

[RDE^{+}07] Audun H. Rikardsen, Ola H. Diserud, J. Malcolm Elliott, J. Brian Dempson, Johannes Sturlaugsson, and Arne J. Jensen. The marine temperature and depth preferences of Arctic charr (Salvelinus alpinus) and sea trout (Salmo trutta), as recorded by data storage tags. Fisheries Oceanography, 16(5):436–447, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reddin:2011:BES

[RDF^{+}11] D. G. Reddin, P. Downton, I. A. Fleming, L. P. Hansen, and A. Mahon. Behavioural ecology at sea of Atlantic salmon (Salmo salar L.) kelts from a Newfoundland (Canada) river. Fisheries Oceanography, 20(3):174–191, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reid:2003:PCZ

[REB^{+}03] Philip C. Reid, Martin Edwards, Gregory Beaugrand, Morten Skogen, and Darren Stevens. Periodic changes in the zooplankton of the North Sea during the twentieth century linked to oceanic inflow. Fisheries Oceanography, 12 (4–5):260–269, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reed:1995:VSE

[Ree95] R. K. Reed. On the variable subsurface environment of fish stocks in the Bering Sea. Fisheries Oceanography, 4(4):317– 323, December 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reum:2013:BAC

[REG^{+}13] Jonathan C. P. Reum, Timothy E. Essington, Correigh M. Greene, Casimir A. Rice, Patrick Polte, and Kurt L. Fresh. Biotic and abiotic controls on body size during critical life history stages of a pelagic fish, Pacific herring (Clupea pallasii). Fisheries Oceanography, 22(4):324–336, July 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reyns:2007:DDP

[REL07] Nathalie B. Reyns, David B. Eggleston, and Richard A. Luettich, Jr. Dispersal dynamics of post-larval blue crabs, Callinectes sapidus, within a wind-driven estuary. Fisheries Oceanography, 16(3):257–272, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rojas:2002:FLD

[REM02] Pablo M. Rojas, Ruben Escribano, and Victor H. Marín. Fish larvae distribution off Mejillones Peninsula (northern Chile) during a coastal upwelling event in Spring 1999: interactions with the cold upwelling plume. Fisheries Oceanography, 11(4):233–244, July 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ravier:2004:LTF

[RF04] Christelle Ravier and Jean-Marc Fromentin. Are the longterm fluctuations in Atlantic bluefin tuna (Thunnus thynnus) population related to environmental changes? Fisheries Oceanography, 13(3):145–160, May 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Royer:2007:ENS

[RF07] F. Royer and J.-M. Fromentin. Environmental noise in spawning areas: the case of Atlantic bluefin tuna (Thunnus thynnus). Fisheries Oceanography, 16(2):202–206, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reddin:2004:THE

[RFD^{+}04] David G. Reddin, Kevin D. Friedland, Peter Downton, J. Brian Dempson, and Conrad C. Mullins. Thermal habitat experienced by Atlantic salmon (Salmo salar L.) kelts in coastal Newfoundland waters. Fisheries Oceanography, 13 (1):24–35, January 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rosengard:2021:CFR

[RFM^{+}21] Sarah Z. Rosengard, Cameron Freshwater, Skip McKinnell, Yi Xu, and Philippe D. Tortell. Covariability of Fraser River sockeye salmon productivity and phytoplankton biomass in the Gulf of Alaska. Fisheries Oceanography, 30(6):666–678, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rosland:1997:DML

[RG97] Rune Rosland and Jarl Giske. A dynamic model for the life history of Maurolicus muelleri, a pelagic planktivorous fish. Fisheries Oceanography, 6(1):19–34, March 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ruiz:2009:BMA

[RGQPN09] Javier Ruiz, Rafael González-Quirós, Laura Prieto, and Gabriel Navarro. A Bayesian model for anchovy (Engraulis encrasicolus): the combined forcing of man and environment. Fisheries Oceanography, 18(1):62–76, January 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rogers:2013:BSM

[RHG^{+}13] Paul J. Rogers, Charlie Huveneers, Simon D. Goldsworthy, James G. Mitchell, and Laurent Seuront. Broad-scale movements and pelagic habitat of the dusky shark Carcharhinus obscurus off Southern Australia determined using popup satellite archival tags. Fisheries Oceanography, 22(2): 102–112, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rogers:2015:LCS

[RHP^{+}15] Paul J. Rogers, Charlie Huveneers, Brad Page, Simon D. Goldsworthy, Micheal Coyne, Andrew D. Lowther, James G. Mitchell, and Laurent Seuront. Living on the continental shelf edge: habitat use of juvenile shortfin makos Isurus oxyrinchus in the Great Australian Bight, Southern Australia. Fisheries Oceanography, 24(3):205–218, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rochette:2012:CHI

[RHRL12] Sebastien Rochette, Martin Huret, Etienne Rivot, and Olivier Le Pape. Coupling hydrodynamic and individualbased models to simulate long-term larval supply to coastal nursery areas. Fisheries Oceanography, 21(4):229–242, July 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rice:1996:PPR

[Ric96] Jake Rice. Population production and regulation in the sea: a fisheries perspective. Fisheries Oceanography, 5(1):70–71, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Richardson:1999:CFE

[RJHC99] K. Richardson, S. H. Jónasdóttir, S. J. Hay, and A. Christoffersen. Calanus finmarchicus egg production and food availability in the Faroe–Shetland Channel and northern North Sea: October–March. Fisheries Oceanography, 8(S1):153– 162, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Romagnoni:2020:ILT

[RKD^{+}20] Giovanni Romagnoni, Kristina Øie Kvile, Knut-Frode Dagestad, Anne Maria Eikeset, Trond Kristiansen, Nils Chr. Stenseth, and Øystein Langangen. Influence of larval transport and temperature on recruitment dynamics of North Sea cod (Gadus morhua) across spatial scales of observation. Fisheries Oceanography, 29(4):324–339, July 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ruzicka:2019:CRP

[RKZHC19] James J. Ruzicka, Stephen Kasperski, Stephani Zador, and Amber Himes-Cornell. Comparing the roles of Pacific halibut and arrowtooth flounder within the Gulf of Alaska ecosystem and fishing economy. Fisheries Oceanography, 28(5):576–596, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Runcie:2019:EAP

[RMH^{+}19] Rosa M. Runcie, Barbara Muhling, Elliott L. Hazen, Steven J. Bograd, Toby Garfield, and Gerard DiNardo. Environmental associations of Pacific bluefin tuna (Thunnus orientalis) catch in the California Current system. Fisheries Oceanography, 28(4):372–388, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Romaine:2002:CEP

[RMM02] Stephen J. Romaine, David L. Mackas, and Michael C. Macaulay. Comparison of euphausiid population size estimates obtained using replicated acoustic surveys of coastal inlets and block average vs. geostatistical spatial interpolation methods. Fisheries Oceanography, 11(2):102–115, March 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Robinson:1994:IOC

[Rob94] Clifford L. K. Robinson. The influence of ocean climate on coastal plankton and fish production. Fisheries Oceanography, 3(3):159–171, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rodionov:2005:STV

[ROB05] Sergei N. Rodionov, James E. Overland, and Nicholas A. Bond. Spatial and temporal variability of the Aleutian climate. Fisheries Oceanography, 14(S1):3–21, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Roger:1994:RAY

[Rog94] Claude Roger. Relationships among yellowfin and skipjack tuna, their prey-fish and plankton in the tropical western Indian Ocean. Fisheries Oceanography, 3(2):133–141, June 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Richards:2016:USD

[ROH16] R. Anne Richards, John E. O’Reilly, and Kimberly J. W. Hyde. Use of satellite data to identify critical periods for early life survival of northern shrimp in the Gulf of Maine. Fisheries Oceanography, 25(3):306–319, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rothschild:1993:GPM

[RP93] B. J. Rothschild and E. V. Patrick. Generation of a phytoplankton maximum in a grazing-extended logistic model. Fisheries Oceanography, 2(3–4):223–230, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reed:2019:RSS

[RPC^{+}19] Daniel Reed, Stéphane Plourde, Adam Cook, Pierre Pepin, Benoit Casault, Caroline Lehoux, and Catherine Johnson. Response of Scotian Shelf silver hake (Merluccius bilinearis) to environmental variability. Fisheries Oceanography, 28(3): 256–272, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reid:1998:OVL

[RPE98] Philip C. Reid, Benjamin Planque, and Martin Edwards. Is observed variability in the long-term results of the continuous plankton recorder survey a response to climate change? Fisheries Oceanography, 7(3–4):282–288, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Castillo:2022:CAE

[RPG^{+}22] Pedro Ramiro Castillo, Cecilia Peña, Daniel Grados, Luis La Cruz, Carlos Valdez, Marissela Pozada-Herrera, and Rodolfo Cornejo. Characteristics of anchoveta (Engraulis ringens) schools in the optimum zone and the physiological stress zone of its distribution between 2011 and 2021. Fisheries Oceanography, 31(5):510–523, September 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Reiss:2000:OLF

[RPT^{+}00] Christian S. Reiss, G. Panteleev, Christopher T. Taggart, J. Sheng, and Brad deYoung. Observations on larval fish transport and retention on the Scotian Shelf in relation to geostrophic circulation. Fisheries Oceanography, 9(3):195– 213, September 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rice:1999:STD

[RQN^{+}99] James A. Rice, John A. Quinlan, Stephen W. Nixon, William F. Hettler, Jr., Stanley M. Warlen, and Petra M. Stegmann. Spawning and transport dynamics of Atlantic menhaden: inferences from characteristics of immigrating larvae and predictions of a hydrodynamic model. Fisheries Oceanography, 8(S2):93–110, December 1999. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ruiz:2018:ADC

[RR18] Javier Ruiz and Margarita M. Rincón. Anchovy dynamics and the colour of climate noise: Assessing climatic scenarios and exploitation strategies in the Gulf of Cádiz. Fisheries Oceanography, 27(6):560–570, November 2018. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rogers:2021:UBM

[RRF^{+}21] Troy A. Rogers, Ana Redondo Rodriguez, Anthony J. Fowler, Mark J. Doubell, Michael J. Drew, Michael A. Steer, Damian Matthews, Charles James, and Bronwyn M. Gillanders. Using a biophysical model to investigate connectivity between spawning grounds and nursery areas of King George whiting (Sillaginodes punctatus: Perciformes) in South Australia’s gulfs. Fisheries Oceanography, 30(1):51–68, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rose:1992:RAL

[RS92] Kenneth A. Rose and J. Kevin Summers. Relationships among long-term fisheries abundances, hydrographic variables, and gross pollution indicators in northeastern U.S. estuaries. Fisheries Oceanography, 1(4):281–293, December 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Raya:2015:DDE

[RS15] Vanesa Raya and Ana Sabatés. Diversity and distribution of early life stages of carangid fishes in the northwestern Mediterranean: responses to environmental drivers. Fisheries Oceanography, 24(2):118–134, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rogachev:1996:CPO

[RSC96] K. A. Rogachev, A. S. Salomatin, and E. C. Carmack. Concentration of pelagic organisms at mesoscale fronts in the western subarctic Pacific: small fish on long waves. Fisheries Oceanography, 5(3–4):153–162, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ralston:2013:IVP

[RSF13] S. Ralston, K. M. Sakuma, and J. C. Field. Interannual variation in pelagic juvenile rockfish (Sebastes spp.) abundance — going with the flow. Fisheries Oceanography, 22(4):288– 308, July 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rooker:2003:IAB

[RSZ^{+}03] Jay R. Rooker, David H. Secor, Vincent S. Zdanowicz, Gregorio De Metrio, and Lidia Orsi Relini. Identification of Atlantic bluefin tuna (Thunnus thynnus) stocks from putative nurseries using otolith chemistry. Fisheries Oceanography, 12 (2):75–84, March 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rosenkranz:2001:EWT

[RTK01] Gregg E. Rosenkranz, Albert V. Tyler, and Gordon H. Kruse. Effects of water temperature and wind on year-class success of Tanner crabs in Bristol Bay, Alaska. Fisheries Oceanography, 10(1):1–12, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rogers:2021:PBI

[RWDA^{+}21] Lauren A. Rogers, Matthew T. Wilson, Janet T. Duffy-Anderson, David G. Kimmel, and Jesse F. Lamb. Pollock and “The Blob”: Impacts of a marine heatwave on walleye pollock early life stages. Fisheries Oceanography, 30(2): 142–158, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rooker:2016:NOP

[RWI^{+}16] Jay R. Rooker, R. J. David Wells, David G. Itano, Simon R. Thorrold, and Jessica M. Lee. Natal origin and population connectivity of bigeye and yellowfin tuna in the Pacific Ocean. Fisheries Oceanography, 25(3):277–291, May 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Rupp:2012:MEB

[RWLP12] David E. Rupp, Thomas C. Wainwright, Peter W. Lawson, and William T. Peterson. Marine environment-based forecasting of Coho salmon (Oncorhynchus kisutch) adult recruitment. Fisheries Oceanography, 21(1):1–19, January 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ruzicka:2011:MBM

[RWP11] James J. Ruzicka, Thomas C. Wainwright, and William T. Peterson. A model-based meso-zooplankton production index and its relation to the ocean survival of juvenile Coho (Oncorhynchus kisutch). Fisheries Oceanography, 20(6):544– 559, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ruggerone:2003:CBA

[RZM^{+}03] G. T. Ruggerone, M. Zimmermann, K. W. Myers, J. L. Nielsen, and D. E. Rogers. Competition between Asian pink salmon (Oncorhynchus gorbuscha) and Alaskan sockeye salmon (O. nerka) in the North Pacific Ocean. Fisheries Oceanography, 12(3):209–219, May 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sagarminaga:2010:STD

[SA10] Y. Sagarminaga and H. Arrizabalaga. Spatio-temporal distribution of albacore (Thunnus alalunga) catches in the northeastern Atlantic: relationship with the thermal environment. Fisheries Oceanography, 19(2):121–134, March 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Spear:2023:CVD

[SADA^{+}23] Adam Spear, Alexander G. Andrews III, Janet Duffy-Anderson, Tayler Jarvis, David Kimmel, and Denise McKelvey. Changes in the vertical distribution of age-0 walleye pollock (Gadus chalcogrammus) during warm and cold years in the southeastern Bering Sea. Fisheries Oceanography, 32 (2):177–195, March 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schwemmer:2009:IWF

[SAG^{+}09] Philipp Schwemmer, Sven Adler, Nils Guse, Nele Markones, and Stefan Garthe. Influence of water flow velocity, water depth and colony distance on distribution and foraging patterns of terns in the Wadden Sea. Fisheries Oceanography, 18 (3):161–172, May 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sepulveda:2018:MBS

[SAH^{+}18] Chugey A. Sepulveda, Scott A. Aalbers, Craig Heberer, Suzy Kohin, and Heidi Dewar. Movements and behaviors of swordfish Xiphias gladius in the United States Pacific Leatherback Conservation Area. Fisheries Oceanography, 27(4):381–394, July 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Saito:2022:OTS

[Sai22] Toshihiko Saito. Optimum timing and size at release of chum salmon: Improving survival by modifying hatchery practices. Fisheries Oceanography, 31(4):416–428, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shigenobu:2017:RCA

[SAO^{+}17] Yuya Shigenobu, Daisuke Ambe, Tsuneo Ono, Ken Fujimoto, Takami Morita, Tadafumi Ichikawa, and Tomowo Watanabe. Radiocesium contamination of aquatic organisms in the estuary of the Abukuma River flowing through Fukushima. Fisheries Oceanography, 26(2):208–220, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shimose:2018:SLI

[SAT^{+}18] Tamaki Shimose, Yoshimasa Aonuma, Toshiyuki Tanabe, Nobuaki Suzuki, and Minoru Kanaiwa. Solar and lunar influences on the spawning activity of Pacific bluefin tuna (Thunnus orientalis) in the south-western North Pacific spawning ground. Fisheries Oceanography, 27(1):76–84, January 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shima:1994:CAI

[SB94] Michiyo Shima and Kevin M. Bailey. Comparative analysis of ichthyoplankton sampling gear for early life stages of walleye pollock (Theragra chalcogramma). Fisheries Oceanography, 3(1):50–59, March 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stein:2004:GCG

[SB04] M. Stein and V. A. Borovkov. Greenland cod (Gadus morhua): modeling recruitment variation during the second half of the 20th century. Fisheries Oceanography, 13(2): 111–120, March 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Solow:2005:TRS

[SB05] Andrew R. Solow and Andrew R. Beet. A test for a regime shift. Fisheries Oceanography, 14(3):236–240, May 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Swain:2006:CHA

[SB06] D. P. Swain and H. P. Benoît. Change in habitat associations and geographic distribution of thorny skate (Amblyraja radiata) in the southern Gulf of St Lawrence: density-dependent habitat selection or response to environmental change? Fisheries Oceanography, 15(2):166–182, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Solow:2007:ENN

[SB07] Andrew R. Solow and Andrew R. Beet. Is the effect of the NAO on North-east Arctic cod, Gadus morhua, recruitment stock-dependent? Fisheries Oceanography, 16(5):479–481, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stratoudakis:2003:CDS

[SBBB03] Y. Stratoudakis, M. Bernal, D. L. Borchers, and M. F. Borges. Changes in the distribution of sardine eggs and larvae off Portugal, 1985–2000. Fisheries Oceanography, 12(1): 49–60, January 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Spencer:2019:BBL

[SBD^{+}19] David M. Spencer, Ian W. Brown, Mark J. Doubell, Christopher J. Brown, Ana Redondo Rodriguez, Shing Yip Lee, Hong Zhang, and Charles J. Lemckert. Bottom boundary layer cooling and wind-driven upwelling enhance the catchability of spanner crab (Ranina ranina) in South-East Queensland, Australia. Fisheries Oceanography, 28(3):317– 326, May 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stabeno:2001:TVP

[SBK^{+}01] P. J. Stabeno, N. A. Bond, N. B. Kachel, S. A. Salo, and J. D. Schumacher. On the temporal variability of the physical environment over the south-eastern Bering Sea. Fisheries Oceanography, 10(1):81–98, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schaefer:2020:MBR

[SBT20] Anne Louise Schaefer, Mary Anne Bishop, and Richard Thorne. Marine bird response to forage fish during winter in subarctic bays. Fisheries Oceanography, 29(4):297–308, July 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Salvanes:2015:SDB

[SBY^{+}15] Anne Gro V. Salvanes, Chris Bartholomae, Dawit Yemane, Mark J. Gibbons, Paul Kainge, Jens-Otto Krakstad, Mathieu Rouault, Arved Staby, and Svein Sundby. Spatial dynamics of the bearded goby and its key fish predators off Namibia vary with climate and oxygen availability. Fisheries Oceanography, 24(S1):88–101, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Spencer:1997:PPV

[SC97] Paul D. Spencer and Jeremy S. Collie. Patterns of population variability in marine fish stocks. Fisheries Oceanography, 6 (3):188–204, October 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sinclair:2005:IES

[SC05] A. F. Sinclair and W. R. Crawford. Incorporating an environmental stock–recruitment relationship in the assessment of Pacific cod (Gadus macrocephalus). Fisheries Oceanography, 14(2):138–150, March 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schirripa:2006:ICS

[SC06] Michael J. Schirripa and J. J. Colbert. Interannual changes in sablefish (Anoplopoma fimbria) recruitment in relation to oceanographic conditions within the California Current System. Fisheries Oceanography, 15(1):25–36, January 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Santana-Cisneros:2021:DMO

[SCAG^{+}21] Mariana L. Santana-Cisneros, Pedro-Luis Ardisson, Ángel F. González, Ismael Mariño-Tapia, Miguel Cahuich-López, Luis Enrique Ángeles-González, Uriel Ordoñez-López, and Iván Velázquez-Abunader. Dispersal modeling of octopoda paralarvae in the Gulf of Mexico. Fisheries Oceanography, 30(6):726–739, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sohn:2010:DDP

[SCDA10] D. Sohn, L. Ciannelli, and J. T. Duffy-Anderson. Distribution and drift pathways of Greenland halibut (Reinhardtius hippoglossoides) during early life stages in the eastern Bering Sea and Aleutian Islands. Fisheries Oceanography, 19(5): 339–353, September 2010. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sun:2020:LDE

[SCF^{+}20] Peng Sun, Qi Chen, Caihong Fu, Yi Xu, Runlong Sun, Jianchao Li, Haiqing Yu, Chi Zhang, Yang Liu, Zhenjiang Ye, Yongjun Tian, and Mikko Heino. Latitudinal differences in early growth of largehead hairtail (Trichiurus japonicus) in relation to environmental variables. Fisheries Oceanography, 29(6):470–483, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schwing:2023:MTI

[Sch23] Franklin B. Schwing. Modern technologies and integrated observing systems are “instrumental” to fisheries oceanography: A brief history of ocean data collection. Fisheries Oceanography, 32(1):28–69, January 2023. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Siegelman-Charbit:2018:PFF

[SCKJ^{+}18] Lia Siegelman-Charbit, J. Anthony Koslow, Michael G. Jacox, Elliott L. Hazen, Steven J. Bograd, and Eric F. Miller. Physical forcing on fish abundance in the southern California Current System. Fisheries Oceanography, 27(5):475–488, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Scott:1995:OFD

[Sco95] Beth Scott. Oceanographic features that define the habitat of Pacific Ocean perch, Sebastes alutus. Fisheries Oceanography, 4(2):147–157, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sullivan:2005:EAO

[SCS05] Mark C. Sullivan, Robert K. Cowen, and Brian P. Steves. Evidence for atmosphere–ocean forcing of yellowtail flounder (Limanda ferruginea) recruitment in the Middle Atlantic Bight. Fisheries Oceanography, 14(5):386–399, September 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Staples:2019:SVP

[SCTB19] Kevin W. Staples, Yong Chen, David W. Townsend, and Damian C. Brady. Spatiotemporal variability in the phenology of the initial intra-annual molt of American lobster (Homarus americanus Milne Edwards, 1837) and its relationship with bottom temperatures in a changing Gulf of Maine. Fisheries Oceanography, 28(4):468–485, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sippel:2007:SMT

[SDHB07] Tim J. Sippel, Peter S. Davie, John C. Holdsworth, and Barbara A. Block. Striped marlin (Tetrapturus audax) movements and habitat utilization during a summer and autumn in the Southwest Pacific Ocean. Fisheries Oceanography, 16(5):459–472, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shuntov:1996:NDA

[SDRL96] V. P. Shuntov, E. P. Dulepova, V. I. Radchenko, and V. V. Lapko. New data about communities of plankton and nekton of the far-eastern seas in connection with climateoceanological reorganization. Fisheries Oceanography, 5(1): 38–44, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sigler:2019:DVM

[SE19] Michael F. Sigler and Katy B. Echave. Diel vertical migration of sablefish (Anoplopoma fimbria). Fisheries Oceanography, 28(5):517–531, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stachura:2014:LNP

[SEM^{+}14] Megan M. Stachura, Timothy E. Essington, Nathan J. Mantua, Anne B. Hollowed, Melissa A. Haltuch, Paul D. Spencer, Trevor A. Branch, and Miriam J. Doyle. Linking Northeast Pacific recruitment synchrony to environmental variability. Fisheries Oceanography, 23(5):389–408, September 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schilling:2020:MSE

[SES^{+}20] Hayden T. Schilling, Jason D. Everett, James A. Smith, John Stewart, Julian M. Hughes, Moninya Roughan, Colette Kerry, and Iain M. Suthers. Multiple spawning events promote increased larval dispersal of a predatory fish in a western boundary current. Fisheries Oceanography, 29(4): 309–323, July 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schaefer:2022:HMU

[SF22] Kurt M. Schaefer and Daniel W. Fuller. Horizontal movements, utilization distributions, and mixing rates of yellowfin tuna (Thunnus albacares) tagged and released with archival tags in six discrete areas of the eastern and central Pacific Ocean. Fisheries Oceanography, 31(1):84–107, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schaefer:2014:MBH

[SFA14] Kurt M. Schaefer, Daniel W. Fuller, and Gabriel Aldana. Movements, behavior, and habitat utilization of yellowfin tuna (Thunnus albacares) in waters surrounding the Revillagigedo Islands Archipelago Biosphere Reserve, Mexico. Fisheries Oceanography, 23(1):65–82, January 2014. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sanz-Fernandez:2021:HCB

[SFGE21] Víctor Sanz-Fernández and Juan Carlos Gutiérrez-Estrada. Historical changes of blackspot seabream (Pagellus bogaraveo) landing patterns in the Strait of Gibraltar from 1983 to 2016: Environmental and legislation effects. Fisheries Oceanography, 30(2):111–126, March 2021. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Suzuki:2020:WMP

[SFK^{+}20] Keita W. Suzuki, Taiki Fuji, Akihide Kasai, Sachihiko Itoh, Shingo Kimura, and Yoh Yamashita. Winter monsoon promotes the transport of Japanese temperate bass Lateolabrax japonicus eggs and larvae toward the innermost part of Tango Bay, the Sea of Japan. Fisheries Oceanography, 29 (1):66–83, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Smith:2016:BBS

[SFL16] Brian E. Smith, Michael D. Ford, and Jason S. Link. Bloom or bust: synchrony in jellyfish abundance, fish consumption, benthic scavenger abundance, and environmental drivers across a continental shelf. Fisheries Oceanography, 25(5): 500–514, September 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sanchez-Garrido:2021:DAS

[SGFR^{+}21] José C. Sánchez-Garrido, Jerome Fiechter, Kenneth A. Rose, Francisco E. Werner, and Enrique N. Curchitser. Dynamics of anchovy and sardine populations in the Canary Current off NW Africa: Responses to environmental and climate forcing in a climate-to-fish ecosystem model. Fisheries Oceanography, 30(3):232–252, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Speirs:2005:MBS

[SGHW05] Douglas C. Speirs, William S. C. Gurney, Michael R. Heath, and Simon N. Wood. Modelling the basin-scale demography of Calanus finmarchicus in the north-east Atlantic. Fisheries Oceanography, 14(5):333–358, September 2005. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schick:2004:BTT

[SGL04] R. S. Schick, J. Goldstein, and M. E. Lutcavage. Bluefin tuna (Thunnus thynnus) distribution in relation to sea surface temperature fronts in the Gulf of Maine (1994–96). Fisheries Oceanography, 13(4):225–238, July 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schonfeld:2022:SDE

[SGL22] Adena J. Schonfeld, James Gartland, and Robert J. Latour. Spatial differences in estuarine utilization by seasonally resident species in Mid-Atlantic Bight, USA. Fisheries Oceanography, 31(6):615–628, November 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Siegel:2005:LTV

[SGN^{+}05] V. Siegel, J. Gröger, T. Neudecker, U. Damm, and S. Jansen. Long-term variation in the abundance of the brown shrimp Crangon crangon (L.) population of the German Bight and possible causes for its interannual variability. Fisheries Oceanography, 14(1):1–16, January 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Somarakis:2006:SHD

[SGS^{+}06] S. Somarakis, K. Ganias, A. Siapatis, C. Koutsikopoulos, A. Machias, and C. Papaconstantinou. Spawning habitat and daily egg production of sardine (Sardina pilchardus) in the eastern Mediterranean. Fisheries Oceanography, 15(4): 281–292, July 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sadorus:2021:MLS

[SGW^{+}21] Lauri L. Sadorus, Esther D. Goldstein, Raymond A. Webster, William T. Stockhausen, Josep V. Planas, and Janet T. Duffy-Anderson. Multiple life-stage connectivity of Pacific halibut (Hippoglossus stenolepis) across the Bering Sea and Gulf of Alaska. Fisheries Oceanography, 30(2):174–193, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sharp:1995:ATN

[Sha95] Gary D. Sharp. It’s about time: new beginnings and old good ideas in fisheries science. Fisheries Oceanography, 4(4): 324–341, December 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shanks:2013:AFD

[Sha13] A. L. Shanks. Atmospheric forcing drives recruitment variation in the Dungeness crab (Cancer magister), revisited.

Fisheries Oceanography, 22(4):263–272, July 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stepputtis:2011:EMS

[SHB^{+}11] Daniel Stepputtis, Hans-Harald Hinrichsen, Uwe Böttcher, Eberhard Götze, and Volker Mohrholz. An example of mesoscale hydrographic features in the central Baltic Sea and their influence on the distribution and vertical migration of sprat, Sprattus sprattus balticus (Schn.). Fisheries Oceanography, 20(1):82–88, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schaber:2012:SCV

[SHG12] Matthias Schaber, Hans-Harald Hinrichsen, and Joachim Gröger. Seasonal changes in vertical distribution patterns of cod (Gadus morhua) in the Bornholm Basin, central Baltic Sea. Fisheries Oceanography, 21(2–3):33–43, March–May 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schilling:2022:CWL

[SHG^{+}22] Hayden T. Schilling, Charles Hinchliffe, Jonathan P. Gillson, Anthony G. Miskiewicz, and Iain M. Suthers. Coastal winds and larval fish abundance indicate a recruitment mechanism for southeast Australian estuarine fisheries. Fisheries Oceanography, 31(1):40–55, January 2022. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shiganova:1998:IBS

[Shi98] T. A. Shiganova. Invasion of the Black Sea by the ctenophore Mnemiopsis leidyi and recent changes in pelagic community structure. Fisheries Oceanography, 7(3–4):305–310, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sogawa:2019:ECS

[SHK^{+}19] Sayaka Sogawa, Kiyotaka Hidaka, Yasuhiro Kamimura, Masanori Takahashi, Hiroaki Saito, Yuji Okazaki, Yugo Shimizu, and Takashi Setou. Environmental characteristics of spawning and nursery grounds of Japanese sardine and mackerels in the Kuroshio and Kuroshio Extension area. Fisheries Oceanography, 28(4):454–467, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stabeno:2005:IPC

[SHM05] Phyllis J. Stabeno, George L. Hunt, Jr., and S. Allen Macklin. Introduction to processes controlling variability in productivity and ecosystem structure of the Aleutian Archipelago. Fisheries Oceanography, 14(S1):1–2, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Salois:2023:SBE

[SHS^{+}23] Sarah L. Salois, Kimberly J. W. Hyde, Adrienne Silver, Brooke A. Lowman, Avijit Gangopadhyay, Glen Gawarkiewicz, Anna J. M. Mercer, John P. Manderson, Sarah K. Gaichas, Daniel J. Hocking, Benjamin Galuardi, Andrew W. Jones, Jeff Kaelin, Greg DiDomenico, Katie Almeida, Bill Bright, and Meghan Lapp. Shelf break exchange processes influence the availability of the northern shortfin squid, Illex illecebrosus, in the Northwest Atlantic. Fisheries Oceanography, 32(5):461–478, September 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Simpson:1992:RSG

[Sim92a] James J. Simpson. Remote sensing and geographical information systems: Their past, present and future use in global marine fisheries. Fisheries Oceanography, 1(3):238– 280, September 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Simpson:1992:RSC

[Sim92b] James J. Simpson. Response of the Southern California current system to the mid-latitude North Pacific coastal warming events of 1982–1983 and 1940–1941. Fisheries Oceanography, 1(1):57–79, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Simpson:1996:CCS

[Sim96] James J. Simpson. Comments on ‘Computer simulations of homeward-migrating Fraser River sockeye salmon: is compass orientation a sufficient direction-finding mechanism in the north-east Pacific ocean?’ by C. G. Dat et al. (1995). Fisheries Oceanography, 5(2):136, June 1996. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See [DLTI95].

Suca:2022:LTP

[SJB^{+}22] Justin J. Suca, Rubao Ji, Hannes Baumann, Kent Pham, Tammy L. Silva, David N. Wiley, Zhixuan Feng, and Joel K. Llopiz. Larval transport pathways from three prominent sand lance habitats in the Gulf of Maine. Fisheries Oceanography, 31(3):333–352, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Suda:2003:SMP

[SK03] Maki Suda and Tatsu Kishida. A spatial model of population dynamics of the early life stages of Japanese sardine, Sardinops melanostictus, off the Pacific coast of Japan. Fisheries Oceanography, 12(2):85–99, March 2003. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sugisaki:2004:DRS

[SK04] Hiroya Sugisaki and Yutaka Kurita. Daily rhythm and seasonal variation of feeding habit of Pacific saury (Cololabis saira) in relation to their migration and oceanographic conditions off Japan. Fisheries Oceanography, 13(S1):63–73, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2004:DPL

[SKHI04] Chiyuki Sassa, Kouichi Kawaguchi, Yuichi Hirota, and Minoru Ishida. Distribution patterns of larval myctophid fish assemblages in the subtropical–tropical waters of the western North Pacific. Fisheries Oceanography, 13(4):267–282, July 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Saito:2011:EJS

[SKHN11] Toshihiko Saito, Toshiki Kaga, Eiichi Hasegawa, and Kazuya Nagasawa. Effects of juvenile size at release and early marine growth on adult return rates for Hokkaido chum salmon (Oncorhynchus keta) in relation to sea surface temperature. Fisheries Oceanography, 20(4):278–293, July 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stabeno:2005:OMA

[SKKS05] P. J. Stabeno, D. G. Kachel, N. B. Kachel, and M. E. Sullivan. Observations from moorings in the Aleutian passes: temperature, salinity and transport. Fisheries Oceanography, 14(S1):39–54, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2002:AVM

[SKKW02] Chiyuki Sassa, Kouichi Kawaguchi, Takahiro Kinoshita, and Chikako Watanabe. Assemblages of vertical migratory mesopelagic fish in the transitional region of the western North Pacific. Fisheries Oceanography, 11(4):193–204, July 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2004:LWL

[SKM04] Chiyuki Sassa, Kouichi Kawaguchi, and Ken Mori. Late winter larval mesopelagic fish assemblage in the Kuroshio waters of the western North Pacific. Fisheries Oceanography, 13(2):121–133, March 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2006:DJM

[SKM06] Chiyuki Sassa, Yoshinobu Konishi, and Ken Mori. Distribution of jack mackerel (Trachurus japonicus) larvae and juveniles in the East China Sea, with special reference to the larval transport by the Kuroshio Current. Fisheries Oceanography, 15(6):508–518, November 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sepulveda:2010:FSM

[SKNLD10] Chugey A. Sepulveda, Ashley Knight, Nicole Nasby-Lucas, and Michael L. Domeier. Fine-scale movements of the swordfish Xiphias gladius in the Southern California Bight. Fisheries Oceanography, 19(4):279–289, July 2010. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Suzuki:2014:MFD

[SKNT14] Keita W. Suzuki, Youta Kanematsu, Kouji Nakayama, and Masaru Tanaka. Microdistribution and feeding dynamics of Coilia nasus (Engraulidae) larvae and juveniles in relation to the estuarine turbidity maximum of the macrotidal Chikugo River estuary, Ariake Sea, Japan. Fisheries Oceanography, 23(2):157–171, March 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Skogen:2005:CLG

[Sko05] Morten D. Skogen. Clupeoid larval growth and plankton production in the Benguela upwelling system. Fisheries Oceanography, 14(1):64–70, January 2005. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2021:IVD

[SKT21] Chiyuki Sassa, Satoshi Kitajima, and Motomitsu Takahashi. Interannual variations in diet of Japanese jack mackerel (Trachurus japonicus) juveniles in the southwestern Sea of Japan in relation to recent growth rate. Fisheries Oceanography, 30 (6):772–786, November 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Lloret:1995:SWM

[SL95] M. Stein and J. Lloret. Stability of water masses-impact on cod recruitment off West Greenland. Fisheries Oceanography, 4(3):230–237, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schick:2009:IPD

[SL09] R. S. Schick and M. E. Lutcavage. Inclusion of prey data improves prediction of bluefin tuna (Thunnus thynnus) distribution. Fisheries Oceanography, 18(1):77–81, January 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Silva:2019:PCF

[SLL19] Claudio Silva, Francisco Leiva, and José Lastra. Predicting the current and future suitable habitat distributions of the anchovy (Engraulis ringens) using the Maxent model in the coastal areas off central-northern Chile. Fisheries Oceanography, 28(2):171–182, March 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sparrevohn:2013:CIR

[SLM13] Claus Reedtz Sparrevohn, Martin Lindegren, and Brian R. Mackenzie. Climate-induced response of commercially important flatfish species during the 20th century. Fisheries Oceanography, 22(5):400–408, September 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Song:2023:CML

[SLZ^{+}23] Liming Song, Tianlai Li, Tianjiao Zhang, Hengshou Sui, Bin Li, and Min Zhang. Comparison of machine learning models within different spatial resolutions for predicting the bigeye tuna fishing grounds in tropical waters of the Atlantic Ocean. Fisheries Oceanography, 32(6):509–526, November 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Snauffer:2014:MDH

[SMA14] Evgeniya L. Snauffer, Diane Masson, and Susan E. Allen. Modelling the dispersal of herring and hake larvae in the Strait of Georgia for the period 2007–2009. Fisheries Oceanography, 23(4):375–388, July 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schabetsberger:2003:PSD

[SMB^{+}03a] R. Schabetsberger, C. A. Morgan, R. D. Brodeur, C. L. Potts, W. T. Peterson, and R. L. Emmett. Prey selectivity and diel feeding chronology of juvenile Chinook (Oncorhynchus tshawytscha) and Coho (O. kisutch) salmon in the Columbia River plume. Fisheries Oceanography, 12(6): 523–540, November 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sibert:2003:HMB

[SMB03b] John R. Sibert, Michael K. Musyl, and Richard W. Brill. Horizontal movements of bigeye tuna (Thunnus obesus) near Hawaii determined by Kalman filter analysis of archival tagging data. Fisheries Oceanography, 12(3):141–151, May 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stoens:1998:NPE

[SMDM98] A. Stoens, C. Menkes, Y. Dandonneau, and L. Memery. New production in the equatorial Pacific: a coupled dynamical– biogeochemical model. Fisheries Oceanography, 7(3–4):311– 316, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sadorus:2014:DPP

[SME^{+}14] Lauri L. Sadorus, Nathan J. Mantua, Timothy Essington, Barbara Hickey, and Steven Hare. Distribution patterns of Pacific halibut (Hippoglossus stenolepis) in relation to environmental variables along the continental shelf waters of the US West Coast and southern British Columbia. Fisheries Oceanography, 23(3):225–241, May 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Springer:1996:BSG

[SMF96] Alan M. Springer, C. Peter Mcroy, and Mikhail V. Flint. The Bering Sea Green Belt: shelf-edge processes and ecosystem production. Fisheries Oceanography, 5(3–4):205–223, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sinclair:2005:DPS

[SMF^{+}05] Elizabeth H. Sinclair, Sue E. Moore, Nancy A. Friday, Tonya K. Zeppelin, and Janice M. Waite. Do patterns of Steller sea lion (Eumetopias jubatus) diet, population trend and cetacean occurrence reflect oceanographic domains from the Alaska Peninsula to the central Aleutian Islands? Fisheries Oceanography, 14(S1):223–242, November 2005. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

StJohn:1992:FRP

[SMH^{+}92] Michael A. St. John, J. Stevenson Macdonald, Paul J. Harrison, Richard J. Beamish, and Edward Choromanski. The Fraser River plume: some preliminary observations on the distribution of juvenile salmon, herring, and their prey. Fisheries Oceanography, 1(2):153–162, June 1992. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2002:HVD

[SMK02] Chiyuki Sassa, H. Geoffrey Moser, and Kouichi Kawaguchi. Horizontal and vertical distribution patterns of larval myctophid fishes in the Kuroshio Current region. Fisheries Oceanography, 11(1):1–10, January 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sasaki:2013:HDB

[SMK^{+}13] Hiroko Sasaki, Hiroto Murase, Hiroshi Kiwada, Koji Matsuoka, Yoko Mitani, and Sei ichi Saitoh. Habitat differentiation between sei (Balaenoptera borealis) and Bryde’s whales (B. brydei) in the western North Pacific. Fisheries Oceanography, 22(6):496–508, November 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Staudinger:2019:ATS

[SMS^{+}19] Michelle D. Staudinger, Katherine E. Mills, Karen Stamieszkin, Nicholas R. Record, Christine A. Hudak, Andrew Allyn, Antony Diamond, Kevin D. Friedland, Walt Golet, Meghan Elisabeth Henderson, Christina M. Hernandez, Thomas G. Huntington, Rubao Ji, Catherine L. Johnson, David Samuel Johnson, Adrian Jordaan, John Kocik, Yun Li, Matthew Liebman, Owen C. Nichols, Daniel Pendleton, R. Anne Richards, Thomas Robben, Andrew C. Thomas, Harvey J. Walsh, and Keenan Yakola. It’s about time: a synthesis of changing phenology in the Gulf of Maine ecosystem. Fisheries Oceanography, 28(5):532–566, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Smith:2021:PIS

[SMS^{+}21] James A. Smith, Barbara Muhling, Jonathan Sweeney, Desiree Tommasi, Mercedes Pozo Buil, Jerome Fiechter, and Michael G. Jacox. The potential impact of a shifting Pacific sardine distribution on U.S. West Coast landings. Fisheries Oceanography, 30(4):437–454, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Scales:2023:FRM

[SMS^{+}23] Kylie L. Scales, Thomas S. Moore II, Bernadette Sloyan, Claire M. Spillman, J. Paige Eveson, Toby A. Patterson, Ashley J. Williams, Alistair J. Hobday, and Jason R. Hartog. Forecast-ready models to support fisheries’ adaptation to global variability and change. Fisheries Oceanography, 32 (4):405–417, July 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stratton:2019:MDC

[SNL19] Mark A. Stratton, Geneviève M. Nesslage, and Robert J. Latour. Multi-decadal climate and fishing predictors of abundance for U.S. South Atlantic coastal fishes and invertebrates. Fisheries Oceanography, 28(5):487–504, September 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stenevik:2012:ESD

[SNV^{+}12] Erling K. Stenevik, Richard D. M. Nash, Frode VikebØ, Petter Fossum, and Kjell Bakkeplass. The effects of survey design and circulation pattern on the perceived abundance of herring larvae: a case study for Norwegian spring spawning herring (Clupea harengus). Fisheries Oceanography, 21(5): 363–373, September 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stratoudakis:2018:GEA

[SOTM^{+}18] Yorgos Stratoudakis, Paulo B. Oliveira, Ana Teles-Machado, José Manuel Oliveira, Maria João Correia, and Carlos Antunes. Glass eel (Anguilla anguilla) recruitment to the River Lis: Ingress dynamics in relation to oceanographic processes in the western Iberian margin and shelf. Fisheries Oceanography, 27(6):536–547, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shepherd:1993:AME

[SP93] J. G. Shepherd and J. G. Pope. Alternative methods for the estimation of immigration to the Icelandic cod stock. Fisheries Oceanography, 2(3–4):254–259, December 1993. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Szuwalski:2013:RSR

[SP13] Cody Szuwalski and André E. Punt. Regime shifts and recruitment dynamics of snow crab, Chionoecetes opilio, in the eastern Bering Sea. Fisheries Oceanography, 22(5):345–354, September 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Surma:2015:PEW

[SP15] Szymon Surma and Tony J. Pitcher. Predicting the effects of whale population recovery on Northeast Pacific food webs and fisheries: an ecosystem modelling approach. Fisheries Oceanography, 24(3):291–305, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Spencer:2008:DID

[Spe08] Paul D. Spencer. Density-independent and densitydependent factors affecting temporal changes in spatial distributions of eastern Bering Sea flatfish. Fisheries Oceanography, 17(5):396–410, September 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schismenou:2016:SCO

[SPG^{+}16] Eudoxia Schismenou, Miquel Palmer, Marianna Giannoulaki, Itziar Alvarez, Kostas Tsiaras, George Triantafyllou, and Stylianos Somarakis. Seasonal changes in otolith increment width trajectories and the effect of temperature on the daily growth rate of young sardines. Fisheries Oceanography, 25(4):362–372, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stephenson:2015:TLR

[SPLS15] Robert L. Stephenson, Michael J. Power, Shawn W. Laffan, and Iain M. Suthers. Tests of larval retention in a tidally energetic environment reveal the complexity of the spatial structure in herring populations. Fisheries Oceanography, 24(6):553–570, November 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sivel:2023:CET

[SPLY23] Elliot Sivel, Benjamin Planque, Ulf Lindstrøm, and Nigel G. Yoccoz. Combined effects of temperature and fishing mortality on the Barents Sea ecosystem stability. Fisheries Oceanography, 32(1):106–120, January 2023. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shepherd:2002:LSS

[SPM02] Travis Shepherd, Fred Page, and Bruce Macdonald. Length and sex-specific associations between spiny dogfish (Squalus acanthias) and hydrographic variables in the Bay of Fundy and Scotian Shelf. Fisheries Oceanography, 11(2):78–89, March 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Santos:2019:BEA

[SPM^{+}19] Régis Santos, Mário Pinho, Octávio Melo, João Gonçalves, Ana Leocádio, Ana Aranha, Gui Menezes, and Eduardo Isidro. Biological and ecological aspects of the deep-water red crab populations inhabiting isolated seamounts to the west of the Azores (Mid-Atlantic Ridge). Fisheries Oceanography, 28(6):723–734, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Seinor:2024:ESA

[SPM^{+}24] Kate Seinor, Steven W. Purcell, Hamish Malcolm, Stephen D. A. Smith, and Kirsten Benkendorff. Extended and spatially asynchronous reproductive periodicity in a harvested, warm-temperate rocky-reef gastropod (Turbinidae). Fisheries Oceanography, 33(1):e12653:1–e12653:??, January 2024. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Springer:1992:RWP

[Spr92] Alan M. Springer. A review: Walleye pollock in the North Pacific — how much difference do they really make? Fisheries Oceanography, 1(1):80–96, March 1992. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Santos:2020:PSM

[SPS^{+}20] Régis Santos, Ana Pabon, Wendell Silva, Hélder Silva, and Mário Pinho. Population structure and movement patterns of blackbelly rosefish in the NE Atlantic Ocean (Azores archipelago). Fisheries Oceanography, 29(3):227–237, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sydeman:2017:PRC

[SPT^{+}17] William J. Sydeman, John F. Piatt, Sarah Ann Thompson, Marisol García-Reyes, Scott A. Hatch, Mayumi L. Arimitsu, Leslie Slater, Jeffrey C. Williams, Nora A. Rojek, Stephani G. Zador, and Heather M. Renner. Puffins reveal contrasting relationships between forage fish and ocean climate in the North Pacific. Fisheries Oceanography, 26(4):

379–395, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic). See corrigendum [JJBCW17].

Springer:1996:SBP

[SPV96] Alan M. Springer, John F. Piatt, and Gus Van Vliet. Sea birds as proxies of marine habitats and food webs in the western Aleutian Arc. Fisheries Oceanography, 5(1):45–55, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stegmann:1999:AMR

[SQW^{+}99] Petra M. Stegmann, John A. Quinlan, Francisco E. Werner, Brian O. Blanton, and Peter Berrien. Atlantic menhaden recruitment to a southern estuary: defining potential spawning regions. Fisheries Oceanography, 8(S2):111–123, December 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sugimoto:1993:SOF

[SR93] Takashige Sugimoto and Kasijan Romimohrtarto. Seminar on oceanography for fisheries. Fisheries Oceanography, 2(2): 106–107, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sims:2002:CTL

[SR02] David W. Sims and Philip C. Reid. Congruent trends in long-term zooplankton decline in the north-east Atlantic and basking shark (Cetorhinus maximus) fishery catches off west Ireland. Fisheries Oceanography, 11(1):59–63, January 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Santora:2009:ISV

[SRCV09] Jarrod A. Santora, Christian S. Reiss, Anthony M. Cossio, and Richard R. Veit. Interannual spatial variability of krill (Euphausia superba) influences seabird foraging behavior near Elephant Island, Antarctica. Fisheries Oceanography, 18(1):20–35, January 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Suca:2018:CLS

[SRM^{+}18] Justin J. Suca, Leif K. Rasmuson, Estrella Malca, Trika Gerard, and John T. Lamkin. Characterizing larval swordfish habitat in the western tropical North Atlantic. Fisheries Oceanography, 27(3):246–258, May 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sakuma:1999:DVD

[SRR99] Keith M. Sakuma, Stephen Ralston, and Dale A. Roberts. Diel vertical distribution of postflexion larval Citharichthys spp. and Sebastes spp. off central California. Fisheries Oceanography, 8(1):68–76, March 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sundermeyer:2005:UCL

[SRR05] Miles A. Sundermeyer, Brian J. Rothschild, and Allan R. Robinson. Using commercial landings data to identify environmental correlates with distributions of fish stocks. Fisheries Oceanography, 14(1):47–63, January 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sakuma:2007:HFP

[SRR07] Keith M. Sakuma, Stephen Ralston, and Dale A. Roberts. High-frequency patterns in abundance of larval Pacific hake, Merluccius productus, and rockfish, Sebastes spp., at a single fixed station off central California. Fisheries Oceanography, 16(4):383–394, July 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Schumacher:1994:UEE

[SS94] J. D. Schumacher and P. J. Stabeno. Ubiquitous eddies of the eastern Bering Sea and their coincidence with concentrations of larval pollock. Fisheries Oceanography, 3(3):182– 190, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sunye:1998:ESV

[SS98] Patricia S. Sunyé and Jacques Servain. Effects of seasonal variations in meteorology and oceanography on the Brazilian sardine fishery. Fisheries Oceanography, 7(2):89–100, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Santos:2019:DEL

[SS19] Régis Vinícius Souza Santos and William Severi. Dynamics of early life-history stages of fish along an estuarine gradient.

Fisheries Oceanography, 28(4):402–418, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stefansdottir:2010:GSD

[SSM^{+}10] Lilja Stefansdottir, Jón Solmundsson, GuDRún Marteinsdottir, Kristján Kristinsson, and Jónas Páll Jonasson. Groundfish species diversity and assemblage structure in Icelandic waters during recent years of warming. Fisheries Oceanography, 19(1):42–62, January 2010. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sabates:2007:AAE

[SSP^{+}07] A. Sabatés, J. Salat, I. Palomera, M. Emelianov, M. L. Fernández De Puelles, and M. P. Olivar. Advection of anchovy (Engraulis encrasicolus) larvae along the Catalan continental slope (NW Mediterranean). Fisheries Oceanography, 16(2): 130–141, March 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Su:2011:ESS

[SSP^{+}11] Nan-Jay Su, Chi-Lu Sun, André E. Punt, Su-Zan Yeh, and Gerard Dinardo. Evaluation of a spatially sex-specific assessment method incorporating a habitat preference model for blue marlin (Makaira nigricans) in the Pacific Ocean. Fisheries Oceanography, 20(6):415–433, September 2011. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Su:2008:ESE

[SSPY08] Nan-Jay Su, Chi-Lu Sun, André E. Punt, and Su-Zan Yeh. Environmental and spatial effects on the distribution of blue marlin (Makaira nigricans) as inferred from data for longline fisheries in the Pacific Ocean. Fisheries Oceanography, 17(6): 432–445, November 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Staby:2013:VDB

[SSR13] Arved Staby, Jantra Srisomwong, and Rune Rosland. Variation in DVM behaviour of juvenile and adult pearlside (Maurolicus muelleri) linked to feeding strategies and related predation risk. Fisheries Oceanography, 22(2):90–101, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stenevik:2003:EVH

[SSSB03] Erling Kåre Stenevik, Morten Skogen, Svein Sundby, and David Boyer. The effect of vertical and horizontal distribution on retention of sardine (Sardinops sagax) larvae in the Northern Benguela — observations and modelling. Fisheries Oceanography, 12(3):185–200, May 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shibata:2017:EDF

[SSW^{+}17] Yasutoki Shibata, Tooru Sakuma, Toshihiro Wada, Yutaka Kurita, Takeshi Tomiyama, Manabu Yamada, Takashi Iwasaki, Takuji Mizuno, and Akibumi Yamanobe. Effect of decreased fishing effort off Fukushima on abundance of Japanese flounder (Paralichthys olivaceus) using an agestructured population model incorporating seasonal coastaloffshore migrations. Fisheries Oceanography, 26(2):193–207, March 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Suzuki:1995:SRL

[ST95] Toshikazu Suzuki and Akira Taniguchi. Sinking rate of loricae of some common tintinnid ciliates. Fisheries Oceanography, 4(3):257–263, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sugimoto:1997:IIV

[ST97] Takashige Sugimoto and Kazuaki Tadokoro. Interannual– interdecadal variations in zooplankton biomass, chlorophyll concentration and physical environment in the subarctic Pacific and Bering Sea. Fisheries Oceanography, 6(2):74–93, July 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sugimoto:1998:IVP

[ST98] Takashige Sugimoto and Kazuaki Tadokoro. Interdecadal variations of plankton biomass and physical environment in the North Pacific. Fisheries Oceanography, 7(3–4):289–299, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Steele:1998:CFR

[Ste98] John H. Steele. From carbon flux to regime shift. Fisheries Oceanography, 7(3–4):176–181, December 1998. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Shimizu:2009:TSL

[STI^{+}09] Yugo Shimizu, Kazutaka Takahashi, Shin-Ichi Ito, Shigeho Kakehi, Hiroaki Tatebe, Ichiro Yasuda, Akira Kusaka, and Tomoharu Nakayama. Transport of subarctic large copepods from the Oyashio area to the mixed water region by the coastal Oyashio intrusion. Fisheries Oceanography, 18(5): 312–327, September 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sharma:2013:RST

[SVEW^{+}13] Rishi Sharma, Luis A. Vélez-Espino, Alex C. Wertheimer, Nathan Mantua, and Robert C. Francis. Relating spatial and temporal scales of climate and ocean variability to survival of Pacific Northwest Chinook salmon (Oncorhynchus tshawytscha). Fisheries Oceanography, 22(1):14–31, January 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Scheuerell:2005:FCI

[SW05] Mark D. Scheuerell and John G. Williams. Forecasting climate-induced changes in the survival of Snake River spring/summer Chinook salmon (Oncorhynchus tshawytscha). Fisheries Oceanography, 14(6):448–457, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Swain:1999:CDA

[Swa99] Douglas P. Swain. Changes in the distribution of Atlantic cod (Gadus morhua) in the southern Gulf of St Lawrence — effects of environmental change or change in environmental preferences? Fisheries Oceanography, 8(1):1–17, March 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sepulveda:2020:IHM

[SWAAB20] Chugey A. Sepulveda, Michael Wang, Scott A. Aalbers, and Jaime R. Alvarado-Bremer. Insights into the horizontal movements, migration patterns, and stock affiliation of California swordfish. Fisheries Oceanography, 29(2):152–168, March 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sexton:2019:SPP

[SWS^{+}19] Stuart C. Sexton, Tim M. Ward, John Stewart, Kerrie M. Swadling, and Charlie Huveneers. Spawning patterns provide further evidence for multiple stocks of sardine (Sardinops sagax) off eastern Australia. Fisheries Oceanography, 28(1):18–32, January 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Stockwell:2001:ACS

[SWZ^{+}01] Dean A. Stockwell, Terry E. Whitledge, Stephan I. Zeeman, Kenneth O. Coyle, Jeffrey M. Napp, Richard D. Brodeur, Alexei I. Pinchuk, and George L. Hunt, Jr. Anomalous conditions in the south-eastern Bering Sea, 1997: nutrients, phytoplankton and zooplankton. Fisheries Oceanography, 10(1): 99–116, March 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Sassa:2009:DMA

[SYT^{+}09] Chiyuki Sassa, Keisuke Yamamoto, Youichi Tsukamoto, Yoshinobu Konishi, and Muneharu Tokimura. Distribution and migration of age-0 jack mackerel (Trachurus japonicus) in the East China and Yellow Seas, based on seasonal bottom trawl surveys. Fisheries Oceanography, 18(4):255–267, July 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Song:2008:EPL

[SZX^{+}08] Li Ming Song, Yu Zhang, Liu Xiong Xu, Wen Xin Jiang, and Jia Qiao Wang. Environmental preferences of longlining for yellowfin tuna (Thunnus albacares) in the tropical high seas of the Indian Ocean. Fisheries Oceanography, 17(4):239–253, July 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takasuka:2006:EDG

[TA06] Akinori Takasuka and Ichiro Aoki. Environmental determinants of growth rates for larval Japanese anchovy Engraulis japonicus in different waters. Fisheries Oceanography, 15(2): 139–149, March 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Taki:2004:DLH

[Tak04] Kenji Taki. Distribution and life history of Euphausia pacifica off northeastern Japan. Fisheries Oceanography, 13(S1): 34–43, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanasichuk:1999:IVA

[Tan99] Tanasichuk. Interannual variation in the availability and utilization of euphausiids as prey for Pacific hake (Merluccius productus) along the south-west coast of Vancouver Island. Fisheries Oceanography, 8(2):150–156, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanasichuk:2002:IIV

[Tan02] R. W. Tanasichuk. Implications of interannual variability in euphausiid population biology for fish production along the south-west coast of Vancouver Island: a synthesis. Fisheries Oceanography, 11(1):18–30, January 2002. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanasichuk:2017:IBB

[Tan17a] R. W. Tanasichuk. An investigation of the biological basis of recruitment, growth and adult survival rate variability of Pacific herring (Clupea pallasi) from British Columbia: a synthesis. Fisheries Oceanography, 26(4):413–438, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Thiaw:2017:EEC

[TAN^{+}17b] Modou Thiaw, Pierre-Amaël Auger, Fambaye Ngom, Timothée Brochier, Saliou Faye, Ousmane Diankha, and Patrice Brehmer. Effect of environmental conditions on the seasonal and inter-annual variability of small pelagic fish abundance off North-West Africa: The case of both Senegalese sardinella. Fisheries Oceanography, 26(5):583–601, September 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tian:2004:MIO

[TAS04] Yongjun Tian, Tatsuro Akamine, and Maki Suda. Modeling the influence of oceanic-climatic changes on the dynamics of Pacific saury in the northwestern Pacific using a life cycle model. Fisheries Oceanography, 13(S1):125–137, December 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Trites:1992:CBG

[TB92] Andrew W. Trites and Michael A. Bigg. Changes in body growth of northern fur seals from 1958 to 1974: density effects or changes in the ecosystem? Fisheries Oceanography, 1 (2):127–136, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Trathan:2003:OVC

[TBB^{+}03] P. N. Trathan, A. S. Brierley, M. A. Brandon, D. G. Bone, C. Goss, S. A. Grant, E. J. Murphy, and J. L. Watkins. Oceanographic variability and changes in Antarctic krill (Euphausia superba) abundance at South Georgia. Fisheries Oceanography, 12(6):569–583, November 2003. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tsai:1998:FAL

[TCC^{+}98] Chu-Fa Tsai, Pei-Yu Chen, Chang-Po Chen, Ming-Anne Lee, Guang-Yaw Shiah, and Kuo-Tien Lee. Fluctuation in abundance of larval anchovy and environmental conditions in coastal waters off south-western Taiwan as associated with the El Niño–Southern Oscillation. Fisheries Oceanography, 6 (4):238–249, February 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takahashi:2012:RGR

[TCL^{+}12] Motomitsu Takahashi, David M. Checkley, Jr., Marisa N. C. Litz, Richard D. Brodeur, and William T. Peterson. Responses in growth rate of larval northern anchovy (Engraulis mordax) to anomalous upwelling in the northern California Current. Fisheries Oceanography, 21(6):393–404, November 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tadokoro:2005:IVN

[TCO^{+}05] Kazuaki Tadokoro, Sanae Chiba, Tsuneo Ono, Takashi Midorikawa, and Toshiro Saino. Interannual variation in Neocalanus biomass in the Oyashio waters of the western North Pacific. Fisheries Oceanography, 14(3):210–222, May 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tian:2009:SAS

[TCS^{+}09] Rucheng C. Tian, Changsheng Chen, Kevin D. E. Stokesbury, Brian J. Rothschild, Qichun Xu, Song Hu, Geoffrey Cowles, Bradley P. Harris, and Michael C. Marino II. Sensitivity analysis of sea scallop (Placopecten magellanicus) larvae trajectories to hydrodynamic model configuration on Georges Bank and adjacent coastal regions. Fisheries Oceanography, 18(3):173–184, May 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Temming:2002:LCC

[TD02] Axel Temming and Ulrich Damm. Life cycle of Crangon crangon in the North Sea: a simulation of the timing of recruitment as a function of the seasonal temperature signal. Fisheries Oceanography, 11(1):45–58, January 2002. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tilburg:2009:HCB

[TDE09] Charles E. Tilburg, Ana I. Dittel, and Charles E. Epifanio. High concentrations of blue crab (Callinectes sapidus) larvae along the offshore edge of a coastal current: effects of convergent circulation. Fisheries Oceanography, 18(3):135– 146, May 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tittensor:2003:MDS

[TDT03] D. P. Tittensor, B. DeYoung, and C. L. Tang. Modelling the distribution, sustainability and diapause emergence timing of the copepod Calanus finmarchicus in the Labrador Sea. Fisheries Oceanography, 12(4–5):299–316, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanaka:2008:VDJ

[TF08] Yuji Tanaka and Peter J. S. Franks. Vertical distributions of Japanese sardine (Sardinops melanostictus) eggs: comparison of observations and a wind-forced Lagrangian mixing model. Fisheries Oceanography, 17(2):89–100, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tiedemann:2017:DUI

[TFB^{+}17] Maik Tiedemann, Heino O. Fock, Patrice Brehmer, Julian Döring, and Christian Möllmann. Does upwelling intensity determine larval fish habitats in upwelling ecosystems? The case of Senegal and Mauritania. Fisheries Oceanography, 26 (6):655–667, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Thompson:2019:EOC

[TGRS^{+}19] Sarah Ann Thompson, Marisol García-Reyes, William J. Sydeman, Mayumi L. Arimitsu, Scott A. Hatch, and John F. Piatt. Effects of ocean climate on the length and condition of forage fish in the Gulf of Alaska. Fisheries Oceanography, 28(6):658–671, November 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Thomson:2011:MTR

[TH11] Richard E. Thomson and Roy A. S. Hourston. A matter of timing: the role of ocean conditions in the initiation of spawning migration by late-run Fraser River sockeye salmon (Oncorhynchus nerka). Fisheries Oceanography, 20 (1):47–65, January 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tracey:2012:EDT

[THH12] Sean R. Tracey, Klaas Hartmann, and Alistair J. Hobday. The effect of dispersal and temperature on the early life history of a temperate marine fish. Fisheries Oceanography, 21(5):336–347, September 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tolimieri:2018:ODS

[THL^{+}18] N. Tolimieri, M. A. Haltuch, Q. Lee, M. G. Jacox, and S. J. Bograd. Oceanographic drivers of sablefish recruitment in the California Current. Fisheries Oceanography, 27(5):458– 474, September 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tadokoro:1996:CCS

[TID^{+}96] Kazuaki Tadokoro, Yukimasa Ishida, Nancy D. Davis, Shoji Ueyanagi, and Takashige Sugimoto. Change in chum salmon (Oncorhynchus keta) stomach contents associated with fluctuation of pink salmon (O. gorbuscha) abundance in the central subarctic Pacific and Bering Sea. Fisheries Oceanography, 5(2):89–99, June 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Thomson:1992:IOC

[TIH^{+}92] K. A. Thomson, W. J. Ingraham, M. C. Healey, P. H. Leblond, C. Groot, and C. G. Healey. The influence of ocean currents on latitude of landfall and migration speed of sockeye salmon returning to the Fraser River. Fisheries Oceanography, 1(2):163–179, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tang:2003:DSV

[TJW^{+}03] Qisheng Tang, Xianshi Jin, Jun Wang, Zhimeng Zhuang, Yi Cui, and Tianxiang Meng. Decadal-scale variations of ecosystem productivity and control mechanisms in the Bohai Sea. Fisheries Oceanography, 12(4–5):223–233, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takikawa:2008:HCF

[TKH08] Tetsutaro Takikawa, Minoru Kitamura, and Naho Horimoto. Horizontal current field, ADCP backscatter, and plankton distribution in Sagami Bay, Japan. Fisheries Oceanography, 17(4):254–262, July 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Terumoto:2022:TCB

[TKM^{+}22] Takayuki Terumoto, Isao Kudo, Koji Miyoshi, Akiyoshi Shinada, and Akira Miyazono. Tight coupling between primary productivity, export production, and the growth of benthic scallops in the coastal region of the Okhotsk Sea along Hokkaido. Fisheries Oceanography, 31(2):205–216, March 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanaka:2011:NST

[TKMS11] Kiyoshi Tanaka, Teruhisa Komatsu, Yutaka Michida, and Sei-Ichi Saitoh. A numerical study on the transport of eggs and larvae of Sergia lucens in Suruga Bay, Japan. Fisheries Oceanography, 20(3):206–218, May 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takasuka:2014:ODP

[TKO^{+}14] Akinori Takasuka, Hiroshi Kuroda, Takeshi Okunishi, Yugo Shimizu, Yuichi Hirota, Hiroshi Kubota, Hideo Sakaji, Ryo Kimura, Shin-Ichi Ito, and Yoshioki Oozeki. Occurrence and density of Pacific saury Cololabis saira larvae and juveniles in relation to environmental factors during the winter spawning season in the Kuroshio Current system. Fisheries Oceanography, 23(4):304–321, July 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takami:2017:EME

[TKW^{+}17] Hideki Takami, Tomohiko Kawamura, Nam-Il Won, Daisuke Muraoka, Jun Hayakawa, and Toshihiro Onitsuka. Effects of macroalgal expansion triggered by the 2011 earthquake and tsunami on recruitment density of juvenile abalone Haliotis discus hannai at Oshika Peninsula, northeastern Japan. Fisheries Oceanography, 26(2):141–154, March 2017. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tomkiewicz:1998:OID

[TLS98] Jonna Tomkiewicz, Klaus M. Lehmann, and Michael A. St. John. Oceanographic influences on the distribution of Baltic cod, Gadus morhua, during spawning in the Bornholm Basin of the Baltic Sea. Fisheries Oceanography, 7(1):48–62, April 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Trites:2007:BFD

[TMM^{+}07] Andrew W. Trites, Arthur J. Miller, Herbert D. G. Maschner, Michael A. Alexander, Steven J. Bograd, John A. Calder, Antonietta Capotondi, Kenneth O. Coyle, Emanuele Di Lorenzo, Bruce P. Finney, Edward J. Gregr, Chester E. Grosch, Steven R. Hare, George L. Hunt, Jr., Jaime Jahncke, Nancy B. Kachel, Hey-Jin Kim, Carol Ladd, Nathan J. Mantua, Caren Marzban, Wieslaw Maslowski, Roy Mendelssohn, Douglas J. Neilson, Stephen R. Okkonen, James E. Overland, Katherine L. Reedy-Maschner, Thomas C. Royer, Franklin B. Schwing, Julian X. L. Wang, and Arliss J. Winship. Bottom-up forcing and the decline of Steller sea lions (Eumetopias jubatus) in Alaska: assessing the ocean climate hypothesis. Fisheries Oceanography, 16(1):46–67, January 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Temperoni:2020:SSV

[TMMM20] Brenda Temperoni, Agueda E. Massa, Patricia Martos, and Marina Marrari. Seasonal and spatial variability in condition of age-0+ Argentine hake Merluccius hubbsi Marini, 1933, in the San Jorge Gulf (Argentina): a bottom-up perspective. Fisheries Oceanography, 29(6):597–611, November 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takeshige:2015:SIC

[TMN^{+}15] Aigo Takeshige, Yoichi Miyake, Hideaki Nakata, Takashi Kitagawa, and Shingo Kimura. Simulation of the impact of climate change on the egg and larval transport of Japanese anchovy (Engraulis japonicus) off Kyushu Island, the western coast of Japan. Fisheries Oceanography, 24(5):445–462, September 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takahashi:2008:DPL

[TMS^{+}08] Masanori Takahashi, Noritaka Mochioka, Sekio Shinagawa, Akihiko Yatsu, and Akinobu Nakazono. Distribution patterns of leptocephali in the Kuroshio–Oyashio transitional region of the western North Pacific. Fisheries Oceanography, 17(3):165–177, May 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tone:2022:MSB

[TNC^{+}22] Kazuki Tone, Yosuke Nakamura, Wei-Chuan Chiang, Hsin-Ming Yeh, Sheng-Tai Hsiao, Chun-Huei Li, Kazuyoshi Komeyama, Masanori Tomisaki, Takamasa Hasegawa, Takashi Sakamoto, Itsumi Nakamura, Yoshitaka Sakakura, and Ryo Kawabe. Migration and spawning behavior of the greater amberjack Seriola dumerili in eastern Taiwan. Fisheries Oceanography, 31(1):1–18, January 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takasuka:2016:GVP

[TNK^{+}16] Akinori Takasuka, Kirara Nishikawa, Hiroshi Kuroda, Takeshi Okunishi, Yugo Shimizu, Hideo Sakaji, Shin-Ichi Ito, Tadashi Tokai, and Yoshioki Oozeki. Growth variability of Pacific saury Cololabis saira larvae under contrasting environments across the Kuroshio axis: survival potential of minority versus majority. Fisheries Oceanography, 25(4): 390–406, July 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takatsu:2002:SDF

[TNM^{+}02] Tetsuya Takatsu, Toshikuni Nakatani, Takanori Miyamoto, Kouji Kooka, and Toyomi Takahashi. Spatial distribution and feeding habits of Pacific cod (Gadus macrocephalus) larvae in Mutsu Bay, Japan. Fisheries Oceanography, 11(2): 90–101, March 2002. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanasichuk:2011:IBB

[TR11] R. W. Tanasichuk and R. Routledge. An investigation of the biological basis of return variability for sockeye salmon (Oncorhynchus nerka) from Great Central and Sproat lakes, Vancouver Island. Fisheries Oceanography, 20(6):462–478, November 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Teixeira:2020:WBE

[TSG^{+}20] Célia M. Teixeira, Priscila M. Silva, Rita Gamito, Sandra Amoroso, Nuno Lopes, Catarina Freitas, Henrique N. Cabral, and José L. Costa. Which are the best environmental conditions for catching the beach-seine target species? A Portuguese case study. Fisheries Oceanography, 29(3):276– 289, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Taguchi:1992:HSV

[TSK^{+}92] Satoru Taguchi, Hiroaki Saito, Hiromi Kasai, Tokihiro Kono, and Yasuhiro Kawasaki. Hydrography and spatial variability in the size distribution of phytoplankton along the Kurile Islands in the western subarctic Pacific Ocean. Fisheries Oceanography, 1(3):227–237, September 1992. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Terazaki:1995:HDS

[TSK^{+}95] Makoto Terazaki, Hiroaki Saito, Hiromi Kasai, Tokihiro Kono, Yasuhiro Kawasaki, and Satoru Taguchi. Horizontal distribution and seasonal variability of the epipelagic chaetognath Sagitta elegans in relation to hydrography in the western subarctic Pacific Ocean. Fisheries Oceanography, 4 (2):158–170, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tsuda:2004:LHE

[TSK04] Atsushi Tsuda, Hiroaki Saito, and Hiromi Kasai. Life histories of Eucalanus bungii and Neocalanus cristatus (Copepoda: Calanoida) in the western subarctic Pacific Ocean. Fisheries Oceanography, 13(S1):10–20, December 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takahashi:2022:LED

[TSK^{+}22] Motomitsu Takahashi, Chiyuki Sassa, Satoshi Kitajima, Mari Yoda, and Youichi Tsukamoto. Linking environmental drivers, juvenile growth, and recruitment for Japanese jack mackerel Trachurus japonicus in the Sea of Japan. Fisheries Oceanography, 31(1):70–83, January 2022. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tu:2012:UCF

[TTC^{+}12] Chen-Yi Tu, Yu-Heng Tseng, Tai-Sheng Chiu, Mao-Lin Shen, and Chih-Hao Hsieh. Using coupled fish behavior– hydrodynamic model to investigate spawning migration of Japanese anchovy, Engraulis japonicus, from the East China Sea to Taiwan. Fisheries Oceanography, 21(4):255–268, July 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tu:2015:ECT

[TTH15] Chen-Yi Tu, Yongjun Tian, and Chih-Hao Hsieh. Effects of climate on temporal variation in the abundance and distribution of the demersal fish assemblage in the Tsushima Warm Current region of the Japan Sea. Fisheries Oceanography, 24(2):177–189, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanaka:2020:OPB

[TTI^{+}20] Yosuke Tanaka, Atsushi Tawa, Taiki Ishihara, Etsuro Sawai, Misato Nakae, Masachika Masujima, and Taketoshi Kodama. Occurrence of Pacific bluefin tuna Thunnus orientalis larvae off the Pacific coast of Tohoku area, northeastern Japan: Possibility of the discovery of the third spawning ground. Fisheries Oceanography, 29(1):46–51, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tanaka:2023:RRE

[TTY^{+}23] Shota Tanaka, Shizuna Togoshi, Naotaka Yasue, Corinne M. Burns, Dominique Robert, and Akinori Takasuka. Revisiting the role of early life growth for survival potential in three clupeoid species. Fisheries Oceanography, 32(2):245–254, March 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takahashi:2005:ETF

[TW05] Motomitsu Takahashi and Yoshiro Watanabe. Effects of temperature and food availability on growth rate during late larval stage of Japanese anchovy (Engraulis japonicus) in the Kuroshio–Oyashio transition region. Fisheries Oceanography, 14(3):223–235, May 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takami:2013:IME

[TWK13] Hideki Takami, Nam-Il Won, and Tomohiko Kawamura. Impacts of the 2011 mega-earthquake and tsunami on abalone Haliotis discus hannai and sea urchin Strongylocentrotus nudus populations at Oshika Peninsula, Miyagi, Japan. Fisheries Oceanography, 22(2):113–120, March 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takahashi:2001:GLE

[TWKW01] M. Takahashi, Y. Watanabe, T. Kinoshita, and C. Watanabe. Growth of larval and early juvenile Japanese anchovy, Engraulis japonicus, in the Kuroshio–Oyashio transition region. Fisheries Oceanography, 10(2):235–247, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tomiyama:2004:ETA

[TY04] Minoru Tomiyama and Shigeaki Yanagibashi. Effect of temperature, age class, and growth on induction of aestivation in Japanese sandeel (Ammodytes personatus) in Ise Bay, central Japan. Fisheries Oceanography, 13(2):81–90, March 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Tyler:1992:CRC

[Tyl92] A. V. Tyler. A context for recruitment correlations: why marine fisheries biologists should still look for them. Fisheries Oceanography, 1(1):97–107, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Takasuka:2021:DDE

[TYO21] Akinori Takasuka, Michio Yoneda, and Yoshioki Oozeki. Density-dependent egg production in chub mackerel in the Kuroshio Current system. Fisheries Oceanography, 30(1): 38–50, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Uye:1999:GVT

[UIU^{+}99] S. Uye, N. Iwamoto, T. Ueda, H. Tamaki, and K. Nakahira. Geographical variations in the trophic structure of the plankton community along a eutrophic–mesotrophic–oligotrophic transect. Fisheries Oceanography, 8(3):227–237, September 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Uchiyama:2020:MBD

[UMK20] Tadayasu Uchiyama, Franz J. Mueter, and Gordon H. Kruse. Multispecies biomass dynamics models reveal effects of ocean temperature on predation of juvenile pollock in the eastern Bering Sea. Fisheries Oceanography, 29(1):10–22, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Uehara:2006:AJY

[UTMS06] Shinji Uehara, Christopher T. Taggart, Takumi Mitani, and Iain M. Suthers. The abundance of juvenile yellowtail (Seriola quinqueradiata) near the Kuroshio: the roles of drifting seaweed and regional hydrography. Fisheries Oceanography, 15(5):351–362, September 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Uye:1992:TFG

[UYF92] Shin-Ichi Uye, Toru Yamaoka, and Toshiro Fujisawa. Are tidal fronts good recruitment areas for herbivorous copepods? Fisheries Oceanography, 1(3):216–226, September 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vanes:1995:SCC

[VAFG95] Anne Gro Vea Sal Vanes, Dag Aksnes, Jan Helge Fosså, and Jarl Giske. Simulated carrying capacities of fish in Norwegian fjords. Fisheries Oceanography, 4(1):17–32, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

VanDerLingen:1998:AAD

[VCB^{+}98] C. D. Van Der Lingen, David M. Checkley, Jr., Manuel Barange, L. Hutchings, and K. Osgood. Assessing the abundance and distribution of eggs of sardine, Sardinops sagax, and round herring, Etrumeus whiteheadi, on the western Agulhas Bank, South Africa, using a continuous, underway fish egg sampler. Fisheries Oceanography, 7(1):35–47, April 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vlietstra:2005:TFA

[VCKH05] Lucy S. Vlietstra, Kenneth O. Coyle, Nancy B. Kachel, and George L. Hunt, Jr. Tidal front affects the size of prey used by a top marine predator, the short-tailed shearwater (Puffinus tenuirostris). Fisheries Oceanography, 14(S1):196–211, November 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Voss:2008:EFI

[VDHF08] R. Voss, M. Dickmann, H.-H. Hinrichsen, and J. Floeter. Environmental factors influencing larval sprat Sprattus sprattus feeding during spawning time in the Baltic Sea. Fisheries Oceanography, 17(3):219–230, May 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vasbinder:2024:SSP

[VFS^{+}24] Kelly Vasbinder, Jerome Fiechter, Jarrod A. Santora, James J. Anderson, Nate Mantua, Steve T. Lindley, David D. Huff, and Brian K. Wells. Size-selective predation effects on juvenile Chinook salmon cohort survival off Central California evaluated with an individual-based model. Fisheries Oceanography, 33(1):e12654:1–e12654:??, January 2024. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Villamor:2011:ECN

[VGPL^{+}11] B. Villamor, C. Gonzalez-Pola, A. Lavín, L. Valdés, A. Lago De Lanzós, C. Franco, J. M. Cabanas, M. Bernal, C. Hernandez, M. Iglesias, P. Carrera, and C. Porteiro. Environmental control of Northeast Atlantic mackerel (Scomber scombrus) recruitment in the southern Bay of Biscay: case study of failure in the year 2000. Fisheries Oceanography, 20(5):397–414, September 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vanderlaan:2014:EIA

[VHCN14] Angelia S. M. Vanderlaan, Alex R. Hanke, Joël Chassé, and John D. Neilson. Environmental influences on Atlantic bluefin tuna (Thunnus thynnus) catch per unit effort in the southern Gulf of St. Lawrence. Fisheries Oceanography, 23 (1):83–100, January 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Voss:1999:VDL

[VHJ99] Rüdiger Voss, Hans-Harald Hinrichsen, and Mike St John. Variations in the drift of larval cod (Gadus morhua L.) in the Baltic Sea: combining field observations and modelling.

Fisheries Oceanography, 8(3):199–211, September 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Villegas-Hernandez:2015:RCA

[VHLM15] Harold Villegas-Hernández, Josep Lloret, and Marta Muñoz. Reproduction, condition and abundance of the Mediterranean bluefish (Pomatomus saltatrix) in the context of sea warming. Fisheries Oceanography, 24(1):42–56, January 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vastano:1992:OAF

[VIS92] Andrew C. Vastano, Lewis S. Incze, and James D. Schumacher. Observation and analysis of fishery processes: larval pollock at Shelikof Strait, Alaska. Fisheries Oceanography, 1 (1):20–31, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Visser:1999:LBS

[VJ99] A. W. Visser and S. H. Jónasdóttir. Lipids, buoyancy and the seasonal vertical migration of Calanus finmarchicus. Fisheries Oceanography, 8(S1):100–106, September 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vaughan:2001:PVP

[VMG01] Shari L. Vaughan, Christopher N. K. Mooers, and Shelton M. Gay III. Physical variability in Prince William Sound during the SEA Study (1994–98). Fisheries Oceanography, 10(S1):58–80, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vestfals:2023:SSD

[VMT^{+}23] Cathleen D. Vestfals, Kristin N. Marshall, Nick Tolimieri, Mary E. Hunsicker, Aaron M. Berger, Ian G. Taylor, Michael G. Jacox, and Brendan D. Turley. Stage-specific drivers of Pacific hake (Merluccius productus) recruitment in the California Current Ecosystem. Fisheries Oceanography, 32(4):352–389, July 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

VabO:1997:IBM

[VN97] Rune VabØ and Leif NØTtestad. An individual based model of fish school reactions: predicting antipredator behaviour as observed in nature. Fisheries Oceanography, 6(3):155–171, October 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Venegas:2019:CIV

[VOB^{+}19] Roberto Venegas, Thomas Oliver, Russell E. Brainard, Mudjekeewis Santos, Rollan Geronimo, and Matthew Widlansky. Climate-induced vulnerability of fisheries in the Coral Triangle: Skipjack tuna thermal spawning habitats. Fisheries Oceanography, 28(2):117–130, March 2019. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vitaliano:2013:BSD

[VPRG13] Joseph Vitaliano, David Packer, Robert Reid, and Vincent Guida. Broad-scale, dense amphipod tube aggregations on the sea bed: implications for resource species that utilize benthic habitats. Fisheries Oceanography, 22(1):61–67, January 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vikebo:2007:IRT

[VSÅO07] F. B. Vikebø, S. Sundby, Bjørn Ådlandsvik, and O. H. Otterå. Impacts of a reduced thermohaline circulation on transport and growth of larvae and pelagic juveniles of arcto-Norwegian cod (Gadus morhua). Fisheries Oceanography, 16(3):216–228, May 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vargas-Yanez:2020:VOM

[VYGT^{+}20] Manuel Vargas-Yáñez, Ana Giráldez, Pedro Torres, María González, Maria del Carmen García-Martínez, and Francina Moya. Variability of oceanographic and meteorological conditions in the northern Alboran Sea at seasonal, inter-annual and long-term time scales and their influence on sardine (Sardina pilchardus Walbaum 1792) landings. Fisheries Oceanography, 29(5):367–380, September 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Vichi:1998:SMM

[VZP98] M. Vichi, M. Zavatarelli, and N. Pinardi. Seasonal modulation of microbially mediated carbon fluxes in the northern Adriatic Sea — a model study. Fisheries Oceanography, 7 (3–4):182–190, December 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ware:1992:JCJ

[War92] Dan Ware. Joint Canada–Japan Fisheries Oceanography Workshop. Fisheries Oceanography, 1(1):108–110, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ware:1995:CHC

[War95] D. M. Ware. A century and a half of change in the climate of the NE Pacific. Fisheries Oceanography, 4(4):267–277, December 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Watanabe:2017:MTI

[Wat17] Kyuji Watanabe. Multi-timescale interactions between pink and chum salmon catch per unit effort in the Bering Sea. Fisheries Oceanography, 26(4):498–506, July 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Woods:1993:DDW

[WB93] John Woods and Wolfgang Barkmann. Diatom demography in winter — simulated by the Lagrangian ensemble method. Fisheries Oceanography, 2(3–4):202–222, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Werner:1999:PON

[WBQL99] Francisco E. Werner, Brian O. Blanton, John A. Quinlan, and Richard A. Luettich, Jr. Physical oceanography of the North Carolina continental shelf during the fall and winter seasons: implications for the transport of larval menhaden. Fisheries Oceanography, 8(S2):7–21, December 1999. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Willette:2001:EPI

[WCP^{+}01] T. M. Willette, R. T. Cooney, V. Patrick, D. M. Mason, G. L. Thomas, and D. Scheel. Ecological processes influencing mortality of juvenile pink salmon (Oncorhynchus gorbuscha) in Prince William Sound, Alaska. Fisheries Oceanography, 10 (S1):14–41, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wyllie-Echeverria:1998:YYV

[WEW98] Tina Wyllie-Echeverria and Warren S. Wooster. Year-to-year variations in Bering Sea ice cover and some consequences for fish distributions. Fisheries Oceanography, 7(2):159–170, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Waring:1993:SWA

[WFRS93] Gordon T. Waring, Carol P. Fairfield, Cynthia M. Ruhsam, and Margaret Sano. Sperm whales associated with Gulf Stream features off the north-eastern USA shelf. Fisheries Oceanography, 2(2):101–105, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wells:2006:CBA

[WGFR06] Brian K. Wells, Churchill B. Grimes, John C. Field, and Christian S. Reiss. Covariation between the average lengths of mature Coho (Oncorhynchus kisutch) and Chinook salmon (O. tshawytscha) and the ocean environment. Fisheries Oceanography, 15(1):67–79, January 2006. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wells:2008:RBO

[WGS^{+}08] Brian K. Wells, Churchill B. Grimes, John G. Sneva, Scott Mcpherson, and James B. Waldvogel. Relationships between oceanic conditions and growth of Chinook salmon (Oncorhynchus tshawytscha) from California, Washington, and Alaska, USA. Fisheries Oceanography, 17(2):101–125, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wells:2007:QEW

[WGW07] Brian K. Wells, Churchill B. Grimes, and James B. Waldvogel. Quantifying the effects of wind, upwelling, curl, sea surface temperature and sea level height on growth and maturation of a California Chinook salmon (Oncorhynchus tshawytscha) population. Fisheries Oceanography, 16(4): 363–382, July 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Walters:1992:MPS

[WHT92] Carl J. Walters, Charles G. Hannah, and Keith Thomson. A microcomputer program for stimulating effects of physical transport processes on fish larvae. Fisheries Oceanography, 1(1):11–19, March 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Willette:2001:FBJ

[Wil01] T. M. Willette. Foraging behaviour of juvenile pink salmon (Oncorhynchus gorbuscha) and size-dependent predation risk. Fisheries Oceanography, 10(S1):110–131, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wilson:2004:BSC

[Wil04] Steven G. Wilson. Basking sharks (Cetorhinus maximus) schooling in the southern Gulf of Maine. Fisheries Oceanography, 13(4):283–286, July 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wyatt:1993:NAT

[WJ93] T. Wyatt and I. R. Jenkinson. The North Atlantic turbine: views of production processes from a mainly North Atlantic perspective. Fisheries Oceanography, 2(3–4):231–243, December 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wilhelm:2015:SNA

[WJM15] Margit R. Wilhelm, Astrid Jarre, and Coleen L. Moloney. Spawning and nursery areas, longitudinal and cross-shelf migrations of the Merluccius capensis stock in the northern Benguela. Fisheries Oceanography, 24(S1):31–45, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wang:2001:NSS

[WJP^{+}01] Jia Wang, Meibing Jin, E. Vincent Patrick, Jennifer R. Allen, David L. Eslinger, Christophee N. K. Mooers, and R. Ted Cooney. Numerical simulations of the seasonal circulation patterns and thermohaline structures of Prince William Sound, Alaska. Fisheries Oceanography, 10(S1):132–148, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wieland:1997:PVD

[WJT97] Kai Wieland and Astrid Jarre-Teichmann. Prediction of vertical distribution and ambient development temperature of Baltic cod, Gadus morhua L., eggs. Fisheries Oceanography, 6(3):172–187, October 1997. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Woodworth-Jefcoats:2020:TEP

[WJW20] Phoebe A. Woodworth-Jefcoats and Johanna L. K. Wren. Toward an environmental predictor of tuna recruitment. Fisheries Oceanography, 29(5):436–441, September 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Watanabe:2003:DCA

[WK03] Hikaru Watanabe and Kouichi Kawaguchi. Decadal change in abundance of surface migratory myctophid fishes in the Kuroshio region from 1957 to 1994. Fisheries Oceanography, 12(2):100–111, March 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wenner:2005:KFI

[WKB^{+}05] E. L. Wenner, D. M. Knott, C. A. Barans, S. Wilde, J. O. Blanton, and Julie Amft. Key factors influencing transport of white shrimp (Litopenaeus setiferus) post-larvae into the Ossabaw Sound system, Georgia, USA. Fisheries Oceanography, 14(3):175–194, May 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wroblewski:1995:WDM

[WKN^{+}95] J. S. Wroblewski, D. W. Kulka, S. Narayanan, A. M. Oake, A. G. Collier, and B. D. Mcgrath. Winter distribution and movements of northern Atlantic cod (Gadus morhua) along the Newfoundland–Labrador continental shelf edge derived from observations on commercial trawlers. Fisheries Oceanography, 4(2):128–146, June 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wang:2018:MMP

[WKR^{+}18] Lifei Wang, Lisa A. Kerr, Nicholas R. Record, Eric Bridger, Benjamin Tupper, Katherine E. Mills, Edward M. Armstrong, and Andrew J. Pershing. Modeling marine pelagic fish species spatiotemporal distributions utilizing a maximum entropy approach. Fisheries Oceanography, 27(6):571– 586, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wilson:2021:IVC

[WL21] Matthew T. Wilson and Ned Laman. Interannual variation in the coastal distribution of a juvenile gadid in the northeast Pacific Ocean: The relevance of wind and effect on recruitment. Fisheries Oceanography, 30(1):3–22, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wang:1998:FAZ

[WLWZ98] Rong Wang, Chaolun Li, Ke Wang, and Wuchang Zhang. Feeding activities of zooplankton in the Bohai Sea. Fisheries Oceanography, 7(3–4):265–271, December 1998. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ward:2006:DHM

[WM06] Peter J. Ward and Ransom A. Myers. Do habitat models accurately predict the depth distribution of pelagic fishes? Fisheries Oceanography, 15(1):60–66, January 2006. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Walker:2000:DVT

[WMD^{+}00] Robert V. Walker, Katherine W. Myers, Nancy D. Davis, Kerim Y. Aydin, Kevin D. Friedland, H. Richard Carlson, George W. Boehlert, Shigehiko Urawa, Yasuhiro Ueno, and Gen Anma. Diurnal variation in thermal environment experienced by salmonids in the North Pacific as indicated by data storage tags. Fisheries Oceanography, 9(2):171–186, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ward:2006:PEN

[WMD^{+}06] Tim M. Ward, Lachlan J. Mcleay, Wetjens F. Dimmlich, Paul J. Rogers, Sam Mcclatchie, Roger Matthews, Jochen Kämpf, and Paul D. Van Ruth. Pelagic ecology of a northern boundary current system: effects of upwelling on the production and distribution of sardine (Sardinops sagax), anchovy (Engraulis australis) and southern bluefin tuna (Thunnus maccoyii) in the Great Australian Bight. Fisheries Oceanography, 15(3):191–207, May 2006. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Watanabe:1999:DVM

[WMK^{+}99] Hikaru Watanabe, Masatoshi Moku, Kouichi Kawaguchi, Kimie Ishimaru, and Akinori Ohno. Diel vertical migration of myctophid fishes (family Myctophidae) in the transitional waters of the western North Pacific. Fisheries Oceanography, 8(2):115–127, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wall:2009:LBE

[WMKR09] Carrie C. Wall, Frank E. Muller-Karger, and Mitchell A. Roffer. Linkages between environmental conditions and recreational king mackerel (Scomberomorus cavalla) catch off west-central Florida. Fisheries Oceanography, 18(3):185– 199, May 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wooster:1993:NPM

[Woo93] Warren S. Wooster. The North Pacific Marine Science Organization (PICES). Fisheries Oceanography, 2(1):40, March 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wooster:1995:PTA

[Woo95] Warren S. Wooster. PICES Third Annual Meeting. Fisheries Oceanography, 4(1):80, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wooster:1997:PFA

[Woo97] Warren S. Wooster. PICES Fourth Annual Meeting. Erratum. Fisheries Oceanography, 6(1):49–50, March 1997. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Welch:1993:VIT

[WP93] David W. Welch and Timothy R. Parsons. δ^{13}C–δ ^{15}N values as indicators of trophic position and competitive overlap for Pacific salmon (Oncorhynchus spp.). Fisheries Oceanography, 2(1):11–23, March 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Werner:1993:IMA

[WPL^{+}93] Francisco E. Werner, Fred H. Page, Daniel R. Lynch, John W. Loder, R. Gregory Lough, R. Ian Perry, David A. Greenberg, and Michael M. Sinclair. Influences of mean advection and simple behavior on the distribution of cod and haddock early life stages on Georges Bank. Fisheries Oceanography, 2(2):43–64, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wiedmann:2012:RPB

[WPN12] Magnus Aune Wiedmann, Ole-Petter Pedersen, and Einar Magnus Nilssen. The recruitment process of the Barents Sea capelin (Mallotus villosus) stock, 2001–2003. Fisheries Oceanography, 21(6):379–392, November 2012. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Williams:2000:PHCb

[WQ00] Erik H. Williams and Terrance J. Quinn II. Pacific herring, Clupea pallasi, recruitment in the Bering Sea and north-east Pacific Ocean, II: relationships to environmental variables and implications for forecasting. Fisheries Oceanography, 9 (4):300–315, December 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Williams:2000:PHCa

[WQI00] Erik H. Williams, Terrance J. Quinn, and Ii. Pacific herring, Clupea pallasi, recruitment in the Bering Sea and north-east Pacific Ocean, i: relationships among different populations.

Fisheries Oceanography, 9(4):285–299, December 2000. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Waluda:2001:RSM

[WRTP01] C. M. Waluda, P. G. Rodhouse, P. N. Trathan, and G. J. Pierce. Remotely sensed mesoscale oceanography and the distribution of Illex argentinus in the South Atlantic. Fisheries Oceanography, 10(2):207–216, June 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Weitkamp:2008:FHM

[WS08] Laurie A. Weitkamp and Molly V. Sturdevant. Food habits and marine survival of juvenile Chinook and Coho salmon from marine waters of Southeast Alaska. Fisheries Oceanography, 17(5):380–395, September 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Winter:2005:ELS

[WSC05] Andreas Winter, Gordon Swartzman, and Lorenzo Ciannelli. Early- to late-summer population growth and prey consumption by age-0 pollock (Theragra chalcogramma), in two years of contrasting pollock abundance near the Pribilof Islands, Bering Sea. Fisheries Oceanography, 14(4):307–320, July 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Williams:2014:IOF

[WSF^{+}14] John G. Williams, Steven G. Smith, Jeffrey K. Fryer, Mark D. Scheuerell, William D. Muir, Tom A. Flagg, Richard W. Zabel, John W. Ferguson, and Edmundo Casillas. Influence of ocean and freshwater conditions on Columbia River sockeye salmon Oncorhynchus nerka adult return rates. Fisheries Oceanography, 23(3):210–224, May 2014. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wilson:2007:APA

[WSP^{+}07] S. G. Wilson, B. S. Stewart, J. J. Polovina, M. G. Meekan, J. D. Stevens, and B. Galuardi. Accuracy and precision of archival tag data: a multiple-tagging study conducted on a whale shark (Rhincodon typus) in the Indian Ocean. Fisheries Oceanography, 16(6):547–554, November 2007. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Wagawa:2016:RBC

[WTK^{+}16] Taku Wagawa, Tsuyoshi Tamate, Hiroshi Kuroda, Shin-Ichi Ito, Shigeho Kakehi, Takeshi Yamanome, and Takuya Kodama. Relationship between coastal water properties and adult return of chum salmon (Oncorhynchus keta) along the Sanriku coast, Japan. Fisheries Oceanography, 25(6):598– 609, November 2016. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Waluda:2004:SSH

[WTR04] C. M. Waluda, P. N. Trathan, and P. G. Rodhouse. Synchronicity in southern hemisphere squid stocks and the influence of the Southern Oscillation and trans polar index. Fisheries Oceanography, 13(4):255–266, July 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Welch:2000:TSR

[WWSE00] D. W. Welch, B. R. Ward, B. D. Smith, and J. P. Eveson. Temporal and spatial responses of British Columbia steelhead (Oncorhynchus mykiss) populations to ocean climate shifts. Fisheries Oceanography, 9(1):17–32, March 2000. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Watanabe:1997:VSG

[WZK97] Yoshiro Watanabe, Hiromu Zenitani, and Ryo Kimura. Variations in spawning ground area and egg density of the Japanese sardine in Pacific coastal and oceanic waters. Fisheries Oceanography, 6(1):35–40, March 1997. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Watanabe:1998:NCC

[WZK^{+}98] Yoshiro Watanabe, Hiromu Zenitani, Ryo Kimura, Chikako Sato, Yutaka Okumura, Hiroya Sugisaki, and Yoshioki Oozeki. Naupliar copepod concentrations in the spawning grounds of Japanese sardine, Sardinops melanostictus, along the Kuroshio Current. Fisheries Oceanography, 7(2):101– 109, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xu:2009:OSS

[XB09] Cailin Xu and Mark S. Boyce. Oil sardine (Sardinella longiceps) off the Malabar Coast: density dependence and environmental effects. Fisheries Oceanography, 18(5):359–370, September 2009. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xu:2020:CES

[XDP^{+}20] Yi Xu, A. Scott Decker, Charles K. Parken, Lynda M. Ritchie, David A. Patterson, and Caihong Fu. Climate effects on size-at-age and growth rate of Chinook salmon (Oncorhynchus tshawytscha) in the Fraser River, Canada. Fisheries Oceanography, 29(5):381–395, September 2020. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xie:1995:GDW

[XH95] Liusen Xie and William W. Hsieh. The global distribution of wind-induced upwelling. Fisheries Oceanography, 4(1):52– 67, March 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xu:2018:EUG

[XMH^{+}18] Haikun Xu, Timothy J. Miller, Sultan Hameed, Larry A. Alade, and Janet A. Nye. Evaluating the utility of the Gulf Stream Index for predicting recruitment of Southern New England–Mid Atlantic yellowtail flounder. Fisheries Oceanography, 27(1):85–95, January 2018. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xia:2023:ITD

[XMW^{+}23] Wentong Xia, Zhongbo Miao, Sai Wang, Kai Chen, Yinglong Liu, and Songguang Xie. Influence of tidal and diurnal rhythms on fish assemblages in the surf zone of sandy beaches. Fisheries Oceanography, 32(5):448–460, September 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xavier:2004:FEI

[XTC^{+}04] J. C. Xavier, P. N. Trathan, J. P. Croxall, A. G. Wood, G. Podestá, and P. G. Rodhouse. Foraging ecology and interactions with fisheries of wandering albatrosses (Diomedea exulans) breeding at South Georgia. Fisheries Oceanography, 13(5):324–344, September 2004. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Xu:2023:SDE

[XWL^{+}23] Min Xu, Yihe Wang, Zunlei Liu, Yong Liu, Yi Zhang, Linlin Yang, Fei Wang, Hui Wu, and Jiahua Cheng. Seasonal distribution of the early life stages of the small yellow croaker (Larimichthys polyactis) and its dynamic controls adjacent to the Changjiang River Estuary. Fisheries Oceanography, 32 (4):390–404, July 2023. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yamamura:2004:TMW

[Yam04] Orio Yamamura. Trophodynamic modeling of walleye pollock (Theragra chalcogramma) in the Doto area, northern Japan: model description and baseline simulations. Fisheries Oceanography, 13(S1):138–154, December 2004. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yamaguchi:2018:USA

[YAM^{+}18] T. Yamaguchi, T. Aketagawa, M. Miyamoto, N. Hirose, and M. Matsuyama. The use of statolith analyses and particletracking experiments to reveal the migratory route of the swordtip squid (Uroteuthis edulis) caught on the Pacific side of Japan. Fisheries Oceanography, 27(6):517–524, November 2018. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yasumiishi:2015:CSO

[YCH^{+}15] Ellen M. Yasumiishi, Keith R. Criddle, Nicola Hillgruber, Franz J. Mueter, and John H. Helle. Chum salmon (Oncorhynchus keta) growth and temperature indices as indicators of the year–class strength of age-1 walleye pollock (Gadus chalcogrammus) in the eastern Bering Sea. Fisheries Oceanography, 24(3):242–256, May 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yoshiki:2015:NSW

[YCS^{+}15] Tomoko M. Yoshiki, Sanae Chiba, Yuka Sasaki, Hiroya Sugisaki, Tadafumi Ichikawa, and Sonia Batten. Northerly shift of warm-water copepods in the western subarctic North Pacific: Continuous plankton recorder samples (2001–2013). Fisheries Oceanography, 24(5):414–429, September 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yang:2019:HBA

[YCS^{+}19] Qiong Yang, Edward D. Cokelet, Phyllis J. Stabeno, Lingbo Li, Anne B. Hollowed, Wayne A. Palsson, Nicholas A. Bond, and Steven J. Barbeaux. How “The Blob” affected groundfish distributions in the Gulf of Alaska. Fisheries Oceanography, 28(4):434–453, July 2019. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yokouchi:2022:LOE

[YIT^{+}22] Katsumi Yokouchi, Hiroshi Ito, Mai Togawa, Kenichi Ueda, and Shigeho Kakehi. Larval occurrence and environmental factors associated with spawning of Pacific oyster Crassostrea gigas in Matsushima Bay, Japan. Fisheries Oceanography, 31(6):641–652, November 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yasuda:1996:LEF

[YK96] Ichiro Yasuda and Daiji Kitagawa. Locations of early fishing grounds of saury in the northwestern Pacific. Fisheries Oceanography, 5(1):63–69, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yonezaki:2008:DCD

[YKB08] Shiroh Yonezaki, Masashi Kiyota, and Norihisa Baba. Decadal changes in the diet of northern fur seal (Callorhinus ursinus) migrating off the Pacific coast of northeastern Japan. Fisheries Oceanography, 17(3):231–238, May 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yasuda:2021:COA

[YKH^{+}21] Tohya Yasuda, Satoshi Kitajima, Akira Hayashi, Motomitsu Takahashi, and Masa aki Fukuwaka. Cold offshore area provides a favorable feeding ground with lipid-rich foods for juvenile Japanese sardine. Fisheries Oceanography, 30(5):455– 470, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yoshimori:1998:MMP

[YKI98] Akira Yoshimori, Tokihiro Kono, and Hitoshi Iizumi. Mathematical models of population dynamics of the kelp Laminaria religiosa, with emphasis on temperature dependence. Fisheries Oceanography, 7(2):136–146, July 1998. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Youcef:2013:SDG

[YLA13] Wahiba Ait Youcef, Yvan Lambert, and Celine Audet. Spatial distribution of Greenland halibut Reinhardtius hippoglossoides in relation to abundance and hypoxia in the estuary and Gulf of St. Lawrence. Fisheries Oceanography, 22(1): 41–60, January 2013. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Ye:1999:DTS

[YMB99] Yimin Ye, H. M. A. Mohammed, and J. M. Bishop. Depth, temperature and salinity preferences of newly recruited penaeid shrimps in Kuwait waters. Fisheries Oceanography, 8 (2):128–138, June 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yemane:2015:STT

[YMK^{+}15] Dawit Yemane, Samuel K. Mafwila, John Kathena, Silvi E. Nsiangango, and Steve P. Kirkman. Spatio-temporal trends in diversity of demersal fish species in the Benguela Current large marine ecosystem region. Fisheries Oceanography, 24 (S1):102–121, March 2015. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yatsu:2021:CEE

[YOIW21] Akihiko Yatsu, Hiroshi Okamura, Taro Ichii, and Kazuyoshi Watanabe. Clarifying the effects of environmental factors and fishing on abundance variability of Pacific saury (Cololabis saira) in the western North Pacific Ocean during 1982– 2018. Fisheries Oceanography, 30(2):194–204, March 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yano:2017:SDA

[YOK^{+}17] Toshikazu Yano, Seiji Ohshimo, Minoru Kanaiwa, Tsutomu Hattori, Masa aki Fukuwaka, Toru Nagasawa, and Sho Tanaka. Spatial distribution analysis of the North Pacific spiny dogfish, Squalus suckleyi, in the North Pacific using generalized additive models. Fisheries Oceanography, 26(6): 668–679, November 2017. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yamashita:2000:RCE

[YOY00] Yoh Yamashita, Tsuguo Otake, and Hideaki Yamada. Relative contributions from exposed inshore and estuarine nursery grounds to the recruitment of stone flounder, Platichthys bicoloratus, estimated using otolith Sr : Ca ratios. Fisheries Oceanography, 9(4):316–327, December 2000. CODEN FIO-CEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yamamoto:2020:CCS

[YOYK20] Masayuki Yamamoto, Hiroaki Omi, Naotaka Yasue, and Akihide Kasai. Correlation of changes in seasonal distribution and catch of red sea bream Pagrus major with winter temperature in the eastern Seto Inland Sea, Japan (1972–2010). Fisheries Oceanography, 29(1):1–9, January 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yasuda:1999:IVJ

[YSW^{+}99] Ichiro Yasuda, Hiroya Sugisaki, Yoshiro Watanabe, Sho-Shiro Minobe, and Yoshioki Oozeki. Interdecadal variations in Japanese sardine and ocean/climate. Fisheries Oceanography, 8(1):18–24, March 1999. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yanagi:1995:RSS

[YTIS95] Tetsuo Yanagi, Hideshi Tsukamoto, Shujo Igawa, and Kohji Shiota. Recruitment strategy of swimming crab, Portunus trituberculatus, in Hiuchi-Nada, Japan. Fisheries Oceanography, 4(3):217–229, September 1995. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yamashita:1996:TSM

[YTY96] Yoh Yamashita, Yoshinari Tsuruta, and Hideaki Yamada. Transport and settlement mechanisms of larval stone flounder, Kareius bicoloratus, into nursery grounds. Fisheries Oceanography, 5(3–4):194–204, September 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yasuda:1994:RBO

[YW94] Ichiro Yasuda and Yoshiro Watanabe. On the relationship between the Oyashio front and saury fishing grounds in the north-western Pacific: a forecasting method for fishing ground locations. Fisheries Oceanography, 3(3):172–181, September 1994. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yasuda:2007:CVK

[YW07] Ichiro Yasuda and Tomowo Watanabe. Chlorophyll-a variation in the Kuroshio Extension revealed with a mixed-layer tracking float: implication on the long-term change of Pacific saury (Cololabis saira). Fisheries Oceanography, 16(5):482– 488, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yatsu:2005:EER

[YWI^{+}05] Akihiko Yatsu, Tomowo Watanabe, Minoru Ishida, Hiroya Sugisaki, and Larry D. Jacobson. Environmental effects on recruitment and productivity of Japanese sardine Sardinops melanostictus and chub mackerel Scomber japonicus with recommendations for management. Fisheries Oceanography, 14(4):263–278, July 2005. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Yatsu:2000:IVS

[YWM^{+}00] Akihiko Yatsu, Tomowo Watanabe, Junta Mori, Kazuya Nagasawa, Yukimasa Ishida, Toshimi Meguro, Yoshihiko Kamei, and Yasunori Sakurai. Interannual variability in stock abundance of the neon flying squid, Ommastrephes bartramii, in the North Pacific Ocean during 1979–1998: impact of driftnet fishing and oceanographic conditions. Fisheries Oceanography, 9(2):163–170, June 2000. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zaitsev:1992:RCT

[Zai92] Yu. P. Zaitsev. Recent changes in the trophic structure of the Black Sea. Fisheries Oceanography, 1(2):180–189, June 1992. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zamon:2001:SPS

[Zam01] Jeannette E. Zamon. Seal predation on salmon and forage fish schools as a function of tidal currents in the San Juan Islands, Washington, USA. Fisheries Oceanography, 10(4): 353–366, December 2001. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhao:2003:RSY

[ZHL^{+}03] X. Zhao, J. Hamre, F. Li, X. Jin, and Q. Tang. Recruitment, sustainable yield and possible ecological consequences of the sharp decline of the anchovy (Engraulis japonicus) stock in the Yellow Sea in the 1990s. Fisheries Oceanography, 12 (4–5):495–501, September 2003. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhang:2014:SRB

[ZHT14] Zane Zhang, John Holmes, and Steven L. H. Teo. A study on relationships between large-scale climate indices and estimates of North Pacific albacore tuna productivity. Fisheries Oceanography, 23(5):409–416, September 2014. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhou:2020:EMO

[ZHX^{+}20] Cheng Zhou, Pingguo He, Liuxiong Xu, Pascal Bach, Xuefang Wang, Rong Wan, Hao Tang, and Yu Zhang. The effects of mesoscale oceanographic structures and ambient conditions on the catch of albacore tuna in the South Pacific longline fishery. Fisheries Oceanography, 29(3):238–251, May 2020. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zang:2022:MAS

[ZJH^{+}22] Zhengchen Zang, Rubao Ji, Deborah R. Hart, Changsheng Chen, Liuzhi Zhao, and Cabell S. Davis. Modeling Atlantic sea scallop (Placopecten magellanicus) scope for growth on the Northeast U.S. Shelf. Fisheries Oceanography, 31(3):271– 290, May 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zenitani:2007:RBD

[ZKT07] Hiromu Zenitani, Naoaki Kono, and Youichi Tsukamoto. Relationship between daily survival rates of larval Japanese anchovy (Engraulis japonicus) and concentrations of copepod nauplii in the Seto Inland Sea, Japan. Fisheries Oceanography, 16(5):473–478, September 2007. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zabel:2011:IBC

[ZLTM11] Richard W. Zabel, Phillip S. Levin, Nick Tolimieri, and Nathan J. Mantua. Interactions between climate and population density in the episodic recruitment of bocaccio, Sebastes paucispinis, a Pacific rockfish. Fisheries Oceanography, 20 (4):294–304, July 2011. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zenitani:1996:SGS

[ZNI96] Hiromu Zenitani, Kaoru Nakata, and Denzo Inagake. Survival and growth of sardine larvae in the offshore side of the Kuroshio. Fisheries Oceanography, 5(1):56–62, March 1996. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zimmermann:2021:FPA

[ZP21a] Mark Zimmermann and Megan M. Prescott. False pass, Alaska: Significant changes in depth and shoreline in the historic time period. Fisheries Oceanography, 30(3):264–279, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zimmermann:2021:PAI

[ZP21b] Mark Zimmermann and Megan M. Prescott. Passes of the Aleutian Islands: First detailed description. Fisheries Oceanography, 30(3):280–299, May 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zainuddin:2008:ATA

[ZSS08] Mukti Zainuddin, Katsuya Saitoh, and Sei-Ichi Saitoh. Albacore (Thunnus alalunga) fishing ground in relation to oceanographic conditions in the western North Pacific Ocean using remotely sensed satellite data. Fisheries Oceanography, 17 (2):61–73, March 2008. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhang:2021:CSH

[ZSY^{+}21] Tianjiao Zhang, Liming Song, Hongchun Yuan, Bo Song, and Narcisse Ebango Ngando. A comparative study on habitat models for adult bigeye tuna in the Indian Ocean based on gridded tuna longline fishery data. Fisheries Oceanography, 30(5):584–607, September 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zorica:2013:ECC

[ZVKŠ13] Barbara Zorica, Ivica VilibiĆ, Vanja ČIkeŠ KeČ, and Jadranka ŠEpić. Environmental conditions conducive to anchovy (Engraulis encrasicolus) spawning in the Adriatic Sea. Fisheries Oceanography, 22(1):32–40, January 2013. CO-DEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhou:2021:SVB

[ZWC^{+}21] Cheng Zhou, Rong Wan, Jie Cao, Liuxiong Xu, Xuefang Wang, and Jiangfeng Zhu. Spatial variability of bigeye tuna habitat in the Pacific Ocean: Hindcast from a refined ecological niche model. Fisheries Oceanography, 30(1):23–37, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhang:2021:ABE

[ZWL21] Chun-Ling Zhang, Zhen-Feng Wang, and Yu Liu. An argo-based experiment providing near-real-time subsurface oceanic environmental information for fishery data. Fisheries Oceanography, 30(1):85–98, January 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhang:2022:SOJ

[ZYT^{+}22] Wenchao Zhang, Zhenjiang Ye, Yongjun Tian, Haiqing Yu, Shuyang Ma, Peilong Ju, and Yoshiro Watanabe. Spawning overlap of Japanese anchovy Engraulis japonicus and Japanese Spanish mackerel Scomberomorus niphonius in the coastal Yellow Sea: a prey–predator interaction. Fisheries Oceanography, 31(4):456–469, July 2022. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zhang:2021:SSJ

[ZYY^{+}21] Wenchao Zhang, Haiqing Yu, Zhenjiang Ye, Yongjun Tian, Yang Liu, Jianchao Li, Qinwang Xing, and Yiqian Jiang. Spawning strategy of Japanese anchovy Engraulis japonicus in the coastal Yellow Sea: Choice and dynamics. Fisheries Oceanography, 30(4):366–381, July 2021. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).

Zaporozhec:1993:PHR

[ZZ93] O. M. Zaporozhec and G. V. Zaporozhec. Preparation of hatchery — reared chum fry for life at sea: osmoregulation dynamics. Fisheries Oceanography, 2(2):91–96, June 1993. CODEN FIOCEN. ISSN 1054-6006 (print), 1365-2419 (electronic).